A Complete Bibliography of Publications in International Journal of Information Security

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

16 September 2024 Version 1.36

Title word cross-reference

(t, n) [QDW09]. 2^{w} [Bae10]. _{D} [MGAS24]. K [BHZ^{+}21, BDD01, MU18, RMSCR19, Rov23, TNA23]. LU [MARK20]. n [CC12]. O(n) [DYDW10]. t [BDD01].

-anonymity [TNA23]. -anonymization [MU18, RMSCR19]. -ary [Bae10]. -based [Yan21]. -database [BDD01]. -Diffie [CC12]. -means [Rov23]. -private [BDD01].

.NET [KKKV07].

/SSL [BJ16].

1 [KJS17]. 100G [GDA22]. 11770 [CH16].

128-bit [CDP22]. 19 [PGMPPC22].

2.0 [AMLH18, MP22]. 2002 [ACM05]. 2003 [BM05, Sne05]. 2004 [DSY06]. 2012 [SKK^{+}17]. 2013 [BJ15]. 25022 [MD23b].

3 [ABM^{+}12]. 3GPP [EWR^{+}09]. 3T’s [RB23].

4.0 [SDG22].

5 [DWU^{+}23]. 5G

[Alg22, KIP22, PH24b, PH24a, PCSA24].

61850 [AEHS24]. 6LoWPAN [BBN24].

802.15.6 [KDM22]. 88 [vOLW05].

95 [ADHN24].

= [TGNA22].

AAIA [YYH^{+}23]. AAnA [SK14]. abnormal [JJJ21]. absence [AvO13]. absolute [AvO13]. abstract [BLM11, DM07, MS14]. abstract-based [BLM11]. abstraction [MLCS16]. Abuse [FGH24]. Abuse-freeness [FGH24]. abused [CYA^{+}18]. AC [MdMF22]. ACAC [MGAS24]. Acceptable [BB04b]. Access [Lop18, SAL17, VHRRMG24, AMZ22, ACF17, AFA^{+}23, ACBC^{+}15, BAB23, BLM11, CF03, CK08, CZ06, GGJ22, HSMY12, IOU^{+}21, JSMG18a, KKR^{+}24, KAC16, KAC17, Kud02, LD17, LRB^{+}10, MG23, MS15, PMDS23, Pen12, PMPGMLLM12, RD16, RSMT24, RT23, SRD^{+}21]. accessibility [SHA20]. Accessible [KB23]. account [BRS06]. Accountable [GGJ22]. accounts [ASN^{+}16]. accreditation [DFBJR18]. Accumulable [SEXY18]. accumulator [KYH18]. accumulators [CHKO12, HRMM20, JCL^{+}18]. accuracy [AIJM24, QLOW09]. accurate [CYA^{+}18, SSD14]. ACE [RSMT24]. achievable [Pla09]. achieve [Pen13]. Achieving [AICC18, IZS08, RSD19]. ACM [BJ15]. Acnn [XT24]. acoustic [CBRY20, HS15]. acquisition [YP12]. across [IKS22]. action [JG15, PKHS23]. action-mask [PKHS23]. actionable [SBG22]. Active [LCPD14, CBRY20, CBC08, DGZFGH13, LTC23]. activities [JJJ21, OBH^{+}20, YKP22]. Activity [Les24, AAZAA23, MGAS24]. Activity-based [Les24]. actual [MTSH18]. actually [BM11]. Ad [LH23, AHT24, Gol12, MS11, SF17, SSSJ24]. ad-hoc [AHT24, MS11]. Adapting [GLP03, Sen14]. Adaptive [GZH^{+}23, KKR^{+}24, PPSS13, MdSC^{+}15].

Adding [CON09]. Additional [YAY^{+}21]. address [ELPB24]. addressing [ASA23]. administration [KK22]. adoption [UGP24, ZJS22]. ADRGAN [PP24b]. ADroid [RHGTSC17]. advanced [AEHS24, PGMLK^{+}13, TS20, TMP13]. Adversarial [BMFS24, KCC^{+}23, LSR^{+}23, AZM24, AVM23, BV24, CSM^{+}24, GZH^{+}23, RC24, TCE23, XZ24, ZTV23]. adversary [ORK23, QDW09]. adverse [SGC22]. advertising [KOSU16, UBK23]. AERAS [FDS^{+}24]. affects [CFBvO09]. African [ETRLG24]. After [PSTS20, NABGJS^{+}24]. against [ANN23, AP22, BHL^{+}21, DdP13, ETRLG24, EMRN17, FTS^{+}20, GYL^{+}07, GI19, KK17, LSR^{+}23, MLYL20, MS11, MYLZ14, MTD^{+}24, Nui12, SK06, SB22, SAH22, TSMH19, VSR15, WWZ^{+}23, YYH^{+}23, BZ20, ZRJ14]. agent [LV10, PKHS23, PDM20, TGNA22, EZ22]. agent-based [PDM20]. agents [SEZ24]. Aggregate [CL13, RSH^{+}24a, RSH^{+}24b]. aggregation [GLMS19, GKS19, LL21, YYH^{+}23]. aggregation-based [GLMS19]. agile [SHOL23]. Agnostic [CMH^{+}24]. agreement [AAV22, CCS07, CL09, CHC24, FGS12, GNS14, KDM22, TK24, ZWQ^{+}17]. ahead [GBC24]. AI [BAG^{+}24, HAS^{+}24, QQK^{+}24, ZP23, ZKP24]. AI-assisted [ZP23]. AI-driven [ZKP24]. AI-enabled [BAG^{+}24]. AiCEF [ZP23]. aided [NT20]. AIHGAT [WGH23]. Alambic [ABFO08]. alert [MdMF22, SGJ19]. alerts [STD21]. algebraic [DS07, KM10, SSVC16]. Algorithm [AP22, JMV01, AT22, Bae10, BHZ^{+}21, ETRLG24, GS15, MB23, MU18, PV22, SAM^{+}24, SSSJ24, SKK^{+}17, TLX09, Zen22]. algorithms [ASA23, BEPL^{+}17, KU16, ML14, WT21]. alignment [KSM10]. all-encompassing [EHM15]. All-or-Nothing [MTW^{+}14]. allocation [SGJ19]. allotROPism [LNX22]. along [VH19]. Alpha [BV24]. alternative [Les24]. AMACs [CL13]. ambients [TZH04]. among [AHX^{+}23]. analyses [ACCS24, TS22]. Analysing [HL04]. Analysis [AMRR17, BR20, CG14, DBMS10, GLP03, SSFB15, Zha23, ZZW^{+}10, AGZA22, ABCC08, AZS24, ASN^{+}16, AAB20, AC08, AAG24, Asf24, ASFG24, BGKZ12, BEPL^{+}17, Bel10, BDMM19, BR23, BBR18, BFT08, BNN04, CPPK15, CF07, DFF^{+}16, DRPW12, FN19, dSFK19, GAdFGM23, GLMS^{+}04, GKBS12, GBC24, HFA24, HLKI15, HS24, IMI18, IOU^{+}21, Inc24, IDHRPCMP15, KW15, KAC17, KKK22a, KA18, LRB^{+}10, MWZ06, MTW^{+}14, MD23b, MOA^{+}24a, NRC15, NAQ24, Oft24, OO20, OT06, PCT22, PDB11, QLOW09, RSV23, SK16, SS05a, SB09, SSM^{+}20, SSL22a, SSE^{+}15, SDG22, SPDR17, SRSM23, Vaj16, VdWZ14, WYL^{+}12, WGH23, XSA13, YWW22, YA22, YL19, ZHZ22]. analysis-based [BR23]. analytics [SCO21]. analyzer [PAI^{+}23]. Analyzing [BJ16, GMLM23, RRI^{+}19, vOLW05, ABK22, SGLC19]. anchor [BB04b]. AND-gate [JSMG18b]. Android [LSV^{+}23, ACS21, AAB20, BSK^{+}23, CAS22, EZLC21, GPS17, IS23, IOU^{+}21, IKS22, IDHRPCMP15, KBY22, KA18, LMMS17, LWL^{+}21, LSG22, LLG22, MGRR19, MS15, RHGTSC17, SA24, STW^{+}23]. Anomalous [AZ22]. Anomaly [Alm24, DGF^{+}17, KCM^{+}15, MD23a, ME23, RHGTSC17, SAM^{+}24, SKH^{+}23]. anomaly-based [RHGTSC17, SAM^{+}24]. Anonymity [GP23, LSWW14, Man21, BSK^{+}20, DFBJR18, HLS18, KNM^{+}24, SS05a, TNA23]. Anonymization [CMH^{+}24, AT22, CR20, HN14, MU18, RMSCR19]. anonymizing [ZO13]. Anonymous [BFG^{+}13, LcSCL^{+}18, SK14, ACHO13, ABFL12, BSV22, BCL09, CPPK15, CHC24, KLPL21, KWCK19, KCB17, Les24, RA24, SF17, SDR20]. ANS [CDP22]. ANS-based [CDP22]. answer [RRI^{+}19, WZ07]. anti [AFF24, GKBS12, KKK22b, AHC^{+}21]. Anti-BlUFf [AHC^{+}21]. anti-malware [KKK22b]. anti-phishing [AFF24]. anti-SPIT [GKBS12]. antiviruses [ASAAS15]. any [DdP13]. AOMDV [MG19]. API [IKS22, You06]. app [CAS22, IKS22]. apples [BBR18]. applicable [QDW^{+}15]. application [ACS21, ANN23, BB22, BEd23, DGF^{+}17, KSB23, Pen12, QQK^{+}24, RSK^{+}24, RSV23, Roe11a, Roe11b, SPM13, VdWZ14, ZZW^{+}10]. application-layer [DGF^{+}17]. Applications [Gri06, AAB20, BCA^{+}10, BAG^{+}24, BR23, BNTW12, BAC24, DJN10, DTK^{+}18, GSS10, HSMY12, HZL^{+}17, KGG09, LS23, SCO21, SNX19, TSZ22, VD24, WYL^{+}12, WCS20, vORM06]. Applied [BJ15]. Applying [MI22]. approach [AIJM24, AuHD^{+}23, AZ22, AŠC24, AV17, ASFG24, BHL^{+}21, BSBK24, CFG17, CMS10, BABB16, DS07, EZLC21, FDS^{+}24, Fra18, GA23, HBH12, JBK21, Kak24, KAC16, KAC17, KCB20, KKK17, KDYS19, LVK18, LTC23, LH23, MGV17, MARK20, MMS16, MLCQ21, MG23, ME23, MSGCDPSS18, MYLZ14, MOA^{+}24b, MP22, NAQ24, NNL20, NA14, PM23, PMW24, PNG^{+}20, SC24, SRK^{+}20, SGSS23, TWP08, VSR15, ZKP^{+}23]. approach-based [SC24]. approaches [JJJ21, JJQ24, LHS^{+}24, PHE^{+}24, SBS23, SAH22, ZO13, ZSN23]. approximate [Rov23]. apps [CSL^{+}23, CSS^{+}24, EZLC21, GPS17, GL24]. APT [AEHS24, SGSS23]. APTs [GMdFLR23]. Arabia [Asf24]. arbitrary [XT24]. archetypes [SEZ24]. architecture [AKG16, BV24, BS22, EHM15, EWR^{+}09, FZ21, FPP^{+}24, LV10, MSP^{+}13, MdAN23, MSKS20, SRSM23, TCE23, Zha23]. architectures [AFA^{+}23, SSL22a, WW07]. area [LCL16, MPG21, RA24, XT24]. areas [BS21]. arguments [ABM^{+}12]. arithmetic [ABB17, KW15]. ARITO [SSD14]. ARM [BZ20]. ARP [MTD^{+}24]. art [AK24b, TDGL23]. artificial [BAC24]. ary [Bae10]. ASICS [BCF^{+}17]. aspects [AICC18, GL24]. assembly [WGH23, Zha23]. assessment [AK23, EAM^{+}23, FJZZ24, HXTP23, JBK21, KB23, NBA^{+}21, Sep23, SIA^{+}24, SFB^{+}24, WHS18, ZTG22]. assessments [HFA24]. Asset [SSR22]. assigned [JTV19]. assisted [CTN24, DYDW10, FAH24, LWL23, PDM20, VPI15, ZP23]. association [OBH^{+}20, VH19]. assumption [HIST09]. assurance [ABN14, LVK18, SKY23]. asymmetric [ZWQ^{+}17]. Ate [ZZH08]. ATNA [ACBC^{+}15]. ATT&CK [YZ24]. attack [AZ19, ASKG24, AEHS24, AYHK18, ASZ^{+}23, Alh24, BB22, BMFS24, BBN24, BHZ^{+}21, CSL^{+}23, DSB19, DRPW12, ELPB24, FAMMZ23, dSFK19, GMS03, GCT24, GZH^{+}23, KAK23, KY24, Lu09, MVA24b, MAZS24, ML14, NABGJS^{+}24, PHS22, PMW24, RSV23, SLGP23, SS05a, SBD23, SSD14, SSL22b, SMV24, SSV22, SDG22, TCT22, WWZ^{+}23, WKWS24, XYZT24, YAY^{+}21, YYH^{+}23, ZXZ^{+}11]. attacker [RMPADF13]. Attacking [SGE02]. Attacks [AKZM20, MLC23, AHT24, ASAAS15, AFF24, AP22, BHL^{+}21, BRS06, BDG23, CK23a, CK23b, CKKK23, CBRY20, BABB16, CSM^{+}24, DGF^{+}17, EB24, ETRLG24, EMRN17, FTS^{+}20, HAAI24, HS15, HFA24, HXTP23, Hub12, JK22, KSB23, KM10, KDYS19, LLWY09, LSR^{+}23, LH23, MVA24a, MLCQ21, MLYL20, MS11, MTD^{+}24, NAQ24, NBA^{+}21, ORK23, Pen11, PSTS20, PNG^{+}20, PPL15, RC24, SGLC19, SB22, SSVC16, TTS^{+}06, VSR15, XT24, XCW^{+}12, YNC22, ZTV23]. Attention [ACS21]. attestation [BFG^{+}13, BCL09, CGL^{+}11, KLPL21]. attitude [AHX^{+}23]. Attribute [KMW24, SRD^{+}21, TNA23, BAB23, GGJ22, JSMG18a, JSMG18b, KLZ^{+}21, NCBG23, ORK23, PMDS23, QLZH15, QDW^{+}15, RD16, ZLZL20]. Attribute-based [SRD^{+}21, BAB23, GGJ22, JSMG18a, JSMG18b, KLZ^{+}21, PMDS23, QLZH15, QDW^{+}15, RD16, ZLZL20]. attributes [JSMG18a, KLZ^{+}21]. auctions [Bra06]. audio [dSFK19]. Audit [CCD^{+}07, BS05]. Audit-based [CCD^{+}07]. auditing [FAH24, SXZC20, TMvM24, WMS^{+}19]. AUTH [RG13]. Authenticated [HM22, AAV22, BBR18, BCF^{+}17, CHMS21, IMI18, KDM22, Lin15, MPS10, Ust11, YLL^{+}18, YRW14, ZWQ^{+}17]. Authenticating [CF07, AZ22]. Authentication [BAB23, DNF^{+}19, EB24, GCSÁBdSS12, AGZA22, AAMH^{+}24, BPW05b, BEY23, Bra22, BJ16, CL13, CHC24, DSB19, DFF^{+}16, EWR^{+}09, Gol12, GTM11, Hal20, HHSS24, HC10, HCN15, HL04, HKO22, IS23, IT05, KKY^{+}23, KWCK19, KML03, KB13, LSWW14, LCPD14, LSR^{+}23, MB16, ML17, MSKD16, MS09, MNC20, MD23b, MP22, PH24b, PH24a, PS17, RSH^{+}24a, RSH^{+}24b, RA24, RG13, SK14, Smi04, SCO21, SDR20, TWP08, TK24, VHT09, WLLW14, YWW22]. Authentication-enabled [BAB23]. authentications [HZL^{+}17]. authenticators [SXZC20]. authorities [JJQ24, LMMO04]. authority [CON09, QLZH15]. authorization [BZV05, KLMM09, MP22, RV03, SSFB15, SK14, WZ07]. authorized [ZZG19]. autoencoder [OPK24]. automata [BCL13, DLR15, LBW05]. Automated [BB22, GLMS^{+}04, JG15, EFB^{+}24, GH05, HBF^{+}24, PP24b, SSM^{+}20, dAKdG10]. Automatic [SDG22, ACMV15, CAS22, KKK22a, PLA^{+}21, ZLGZ19]. automatically [KM07, XCW^{+}12]. Automating [AKA24, SNX19]. automation [CSL^{+}23]. autoML [PAI^{+}23]. autonomous [AG23]. AutoPKI [HBF^{+}24]. availability [Bel10]. averaging [WKWS24]. AVL [RBD02]. AVOA [ETRLG24]. aware [Alg22, DK24, DGF^{+}17, KKR^{+}24, KJG^{+}11, MBRPS18, MMF^{+}24, RSPMB16, SRK^{+}20, TGNA22, Vaj16, YKP22]. awareness [AYM24, MPS14, MOA^{+}24a, Oft24]. AWS [AMZ22].

BA [MVA24a]. BA-flag [MVA24a]. back [KNL16]. bad [HAS^{+}24]. balance [MYLZ14]. BAPRP [LH23]. BAR [KCB17]. base [QLH^{+}24]. Based [LLW^{+}16, AK23, AYHK18, ACB14, ASZ^{+}23, AuHD^{+}23, ADHN24, Alm24, EOC^{+}24, AZ22, AK24a, AKA24, AFF24, AC08, AAG24, BV24, BHKM23, BMFS24, BFP03, BFPP07, BR23, BS22, BAB23, BBR18, BLM11, CDP22, CCD^{+}07, CSC^{+}23, CMS10, CK24, CCS07, CHZ16, CKPM24, CFBvO09, BABB16, CSL^{+}23, CSS^{+}24, CG23, CK08, Dan07, Des09, EZLC21, EMRN17, FZ21, FAMMZ23, FGH24, FGS12, GCH^{+}19, GPS17, GLMS19, GGJ22, GDA22, GBDJ14, GMSS23, GTM11, HJDC15, Ham23, HHSS24, HS09, HC10, HRL09, HKO22, INS21, IMI18, IT05, JCL^{+}18, JSMG18a, JSMG18b, JBK21, JJQ24, Kak24, KG11, KLZ^{+}21, KAK22, KBY22, KKY^{+}23, KKK17, KM22, KML03, KLMM09, Kud02, KB22, KNL16, LHS^{+}24, LMG17, LKH09, LLH21, Les24, sLC05, LH15, LMD17, LD17, LWA21, LP11, LBZ^{+}10, LMMO04, LSR^{+}23, LSG22, LSV^{+}23, LLG22]. based [MPS10, MGV17, MPS14, MP15, MS09, MAZS24, MMS16, MKS^{+}23, MdMF22, MLYL20, MK21, MS11, MD23b, MUiH^{+}21, MS14, MOA^{+}24a, ME23, Moh23, MLC23, MB23, MRW02, MSGCDPSS18, MOA^{+}24b, MFES04, MTD^{+}24, NAM06, NSNK06, NNL20, NMBB12, OT06, PMDS23, PDM20, Pen11, PGMPPC22, PPL15, QLZH15, QLH^{+}24, QDW09, QDW^{+}15, RSH^{+}24a, RSH^{+}24b, RSK^{+}24, RA24, RD16, RCC^{+}24, RBEH15, RG13, RMSCR19, Rov23, RV03, RHGTSC17, RC24, SAM^{+}24, SM24, SC24, SJ09, SBB19, SV11, SPM13, SKLP20, SSP14, SS05a, SBD23, SRK^{+}20, SA24, SFT^{+}24, SSL22a, STW^{+}23, SK14, SRD^{+}21, SSSJ24, SS05b, SdHZ16, SCO21, SSJ22, SHW23, SAT09, SDR20, SGE02, SPDR17, TSU^{+}24, TLX09, TDRR20, TND^{+}15, Ust11, VBAG24, VGG24, VD24, VH19, WR08, WZZ^{+}24, WCS20, XT24, XCW^{+}12, XSA13, XYZT24, Yan21, YA22, YL20]. based [YSM10, ZBC23, ZGC07, ZWQ^{+}17, ZLGZ19, ZLZL20, ZTG22, ZGRS23, ZVH15, ZKP^{+}23, ZIR24, ZR24]. basic [BCJ^{+}11, PHE^{+}24]. BASPED [CTN24]. batch [Pen13, RSH^{+}24a, RSH^{+}24b]. batch-based [RSH^{+}24a, RSH^{+}24b]. BATMAN [SSSJ24]. Bayes [Sen14]. Bayesian [ETAHCR08, GBG18, SGSS23]. be [ASN^{+}16, EZLC21]. beamforming [KCC^{+}23]. bee [SS17]. beehives [SS17]. behaves [ASN^{+}16]. Behavior [BABB16, AZ22, MLCS16, XCW^{+}12]. Behavior-based [BABB16, XCW^{+}12]. behavioral [BEY23, CSC^{+}23, KLMM09, SCO21]. behavioral-based [CSC^{+}23]. Behaviour [PPL15, AHX^{+}23, IS23, LCPD14, ZTV23]. behaviours [SCY^{+}24]. belief [QLH^{+}24, SGSS23]. Bello [PHE^{+}24]. benchmark [BB22]. benefit [DRPW12, Zen22]. benefits [Rus04]. better [RAC16]. between [ACS21, Auf20, FGS12, LKH09]. beyond [PCSA24, SB22]. bibliometric [AAM23, AZS24, Dwi24]. bidimensional [KCM^{+}15]. big [CLG23]. bilayer [MLCS16]. BILSTM [NK24]. binary [ZSN23]. bindings [MSKD16]. Bio [ZZW^{+}10]. Bio-Inspired [ZZW^{+}10]. Biometric [CMH^{+}24, Pla09, BCA^{+}10, HCN15, IT05, LSR^{+}23]. biometrics [BEY23, BCA^{+}10, SCO21]. Bipartite [YOV09]. birth [VMCR23]. birthmark [XCW^{+}12]. bisection [UMN^{+}20]. bit [CDP22, GSS10, SBB19, SKLP20, YOV09]. bit-length [YOV09]. Bitcoin [AMLH18, BSK^{+}20, FTS^{+}20, ML17, PSDSNAHJ19, Sat20]. BitCracker [AB22]. BitLocker [AB22]. BitM [TCT22]. bits [BR18]. Black [DTK^{+}18, GZH^{+}23, LSR^{+}23]. Black-box [DTK^{+}18, GZH^{+}23, LSR^{+}23]. blackhole [LH23]. blank [WPD18]. BlindIdM [NA14]. BLISS [VD24]. block [CYK09, KM10, Lu09, MARK20]. Blockchain [EOC^{+}24, CTN24, CK24, MdAN23, Moh23, RSK^{+}24, SSSJ24, AKR24, AAM23, AHC^{+}21, BBG24, FAH24, FPP^{+}24, FGH24, FYF22, GGJ22, GdSdC24, HFA24, LLH21, LWA21, LWL23, MVA24a, MVA24b, MAZS24, MLYL20, MG23, PGMPPC22, PP24b, SHD^{+}24, SSV22, VD24, YSD^{+}20, YL20, ZWX20, ZLJW20]. blockchain-assisted [FAH24]. Blockchain-based [EOC^{+}24, CK24, RSK^{+}24, SSSJ24, FGH24, LLH21, LWA21, PGMPPC22, VD24, YL20]. blockchain-enabled [LWL23, ZWX20]. Blockchain-oriented [Moh23]. blockchains [GAdFGM23]. blocks [CMR06]. BLoCNet [BCGL23]. Bloom [MB16]. BlUFf [AHC^{+}21]. board [MMF^{+}24]. body [RA24]. bombing [SMV24]. bootstrapping [EWR^{+}09]. botnet [AKG16, FN19, HGH23, LHS^{+}24, NNL20]. botnets [AMLH18, FN19, KA18, LL14]. bounded [LVK18]. box [CKW19, DTK^{+}18, GZH^{+}23, LSR^{+}23]. Brazilian [MdAN23]. breaches [MRCK24]. Breaking [Dan07, RSMA19]. Bring [ACMV15]. Broadcast [GSP^{+}16, KCB17, LLW^{+}16, CL09, Man21, PPSS13]. browser [PSTS20, TSMH19, TCT22]. Browser-in-the-Middle [TCT22]. BRSIM [BP08]. BRSIM/UC [BP08]. Bu [AKT23].

Bu-Dash [AKT23]. bug [Alq24]. Building [LD07]. Business [ANN23, INS21, KAC16, KBH07]. Business-layer [ANN23]. bypassing [BV24].

C [CHC24]. C-V2X [CHC24]. C2 [HAAI24]. C2-Eye [HAAI24]. cache [TTS^{+}06]. caching [ACB14]. CADS [FAMMZ23]. CADS-ML [FAMMZ23]. CADS-ML/DL [FAMMZ23]. calculator [CR24]. calculus [BNN04]. call [CKPM24, LWL^{+}21, STW^{+}23, YAY^{+}21]. call-based [CKPM24]. calls [IKS22]. campaigns [AYM24]. Canada [Lev07]. candidates [ABM^{+}12]. can’t [JTV19]. capabilities [ASAAS15, Lop18]. capability [CL08, Fon08]. capacity [SJ09]. capture [MR03]. card [ABFL12, DMDD16, GdKGV14, MUiH^{+}21, MS14, MLM19]. card-based [MUiH^{+}21, MS14]. cardiac [NBA^{+}21]. cards [MPP14]. carried [QLOW09]. cascade [WPD18]. cascade-instantiable [WPD18]. case [BSCZ11, BDG21, CSM^{+}24, GLMS^{+}04, IS23, SKY23, SKK^{+}17, TS20, TNA23, ZD22, ZHZ22]. cases [Pla09]. cash [LCL14]. Cashing [PPAHC24]. CASSANDRA [MCD11]. cast [YYK^{+}18]. Causal [BR23]. CCA [PPSS13]. CDH [HIST09]. center [NABGJS^{+}24]. centered [GMGM21]. centralized [AM23]. centric [AFA^{+}23, CAS22]. ceremonies [MdSC^{+}15]. Certificate [MFES04, CLPP11, JJQ24, RSH^{+}24a, RSH^{+}24b, Rus04, SVKV21, Ust11]. certificate-based [Ust11]. certificate-free [CLPP11]. certificate-less [RSH^{+}24a, RSH^{+}24b]. Certificateless [RSD19, ZWQ^{+}17, CFG17, Den08, FGS12, SSP14]. certificates [PGMPPC22, ZJS22]. certification [BCF^{+}17, LMMO04]. certified [IZS08]. cFEM [MKS^{+}23]. chain [FAH24, HAAI24, MGAS24]. chained [DZW^{+}18]. chains [AHC^{+}21]. Chaintegrity [ZWX20]. challenge [LWA21, MLYL20]. challenge-based [LWA21, MLYL20]. challenged [KB23]. Challenges [TSZ22, GBC24, HC10, QQK^{+}24, RCC^{+}24, UGP24]. changes [YAY^{+}21]. channel [CBRY20, dSFK19, HS15, KDYS19, MSKD16, MS09, MTW^{+}14, SB22, VBAG24]. channels [AZM24, DHW11, KK17, KML03, MPG21]. chaos [KAK22]. chaos-based [KAK22]. chaotic [PV22, ZMS22]. characteristics [BSK^{+}23]. characterization [GMdFLR23, MSN02, NLS24, VHRRMG24]. charging [ACT23]. cheating [CG21]. check [KO02]. checker [BMV05]. checking [AC08, GKBS12, HL04, MS15, NST09, YAM^{+}15]. checks [TRT^{+}24]. checksums [GKKT10]. CHES [DSY06]. chip [KKY^{+}23]. choice [Tor20]. ChoKIFA [BHL^{+}21]. chosen [LLW^{+}16]. chosen-ciphertext [LLW^{+}16]. chrome [PSOMS22]. CIED [NBA^{+}21]. CIP [Yan21]. cipher [Lu09, TTS^{+}06]. ciphers [CYK09, IMI18, KM10]. Ciphertext [JSMG18a, JSMG18b, LMG17, LLW^{+}16, XZ21]. Ciphertext-policy [JSMG18a, JSMG18b]. ciphertexts [JSMG18b, PPSS13]. circuits [AK24a, ZXZ22]. civil [MdAN23]. CKKS [PHE^{+}24]. class [DS07, DJS^{+}24, GDA22, MLCS16]. classes [ASA23]. Classification [SAL17, AIJM24, Alq24, CBC08, GDA22, JBK21, KEA24, LTC23, LWL^{+}21, PMW24, PLA^{+}21, RSK^{+}24, ZSN23]. classifiers [PKHS23]. classify [BEPL^{+}17]. Classifying [SAH22]. click [ALPW13, CFBvO09, MBHT17, SGLC19]. click-based [CFBvO09]. clickbait [BV24]. client [ANN23, Bra22]. client-side [ANN23]. clients [SPDR17]. close [DHW11, VHRRMG24]. closed [AKA24]. Cloud [Abb13, ABN14, Ano14, BMP^{+}14, BS22, CMMPS15, CSL^{+}23, CSS^{+}24, DDX19, ELPB24, FZ21, FAMMZ23, FSG^{+}14, GGJ22, GMH14, HK19, KNL16, LZQ^{+}18, LS23, Man21, MSGCDPSS18, MOA^{+}24b, NT20, PDM20, PPL15, SKY23, UGP24, VH19, WMS^{+}19, YAM^{+}15, ZLZL20, Alg22]. cloud-based [BS22, CSL^{+}23, CSS^{+}24, FAMMZ23, KNL16]. cloud-outsourcing [ZLZL20]. clouds [AAMH^{+}24]. cluster [MKS^{+}23]. clustering [BT07, DK24, PLA^{+}21, RMSCR19, SKK^{+}17, STD21, TBGB20]. clustering-based [RMSCR19]. CNN [ETRLG24, SSL22b]. co [BMP^{+}14, NMH^{+}24, PPL15]. co-locations [NMH^{+}24]. co-residence [PPL15]. co-resident [BMP^{+}14]. code [BSK^{+}23, DMDD16, HZL^{+}17, KKKV07, LS23, ZR24]. codebook [SFT^{+}24]. codes [BHKM23, CL13, Nui12]. coding [EMRN17, TWP08]. coding-enabled [EMRN17]. coercion [CLO^{+}24]. coercion-resistant [CLO^{+}24]. Coin [KM22]. Coin-based [KM22]. Collaborative [WKWS24, EOC^{+}24, BEd23, LWA21, LWL23, MLYL20, XZ24, XYZT24]. collect [PD21]. collection [BSV22, SV11]. Collision [MS09, CHKO12, Hal20, KW15]. collision-resistant [CHKO12]. collusion [DdP13, LL21, Nui12]. collusion-resistant [LL21]. collusion-secure [Nui12]. Colored [SSFB15]. coloring [RS18]. colors [PM23]. Coloured [BKBB20, BBB20]. column [CMS10]. combinations [KAC17]. combinatorial [CMR06]. combined [CLG23]. Combining [SCO21, GMSS23, SSE^{+}15]. command [HAAI24, SNX19]. commerce [ABFO08, GLP03, ZGC07]. commercial [NVB^{+}02]. commit [ZR24]. commitment [GSS10, HCN15]. commitments [WCS20]. Commix [SNX19]. commodity [TND^{+}15]. commodity-based [TND^{+}15]. common [BSBK24, DMP13, OPK24, PHE^{+}24]. communication [CPPK15, DWU^{+}23, GBG18, KCB17, LMD17, LLG22, PBI^{+}23, VdWZ14]. communication-efficient [LMD17]. communications [FZ21, KAK22, KIP22, MG19, MB16]. comparative [ACCS24]. Comparing [BBR18, GMLM23]. comparison [AHX^{+}23, SDW23, SRD^{+}21]. comparisons [AL05]. compatible [LI07]. competency [ST24]. Complete [ABCC08, BB04a, LCL14, MSN02]. completeness [WHS18]. complex [SDG22]. compliance [ADHN24, CCD^{+}07, JJQ24, TRT^{+}24]. compliance-based [JJQ24]. Compliant [DVB02, GdSdC24, PSOMS22]. components [AV17]. composability [Vaj16]. composable [BFS^{+}13]. comprehensive [AB24, BF13, CKKK23, GCT24, GBC24, HGH23, KSB23, KC24, LRB^{+}10, Moh23, SA24, SB09]. compressing [MP16]. compression [CDP22, RSMA19]. compromised [SPDR17]. computation [ABB17, BNTW12, DDX19, PCK22a, Pen13, SZ24]. Computational [BP04, LCL16]. computationally [PMW24]. Computations [KM22, BGP07b, KU16]. compute [MCD11]. Computer [Lan01, AZS24, BCEM04, LL14, vORM06]. Computing [BJ15, Ano14, GRV24, KKR^{+}24, LI07, MOA^{+}24b, PDM20, UGP24, Pri04]. concept [Sen14]. conceptual [JGK14, TV24]. Concise [WCS20]. Concrete [BLM11, BCL09, RSD19]. Concrete- [BLM11]. concurrent [ASAAS15]. condition [ANN23]. Conditional [BDHK08, BMFS24]. conditions [BDPV14, sLC05, SGC22]. Conference [CL09]. confidential [HN14]. confidentiality [BB04a, KNL16, XZ21]. configuration [ABR16, ABCC08]. confinement [Fon08, SPM13, SS05b]. confirmation [PSDSNAHJ19]. conflict [RV03]. conjunctive [FRG19, LZQ^{+}18]. connected [MGAS24, MLCQ21, Yan21]. connection [BS22, HAAI24, LSWW14, SS05a]. connection-based [SS05a]. consensus [Kak24]. considering [HXTP23, TTS^{+}06]. constant [BGP07b, PPSS13]. constant-round [BGP07b]. constant-size [PPSS13]. constants [BR17]. constrained [SS17]. constraints [HS24]. construct [YSM10]. construction [Ala17, ASZ^{+}23, BGKZ12, NSNK06, YYK^{+}18]. constructions [HSMY12, KME^{+}16, ZVH15]. consumer [LKH09]. consumption [RBEH15]. consumption-based [RBEH15]. contacts [DMP13]. containing [IT05]. Content [BFP03, HC10, LKH09, Pla09, SSJ22, ZP23, ZKP24]. Content-based [BFP03, HC10, SSJ22]. context [DGF^{+}17, GMGM21, HS09, JGK14, KKR^{+}24, MMF^{+}24, SRK^{+}20, SCO21, ULEY24]. context-aware [DGF^{+}17, KKR^{+}24, MMF^{+}24, SRK^{+}20]. context-centered [GMGM21]. context-sensitive [HS09]. contextual [CCB08]. Continuous [Yan21, BEY23, EG18, GGP21]. continuous-time [GGP21]. contourlet [SJ10]. contract [FGH24, IZS08, JT24, MdAN23]. contracts [FPP^{+}24, GYL^{+}07, PP24b]. Control [LMMS17, SAL17, AMZ22, AM23, ACCS24, ACF17, AFA^{+}23, AJC^{+}09, Ano11b, ACBC^{+}15, BAB23, BLM11, BNN04, CF03, CCD^{+}07, CR24, CK08, CZ06, GGJ22, HAAI24, HS09, HSMY12, KKR^{+}24, KAC16, KAC17, KLMM09, Kud02, KB22, KNL16, LRB^{+}10, MGAS24, MSKS20, MG23, MS15, Pen12, QLH^{+}24, RSMT24, RT23, RM12, SRD^{+}21, TZH04, ZM07]. Controlled [BB04a, BFG^{+}13, SM10]. controller [MPG21]. conversations [DBMS10]. convertible [LL21, Lin15]. convex [ALPW13]. Convey [SSR22]. cookie [ACB14]. cookie-based [ACB14]. cooperation [AGIK07]. coordinated [WKWS24]. coprocessors [Smi04]. core [KKK22b]. correcting [KM07]. Correction [BKBB20, CK23a, LSV^{+}23, PH24b, RSH^{+}24a]. correlation [AZM24, MdMF22]. corruption [BHKM23]. corruptions [KLPL21]. cost [DRPW12, DYDW10, Zen22]. cost-benefit [DRPW12, Zen22]. counter [BF13]. counter-measure [BF13]. counterfeit [AHC^{+}21]. Countermeasure [MLC23, GBDJ14, MTW^{+}14, Pen11]. countermeasures [ACT23, Bae10, HFA24, JK22, PSTS20, SB22, YNC22]. Counting [KM10]. country [SKK^{+}17]. cover [SJ10]. COVERAGE [AGIK07]. covert [DHW11, MPG21, VBAG24]. COVID [PGMPPC22]. COVID-19 [PGMPPC22]. CPU [EKY^{+}24]. CPU-optimized [EKY^{+}24]. Cracker [SGE02]. create [SKB23]. credential [AYHK18, ABFL12]. credentials [AFA^{+}23, SF17]. critical [ELPB24, Oft24, EEB^{+}15]. cross [IS23, PMPGMLLM12]. cross-layer [PMPGMLLM12]. cross-setting [IS23]. CRT [OT06]. CRT-based [OT06]. CRUST [GW09]. Cryptanalysis [ALPW13, DSY06, JT21, MS11]. Cryptanalytic [CJMS19]. Crypto [You06, PPAHC24]. cryptocurrency [FYF22]. cryptographic [ARMLS06, ASFG24, BPW05b, BGK08, BDH^{+}10, DFF^{+}16, DSRHC16, GW09, JCL^{+}18, KU16, Küs05, MR03, MS14, MSGCDPSS18, ZJS22]. Cryptographically [BCJ^{+}11]. cryptography [ALOW15, Bra22, HS24, KIP22, LP11, SDR20]. cryptography-steganography [HS24]. cryptologic [DVB02]. cryptosystem [HCN15, IT05, KAK22, MS11, PHE^{+}24, SGE02]. cryptosystems [KG11, NMBB12, OT06]. Cryptoviral [You06]. CSOC [SGJC18, SGJ19, SGC22]. CT [FYF22]. CT-GCN [FYF22]. cuckoo [VH19]. current [HC10]. Curve [JMV01, KIP22, SDR20]. curves [HMCD04, KSZ07]. cut [UMN^{+}20]. Cyber [AG23, HXTP23, KMMG23, MBH^{+}21, SSD14, SIA^{+}24, TDGL23, ASKG24, AK23, EOC^{+}24, AYM24, AKA24, CK23a, CK23b, CG23, ETRLG24, FPP^{+}24, HSLB24, LWL23, NABGJS^{+}24, NBA^{+}21, OCG24, PAI^{+}23, PCK22b, RV19, SDW23, SBCP21, SCY^{+}24, SBG22, SB22, SAH22, TV24, XYZT24, YNC22, ZP23, ZKP24]. Cyber-attack [SSD14, NABGJS^{+}24]. cyber-attacks [NBA^{+}21]. Cyber-physical [HXTP23, ASKG24, AK23, CK23a, CK23b, LWL23, PCK22b, XYZT24]. cyber-security [CG23, NABGJS^{+}24]. Cyberattack [RE24, TBGB20]. cyberbullying [KB22]. Cybersecurity [FDS^{+}24, RCC^{+}24, Asf24, EAM^{+}23, MSSO^{+}24, MOA^{+}24a, MOA^{+}24b, PCSA24, SB22, VSN22, YZ24]. Cyberspace [VHRRMG24]. cycle [AYHK18].

D2D [KIP22]. D3FEND [YZ24]. damage [WGMB13]. DAMe [PMPGMLLM12]. DAPP [KKK22a]. dark [JKLJ24, ZD22]. darknet [OBH^{+}20]. Dash [AKT23]. Data [LMMS17, PCK22b, SV11, SAL17, VdWZ14, AT22, AaBZS23, AAZAA23, ACF17, ABN14, AAB20, BS05, BNTW12, BDH^{+}10, CR20, CTM^{+}16, CTN24, CK24, CKS21, CLG23, CSM^{+}24, EMRN17, EAH^{+}07, GSAMCA18, GKS19, HJDC15, HK19, HXTP23, HH16, JK22, JJJ21, KNM^{+}24, Küs05, KY24, LLH21, LL21, LD17, LS23, Man21, MSKS20, MSGCDPSS18, MOA^{+}24b, NT20, OO20, OSSK16, OBH^{+}20, RT23, SBS23, SAC22, SSL22a, SRD^{+}21, SBG22, SMMN12, SXZC20, TS22, TNA23, VH19, XZ21, YAM^{+}15, ZO13, ZLZL20, ZTG22, Zha23]. data-leakage [RT23]. Database [KSM10, BDD01, DA21]. databases [BW08, CMS10, EAH^{+}07, HN14]. dataset [ASA23, AŠC24, BCGL23, KNSS24, NRC15]. dataset-independent [BCGL23]. datasets [GBC24, MU18, SKB23]. day [DGF^{+}17, SLGP23]. DDoS [AZ19, KSB23, KCB20, SRSM23]. de-correlation [AZM24]. DEALER [MPP21]. dealership [GSP^{+}16]. Deceiving [CSM^{+}24]. Decentralised [SHD^{+}24, PC19]. decentralized [LLWY09, LSG22, LSV^{+}23, MPP21, SVKV21]. deception [CK23a, CK23b]. decidability [Küs05]. decision [HSLB24, PHS22, WPD18, ZGRS23]. decision-makers [HSLB24]. deck [HSLB24]. decomposition [MARK20]. dedicated [ZZG19]. deduplication [NT20]. Deep [ME23, AAZAA23, AZM24, AKA24, BCGL23, EFB^{+}24, ETRLG24, HNHY23, Inc24, JT24, Kak24, KBY22, KCB20, KNSS24, MR24, MK21, OPK24, RSK^{+}24, RE24, SSJ22, WKWS24]. deep-autoencoder [OPK24]. deepfake [AVM23]. Defeating [DSB19, ZHZ22]. defence [AB24, GCT24]. defences [HXTP23, PNG^{+}20]. defending [SK06]. Defense [WWZ^{+}23, AG23, BB22, ETRLG24, KK22, MYLZ14, RE24, TCE23, VSR15, WR08, ZRJ14]. defense-in-depth [AG23]. defenses [BDG23, LLWY09]. defensive [AVM23, KCC^{+}23]. defined [KSB23, MdMF22, MTD^{+}24, PVGG24, RE24, SKH^{+}23]. Definition [TMP13]. degree [SSVC16]. Delay [ZO13]. Delay-sensitive [ZO13]. delays [SDW23]. delegatability [HYWS11]. Delegation [CK08, BGTCCBB10, CON09, GOBdlC11, PMDS23, WZ07, ZZG19]. delegations [RV03]. deniable [GCH^{+}19, JT21]. Denial [LLWY09, WR08, TGS17]. denial-of-service [WR08]. dependence [HS09]. dependencies [ACF17, MGAS24]. Dependency [AK23, ASKG24, CMS10]. Dependency-based [AK23, CMS10]. deploying [GDA22]. Deployment [ACBC^{+}15, HIDFGHR19, AAV22, Das12]. deployments [PCT22]. depth [AG23, PCT22, SAT09]. description [SM10]. descriptions [CAS22]. Design [DNF^{+}19, KAK22, MBRPS18, SRSM23, BSCZ11, CMR06, CFBvO09, CSS^{+}24, FN19, KMMG23, LMMO04, NCBG23, SBD23]. designated [HSMW08, HYWS11, RSD19]. designed [BAG^{+}24, KCM^{+}15, KC24]. Designing [AV17, PDM20]. desynchronization [Tip22]. detect [AFF24, BABB16, EZLC21, JJJ21, MLCQ21]. Detecting [DGF^{+}17, AGIK07, BHKM23, BMP^{+}14, HAAI24, KDYS19, NAQ24, PPL15, SGLC19, vORM06]. Detection [ABR16, CPPK15, sLC05, NLS24, RC24, SAL17, AZ19, AHT24, ATS23, ASAAS15, ASA23, Alg22, Alh24, AuHD^{+}23, ASN^{+}16, Alm24, AZ22, AK24b, AKA24, AAG24, AVM23, ASFG24, BV24, BSK^{+}23, BMFS24, BEPL^{+}17, BHL^{+}21, BDMM19, BBN24, BAG^{+}24, BR23, BT07, BSBK24, BCGL23, CK23a, CK23b, CSC^{+}23, CL08, CLG23, CKPM24, CG21, DPMC24, DTK^{+}18, DJS^{+}24, DGF^{+}17, EFB^{+}24, FAMMZ23, GPS17, GDA22, GBC24, Ham23, HGH23, HNHY23, IKS22, JT24, Kak24, KAK23, KKK22a, KBY22, KA18, KKK17, KJG^{+}11, KB22, KNSS24, KSM10, LHS^{+}24, LWA21, LWL23, MD23a, MR24, MKS^{+}23, McH01, MLYL20, MLCS16, MMF^{+}24, MK21, MI22, ME23, MB23, MP22, NK24, NNL20, OPK24, PM23, PAI^{+}23, PMW24, PDM20, PP24b, PV22, QLH^{+}24, RSV23, RB23, RBEH15, RHGTSC17, SAM^{+}24, SBS23, SM24, SC24, SLGP23, Sen14]. detection [SBD23, SRK^{+}20, SGSS23, SA24, STW^{+}23, SMV24, SF17, SKB23, SKH^{+}23, SSJ22, STD21, TBGB20, TKKO20, TSU^{+}24, TLX09, VLGL24, VL13, WGH23, WZZ^{+}24, WKWS24, WAB^{+}09, XT24, ZKP^{+}23, ZXZ^{+}11, ZTV23, ZIR24, ZGK07]. detection-based [AZ22]. detector [SBD23]. detectors [AvO13]. deterministic [GS15, RMSCR19]. DeTRACT [SVKV21]. Developing [KBJ22, SK16]. Development [KK17, CSS^{+}24, GRV24, INS21, LNX22]. device [ATN^{+}24, GA23, KWCK19, PH24b, PH24a, ACMV15]. Devices [LMMS17, ABK22, BS22, FBFGEM21, GCSÁBdSS12, IDHRPCMP15, KC24, LKH09, LCPD14, LSG22, MR03, MNC20, NBA^{+}21, TKKO20, UBK23, WLT^{+}24, ZLZL20, LSV^{+}23]. DFTMicroagg [AT22]. diagnosis [RSK^{+}24]. dictatorship [Tor20]. differences [RC24]. Differential [dSFK19, MTW^{+}14, SAC22, ZHZ22]. differentially [DA21, NMH^{+}24, PHS22]. Diffie [CC12, DS07]. diffusion [GGP21]. Digital [JMV01, SSM^{+}20, SSR22, ASF04, GA23, HG24, JG15, MSSO^{+}24, PGMPPC22, TSZ22, ZTG22]. dilemmas [NABGJS^{+}24]. dimensional [SGSS23]. dimensionality [PMW24]. direct [BCL09, KME^{+}16, KLPL21]. directed [Oft24]. directions [TDGL23]. disassembling [BSK^{+}23]. disclosure [KPM12, ORK23, RAC16, TNA23]. discounts [DFBJR18]. discovery [CYA^{+}18, DMP13, JM17, SAM^{+}24, ZR24]. discrete [HMCD04, SBD23]. Discretionary [Fon08]. discrimination [ZZG19]. diseases [RSK^{+}24]. Disposable [HRMM20]. Disposal [SSM^{+}20]. dissociation [VH19]. distillation [KCC^{+}23]. distinguish [HTM11]. Distributed [AFA^{+}23, RS18, ASF04, BEd23, Das12, DHS04, HN14, KKK17, Pen12, RSH^{+}24a, RSH^{+}24b, TLX09, WZ07, WLT^{+}24]. Distributing [VSM06, AYHK18]. distribution [ASF04, BFPP07, CMR06, GP17, HG24, MARK20, PSA23, SFT^{+}24, YYK^{+}18]. diversification [BB22]. diversity [AAG24]. DL [FAMMZ23]. DNA [IHNT02]. DNS [BDG21, SPDR17, ZTV23]. Do [BM11, EZLC21]. document [CF03]. documents [BFP03, BFPP07]. does [GSM^{+}11]. DoH [NLS24]. Dolev [BPW05b, BP08]. domain [CON09, CYA^{+}18, KCB20, LSR^{+}23, OCG24, PGMLK^{+}13, SCL^{+}18, SBG22, ST24, YA22]. domain-specific [LSR^{+}23, ST24]. DomainProfiler [CYA^{+}18]. domains [GMH14, ZMS22]. dominant [PM23]. DOMtegrity [TSMH19]. don’t [JTV19]. Double [ACHO13, PSDSNAHJ19, PS17]. Double-authentication-preventing [PS17]. Double-spending [PSDSNAHJ19]. Double-trapdoor [ACHO13]. down [Auf20, JTV19]. download [NRC15]. DPA [Bae10]. DREAD [ZTG22]. drift [Sen14]. drive [NRC15]. drive-by [NRC15]. driven [AMZ22, ZKP24]. DRM [LKH09]. DroidRista [AAB20]. dropping [VGG24]. DSA [MR04]. DTE [LH15]. DTLS [TGS17]. dual [AT22]. dual-level [AT22]. DUEF [CR20]. DUEF-GA [CR20]. during [CBC08]. Dynamic [AAMH^{+}24, BDMM19, CR24, DFBJR18, HK19, KYH18, Roe11a, ZM07, ANN23, AKT23, ASFG24, CON09, CMS10, CZ06, CHC24, HRMM20, KK17, KCB20, RV19, SSE^{+}15, YYK^{+}18, Roe11b]. dynamicity [AICC18]. dynamics [MRW02].

E-business [KBH07]. E-cash [LCL14].

e-commerce [GLP03, ZGC07].

e-government [GOBdlC11]. e-healthcare [QDW^{+}15]. e-services [DDPS02]. e-voting [MS11, Pen11, ZWX20]. Early [EKY^{+}24, RSV23, Alm24, ASFG24, BHL^{+}21, CYA^{+}18]. eavesdropping [CPPK15]. ECDSA [JMV01]. eCK [Ala17]. eCK-secure [Ala17]. ecosystems [HLA^{+}24, SIA^{+}24]. edge [GRV24, SSL22a]. edge-based [SSL22a]. Edit [LBW05, BCL13]. Editorial [GMO01, Gol08]. editors [ANS^{+}12, CM16, TCS^{+}20, CHM18, DMRS07, JZ11, YSM16, ZLL12]. education [AHX^{+}23, AYM24, KMMG23]. eduroam [PMPGMLLM12]. eduroam/DAMe [PMPGMLLM12]. effect [BR23, Zha23]. Effective [AuHD^{+}23, ASZ^{+}23, BSBK24, MGRR19, ZGC07]. Effectiveness [SK16, KC24, SGJC18, SGC22]. effects [SV11]. efficiency [HMCD04, MGV17, PDB11, Pen13]. Efficient [CLG23, HRL09, HH16, HYWS11, JCL^{+}18, KLZ^{+}21, KJ14, KU16, KSZ07, LMG17, LBZ^{+}10, MP16, MRG23, NP10, PCK22a, RD16, SMMN12, AKMW20, Bae10, BR18, BS21, BCD^{+}13, CKS21, CLO^{+}24, DK24, EMRN17, FAMMZ23, HN14, HYWS12, KKK22b, LMD17, LCL14, LWL^{+}21, MG19, MCD11, NT20, PMW24, Pen11, PV22, TK24, YYH^{+}23]. effort [SGJ19]. eID [RLEM18]. eIDs [HRMM20]. Elastic [CYK09]. electric [ACT23]. Electronic [GSM^{+}11, ABFO08, BDHZ15, BSV22, CLO^{+}24, DJN10, GP23, KO02, NBA^{+}21]. Elias [BR18]. elicit [SK16]. elicitation [FFG20]. elliptic [HMCD04, KIP22, KSZ07, SDR20, JMV01]. Email [SMV24, AIJM24, IZS08, WR15]. emails [SSM^{+}20]. embedded [VLGL24]. Embedding [BDH^{+}10, KEA24, SJ09]. EMBLEM [SKLP20]. empirical [ADHN24]. employee [Sep23]. employees [ADHN24]. enabled [AAV22, BAG^{+}24, BAB23, EMRN17, LWL23, NBA^{+}21, ZWX20, SHA20]. Enabling [BS05, JAYZ21]. encapsulation [MPS14, SKLP20]. encoding [SKLP20, ZSN23]. encompassing [EHM15]. Encouraging [YM19]. encrypted [ABK22, CTM^{+}16, CKS21, DBMS10, GSAMCA18, HH16, KJ14, OO20, OSSK16, SEXY18, WPD18, WR15]. encryption [ARMLS06, AKMW20, BZ03, CDP22, CTN24, CC12, CHMS21, Den08, DDX19, EHSS14, EKY^{+}24, FRG19, FGS12, GMMV05, GSP^{+}16, HK19, IMI18, JCL^{+}18, JSMG18a, JSMG18b, KME^{+}16, KLZ^{+}21, LMG17, LZQ^{+}18, Lin15, MRG23, PMDS23, PP24a, PPSS13, QLZH15, QDW^{+}15, SSP14, SRD^{+}21, STD21, XZ21, YP06, ZZG19, ZVH15, ZMS22, LLW^{+}16, Dan07]. end [BB04b, KNSS24]. end-to-end [KNSS24]. ended [Küs05]. Energy [DK24, RCC^{+}24, RBEH15]. Energy-efficient [DK24]. enforce [RSMT24]. enforced [BM11, GGJ22]. enforcement [ACMV15, AICC18, BCL13, DLR15, LBW05]. enforcing [BB04a]. engine [LNX22]. engineering [AZ19, AHT24, GMMZ06, MLM19, SK16, BZ20]. enhance [Inc24, SCO21]. Enhanced [ABN14, GGP21, YAM^{+}15, KBJ22, LV10, MSP^{+}13, PVGG24, PV22, SDR20, WLT^{+}24, JTV19]. Enhancing [ASKG24, ACB14, AK24a, CLW^{+}11, DA21, EFB^{+}24, HHSS24, KLMM09, MLYL20, XZ21]. ensemble [AIJM24, AuHD^{+}23, BSBK24, SM24, SC24, SSSJ24, WKWS24, YA22, ZBC23]. Ensuring [SAT09, SGJ19, TSMH19]. enterprise [WYL^{+}12]. entity [ZP23]. Entropy [BEPL^{+}17, HLKI15, RC24, SV11]. environment [BEd23, CLW^{+}11, CLG23, DJS^{+}24, GOBdlC11, INS21, KAK23, LRB^{+}10, MI22, NK24, NVB^{+}02, YL20, ZBC23]. environmental [SHA20]. environments [EOC^{+}24, AKMW20, FSG^{+}14, GYL^{+}07, KKR^{+}24, LI07, PPL15, WAB^{+}09]. epidemics [AGIK07]. equations [KM10, SSVC16]. Erratum [Roe11b]. Error [KKY^{+}23, RMPADF13]. errors [ABR16]. escalation [FTS^{+}20, KY24, YAY^{+}21]. Escrow [ARMLS06, YSM10]. Escrow-free [ARMLS06]. ESORICS [Sne05]. Establishing [BS22, Abb13]. establishment [BVS07, Das12, YRW14]. Estimating [AD08, KEA24, WHS18]. estimation [RMPADF13]. Ethereum [JT24]. EUFORBIA [BFP03]. Evading [PKHS23, XCW^{+}12]. Evaluating [HS24, JJJ21, MUiH^{+}21, PHS22, ZIR24]. Evaluation [AvO13, BDG23, AZ19, AMZ22, AKG16, BB04a, CR20, HIDFGHR19, PP24a, RHL17, SKY23, SK06, SNX19, TKKO20, TND^{+}15]. evenhandedness [IZS08]. event [AMZ22, SBD23]. event-driven [AMZ22]. events [RHGTSC17]. everlasting [GP23]. evidence [GA23, SdHZ16]. evidence-based [SdHZ16]. evident [GMSS23]. evolution [TV24]. EXAM [LRB^{+}10]. example [PHE^{+}24]. examples [BV24, XZ24]. exchange [Ala17, BBR18, BCF^{+}17, BFS^{+}13, CHZ16, DGZFGH13, ETAHCR08, HM22, HYWS12, MPS10, MPP14, MPP21, MSN02, SEXY18, Ust11, YLL^{+}18]. excursus [Fra18]. executable [BHKM23]. Execution [SK06, BHKM23, MLO^{+}04, MUiH^{+}21]. exercise [CG23, ZP23, ZKP24]. exfiltration [ZTV23]. exfiltrator [ZTV23]. exhaustive [KJS17]. Expediting [CSS^{+}24]. experience [MdAN23]. experimental [AMZ22, BCD^{+}13, DSB19]. experiments [QLOW09]. explainable [QQK^{+}24]. explicit [KAC16]. exploitation [JM17, MLM19, SEZ24, SNX19]. exploiting [ACF17]. exploits [ZRJ14]. Exploring [Inc24, HS15]. exponentiation [Bae10]. expression [WR15]. expressive [RD16]. extended [ACF17, AKT23, BGTCCBB10, Lop18, MS15, QLH^{+}24]. Extending [AMZ22, KKKV07]. extensibility [PHE^{+}24]. extension [JSMG18a, ST24]. extensions [PSTS20, PSOMS22, TSMH19]. external [CLW^{+}11]. extortion [You06]. extractable [ABPR24, KKK22b, WT16]. extraction [GA23, KNSS24, MKS^{+}23]. Eye [HAAI24].

FA [Bae10]. FaaS [TMvM24]. faceted [BFT08, QLOW09]. facets [AKG16]. factor [AGZA22, CG14, ML17, MNC20]. failures [SK06]. Fair [KM03, DGZFGH13, GRV05, HYWS12, SEXY18, SGJ19]. fairness [TMvM24, TMvM24]. fake [RB23]. false [HXTP23]. falsification [AHT24]. familiarity [ZGC07]. familiarity-based [ZGC07]. family [AMP12, MSN02, PLA^{+}21]. Fast [Rov23, ZLZL20]. fasttext [Alq24]. fault [CL09, GKS19, RBD02]. fault-tolerance [CL09]. fault-tolerant [GKS19, RBD02]. FCS [ACM05]. FCS/VERIFY [ACM05]. FDCO [ZLZL20]. Feature [AHT24, KCB20, AZ19, KNSS24, MKS^{+}23, SBS23, YP06, ZIR24]. features [AVM23, GDA22, OPK24, ZJS22]. Federated [Alh24, Ham23, TGNA22, WW07, YYH^{+}23]. federating [PMPGMLLM12]. federations [SHD^{+}24]. few [WZZ^{+}24]. few-shot [WZZ^{+}24]. Fiat [HHSS24]. file [BLS^{+}23, sLC05, ZLGZ19, ZLGZ19]. file-based [sLC05]. filter [MB16, VH19, ZIR24]. filter-based [ZIR24]. filtering [BFP03, HBK22]. find [SSVC16]. fine [KLMM09, RT23]. fine-grained [KLMM09, RT23]. fingerprint [CC10, Nui12]. fingerprinting [ATN^{+}24]. finite [BCL13]. firewall [ABR16, KNL16, MWZ06]. firewalls [HBK22]. firmware [WLT^{+}24, YL20]. first [LM06, ZD22]. first-order [LM06]. fixed [HMCD04]. FL [TGNA22]. flag [MVA24a]. Flexible [JSMG18b, ORK23, QDW^{+}15, SBG22]. floating [ABB17, KW15]. floating-point [ABB17]. Flood [MLC23]. flooding [BHL^{+}21, JK22]. Flow [HS09, SdHZ16, ACCS24, AAB20, BMP^{+}14, BNN04, MP22, ZM07]. Flow-based [SdHZ16]. Flow-sensitive [HS09]. Flowchart [SM10]. flows [IKS22]. fly [CL13]. focusing [KKK22b, YAY^{+}21]. fog [AAV22, KKR^{+}24, RA24, KAK23]. fog-based [RA24, Moh23]. fog-enabled [AAV22]. fog-IoT [KAK23]. FOO [ZLJW20]. forecasting [ZKP24]. Forensic [JKLJ24, GA23, GMGM21, GRV24]. forest [PP24a]. form [DdP13]. Formal [ACMV15, GKBS12, Yon18, dAKdG10, BKBB20, BBB20, CMN^{+}18, KAC17, NSNK06, PDB11, RV03, VdWZ14]. formalism [GBDJ14]. formalization [MS14]. format [ZLGZ19]. format-based [ZLGZ19]. forward [CDF^{+}13, KME^{+}16]. forward-secure [KME^{+}16]. Fostering [ADHN24]. FOTB [YL20]. four [Dan07]. FPGA [RA24]. Fractional [BCA^{+}10]. fragmentation [LS23]. frame [SHW23]. framework [AZ19, Abb13, ADHN24, AAB20, EOC^{+}24, AVM23, ABFL12, BBG24, BMFS24, BFT08, CR20, CK24, EAM^{+}23, EFB^{+}24, FAH24, HAAI24, JAYZ21, KMMG23, KAK23, KNL16, LWA21, LWL23, MB16, MLO^{+}04, Moh23, PDM20, RSMT24, RV19, RV03, SVKV21, SSM^{+}20, SKY23, Vaj16, VdWZ14, WR08, WHS18, XSA13, Yan21, YL20, ZP23, ZRJ14, ZGRS23]. frameworks [YZ24]. free [ARMLS06, BGKZ12, CLPP11, YRW14]. freeness [FGH24, HIST09]. frequency [ATN^{+}24, YA22]. friendly [CLPP11]. Fujisaki [GMMV05]. full [Bra06, RSH^{+}24a, RSH^{+}24b]. Fully [CDF^{+}13, MLCQ21]. function [AGZA22, BGKZ12, KKY^{+}23, LWL^{+}21, SB14, STW^{+}23, TK24]. functional [WCS20]. functionality [SPM13]. functionality-based [SPM13]. functions [AMZ22, AMP12, EG18, GKKT10, MS09, PHE^{+}24, SM10]. fusion [HS24]. FusionGuard [Alm24]. future [CYA^{+}18, LMMP06, PSA23, TDGL23]. fuzzy [DK24, HCN15, TMM^{+}19].

G [DWU^{+}23, DWU^{+}23]. G/6 [DWU^{+}23]. GA [CR20]. Gait [HCN15]. Game [wLW05, ETRLG24, LVK18, MTD^{+}24, PNG^{+}20, VGG24, WYL^{+}12]. game-based [VGG24]. game-theoretic [LVK18, PNG^{+}20]. games [MB23]. Gaming [VBAG24]. gap [Auf20]. garbled [ZXZ22]. gas [ADHN24]. gate [AK24a, JSMG18b]. GaTeBaSep [MTD^{+}24]. GCN [FYF22]. GDPR [GdSdC24]. general [BCD^{+}13, Pen11]. generalised [BKMR08]. generalization [DJN10]. Generalized [BR18, KMR09]. Generalizing [KG11, AMP12]. generated [AZM24, IHNT02]. Generating [ESSP24, LC04, XZ24, BV24]. generation [AKZM20, AMLH18, BHZ^{+}21, CAS22, HG24, KJG^{+}11, KSZ07, MR04, YKP22, ZP23, ZKP24]. Generative [HAS^{+}24, XZ24]. Generic [Ala17, ZVH15, EWR^{+}09, FAH24, KME^{+}16]. genetic [AŠC24]. geographic [ABN14]. GeoProof [ABN14]. geospatial [HK19]. gesture [GCSÁBdSS12]. Global [Pri04]. goal [Oft24]. goal-directed [Oft24]. goals [RGL16]. good [HAS^{+}24]. Gossamer [DDZ22]. gossip [ML14]. government [GOBdlC11, UGP24]. GPU [VPI15]. GPU-assisted [VPI15]. GPUs [AB22]. GQ [HRL09]. graduates [AHX^{+}23]. grained [KLMM09, RT23]. graph [ASKG24, BCEM04, CR20, LHS^{+}24, NNL20, RV03, STW^{+}23, WGH23, WWZ^{+}23, YL19]. graph-based [LHS^{+}24, NNL20]. graph-theoretical [BCEM04]. graphical [AKT23, CG14, CFBvO09, GTM11, IS23, WLLW14]. graphs [ELPB24, HS09, HLKI15, KB13, LWL^{+}21, SDG22]. gray [SM24]. grey [CKW19]. grey-box [CKW19]. grid [AKZM20, AT22, AEHS24, CLPP11, DPMC24, FJZZ24, KLMM09, LP11, PSA23, PNG^{+}20]. Gridchain [PSA23]. grids [MD23a, SAH22]. Groth [ABPR24]. ground [AvO13]. Group [EHSS14, TDRR20, BVS07, CHC24, DFBJR18, GNS14, HM22, IDHRPCMP15, PBI^{+}23, RBD02, YLL^{+}18, ZWQ^{+}17]. groups [PJ10]. guarantee [ABCC08, Bel10]. guarantees [LSWW14]. Guard [ZLGZ19]. Guardian [PP24b]. guessers [PCT22]. guest [ANS^{+}12, CM16, TCS^{+}20, CHM18, DMRS07, JZ11, YSM16, ZLL12].

Haar [KCM^{+}15]. hacking [EZ22]. hand [GCSÁBdSS12]. Handling [WZ07, AŠC24]. hands [CKKK23]. hands-on [CKKK23]. handshake [TGS17]. hard [EAH^{+}07]. hard-to-reverse [EAH^{+}07]. hardening [DMDD16, DRPW12, MRW02, Zen22]. hardware [BSCZ11, DYDW10, KBJ22, KKK22b]. hardware-assisted [DYDW10]. harvesting [BR18]. hash [AMP12, BGKZ12, GKKT10, MS09, MFES04, SB14]. HashEdDSA [BST21]. hashing [BDPV14, CHKO12, CJMS19, MP16]. Hate [AK24b]. HBA [SC24]. health [BDHZ15, KNM^{+}24, QLZH15, SAC22, SHA20, SDR20]. health/accessibility [SHA20]. healthcare [BEd23, FDS^{+}24, MG23, PH24b, PH24a, QDW^{+}15, RSH^{+}24a, RSH^{+}24b, SIA^{+}24, SFB^{+}24]. Hellman [CC12, DS07]. heterogeneous [RLEM18, SAT09, WGH23, WKWS24]. HIBE [ZZG19]. hidden [ABB17, GSS10, SSVC16]. Hiding [GI19, XZ21, ZXZ22]. Hierarchical [BKBB20, BBB20, NAM06, FTS^{+}20, JCL^{+}18, LLW^{+}16]. High [BNTW12, AM23, BB04b, KKK22b, Pen13]. high-efficient [KKK22b]. high-end [BB04b]. High-performance [BNTW12]. high-rate [AM23]. higher [PDB11]. Highly [PGMPPC22, AAB20]. hijacking [SGLC19]. historical [TV24]. history [TDRR20]. HiveSec [SS17]. Hoc [LH23, AHT24, Gol12, MS11, SF17, SSSJ24]. home [AZ22, FBFGEM21, SRD^{+}21, WLT^{+}24]. homes [BAB23, HKO22]. homology [WGH23]. homomorphic [DDX19, EMRN17, MMS16, PP24a, Rov23, SEXY18, STD21, SXZC20]. Homomorphico [PHE^{+}24]. HoneyCirculator [AYHK18]. honeypot [EFB^{+}24, LSG22, LSV^{+}23]. honeytoken [AYHK18]. hop [BT07]. host [CLW^{+}11, KLPL21]. hotels [DZW^{+}18]. HTTP [MLC23]. HTTPScout [MLC23]. hull [ALPW13]. human [AAZAA23, ALPW13, MBHT17]. hunger [MB23]. HWMSNs [RSH^{+}24a, RSH^{+}24b]. Hybrid [KML03, BBN24, BCGL23, CC10, GMMV05, MMS16, MB23, SGSS23, TCE23]. Hydras [PC19]. hygiene [SCY^{+}24]. hyperparameter [GZH^{+}23].

IC [AHC^{+}21]. ICS [ESSP24]. ICT [HLA^{+}24]. ID [FGS12, WCS20, ZZG19]. ID-based [FGS12, WCS20]. identification [ATN^{+}24, AAZAA23, ASZ^{+}23, ALPW13, CL08, FYF22, GBG18, GI19, LcSCL^{+}18, MGV17, NCBG23, NRC15, PLA^{+}21]. Identifier [ZZG19]. identifiers [IHNT02]. identifying [KGG09, SPDR17]. Identity [CCS07, LP11, LLW^{+}16, CHZ16, FPP^{+}24, GOBdlC11, GMLM23, HRL09, JCL^{+}18, LMG17, LBZ^{+}10, MPS14, NA14, ORK23, SHD^{+}24, TMP13, Ust11, VdWZ14, YSM10, ZWQ^{+}17]. Identity-Based [LLW^{+}16, CCS07, LP11, CHZ16, HRL09, JCL^{+}18, LMG17, LBZ^{+}10, MPS14, Ust11, YSM10, ZWQ^{+}17]. IDS [ATS23, BBN24, ZBC23]. IDSIC [CL08]. IEC [AEHS24, CH16, MD23b]. IEEE [KDM22]. If [ASN^{+}16, MBHT17]. IHE [ACBC^{+}15]. II [BB04b]. IIJ [SKK^{+}17]. image [HC10, RC24, ZMS22]. images [HLKI15, MBHT17]. imbalanced [ASA23, AŠC24]. immune [Hub12, MdMF22]. immutable [SVKV21]. impact [AHT24, SSD14, VBAG24, ZIR24]. implantable [NBA^{+}21]. implement [ABFL12]. implementation [Auf20, BSCZ11, DNF^{+}19, IDHRPCMP15, KAK22, MBRPS18, MFES04, PVGG24, RA24, TKKO20]. implementations [RSMA19, UMN^{+}20]. implemented [MS15]. Implementing [ALOW15, BGK08]. implications [HMCD04]. implies [EHSS14]. improve [YM19]. Improved [HLS18, ABN14, CLO^{+}24, LS23, MLCS16, WT21, ZLJW20]. improvement [AaBZS23, FJZZ24, YWW22, Yan21]. Improving [AIJM24, CH16, KLPL21, TTS^{+}06, BJ16, TGS17, TG05]. IMS [VL13]. in-depth [PCT22]. in-VM-assisted [PDM20]. inappropriate [LTC23]. incentive [LI07]. incentives [MPP21]. INCHAIN [FPP^{+}24]. incidents [SDW23]. including [NMH^{+}24]. incomplete [BW08]. Incompleteness [SAL17]. Inconsistency [SAL17, ACS21]. incorporating [BCF^{+}17]. incremental [TBGB20, ZKP^{+}23]. independent [BCGL23]. indifferentiability [BGKZ12]. indistinguishability [SXZC20]. individual [Man21]. inductive [MP15]. industrial [CR24, QLH^{+}24, RM12, WZZ^{+}24]. industry [SDW23, SDG22]. industry-level [SDW23]. inference [AJC^{+}09, JG15, KMW24, WWZ^{+}23]. influences [AHX^{+}23]. inform [PCT22]. Information [ACCS24, CMH^{+}24, KBH07, TND^{+}15, UGP24, ASF04, AD08, ADHN24, AFA^{+}23, ABFL12, BDHZ15, BB04a, BDD01, CMMPS15, DdP13, DBMS10, Dwi24, EFH09, GI19, HS09, HSHM24, INS21, IKS22, IHNT02, KPM12, LLH21, LTC23, LMD17, LcSCL^{+}18, MU18, NAM06, PD21, Pla09, ST24, TKKO20, WHS18, ZM07]. information-centric [AFA^{+}23]. information-distribution [ASF04]. Information-theoretically [TND^{+}15]. informed [AG23, ASFG24]. infrastructure [CLPP11, HBF^{+}24, Oft24, SSV22, TS20, TMP13, Gri06]. infrastructures [ACT23, EEB^{+}15]. infringement [SCY^{+}24]. Inheritance [SSR22]. injection [DSB19, DTK^{+}18, HXTP23, Kak24, NAQ24, SEZ24, SNX19]. insecure [KM07]. insecurity [Sat20]. Insider [SSV22, YP12, ASA23, MLYL20, Sep23, YA22]. inspection [BDF04]. inspired [SS17, ZZW^{+}10]. install [VGG24]. instance [Sen14]. instance-weighted [Sen14]. instances [BMP^{+}14, JG15]. instantiable [WPD18]. instantiations [CYK09]. Instruction [DM07, WGH23]. Instruction-level [DM07]. instructions [EKY^{+}24]. insurance [FPP^{+}24, GYL^{+}07, TDGL23]. integers [DDX19]. integrated [JT24, VD24]. Integrating [GRV24, Ust11, ZKP24, ZGK07]. integration [AKR24, VCD21]. Integrity [IMI18, LZQ^{+}18, CK24, EEB^{+}15, LVK18, SXZC20, TSMH19, YAM^{+}15, EEB^{+}15]. integrity-checking [YAM^{+}15]. Integrity-OrBAC [EEB^{+}15]. Integrity-verifiable [LZQ^{+}18]. intelligence [AKA24, BAC24, HLA^{+}24, KKK17, MPP21, RV19, SBCP21, VSN22, ZKP24]. intelligence-based [KKK17]. Intelligent [CSC^{+}23, MOA^{+}24b, AAMH^{+}24, AFF24, ZGRS23]. intended [GCT24]. intent [LLG22, EZLC21]. intent-based [LLG22]. IntentAuth [LLG22]. inter [LLG22]. inter-process [LLG22]. Interacting [vOLW05]. interactions [Sep23]. Interactive [MS09, CDF^{+}13, CHZ16, CL09, EHSS14, HHSS24, JT21, KBJ22, MS09]. interest [BHL^{+}21, JK22]. interface [CFBvO09]. intermediaries [DGZFGH13]. internal [Sep23]. International [Dwi24, KBH07]. internet [AKR24, WZZ^{+}24, AAMH^{+}24, Alh24, AHX^{+}23, CK24, DHS04, EB24, Ham23, MOA^{+}24b, OPK24, SM24, VD24]. InterPlanetary [BLS^{+}23]. interpretability [ZGRS23]. interpretable [CLG23]. interpretation [DM07]. interruption [AKZM20]. intersection [MCD11, SZ24]. introspection [AYHK18]. Intrusion [BT07, DPMC24, McH01, SBS23, WAB^{+}09, ATS23, Alg22, AuHD^{+}23, AAG24, BR23, BCGL23, CL08, CLG23, DJS^{+}24, GDA22, Ham23, KKK17, KNSS24, KSM10, LWA21, LWL23, MR24, MKS^{+}23, MLYL20, MMF^{+}24, MK21, ME23, MB23, NK24, OPK24, PMW24, PV22, QLH^{+}24, RBEH15, SM24, SC24, SRK^{+}20, SSJ22, STD21, TBGB20, TSU^{+}24, TLX09, VL13, WZZ^{+}24, WKWS24, ZKP^{+}23, ZIR24, ZGK07]. intrusions [MYLZ14]. invariants [XYZT24]. inversion [PHS22]. inverting [YYH^{+}23]. Investigating [ASAAS15]. Investigation [KSB23, PHE^{+}24, JKLJ24, PSA23, SGJ19]. investigations [GMGM21, JG15]. involvement [ZBD06]. IoT [LSV^{+}23, VD24, AAV22, AMZ22, AKMW20, Auf20, BAG^{+}24, BR23, BS22, CSC^{+}23, FFG20, GA23, GMGM21, GRV24, HHSS24, HBF^{+}24, Inc24, KKR^{+}24, KAK23, KAK22, KKK22b, KNSS24, LSG22, MAZS24, MNC20, NK24, NNL20, OBH^{+}20, PCSA24, PV22, QQK^{+}24, RCC^{+}24, SHA20, SHOL23, TKKO20, WLT^{+}24, YL20, ZBC23]. IoT-based [RCC^{+}24]. IoT-enabled [SHA20]. IoT-oriented [KKK22b]. IoTvulCode [BAG^{+}24]. IoUT [Moh23]. IoV [AGZA22]. IP [BABB16, RS18, WAB^{+}09]. IPFS [BLS^{+}23, PC19]. IPsec [TG05]. IPTV [KOSU16, SCY^{+}24]. ISA [ADHN24]. ISA-95-based [ADHN24]. ISC [BM05]. ISHO [SC24]. ISM [MdMF22]. ISM-AC [MdMF22]. ISO [CH16, DFF^{+}16, MD23b]. ISO-standards-track [DFF^{+}16]. ISO/IEC [CH16, MD23b]. isogeny [HM22]. isolation [JAYZ21]. isomorphisms [GMS03]. ISP [KCB20]. Issue [KBH07, Ano11b, ACM05, BJ15, BM05, Daw04, Pri04, Sne05, YSD^{+}20]. issues [EB24, FSG^{+}14, HC10].

Java [MLM19]. join [TDRR20]. Journal [KBH07, Dwi24].

K-anonymity [KNM^{+}24]. KAMU [PGMLK^{+}13]. Kasahara [CHZ16]. Keeping [BW08]. kerberized [PMPGMLLM12]. Kerberos [BCJ^{+}11, MPP14, PGMLK^{+}13]. kernel [KY24, YAY^{+}21]. KEVM [PP24b]. Key [AKMW20, BRS06, Gri06, MARK20, NVB^{+}02, AAV22, Ala17, AK24a, ALOW15, BCJ^{+}11, BZ03, BBR18, BVS07, BCF^{+}17, Bra22, BFS^{+}13, CMR06, CTN24, CCS07, CC12, CHZ16, CL09, CH16, CHC24, DJN10, Das12, EHSS14, FRG19, FGS12, GNS14, GP17, Hal20, HBF^{+}24, HM22, KDM22, Lu09, MPS10, MPS14, MPP14, MSN02, PP24a, RHL17, RBD02, SKLP20, SSP14, TMM^{+}19, TGS17, TK24, Ust11, YLL^{+}18, YRW14, YYK^{+}18, YSM10, ZWQ^{+}17]. Key-updatable [AKMW20]. Keyboard [HS15, LSR^{+}23]. keyboard-based [LSR^{+}23]. keyboards [CSM^{+}24]. keys [AZM24, BR20, BS22, BDH^{+}10, GW09, IT05, LC04, PPSS13, Pla09]. keystroke [CF07, MRW02]. Keyword [VH19, AKMW20, FRG19, LZQ^{+}18, OO20, OSSK16]. Keyword-based [VH19]. kNN [OO20]. knowledge [AHX^{+}23, HHSS24, SBG22, WCS20, YP12].

labels [ZM07]. language

[HNHY23, SIA^{+}24, TS20]. Large

[TLX09, Das12, RLEM18, SMMN12, SDG22, ZWX20]. Large-scale [TLX09, Das12, RLEM18, SDG22, ZWX20]. LASSI [AAV22]. lateral [SKB23]. Lattice [GCH^{+}19, SSP14, BBR18, WT21]. Lattice-based [GCH^{+}19, SSP14, BBR18]. lattices [KJ14]. law [MdAN23]. layer [ANN23, DGF^{+}17, KSB23, PMPGMLLM12, QQK^{+}24, WR08]. LDES [SBD23]. leakage [AD08, DdP13, DBMS10, RT23, WT16, XT24, XZ21]. leakage-resilient [WT16, XZ21]. leakages [MRCK24]. learner [HNHY23]. Learning [MLC23, AZ19, AAZAA23, Alh24, AZM24, AKA24, AFF24, BBG24, BCGL23, CSM^{+}24, CG23, EFB^{+}24, ETRLG24, EZ22, GSAMCA18, GBG18, Ham23, Inc24, JT24, Kak24, KBY22, KA18, KCB20, KB22, KNSS24, LL22, LTC23, LH23, MR24, MK21, ME23, MP22, NCBG23, NK24, OBH^{+}20, PM23, RSK^{+}24, RE24, SBS23, SM24, SLGP23, SRK^{+}20, SMV24, SKB23, SKH^{+}23, SEZ24, SSE^{+}15, TSU^{+}24, TLX09, TMvM24, TGNA22, ULEY24, WZZ^{+}24, YYH^{+}23, YA22, ZBC23]. learning-based [AFF24, Ham23, KBY22, KB22, SRK^{+}20, TSU^{+}24]. least [PSOMS22]. ledger [WLT^{+}24]. left [Bae10]. legacy [KSB23]. legal [Lev07]. length [YOV09]. Less [BFS^{+}13, RSH^{+}24a, RSH^{+}24b]. let [EZLC21]. level [AT22, AHX^{+}23, DM07, MG23, SDW23, SCL^{+}18, SGJC18, SGC22, ZTG22]. Leveraging [RV19, GDA22, GMdFLR23, KKR^{+}24, KNM^{+}24]. Lib2Desc [CAS22]. libraries [CAS22]. library [BPW05b, PHE^{+}24]. lifecycle [ZKP24]. LightGBM [PV22]. Lightweight [MNC20, RBEH15, AAV22, DJS^{+}24, KAK22, OPK24, PDM20]. like [ASN^{+}16]. Limits [BP08]. line [BCD^{+}13, LMMO04, SZ24]. line/on [BCD^{+}13]. linear [SBD23]. Link [DMDD16]. Link-time [DMDD16]. linkability [BFG^{+}13]. Linkable [GP17]. list [Des09]. list-based [Des09]. Listega [Des09]. literature [AKR24, ATS23, EB24, MVA24b, PCSA24, SAC22, ZR24]. local [LWL^{+}21, PSA23]. localization [ZXZ^{+}11]. Locating [XYZT24]. location [EG18, NMH^{+}24, RA24]. locations [NMH^{+}24]. lock [AMZ22]. Lockmix [BSK^{+}20]. lockpicking [GdKGV14]. log [HBH12, YKP22]. logarithm [HMCD04]. logarithmic [HM22]. logarithmic-order [HM22]. logging [GMSS23]. logic [LM06, SBD23, SdHZ16]. login [KPM12]. logs [LL22, SKB23]. long [VLGL24, ZKP^{+}23]. long-term [VLGL24]. look [Auf20, ZD22]. looks [ASN^{+}16]. lookup [PP24a]. loss [MU18]. lottery [GSS10]. Low [BGP07b, GS15, MU18, SSVC16]. low-deterministic [GS15]. Low-randomness [BGP07b]. LPN [RG13]. LSTM [ETRLG24]. LTE [LSWW14]. LWE [SKLP20]. LWE-based [SKLP20].

m [LNX22]. MAC [EMRN17, SHW23]. MAC-based [EMRN17]. Machine [MLC23, MP22, AZ19, CSM^{+}24, GSAMCA18, Inc24, KA18, KB22, LL22, LH23, MS14, PM23, SBS23, SLGP23, SMV24, SKB23, SKH^{+}23, SSE^{+}15, TSU^{+}24, TMvM24, ULEY24]. machine-learning [SSE^{+}15]. machines [MLCS16, PDM20, vOLW05]. Maintaining [SGC22]. makers [HSLB24]. Making [BR17]. MalFamAware [PLA^{+}21]. MALICIA [NRC15]. Malicious [BSK^{+}23, GAdFGM23, NCBG23, NLS24, AKZM20, ABB17, BRS06, CMS10, EFB^{+}24, GPS17, KK22, LNX22, RHGTSC17, TSMH19, WGMB13]. Malware [HLKI15, MLCS16, ZSN23, BEPL^{+}17, BDMM19, BSBK24, CSC^{+}23, GAdFGM23, KBY22, KKK22b, LWL^{+}21, MI22, OBH^{+}20, PKHS23, PDM20, PC19, PLA^{+}21, SGSS23, SA24, STW^{+}23, SKK^{+}17, TKKO20, VGG24, VPI15, VLGL24, WGH23, ZRJ14]. manageability [TG05]. Management [CF03, AKR24, ASF04, EOC^{+}24, AG23, BF13, CH16, GMMZ06, GMLM23, GH05, LLWY09, Lop18, LcSCL^{+}18, NA14, PGMPPC22, RHL17, RLEM18, RBD02, SHA20, SSN15, SIA^{+}24, TMP13, VdWZ14, WPD18, WLT^{+}24, WW07, dAKdG10, vOLW05]. managing [AMLH18]. mandatory [DLR15]. MANET [SSSJ24, Alg22]. MANET-Cloud [Alg22]. manipulation [PD21]. manufacturing [TRT^{+}24]. map [ZGRS23, MD23b]. MAPAS [KBY22]. mapping [OBDM^{+}24]. maps [ZMS22]. Maritime [HSLB24, MSSO^{+}24, OCG24, YZ24]. marketplaces [ZTG22]. Markov [ABB17, RHL17]. mask [PKHS23]. masking [Rov23]. masquerade [Sen14]. Mastering [KNM^{+}24]. matching [FHV18, SBB19]. mathematical [PHE^{+}24]. MaX [BFP03]. maximum [AD08, BHZ^{+}21]. McEliece [NMBB12]. McEliece-based [NMBB12]. mean [BM11]. means [PHE^{+}24, Rov23]. measure [BF13, CLO^{+}24, ORK23, SGJC18]. measurement [RMPADF13]. Measuring [RGL16, SBCP21]. mechanism [ACF17, AB24, IOU^{+}21, KY24, LI07, MVA24a, PD21, RE24, TKKO20, YAY^{+}21]. mechanisms [BV24, CSL^{+}23, LBW05]. media [AK24b, KB22, RAC16, ZLGZ19]. mediated [Hal20, VSM06]. mediation [Asf24]. Medical [Alh24, KC24, LLH21, RSH^{+}24a, RSH^{+}24b]. MEDICALHARM [KC24]. mediums [GCT24]. meet [FN19]. meets [AB22, LLBL18]. member [TDRR20]. membership [WWZ^{+}23]. memorability [YM19]. memory [BHKM23, CJMS19, NAQ24, SV11, ZKP^{+}23, vOLW05]. memory-based [SV11]. mercurial [WCS20]. merging [CMR06]. merits [HS24]. Merkle [MFES04]. mesh [EHM15, SSN15]. Message [ANS^{+}12, CM16, CHM18, JZ11, TCS^{+}20, YSM16, ZLL12, CL13, GP17, MS09]. messages [AM23]. messaging [Man21]. metadata [ACS21, KKK22b]. metamorphic [LNX22]. metering [GLMS19, MSP^{+}13, RDK18, SSV22, TS20]. Method [SAL17, BHKM23, CYK09, FFG20, IT05, KGG09, KB23, MP15, MKS^{+}23, PJ10, QLH^{+}24, SKLP20, Sep23, SPDR17, WGH23, WZZ^{+}24, WHS18, YKP22]. methodologies [GRV24]. methodology [AvO13, Des09, GMMZ06, GMGM21, SGJC18, SGJ19]. methods [BSK^{+}23, CMN^{+}18, KEA24, MD23a, SBS23, TS22]. metrics [BDG23, KEA24]. metro [JAYZ21]. microaggregation [SMMN12]. microblogging [ASN^{+}16]. microscope [PCT22]. Microsoft [ELPB24, You06]. Middle [TCT22]. middleware [SRD^{+}21]. military [SBG22]. minimisation [VdWZ14]. Minimizing [KPM12, ZBD06]. mining [BDMM19, BNTW12, HBH12, MVA24a, MVA24b, Sat20, ZR24]. mIoMT [PH24b, PH24a]. misbehavior [SF17]. MITF [SKK^{+}17]. Mitigating [YA22, SRSM23]. Mitigation [KY24, AGL24, AHC^{+}21, BHL^{+}21, BBN24, EKY^{+}24, KCB20, KB22, MR24, SMV24, SSV22, SDG22, YP12]. mitigations [KCC^{+}23]. MITRE [YZ24]. mix [BSK^{+}20, Dan07, MS11, Pen11]. mix-based [MS11, Pen11]. mix-related [Dan07]. ML/DL [FAMMZ23]. MLChain [BBG24]. MLSTL [TSU^{+}24]. MLSTL-WSN [TSU^{+}24]. mmWave [KCC^{+}23]. Mobile [EWR^{+}09, FN19, HGH23, LMMS17, ABK22, ACB14, AAZAA23, AKG16, AMRR17, CSL^{+}23, CSS^{+}24, CF07, GSM^{+}11, GCSÁBdSS12, HIDFGHR19, HCN15, HRMM20, HZL^{+}17, IDHRPCMP15, LCPD14, LH15, LL14, LV10, MTSH18, SSSJ24, UBK23, VBAG24, WGMB13, ZLZL20, BT07, SSE^{+}15]. mobile-phone [WGMB13]. Mobile-Sandbox [SSE^{+}15]. MobileTrust [LV10]. Model [CMH^{+}24, SSR22, ASF04, Ala17, ASZ^{+}23, ABB17, AC08, BBG24, BMV05, BGTCCBB10, BS22, BCEM04, DJS^{+}24, EEB^{+}15, FBFGEM21, FYF22, GMMZ06, Gol12, GYL^{+}07, GKBS12, HNHY23, HL04, INS21, JT24, JGK14, Kud02, LL22, Les24, LTC23, LH15, MdSC^{+}15, MGAS24, MdMF22, NK24, NSNK06, PHS22, QLOW09, RSD19, SK16, SPM13, SSP14, SKH^{+}23, SHW23, Tip22, TND^{+}15, YLL^{+}18, ZGC07, ZTG22, EZ22, ZZW^{+}10]. Model-Agnostic [CMH^{+}24]. model-checking [AC08]. model-oriented [SK16]. Modeling [CCB08, DLR15, SDW23, ACMV15, BEd23, EZ22, GMLM23, KAC17, KC24, MB23, Sep23, Yon18]. models [ASKG24, ABB17, Auf20, CSM^{+}24, Den08, DRPW12, FGS12, GGP21, KAC16, KMW24, KB22, KCC^{+}23, MR24, SM24, ST24, VLGL24, XZ24]. moderated [Asf24]. modern [HLA^{+}24, KC24]. modes [BDPV14]. Modification [PDB11]. modified [KJS17, ZTV23]. modular [MSKD16, OCG24, YOV09]. modules [KKK22a]. monitor [SGJC18]. monitoring [EKY^{+}24, HBH12, IOU^{+}21, LHS^{+}24, SHA20, SDR20, vORM06]. motivate [Fra18]. mouse [LSR^{+}23, YA22]. mouse- [LSR^{+}23]. movement [SKB23]. moving [BB22, DA21, HXTP23]. MPC [BST21, KBJ22, RSMA19]. MQTT [PV22, ZBC23]. MQTT-IoT [PV22]. MSOffice [VLGL24]. MSOffice-embedded [VLGL24]. Multi [BT07, EHM15, KWCK19, LS23, OSSK16, SFT^{+}24, YYK^{+}18, ZZW^{+}10, EOC^{+}24, BNTW12, BFT08, CON09, Das12, DJS^{+}24, FAMMZ23, GMH14, HNHY23, Kak24, KM03, LWL^{+}21, MNC20, PGMLK^{+}13, PP24a, PBI^{+}23, QLZH15, QLOW09, RSK^{+}24, SKLP20, SGSS23, WR08, Zen22]. multi-attack [FAMMZ23]. multi-authority [QLZH15]. multi-bit [SKLP20]. Multi-cast [YYK^{+}18]. multi-class [DJS^{+}24]. multi-classification [LWL^{+}21]. Multi-cloud [LS23]. Multi-device [KWCK19]. multi-diagnosis [RSK^{+}24]. multi-dimensional [SGSS23]. multi-domain [CON09, PGMLK^{+}13]. multi-domains [GMH14]. multi-faceted [BFT08, QLOW09]. multi-factor [MNC20]. Multi-hop [BT07]. multi-key [PP24a]. Multi-keyword [OSSK16]. multi-language [HNHY23]. multi-layer [WR08]. Multi-Net [ZZW^{+}10]. multi-objective [PBI^{+}23, Zen22]. Multi-operator [EHM15]. multi-organizational [EOC^{+}24]. Multi-party [SFT^{+}24, BNTW12, KM03]. multi-phase [Das12]. multi-view [Kak24]. multiagent [ZGC07]. multicast [DWU^{+}23, MP15, PJ10, TWP08]. multicast-based [MP15]. multiclass [KEA24]. multicoupon [HIDFGHR19]. multifaceted [JBK21]. multifactor [IT05]. multimedia [MRCK24]. multiple [CC12, HMCD04, WMS^{+}19]. multiple-key [CC12]. multiplication [YOV09]. multiplications [HTM11]. multipliers [YOV09]. multiresolution [VSR15]. multiset [BA16]. multisignatures [HRL09]. must [ASN^{+}16]. mutable [MGAS24]. Mutual [HZL^{+}17, MNC20].

naive [Sen14]. named [JK22, ZP23]. names [CYA^{+}18]. Narrow [Tip22]. natural [SIA^{+}24]. NDCG [BDG23]. NDN [BHL^{+}21]. need [JTV19]. Negative [EFH09, EAH^{+}07, DNF^{+}19]. negotiation [KLMM09]. negotiation-based [KLMM09]. Neighbor [SAM^{+}24]. neighborhood [LWL^{+}21]. Net [ZZW^{+}10, WYL^{+}12]. Nets [BKBB20, BBB20, SSFB15]. Network [GPS17, MR24, AHT24, ATS23, ABK22, ABCC08, AuHD^{+}23, AAG24, BB22, BMFS24, BMP^{+}14, CLG23, CL09, DK24, EMRN17, GDA22, GBG18, GH05, JJJ21, JKLJ24, LHS^{+}24, LBZ^{+}10, wLW05, MARK20, MKS^{+}23, ME23, MYLZ14, MPG21, PAI^{+}23, PMW24, PVGG24, PMPGMLLM12, RA24, RSV23, RS18, SAM^{+}24, SGSS23, SSSJ24, SAT09, SRSM23, TLX09, TCE23, WYL^{+}12, WKWS24, YKP22, Zen22, Zha23, dAKdG10]. Network-based [GPS17]. Networked [MR03, BHZ^{+}21, SRD^{+}21]. networking [AFA^{+}23, JK22, MdMF22]. networks [ACB14, BBN24, BDG23, CMR06, CPPK15, Das12, DRPW12, DWU^{+}23, EHM15, EMRN17, FN19, Gol12, GDA22, GGP21, GBG18, KSB23, KIP22, LTC23, LH23, MLYL20, MTD^{+}24, PH24b, PH24a, RSH^{+}24a, RSH^{+}24b, RE24, RBEH15, RMSCR19, SS17, SSN15, SF17, SKH^{+}23, SDG22, TK24, WWZ^{+}23, YRW14, ZXZ^{+}11, BT07]. neural [BDG23, SAM^{+}24, WWZ^{+}23, WKWS24]. neuro [DK24]. neuro-fuzzy [DK24]. news [RB23]. next [AMLH18, EZLC21, EZLC21]. next-generation [AMLH18]. Next-Intent [EZLC21]. NICs [CBC08]. nightmare [EZLC21]. NIKE [FJZZ24]. NIZK [WT16]. NLP [BV24, MI22]. NLP-based [BV24]. no [BB04b, CMN^{+}18]. Node.js [KKK22a]. nodes [BHZ^{+}21]. noise [EG18]. noisy [BDH^{+}10, KML03]. Non [BV24, CDF^{+}13, CHZ16, CL09, EHSS14, GRV05, HHSS24, HYWS11, JT21, KBJ22, KM03, MP16, RMSCR19, SB22]. Non-Alpha-Num [BV24]. non-compressing [MP16]. non-cyber [SB22]. non-delegatability [HYWS11]. non-deterministic [RMSCR19]. non-interactive [CDF^{+}13, CHZ16, CL09, EHSS14, HHSS24, JT21, KBJ22]. non-repudiation [GRV05, KM03]. noninterference [BP04]. nonlinear [MB23]. notaries [MdAN23]. note [ZZH08]. Nothing [MTW^{+}14]. notifications [RSMT24]. notions [BCL09, BFS^{+}13]. Novel [Alg22, BV24, GCT24, KK22, KNL16, Lin15, MK21, MG23, MB23, NNL20, PVGG24, SGSS23, SHW23, WGH23]. NSDroid [LWL^{+}21]. nuclear [Alm24]. Num [BV24]. number [SBD23]. numerical [SMMN12, TNA23, ZO13].

OAuth [MP22]. Obfuscated [VCD21, BR20, VLGL24]. obfuscation [OSSK16, SXZC20]. object [HS09, SRD^{+}21, SS05b]. object-sensitive [HS09]. objective [PBI^{+}23, Zen22]. objects [DA21, RT23]. oblivious [BA16, HSMY12, LD17, TND^{+}15]. observatory [VSN22]. observer [YAY^{+}21]. obtain [Bra06]. OCFSDA [OPK24]. OCPP [ACT23]. Octopus [BLS^{+}23]. Off [BCD^{+}13, CJMS19, FAH24]. off-chain [FAH24]. Off-line [BCD^{+}13]. Off-line/on-line [BCD^{+}13]. office [AYM24]. officer [HSLB24]. Offline [MWZ06, LMG17, LBZ^{+}10]. OFMC [BMV05]. Ohgishi [CHZ16]. oil [ADHN24]. Okamoto [GMMV05]. Olympus [GdSdC24]. on-board [MMF^{+}24]. on-chain [FAH24]. on-line [BCD^{+}13, LMMO04]. on-the-fly [CL13]. one [GMS03, GDA22, MLCS16, YRW14]. one-class [GDA22, MLCS16]. one-pass [YRW14]. onion [CDF^{+}13, CFG17]. OnionDNS [SCL^{+}18]. Online [SSR22, AK24b, KB22, LMG17, LD07, Les24, LBZ^{+}10, PLA^{+}21, QLH^{+}24, SCY^{+}24]. online/offline [LMG17, LBZ^{+}10]. only [LD17]. Opacity [BKMR08]. open [AAG24, BAC24, EB24, ESSP24, Küs05, RM12, SVKV21, SKB23, VSN22]. open-ended [Küs05]. open-source [AAG24, SKB23, VSN22]. OpenFHE [PHE^{+}24]. opening [EHSS14]. operational [SGJC18, SGC22]. operations [BA16, HTM11, NRC15, VHRRMG24]. operator [EHM15]. operators [Oft24]. opportunities [AK24b]. Optimal [DRPW12, EG18, AK24a, GYL^{+}07, VSR15]. optimisation [PDB11]. optimistic [DGZFGH13, HYWS12, SEXY18]. Optimization [HBK22, MU18, DK24, ETRLG24, GZH^{+}23, KLZ^{+}21, LL14, PBI^{+}23, TGNA22, Zen22]. Optimized [OPK24, BMFS24, EKY^{+}24]. optimizer [SM24]. optimizer-based [SM24]. oracle [HYWS11, Yon18]. oracles [HSMW08, HYWS12]. OrBAC [EEB^{+}15, GBDJ14]. order [HM22, LM06, PSTS20]. organization [INS21]. organizational [EOC^{+}24, Asf24]. organizations [SFB^{+}24]. oriented [KKK22b, Moh23, SK16, WW07]. OSBE [HSMY12]. OSINT [BAC24]. OSN [RT23]. OSPU [Zha23]. OT [HSLB24]. Outbound [Smi04]. Outsourced [CG21, FHV18, CTM^{+}16, CTN24, LD17]. outsourcing [AL05, GYL^{+}07, HJDC15, KU16, TRT^{+}24, ZLZL20]. overlays [MLCQ21]. overview [BMP05, TS22]. Own [ACMV15].

P2ADF [KAK23]. P2P [MLCQ21]. P3 [BS22, YL19]. PAASO [TGNA22]. packers [AGL24]. packet [JK22, AM23]. packets [vORM06]. packing [BEPL^{+}17]. pads [dSFK19]. page [PD21, TSMH19]. PageDumper [PD21]. pages [PM23, SGLC19]. Paillier [NSNK06, DJN10]. Paillier-based [NSNK06]. pairing [MSGCDPSS18, YRW14, ZZH08]. pairing-based [MSGCDPSS18]. pairing-free [YRW14]. pairings [BCL09, CCS07]. papers [ACM05]. paradigm [BCD^{+}13, GRV24]. parallel [HBK22]. parallelism [SBB19]. parallelizable [MP16]. parameter [DTK^{+}18]. parameters [NMBB12]. parazoa [AMP12]. parking [BS21]. partial [CKW19, MRCK24, XZ21]. particle [TGNA22]. parties [HZL^{+}17, KPM12]. party [BNTW12, CAS22, KM03, LCL16, MR04, SFT^{+}24, SZ24]. pass [YRW14]. passenger [AG23]. Passive [CBC08, SS05a, ALOW15]. Passive-attack [SS05a]. Password [CHMS21, MPS10, MRW02, PCT22, AŠC24, AKT23, CG14, CJMS19, GTM11, HKO22, KJS17, KB23, Lop18, MSKD16, SB09, WLLW14, YM19, YL19]. Password-authenticated [CHMS21, MPS10]. password-based [HKO22]. passwords [CFBvO09, JTV19, YL19]. past [JG15, LMMP06]. patch [ZR24]. patching [JM17]. paths [BHZ^{+}21, ELPB24]. pathway [JKLJ24]. PatrIoT [SHOL23]. Pattern [IS23, FHV18, NLS24, OSSK16, SBB19, TS20, YA22]. patterns [CFBvO09]. pay [BS21, Roe11a, Roe11b, DZW^{+}18]. pay-by-phone [BS21]. pay-TV [Roe11b, Roe11a, DZW^{+}18]. payment [Les24]. payments [Les24, PPAHC24]. PBAC [Kud02]. PDGuard [MSKS20]. peer [BLS^{+}23]. peer-to-peer [BLS^{+}23]. PEKS [AKMW20]. Penalizing [MRCK24]. pentesting [HAS^{+}24]. Pepal [MRCK24]. perception [HSLB24, MTSH18]. Perceptions [Asf24, NABGJS^{+}24]. Perceptual [SB14]. Perfect [Hub12]. performance [AZ19, BBR18, BNTW12, DDZ22, HIDFGHR19, IDHRPCMP15, SK16, SSN15, VLGL24]. perimeter [PVGG24]. Periodicity [JM17]. permissions [ACS21, SA24, ZSN23]. permissions-based [SA24]. permutations [BR17]. persistent [AEHS24]. Persona [SSR22]. personal [IHNT02, MSKS20, QLZH15]. personalised [GL24]. personality [PPL15]. personalized [CG23]. perspective [AAM23, AAG24]. perspective-retrospective [AAG24]. perspectives [INS21, KC24]. perturbation [WWZ^{+}23]. Petri [BKBB20, SSFB15, BBB20]. phase [Das12]. Phish [PM23]. Phish-Sight [PM23]. phishing [AFF24, ETRLG24, FYF22, GCT24, PM23]. phone [BS21, CF07, GSM^{+}11, HCN15, WGMB13]. phylogeny [BDMM19]. physical [ASKG24, AGZA22, AK23, CK23a, CK23b, HXTP23, LWL23, PCK22b, SM10, XYZT24]. physically [BR17, BR20, KKY^{+}23, TK24]. PIN [dSFK19]. PINs [MTSH18]. PIOAs [Yon18]. Pioneering [PP24b]. PIR [DYDW10]. pirates [Nui12]. PiSHi [MBHT17]. PKI [BB04b, Daw04, LC04, LMMP06, LMMO04, SVKV21, VSM06]. placement [AK24a]. PLAID [DFF^{+}16]. Plaintext [MPS14]. plane [AM23]. plant [Alm24]. platform [IDHRPCMP15, MBRPS18, MG23]. platforms [KPM12, RHGTSC17]. play [VGG24]. playbooks [ESSP24]. playground [PC19]. PLC [JAYZ21]. please [PSTS20]. point [ABB17, KW15]. poisoning [BDG21, CSM^{+}24]. policies [ABCC08, ADHN24, ACMV15, AICC18, BCL13, BBB20, CF03, CCB08, KAC17, LBW05, LRB^{+}10, RRI^{+}19, SB09, SRD^{+}21, BKBB20]. Policy [SAL17, BZV05, GMH14, JSMG18a, JSMG18b, KNL16, MS15, TG05, XSA13, dAKdG10]. PolicyUpdater [CZ06]. pollution [EMRN17, KKK22a]. Polly [SGE02]. polymorphic [KJG^{+}11]. polynomial [PCK22a, SGE02, TND^{+}15]. polynomial-based [SGE02]. polynomials [GMS03]. popular [FBFGEM21]. Portfolio [LL14]. position [AHT24]. post [TSZ22]. post-quantum [TSZ22]. posteriori [ACBC^{+}15]. postfix [YL19]. postgraduates [AHX^{+}23]. Posthumously [SSR22]. potential [Inc24, WGMB13]. power [AKZM20, AEHS24, Alm24, DPMC24, HG24, MD23a]. powered [EFB^{+}24]. PPAM [PH24b, PH24a]. PPAM-mIoMT [PH24b, PH24a]. practicability [IDHRPCMP15]. Practical [DDX19, LLW^{+}16, ALOW15, GKS19, HRMM20, KOSU16, KBY22, LCL14, MI22, Pen12, SHOL23, VHT09, WR15]. practice [PPAHC24]. practices [LD07, SCY^{+}24]. Pragmatic [KDM22]. pre [CMR06, KMW24, MARK20, Pen13]. pre-computation [Pen13]. pre-distribution [CMR06, MARK20]. pre-trained [KMW24]. precise [AAB20, IKS22]. prediction [KCC^{+}23, NK24]. predictive [LL22]. Preface [Ano11a, ACM05, BGP07a, BM05, Daw04, DV08, DMRS07, Pri04, Sne05, Wai04, ZL06]. prefix [BGKZ12, YL19]. presence [ZTV23]. present [LMMP06]. preservation [LLH21, RAC16]. preserve [EEB^{+}15]. preserved [JSMG18a, LD17]. Preserving [CMH^{+}24, ABFO08, BEY23, BGK08, BLS^{+}23, BBG24, BSK^{+}20, BS21, CTM^{+}16, DFBJR18, GGJ22, GKS19, HLS18, HRMM20, KAK23, KOSU16, KB13, KNL16, MB16, MCD11, NST09, NA14, PH24b, PH24a, QLZH15, RSH^{+}24a, RSH^{+}24b, RDK18, SZ24, SSL22a, STD21, TMvM24, Tor20, WMS^{+}19, YWW22, Yan21]. Press [VGG24]. Preventing [RAC16, PS17, YP12]. prevention [Alg22, LH23, MVA24a, MAZS24, PSDSNAHJ19, VL13, YAY^{+}21]. primitives [MP16, SM10]. principle [Bel10]. Principles [CGL^{+}11]. prioritization [SDG22]. Privacy [AaBZS23, BEY23, KB13, MB16, MSP^{+}13, NST09, QLZH15, RDK18, SZ24, SSL22a, UBK23, ACCS24, ABFO08, AZS24, AJC^{+}09, AMRR17, BGK08, BLS^{+}23, BBG24, BSK^{+}20, BCA^{+}10, BDHZ15, BS21, Bra06, CR20, CTM^{+}16, EG18, EAH^{+}07, GL24, GGJ22, GLMS^{+}04, GGP21, GKS19, GYL^{+}07, GP23, HRMM20, KNM^{+}24, KAK23, KOSU16, KMW24, KBH07, KNL16, LD07, LD17, LcSCL^{+}18, MCD11, MBRPS18, NA14, OAJ23, PH24b, PH24a, PGMLK^{+}13, PSA23, RSH^{+}24a, RSH^{+}24b, RA24, RSPMB16, RMPADF13, SAC22, ST24, STD21, Tip22, TMvM24, Tor20, TGNA22, UGP24, ULEY24, WMS^{+}19, YWW22, Yan21, YSD^{+}20, YAM^{+}15, ZHZ22, ZZG19]. privacy-aware [MBRPS18, RSPMB16, TGNA22]. Privacy-enhanced [MSP^{+}13]. privacy-preserved [LD17]. Privacy-preserving [BEY23, KB13, MB16, NST09, QLZH15, RDK18, SZ24, ABFO08, BGK08, BBG24, BSK^{+}20, BS21, CTM^{+}16, GKS19, HRMM20, KAK23, KOSU16, MCD11, NA14, PH24b, PH24a, RSH^{+}24a, RSH^{+}24b, STD21, TMvM24, Tor20, WMS^{+}19, YWW22, Yan21]. Private [BA16, DMP13, BDD01, BGP07b, DA21, HSHM24, KW15, LMD17, NMH^{+}24, PHS22, PGMPPC22, TMM^{+}19, VH19, WR15]. privilege [FTS^{+}20, KY24, PMDS23, PSOMS22, YAY^{+}21]. Probabilistic [DHW11, BP04, GYL^{+}07, JJQ24, MCD11, PJ10]. probability [BHZ^{+}21]. problem [CC12, GOBdlC11, GMS03, GP17, KG11, RG13, TWP08, YSM10]. problems [HMCD04]. process [BDMM19, HBH12, LLG22, NAQ24, Yan21]. processes [RHL17]. processing [MSKS20, SIA^{+}24]. processor [KKK22b, TKKO20]. production [AKA24]. products [BAG^{+}24]. profiles [EFB^{+}24]. profiling [LCPD14, NLS24, YA22]. program [BDF04, HS09]. programmability [Yon18]. programmable [Smi04]. programme [OCG24]. programming [AŠC24, RRI^{+}19, WZ07]. programs [KM07, WGMB13]. progressive [WPD18]. project [BFP03]. proof [HLS18, WLLW14]. proofs [BCJ^{+}11]. propagation [IKS22]. property [TS20]. Proposal [IHNT02, IT05, FGH24]. proposals [BJ16]. protect [TS22]. protected [BJ16]. Protecting [EAH^{+}07, MSSO^{+}24, SAH22]. Protection [AM23, BZ20, AJC^{+}09, GKBS12, KK17, KMW24, MGRR19, MOA^{+}24b]. protections [VCD21]. protocol [AAV22, AGZA22, Ala17, ALPW13, Bel10, BFT08, CL09, CHC24, DFF^{+}16, DGZFGH13, FJZZ24, HL04, HYWS12, KIP22, LSWW14, LH23, MG19, ML17, MNC20, MTD^{+}24, RG13, RGL16, SAM^{+}24, SSSJ24, SDR20, YRW14, YAM^{+}15, ZBC23]. Protocols [DHS04, LSG22, LSV^{+}23, AC08, ABFL12, BGK08, BMV05, BBR18, Bra22, BNN04, CCS07, DVB02, DS07, EB24, Fra18, GLP03, GRV05, Hal20, HLS18, HHSS24, KM03, Küs05, LCL16, MP15, MUiH^{+}21, MSN02, MS14, QQK^{+}24, SSFB15, Ust11, Vaj16, VdWZ14]. prototype [KKK22a, TKKO20]. Provable [YRW14, NSNK06]. Provably [FRG19, LCL14, RG13, VHT09, YYK^{+}18]. provenance [Abb13]. provided [TS22]. providing [PGMLK^{+}13]. Provision [Kud02]. Provision-based [Kud02]. provisioning [TGS17]. proxy [LL22, Lin15]. pseudo [HIST09]. pseudo-freeness [HIST09]. PSO [TGNA22]. Public [Bra22, SXZC20, AKMW20, ALOW15, BCJ^{+}11, BZ03, CTN24, DJN10, EHSS14, FRG19, GW09, GMdFLR23, HBF^{+}24, LC04, Pen12, SSP14, Gri06]. public-key [AKMW20, ALOW15, BCJ^{+}11, DJN10, EHSS14, FRG19, SSP14]. Publisher [PH24b]. publishing [AAM23, SAC22]. PUF [AHC^{+}21]. push [BFPP07]. push-based [BFPP07]. pushdown [BCL13]. puzzle [WR08]. puzzle-based [WR08]. PVSS [Pen12].

QCA [AK24a]. QCA-based [AK24a]. QR [HZL^{+}17]. QR-code [HZL^{+}17]. quality [SBCP21]. quantitative [JBK21, SKY23]. quantum [Sat20, TSZ22]. queries [HSHM24, MRG23]. query [BB04a, KCM^{+}15]. queue [BF13]. QUIC [CKKK23]. quick [WT21]. quotient [PCK22a].

R [SKLP20]. race [ANN23, sLC05]. racer [ANN23]. Radio [ATN^{+}24]. RADS [SKH^{+}23]. RAM [LD17]. RAM-based [LD17]. RAMA [SFB^{+}24]. ramp [LMD17]. Random [ASA23, PP24a, Tor20, BR17, BR18, Das12, HSMW08, HYWS11, HYWS12, UMN^{+}20, VSR15, Yon18]. randomization [PCK22b]. Randomized [ML14]. randomness [BGP07b, Inc24]. range [HSHM24, KMMG23, MRG23]. rank [BDG23]. ranked [OO20]. ranking [JJQ24]. ransom [PPAHC24]. ransomware [Alm24, ASFG24, CKPM24, CMN^{+}18, EKY^{+}24]. ratcheting [GMSS23]. rate [AM23]. rational [ETAHCR08]. RBAC [BGTCCBB10]. re [GI19, Dan07]. Re-encryption [Dan07]. re-identification [GI19]. reacting [AGIK07]. reactive [BDHK08]. Reactively [BPW05a]. reader [GdKGV14]. Real [CK23b, CKPM24, GDA22, KCB17, RLEM18, SBS23, SKH^{+}23, TSZ22, CK23a]. Real-time [CK23b, CKPM24, GDA22, KCB17, SBS23, SKH^{+}23, CK23a]. real-world [TSZ22]. realistic [BB22, HS15]. realization [KAC16]. realize [AKMW20]. realizing [RD16, ZZG19]. really [BM11]. reasoning [GMMZ06]. receivers [AP22]. reciprocal [BDG23]. recognition [GCSÁBdSS12, NLS24, ZP23]. Recommendation [PBI^{+}23, OAJ23]. recommendations [YNC22]. recommender [ABFO08]. record [QLZH15]. records [QDW^{+}15]. recovery [CMS10, JAYZ21, NVB^{+}02]. rectangle [Lu09]. recursive [Rus04]. Redistributing [LKH09]. reducing [AYM24]. reduction [KKY^{+}23, PMW24, WT21]. redundancy [JAYZ21]. references [ZD22]. refinement [dAKdG10]. reflects [PPL15]. refunds [KO02]. regions [EG18]. regression [MB23]. Regular [Bae10, AYM24, WR15]. regular-expression [WR15]. regulated [BS21]. reinforcement [AFF24, EZ22, ME23, SRK^{+}20, SEZ24]. rekeying [PJ10, TDRR20]. Related [Lu09, AaBZS23, Dan07]. Related-key [Lu09]. relation [MS11, Pen11]. Relations [FGS12]. relationships [CTM^{+}16, RAC16]. relaxed [BFS^{+}13]. release [DA21, NAM06]. reliable [ABCC08, AvO13]. relocation [KY24]. remanence [SV11]. remedy [SCY^{+}24]. remote [AYM24, CGL^{+}11, GW09, KK22, MMS16, YAM^{+}15]. renewable [BDH^{+}10, RCC^{+}24]. reordering [WT21]. Repairing [GLMS19]. repeated [KLZ^{+}21, vORM06]. Replacement [XCW^{+}12]. reporting [MPP21, SDW23, SBG22]. reports [Alq24]. representations [EFH09]. repudiation [GRV05, KM03]. reputation [LI07, SSV22, SdHZ16]. Requirements [GMMZ06, WW07, CSL^{+}23, FFG20, SK16]. resampling [ASA23]. Rescue [BST21]. Research [Gri06, AZS24, BAC24, GDA22, NABGJS^{+}24, RM12, SHOL23]. residence [PPL15]. resident [BMP^{+}14]. resilience [ASKG24, SAH22, TV24]. resilient [MR03, WT16, XZ21]. resistance [TGS17]. resistant [CHKO12, CLO^{+}24, LL21, SCL^{+}18, MS09]. ResNet152 [SC24]. resolutions [RV03]. Resolving [CTM^{+}16]. resource [LVK18, SS17]. resource-bounded [LVK18]. resource-constrained [SS17]. resources [GMdFLR23, TZH04]. response [SSD14]. responsibility [Asf24]. Restricted [SF17]. Restricting [RT23]. results [BCD^{+}13, DLR15]. retention [GSM^{+}11]. retraining [KCC^{+}23]. retrieval [BDD01, HSHM24, LMD17]. retrospective [AAG24]. reusable [KO02]. reveal [GSM^{+}11]. Reverse [MLM19, EAH^{+}07, BZ20]. reversed [KYH18]. review [AKR24, ATS23, BAC24, CKKK23, DPMC24, EB24, HG24, KSB23, LHS^{+}24, MVA24b, OAJ23, PCSA24, RB23, SAC22, SA24, TV24, UGP24, VMCR23, ZR24]. revisited [ABPR24, BZ03, BRS06, BVS07, BCD^{+}13, CHZ16, GMMV05, RSD19, RDK18]. Revisiting [CKKK23]. revocable [JCL^{+}18, Lin15]. revocation [HRMM20, MFES04, NST09, QLZH15, ZJS22]. revoke [NP10]. revoked [RSMT24]. revoking [DdP13]. RFID [ALOW15, SV11, Tip22]. rhythm [VGG24]. rich [RMSCR19]. right [Bae10, Lop18]. right-to-left [Bae10]. rights [ASF04, LKH09, RSMT24]. Rigorous [GH05]. ring [GCH^{+}19, JT21]. rings [PCK22a]. rise [AZS24]. Risk [INS21, MYLZ14, NBA^{+}21, AK23, EOC^{+}24, AG23, EAM^{+}23, HXTP23, MTSH18, ORK23, RV19, SSD14, SFB^{+}24, SCO21, SDG22, TNA23, WHS18, ZTG22]. risk-based [SCO21]. risk-level [ZTG22]. risks [AYM24, GYL^{+}07, HSLB24, JBK21]. RL [PKHS23]. RLET [DJS^{+}24]. road [GBC24]. Robotics [YNC22]. Robust [AŠC24, GKBS12, KNM^{+}24, MS11, SRK^{+}20]. robustness [ZWX20]. role [BSK^{+}23, CK08, QQK^{+}24, ZVH15]. role-based [CK08, ZVH15]. ROM [MLM19]. RORI [GBDJ14]. RORI-based [GBDJ14]. round [ABM^{+}12, BGP07b, GNS14, NSNK06]. rounds [Lu09]. Routing [SSN15, BT07, CDF^{+}13, CFG17, KCB17, LH23, MG19]. RPCAE [Lin15]. rPIR [LMD17]. RSA [LC04, MPS10]. rule [KAC16, OBH^{+}20, QLH^{+}24]. rules [ABCC08, HBK22, KAC17]. run [AGL24, LBW05, PD21]. run-time [AGL24, LBW05, PD21]. runtime [DLR15, KDYS19]. RWArmor [ASFG24].

safeguarding [BCA^{+}10]. safer [CMMPS15]. safety [AaBZS23]. safety-related [AaBZS23]. sail [BB04b]. SAINT [VSN22]. Sakai [CHZ16]. saliency [ZGRS23]. salp [PV22]. SAML [EWR^{+}09]. samples [SSVC16]. Sandbox [SSE^{+}15]. sanitization [ZLGZ19]. SARSA [ME23]. SARSA-based [ME23]. SAS [KJG^{+}11]. SAT [AC08]. SAT-based [AC08]. satellite [KW15]. satisfaction [KC24]. Saudi [Asf24]. saving [YZ24]. SCADA [Ano11b, HBH12]. scalability [PVGG24, TGS17]. Scalable [HKO22, RMSCR19, HRMM20, KCB17, MK21, SB14, YYK^{+}18]. scale [Das12, RLEM18, SDG22, TLX09, ZWX20]. scan [AvO13]. scanning [CBC08]. scattered [SBG22]. scenario [ASZ^{+}23, SRD^{+}21]. scenarios [HS15, PGMLK^{+}13]. scheduling [PNG^{+}20]. scheme [AaBZS23, AKT23, BZ03, BDD01, BCL09, CG14, CMR06, CC10, Das12, DSY06, EMRN17, FTS^{+}20, HCN15, IS23, JT21, KOSU16, LLH21, Lin15, MD23b, NT20, PDB11, QDW09, RSH^{+}24a, RSH^{+}24b, RHL17, RA24, RSD19, SHA20, TK24, WLT^{+}24, WLLW14, YWW22, YYH^{+}23, ZLZL20, ZLJW20]. schemes [BPW05a, BR18, CJMS19, DdP13, Dan07, DHS04, Den08, GP17, HSMY12, HYWS11, NP10, PBI^{+}23, TMM^{+}19]. score [SSV22]. scoring [OSSK16]. SDN [AM23, AB24, LWL23, MR24, MLCQ21, Moh23, MLC23, SHW23]. SDN-assisted [LWL23]. SDN/Fog [Moh23]. SDN/Fog-based [Moh23]. SDSI [LM06]. SE [MG19, SC24]. SE-AOMDV [MG19]. seafarers [MOA^{+}24a]. SealFSv2 [GMSS23]. search [AKMW20, CKS21, DK24, FRG19, HH16, KJS17, OO20, OSSK16]. searchability [HJDC15]. searchable [CTN24, CHMS21, HK19, LZQ^{+}18, MRG23]. searches [WR15]. searching [VH19]. seas [MSSO^{+}24]. Seberry [BZ03]. second [ABM^{+}12]. secrecy [CDF^{+}13, Hub12]. secret [CG21, DdP13, GMS03, HJDC15, LMD17, PPSS13, QDW09, SFT^{+}24]. secrets [BW08, JTV19]. Secure [ABB17, AL05, BVS07, CKW19, CKS21, CSL^{+}23, DZW^{+}18, FTS^{+}20, HN14, HSMW08, KW15, KM22, LL21, LCL16, MSKD16, PMDS23, SJ09, SBB19, TRT^{+}24, UMN^{+}20, ZXZ^{+}11, ASF04, Ala17, BPW05a, BSK^{+}20, BDF04, BNTW12, BT07, CSS^{+}24, DdP13, DWU^{+}23, DDX19, FZ21, FRG19, GA23, KME^{+}16, KAK22, KU16, KSZ07, LCL14, LLW^{+}16, MG19, MB16, MLO^{+}04, MBRPS18, MPP14, MD23b, MSKS20, MG23, Moh23, NT20, NMBB12, Nui12, OO20, PJ10, PBI^{+}23, QDW^{+}15, RG13, SJ10, SK16, Smi04, TND^{+}15, WPD18, YL20, YYK^{+}18, YAM^{+}15]. secured [BS22, DJS^{+}24, DK24, EHM15]. Securesdp [PVGG24]. Securing [AYM24, DDPS02, LLG22, MPG21, ZBC23, BFPP07, CC10, PDM20, PH24b, PH24a, PCK22b, SIA^{+}24]. Security [AGZA22, Alq24, Ano14, BCL13, BDHZ15, FJZZ24, FSG^{+}14, GMH14, HJDC15, OO20, OT06, PNG^{+}20, SB22, WLT^{+}24, YWW22, ZZW^{+}10, AK23, ACB14, ABCC08, ADHN24, AZS24, ANN23, AK24a, AHX^{+}23, AŠC24, ABM^{+}12, AYM24, Ano11b, AC08, Auf20, AICC18, BCJ^{+}11, BGKZ12, BMV05, Bel10, BKBB20, BBB20, BCF^{+}17, BFT08, BCL09, BFS^{+}13, BNN04, BCEM04, CDP22, CAS22, CLW^{+}11, CKKK23, CFBvO09, CG23, CYK09, CLPP11, CCB08, DSB19, DM07, Den08, DRPW12, Dwi24, EHM15, FGS12, GL24, GS15, GLMS^{+}04, GYL^{+}07, GMLM23, GRV05, GH05, HS15, HHSS24, HSLB24, HMCD04, INS21, Inc24, JBK21, JGK14, KBJ22, KMMG23, KLPL21, KIP22, KKKV07, KBH07, KLMM09, KMR09, KCC^{+}23, KC24, LHS^{+}24, Lan01, LBW05, LP11, LV10, LS23, wLW05, MBH^{+}21, MdSC^{+}15, MP15, MdMF22, MSGCDPSS18, MTD^{+}24]. security [NABGJS^{+}24, NSNK06, OBDM^{+}24, OCG24, PVGG24, PDB11, QQK^{+}24, RM12, RG13, RGL16, SK16, SS17, SCY^{+}24, SKY23, SSN15, SBG22, SM10, SAT09, TG05, UGP24, ULEY24, VSM06, VBAG24, VD24, WYL^{+}12, WHS18, YNC22, YLL^{+}18, YRW14, YSD^{+}20, YP06, YM19, ZKP24, Zha23, ZGRS23, ZM07, dAKdG10, vOLW05, BJ15, Pri04, KBH07]. security-centric [CAS22]. Security-enhanced [WLT^{+}24]. security-mediated [VSM06]. security-sensitive [HS15]. SEDS [NT20]. segment [SZ24]. seizure [SCL^{+}18]. seizure-resistant [SCL^{+}18]. selected [ACM05]. Selecting [NMBB12]. selection [GBDJ14, OPK24, SBS23, SJ10, SKY23, SSVC16, ZGC07, ZIR24]. selective [ZZG19]. selective-ID [ZZG19]. self [FPP^{+}24, GMLM23, KK22, MVA24a]. self-defense [KK22]. self-prevention [MVA24a]. self-sovereign [FPP^{+}24, GMLM23]. Selfish [MVA24b, MVA24a]. SELinux [XSA13]. sellers [ZGC07]. semantic [YKP22]. semantic-aware [YKP22]. semantics [KJG^{+}11]. semi [Zha23]. semi-network [Zha23]. sensitive [HS15, HS09, RAC16, STW^{+}23, TS22, ZO13]. sensor [AAMH^{+}24, AAZAA23, CMR06, Das12, DK24, KCM^{+}15, LBZ^{+}10, MARK20, OBH^{+}20, RSH^{+}24a, RSH^{+}24b, RBEH15, TK24, YRW14, ZXZ^{+}11]. sensors [CSC^{+}23, CLW^{+}11, MTSH18]. sequence [AL05, BZV05, BSK^{+}23, KSM10]. sequential [BDPV14]. series [KKK22b]. server [Bra22, MYLZ14, NT20]. server-aided [NT20]. service [BSK^{+}20, LLWY09, NA14, WR08, WW07, TGS17, TMvM24]. service-oriented [WW07]. services [AV17, DDPS02, EWR^{+}09, GGJ22, KBH07, MdAN23, MS15, PMPGMLLM12, SHA20, SKY23, VL13]. session [LL22, SCO21]. sessions [ACB14]. set [BA16, MCD11, RRI^{+}19, WZ07, BMP05]. Sets [SAL17, WCS20]. setting [IS23]. settings [AVM23]. severity [CR24]. SFCGDroid [STW^{+}23]. SHA [ABM^{+}12, KJS17]. SHA-1 [KJS17]. SHA-3 [ABM^{+}12]. Shamir [HHSS24]. shareable [AKA24]. shared [BS05]. sharing [CMMPS15, CG21, HJDC15, KNM^{+}24, LKH09, LMD17, QDW09, SFT^{+}24, SSL22a]. sharing-based [HJDC15, LMD17]. ships [AG23]. Short [Nui12, VLGL24, JSMG18b, LMG17, SK14, ZKP^{+}23]. Short- [VLGL24]. short-term [ZKP^{+}23]. shortcut [MLM19]. shot [SLGP23, WZZ^{+}24]. Shoulder [WLLW14]. Shoulder-surfing-proof [WLLW14]. shuffle [DYDW10, Pen13]. shuffles [NSNK06]. shuffling [PDB11]. side [ANN23, CBRY20, dSFK19, HS15, KDYS19, MTW^{+}14, SB22]. side-channel [CBRY20, dSFK19, KDYS19, MTW^{+}14, SB22]. SIEM [VMCR23]. Sight [PM23]. Signature [JMV01, TMM^{+}19, AAG24, BPW05a, DHS04, EHSS14, GP17, HSMW08, HYWS11, JT21, KJG^{+}11, LBZ^{+}10, LWL^{+}21, RSH^{+}24a, RSH^{+}24b, RD16, RSD19, Yon18, ZBD06]. signature-based [AAG24]. signatures [ACHO13, BCD^{+}13, GCH^{+}19, IDHRPCMP15, IT05, KJ14, MR04, PS17, SEXY18, SK14, WPD18, YSM10]. signcryption [LMG17, RD16]. signers [BRS06]. signing [FGH24, IZS08, TSZ22]. SilentKnock [VHT09]. simple [BR20, MD23b]. Simplified [BCL09]. simulatability [BDHK08]. simulatable [BPW05b]. Simulating [SEZ24]. Simulation [ABPR24, BB22, SB09, WT16]. simulation-extractable [WT16]. simultaneous [Man21]. sinkhole [BBN24]. sinking [YZ24]. size [DFBJR18, PPSS13]. Skype [DBMS10]. SLE [vOLW05]. slicing [MS15]. small [MP16]. smart [AKZM20, AT22, AMZ22, AZ22, BAB23, DMDD16, DJS^{+}24, FPP^{+}24, FJZZ24, FBFGEM21, GdKGV14, GLMS19, HKO22, JT24, MSP^{+}13, MPP14, MGAS24, MdAN23, MMF^{+}24, PNG^{+}20, PP24b, RCC^{+}24, RDK18, SSL22b, SRD^{+}21, SSV22, SAH22, WLT^{+}24, Yan21, GAdFGM23]. SmartiPhish [AFF24]. smashing [MGRR19]. SMOTETomek [TSU^{+}24]. smudge [SSL22b]. SNARK [ABPR24]. SOAP [DDPS02]. social [AK24b, DMP13, FN19, KPM12, KB22, LTC23, MAZS24, RAC16, Tor20]. SOCs [Oft24]. software [AKA24, BAG^{+}24, GBC24, JM17, KSB23, LNX22, MLO^{+}04, MdMF22, MTD^{+}24, PVGG24, RE24, SK16, SK06, SKH^{+}23, VCD21, XCW^{+}12, ZGK07]. software-defined [KSB23, MdMF22, MTD^{+}24, PVGG24, RE24, SKH^{+}23]. solution [AFF24, MCD11, PMPGMLLM12, SK16, SFB^{+}24]. solutions [HIDFGHR19]. Solving [GOBdlC11, GP17, HMCD04]. some [BR20]. SonarSnoop [CBRY20]. sound [BCJ^{+}11, BDPV14]. soundness [BP08]. source [AKA24, AAG24, BAC24, SKB23, VSN22, ZR24]. sources [CF03]. sovereign [FPP^{+}24, GMLM23]. Space [MBH^{+}21, PCK22b]. spam [AIJM24, ASN^{+}16, BABB16, EFB^{+}24, SSM^{+}20]. spammer [ASN^{+}16]. Spanish [NABGJS^{+}24]. Spark [MAZS24]. Spark-blockchain [MAZS24]. sparrow [DK24]. sparse [ACB14]. Spatial [AZM24, AVM23, DA21, ZMS22]. spatio [SKK^{+}17]. spatio-temporal [SKK^{+}17]. spatiotemporal [ZMS22]. Special [Ano11b, BJ15, KBH07, YSD^{+}20, ACM05, BM05, Daw04, Pri04, Sne05]. specific [KME^{+}16, LSR^{+}23, ST24]. specification [TS20]. specifications [ZGK07]. Specifying [BGK08]. spectrum [CSC^{+}23]. speech [AK24b]. spending [PSDSNAHJ19]. spirit [PSOMS22]. SPIT [GKBS12]. SPKI [LM06]. SPKI/SDSI [LM06]. sponge [AMP12]. spoofed [AFF24]. spoofing [Hub12, MTD^{+}24]. spotlight [ACT23]. spread [WGMB13]. SpyDetector [KDYS19]. spyware [Alm24]. SQL [DSB19, Kak24, SEZ24]. squaring [HTM11]. SRAM [KKY^{+}23]. SRAM-based [KKY^{+}23]. SRU [CLG23]. SSH [LL22]. SSL-protected [BJ16]. SSO [PMPGMLLM12]. SSPFA [MGRR19]. stabilizers [BR20]. Stack [BDF04, MGRR19]. stacking [AIJM24, AuHD^{+}23]. stacking-based [AuHD^{+}23]. stage [AEHS24]. stages [ASZ^{+}23]. standard [Ala17, CH16, RSD19, SSP14]. standards [DFF^{+}16, ESSP24]. star [RS18]. Start [ELPB24]. State [vOLW05, ASKG24, AK24b, PPAHC24, TDGL23]. Stateful [LMMS17, YLL^{+}18]. Static [SS05b, AAB20, ASFG24, CMS10, MS15, SCO21, SSE^{+}15, ZM07]. static-informed [ASFG24]. statistical [NLS24]. status [HC10]. Std [KDM22]. steal [SGLC19]. Stealing [MTSH18]. stealth [KK22]. steganography [Des09, HS24, SJ09, SJ10]. steps [ASZ^{+}23]. stepwise [BMFS24]. sticky [SRD^{+}21]. Stochastic [WYL^{+}12]. storage [FAH24, GW09, GMSS23, LZQ^{+}18, NT20, YAM^{+}15]. storage-based [GMSS23]. storages [CTM^{+}16]. storing [HK19]. STORK [RLEM18]. STR [IHNT02]. strategies [wLW05]. strategy [SFT^{+}24]. stream [GDA22, HL04, IMI18, TWP08]. stream-based [GDA22]. streaming [SCY^{+}24, ZO13, ZHZ22]. Street [MD23b]. strength [KB23, RGL16]. STRIDE [ZTG22]. STRIDE/DREAD [ZTG22]. string [KCM^{+}15]. Strong [CHKO12, XZ21, HYWS11, YLL^{+}18]. stronger [PDB11]. Strongly [WT16]. Structural [KA18, GI19]. structure [QDW09, TTS^{+}06]. structures [KN07, Küs05, RD16, Rus04]. study [ADHN24, BSCZ11, BKBB20, BBB20, BDG21, BF13, CKKK23, CSM^{+}24, DSB19, DSRHC16, Dwi24, GL24, GLMS^{+}04, IS23, KB22, KC24, OBDM^{+}24, OBH^{+}20, SCY^{+}24, SKY23, SKK^{+}17, TS20, VGG24, XSA13, ZD22, ZHZ22]. Studying [WT21]. style [BPW05b, BP08]. subject [LcSCL^{+}18]. subjective [SdHZ16]. subset [FRG19, KG11]. subspace [RG13]. substation [AEHS24]. substitution [AP22, BRS06]. subsystem [VSN22]. subtree [Roe11a, Roe11b]. Sufficient [BDPV14]. suffix [BGKZ12]. suffix-free-prefix-free [BGKZ12]. suitable [Pla09]. sum [KG11]. supersingular [HM22]. Supervised [GSAMCA18, CSM^{+}24, SKB23]. supply [AKZM20, AHC^{+}21, HAAI24]. support [CON09, MLCS16, ZGRS23]. supporting [ARMLS06, HSHM24, JSMG18a, JSMG18b, MRG23, RV03]. supports [WR15]. surface [JKLJ24, ZD22]. surfing [WLLW14]. surveillance [Lev07, LcSCL^{+}18, MBRPS18, RSPMB16]. Survey [MOA^{+}24a, ZMS22, ATN^{+}24, ABK22, AGL24, AK24b, AB24, Den08, FSG^{+}14, GCT24, HGH23, JK22, MD23a, MAZS24, QQK^{+}24, TSZ22, UBK23, VHRRMG24]. Survey-based [MOA^{+}24a]. surveys [AAM23, Ano22]. suspect [GBG18]. sustainable [DJS^{+}24]. Swarm [HLA^{+}24, KKK17, PV22, TGNA22]. Swarm-intelligence [HLA^{+}24]. swarms [SS17]. switching [SHW23]. symbolic [BHKM23, BMV05, BFT08]. Symmetric [BPW05b, Bra22, HK19, MRG23]. Symposium [BJ15]. Synthesizing [NMH^{+}24]. Sysmon [SKB23]. System [JTV19, ATS23, AMZ22, ABFO08, Alg22, AZ22, Ano11b, AKA24, BLS^{+}23, BEd23, BFP03, BFPP07, BR23, BS21, BSV22, BCGL23, CL08, CLO^{+}24, CKPM24, CZ06, DJN10, EZLC21, ETRLG24, FAMMZ23, GLMS19, GdSdC24, Ham23, IZS08, KO02, KJS17, KBY22, KNSS24, LKH09, LCL14, MGV17, Man21, MMF^{+}24, MK21, ME23, MFES04, MS15, NK24, PGMPPC22, PV22, RM12, RLEM18, RBEH15, SAM^{+}24, SBS23, SC24, SBD23, SRK^{+}20, SSD14, SSJ22, TS20, VBAG24, VD24, VL13, YAY^{+}21, ZWX20, ZTG22, vORM06, DNF^{+}19]. System-assigned [JTV19]. system-based [EZLC21, MGV17]. Systematic [VMCR23, AKR24, ATS23, BAC24, EB24, HS24, MVA24b, OAJ23, OBDM^{+}24, PCSA24, SAC22, SK16, TS22, UGP24, ZR24]. systems [ASF04, ASKG24, AK23, Alm24, AMRR17, AAG24, BDHZ15, BHZ^{+}21, BB04a, BCF^{+}17, BKMR08, CK23b, CR24, DSRHC16, GMH14, GMLM23, HSLB24, HZL^{+}17, Hub12, JAYZ21, LVK18, LLH21, LWL23, LV10, MGAS24, MB23, OAJ23, OBDM^{+}24, PH24b, PH24a, PCK22b, QLH^{+}24, RSPMB16, Roe11a, SK16, SV11, SS05a, SDR20, SB22, TBGB20, TMvM24, XYZT24, ZGC07, ZGRS23, ZIR24, dAKdG10, CK23a, Roe11b].

table [PD21]. tables [PP24a]. TabNet [CLG23]. Tacit [JTV19]. tag [EMRN17]. tagging [GP17]. tags [ACHO13, ALOW15, LL21]. taint [IKS22]. Taking [BRS06, KNL16]. tallying [MMS16]. Talos [CMN^{+}18]. tamper [GMSS23]. tamper-evident [GMSS23]. tampering [DTK^{+}18]. target [BB22, BHZ^{+}21, HXTP23]. targeted [GBG18, KOSU16, MLCQ21, UBK23]. targets [GCT24, XYZT24]. task [Oft24, Yon18]. task-PIOAs [Yon18]. taxonomies [RB23]. taxonomy [CSL^{+}23, GCT24]. TCP [FZ21]. TCP/UDP [FZ21]. TCP/UDP-based [FZ21]. technical [GMdFLR23]. technique [GLP03, Rov23, SSVC16, WT21]. Techniques [TG05, ATN^{+}24, AIJM24, AGL24, BMP^{+}14, CH16, GCT24, Inc24, KK22, KA18, LSR^{+}23, MI22, OAJ23, RE24, SMMN12, SSE^{+}15, ZMS22]. technologies [KBH07, SAT09, WLT^{+}24, YSD^{+}20]. technology [LH15, MVA24a, SSV22, VMCR23]. telecare [LLH21]. telemetry [NBA^{+}21]. telemetry-enabled [NBA^{+}21]. telephony [AMRR17, BABB16]. tell [MBHT17]. templates [CC10, HTM11]. temporal [AVM23, DA21, SBD23, SKK^{+}17, TS20]. Temporarily [GSS10]. tenant [ELPB24]. TENET [TCE23]. term [VLGL24, ZKP^{+}23]. TermID [KKK17]. terms [MUiH^{+}21]. test [LLBL18]. testbed [RM12]. testing [ANN23, AV17]. tests [CSL^{+}23]. tetromino [VBAG24]. tetromino-based [VBAG24]. text [ZSN23]. their [Auf20, HSMY12, KAC17, PNG^{+}20, SB22, WGMB13]. theoretic [LVK18, PNG^{+}20]. theoretical [BCEM04]. theoretically [TND^{+}15]. Theory [Rus04, CG23, ETRLG24, MTD^{+}24]. theory-based [CG23, MTD^{+}24]. there [ACS21]. Things [AAMH^{+}24, Alh24, CK24, EB24, Ham23, MOA^{+}24b, OPK24, SM24, VD24]. thinking [ELPB24]. third [CAS22, KPM12]. third-party [CAS22]. threat [AEHS24, ASA23, AG23, AKA24, Auf20, GMLM23, IOU^{+}21, KC24, MdSC^{+}15, MPP21, PAI^{+}23, RV19, SBCP21, Sep23, SIA^{+}24, SHOL23, YP12, YA22, ZKP24]. threat-informed [AG23]. threats [ACT23, SAH22, VSN22]. three [LCL16, NSNK06, Nui12, AGZA22]. Three-factor [AGZA22]. three-party [LCL16]. three-round [NSNK06]. Threshold [Pen12, WMS^{+}19, BCD^{+}13, HKO22, JSMG18b, QDW09]. Thresholdizing [BST21]. ThunderSecure [GDA22]. Time [KME^{+}16, AGL24, BDG21, CK23b, CJMS19, CKPM24, DMDD16, GDA22, GGP21, KKK22b, KCB17, LLBL18, LKH09, LBW05, MUiH^{+}21, NAM06, PD21, RB23, SBS23, SKH^{+}23, Vaj16, CK23a]. time-aware [Vaj16]. time-based [LKH09, NAM06]. time-memory [CJMS19]. Time-specific [KME^{+}16]. Timed [BKBB20, BBB20, GLMS^{+}04]. timing [DHW11, KK17]. TL [NK24]. TL-BILSTM [NK24]. TLS [BJ16]. TLS- [BJ16]. TLS-/SSL-protected [BJ16]. Token [GTM11]. Token-based [GTM11]. tokenization [DSRHC16]. tolerance [CL09, SSD14]. tolerant [GKS19, RBD02]. toll [BSV22]. tool [KK22, SB09, SKB23, VSN22]. tools [AV17, HFA24]. top [FPP^{+}24, SCL^{+}18]. top-level [SCL^{+}18]. Topology [ZXZ22]. Topology-hiding [ZXZ22]. Tor [JKLJ24, ZD22]. toxicity [AK24b]. toys [Yan21]. trace [NP10, XT24]. traceable [ACHO13, Man21, SK14]. traceback [RS18]. traces [NMH^{+}24]. Tracing [TV24, DdP13, LCL14, Roe11a, Roe11b]. Track [Ano22, BJ15, DFF^{+}16]. Tracking [ZJS22]. trade [CJMS19]. trade-off [CJMS19]. tradeoff [SSN15]. traditional [GLP03]. traffic [ABK22, JJJ21, NLS24, PAI^{+}23, RSV23, SBS23, SPDR17, ZHZ22]. train [JAYZ21]. trained [KMW24]. training [CG23, FDS^{+}24, GSAMCA18, KMMG23, OCG24]. traits [BSBK24]. transactions [BLS^{+}23, CMS10, FYF22, PSDSNAHJ19, SK06]. transfer [Alh24, HBF^{+}24, NK24]. transform [KCM^{+}15, SJ10, ZMS22]. transformation [HHSS24]. transformation-based [HHSS24]. Transformational [KM07]. transformations [BZV05, BDF04]. Transforms [MTW^{+}14]. transition [BKMR08]. transparent [SVKV21]. trapdoor [ACHO13]. TRBAC [BKBB20, BBB20]. tree [ABR16, BDPV14, DRPW12, MFES04, PHS22]. trees [KB13, RBD02]. Trends [AAM23, GBC24, TDGL23]. triage [TBGB20]. triangle [LCL16]. Trojan [BSCZ11]. truly [BR18]. Trust [AKR24, BEd23, KN07, MAZS24, Abb13, Alg22, DK24, FFG20, FBFGEM21, GMMZ06, HBF^{+}24, KBH07, LD07, LLWY09, LV10, QLOW09, RSPMB16, SHA20, WW07, ZGC07]. trust-aware [Alg22, DK24]. TrUStAPIS [FFG20]. Trusted [LH15, CKW19, KBJ22]. Trustworthy [ULEY24, MBRPS18]. truth [AvO13]. TTP [ZBD06]. TTS [DSY06]. Tuning [GNS14]. Turkish [MOA^{+}24a]. turnkey [SKB23]. TV [Roe11b, DZW^{+}18, Roe11a]. tweets [EFB^{+}24]. twice [YOV09]. twins [HG24]. Two [AEHS24, LTC23, MR04, ML17, CG14, GNS14, MS09, MG23, SSFB15, SZ24, ST24]. two-channel [MS09]. Two-factor [ML17, CG14]. two-level [MG23].

Two-model [LTC23]. Two-party [MR04, SZ24]. two-round [GNS14]. Two-stage [AEHS24]. type [EZLC21, FN19]. typing [DM07, KM07]. typology [DWU^{+}23, RB23].

ubiquitous [DJS^{+}24, LI07]. UC [BP08]. UDP-based [FZ21]. ugly [HAS^{+}24]. UMTS [LSWW14]. UMTS/LTE [LSWW14]. unauthorized [YP12]. Uncertain [AJC^{+}09]. uncertainty [SdHZ16]. unclonable [AGZA22, BR17, KKY^{+}23, SM10, TK24]. uncovering [VSN22]. undergraduates [AHX^{+}23]. Understanding [LM06, MS15, RSPMB16, WGMB13, YL19]. undetectable [VHT09]. unforgeability [Hal20, Yon18]. unification [KM07]. unified [GA23]. uniform [KAC16]. unifying [BCD^{+}13]. unit [BHKM23]. unit-based [BHKM23]. universal [AKT23, HSMW08, Vaj16, ZWX20, ZXZ22, Dan07]. unknown [BEPL^{+}17]. unlinkable [BSV22]. Unlock [IS23]. unlocking [IS23]. unmanaged [KKKV07]. Unmasking [BSBK24]. Unpicking [DFF^{+}16]. unsupervised [KNSS24]. untrusted [GW09]. updatable [ABR16, AKMW20]. update [JSMG18a, YL20]. updates [AKMW20]. uploaders [WMS^{+}19]. URL [KCM^{+}15]. usability [MD23b]. Usage [LMMS17, RSMT24, SF17, YA22]. USB [CSM^{+}24]. use [LMMO04, Pla09, SK06, SS05b]. use-based [SS05b]. useful [DHS04]. User [AAZAA23, CFBvO09, CLPP11, Hal20, BFG^{+}13, CON09, IS23, KC24, MGV17, MTSH18, MD23b, PGMLK^{+}13, TK24, XSA13, YL19]. user-controlled [BFG^{+}13]. User-friendly [CLPP11]. User-mediated [Hal20]. users [AZ22, CF07, KOSU16, KB23, YM19, YA22]. uses [GAdFGM23]. Using [BBB20, GBG18, HTM11, RSMT24, RBD02, SJ10, Sen14, STD21, TZH04, AIJM24, AMZ22, AGZA22, ACB14, AAZAA23, AMLH18, AuHD^{+}23, AZM24, Alq24, AFA^{+}23, AG23, AK24a, AGIK07, AHC^{+}21, AVM23, AICC18, ACBC^{+}15, BEY23, BBG24, BMP^{+}14, BCL13, BDMM19, BS22, BCGL23, BDG23, CAS22, CLW^{+}11, CF07, DGF^{+}17, ELPB24, EWR^{+}09, EKY^{+}24, GKKT10, GGJ22, GSAMCA18, GBDJ14, GKBS12, HLKI15, HCN15, HL04, IDHRPCMP15, JCL^{+}18, JJQ24, KK22, KJS17, KA18, KIP22, KNSS24, KSM10, KCC^{+}23, KY24, LL22, LM06, LWL^{+}21, MR24, MB16, MARK20, MSKD16, MP15, MLYL20, MP16, MLCS16, MD23b, NCBG23, NAQ24, NLS24, OBH^{+}20, PMDS23, PKHS23, PM23, PP24a, PP24b, PV22, PBI^{+}23, PHE^{+}24, QLZH15, RHL17, RSV23, RE24, RT23, RRI^{+}19, SAM^{+}24, SBS23, SSFB15, SBD23, SSD14, SGSS23, SSL22b, SHD^{+}24, SMV24]. using [SIA^{+}24, SSV22, SSSJ24, SKH^{+}23, SEZ24, TBGB20, TKKO20, TSU^{+}24, TK24, VH19, WGH23, WKWS24, WLT^{+}24, Yon18, You06, ZP23, ZBC23, ZLJW20, ZMS22, vOLW05, BKBB20]. utilising [BAC24, LCPD14]. utilities [DA21]. Utility [CMH^{+}24, CR20, GGP21]. Utility-Preserving [CMH^{+}24]. utilized [IKS22]. utilizing [SC24, ZSN23].

V2X [CHC24]. validating [BS22]. Validation [ST24, AGIK07, ZBD06, dAKdG10]. valley [Auf20]. Value [IKS22]. Value-utilized [IKS22]. VANET [AaBZS23, YWW22]. variants [BCD^{+}13]. variation [LNX22]. various [RSK^{+}24, TS22]. Vaudenay [Tip22]. vector [MMS16, MLCS16]. vector-based [MMS16]. vehicle [ACT23]. vehicles [AKR24, MMF^{+}24]. vehicular [AHT24, LH23, MG19, MB16, SF17]. verifiability [ZWX20]. Verifiable [NSNK06, AFA^{+}23, KU16, LZQ^{+}18, PCK22a, TMvM24]. Verifiably [WPD18, KJ14, SEXY18]. verification [BMP05, CKW19, CL13, DS07, GMH14, PH24b, PH24a, Pen12, Pen13, Yon18, Zha23]. verifier [HSMW08, HYWS11, RSD19]. VERIFY [ACM05]. Verifying [MP15]. version [AKT23, SBD23]. versions [ABPR24]. versus [Bra22, LSR^{+}23, MTSH18, VLGL24]. very [Bae10]. via [AaBZS23, CSM^{+}24, IOU^{+}21, MTSH18, MS14, ZHZ22]. victims [CMN^{+}18]. video [LcSCL^{+}18, MBRPS18, RSPMB16, SB14]. view [Kak24, MD23b]. views [KMR09]. violation [GYL^{+}07]. virtual [PDM20, RM12, SHW23]. visible [LC04]. Visualization [XSA13, JJJ21, YL19]. Visualization-based [XSA13]. visualized [HLKI15]. visualizing [SBCP21]. visually [KB23]. VM [PDM20]. voice [WAB^{+}09]. VoIP [VL13]. voltage [MPG21]. volumes [SMMN12]. voting [CLO^{+}24, DJN10, GP23, MMS16, MS11, Pen11, SM24, ZWX20, ZLJW20]. Vulnerabilities [HFA24, AKA24, BHKM23, DTK^{+}18, SNX19, YNC22]. Vulnerability [ZR24, AV17, BAG^{+}24, CR24, ESSP24, GBC24, JT24, JM17, KEA24, KKK22a, MLM19, MP22, PP24b, SEZ24, ZRJ14, EZLC21]. vulture [ETRLG24].

wallet [FTS^{+}20]. warning

[Alm24, ABFL12]. Waste [SSM^{+}20].

Watermarking [Fra18, BMP^{+}14]. wavelet [KCM^{+}15]. weaknesses [DDZ22]. wearable [RSH^{+}24a, RSH^{+}24b, SAC22, SDR20]. Web [EZ22, AYHK18, ANN23, AV17, DTK^{+}18, GLMS^{+}04, JKLJ24, LL22, MSKD16, PM23, RSV23, SGLC19, SCO21, TSMH19, ZD22, BFP03, BJ16, EWR^{+}09, IOU^{+}21, KGG09, SNX19]. web-based [AYHK18]. webshell [HNHY23]. website [AFF24, NCBG23]. WebView [IOU^{+}21]. weighted [BMFS24, Sen14, WKWS24]. wildcard [CKS21, HH16]. windows [ASFG24]. WinRAR [YP06]. Wireless [BT07, AZM24, CMR06, CBC08, DK24, EHM15, EMRN17, KKK17, LBZ^{+}10, MARK20, RSH^{+}24a, RSH^{+}24b, RA24, RBEH15, SS17, SSN15, TK24, YRW14, ZXZ^{+}11]. Wirelessly [GdKGV14]. within [KCB20]. without [GW09, HSMW08, HYWS11, HYWS12, SXZC20, YSM10, ZXZ22]. wolf [SM24]. word [KEA24]. workflow [ARMLS06]. working [AHX^{+}23, AYM24]. workplace [Lev07]. world [TSZ22, ZD22]. worm [AGIK07, KJG^{+}11]. worms [vORM06]. wrinkle [BDG21]. Write [LD17, JTV19]. Write-only [LD17]. WSN [TSU^{+}24]. WSNs [TSU^{+}24].

X.509 [ZJS22]. XACML [CON09, RRI^{+}19]. XAIoT [QQK^{+}24]. Xen [PPL15]. Xen-based [PPL15]. XML [BFPP07, CF03, KMR09]. XOR [BP08, BGP07b]. XQuery [DTK^{+}18]. XTEA [Lu09].

Yao [BPW05b, BP08]. Yao-style [BPW05b, BP08]. yoking [HLS18].

Zephyrus [GAdFGM23]. zero [DGF^{+}17, HHSS24, PSDSNAHJ19, SLGP23, WCS20]. zero-confirmation [PSDSNAHJ19]. zeroday [DGF^{+}17, SLGP23]. zero-knowledge [HHSS24, WCS20]. zero-shot [SLGP23]. Zheng [BZ03]. zk [ABPR24]. zk-SNARK [ABPR24]. ZombieCoin [AMLH18].

References

Alzaidi:2020:DHP

[AAB20] Areej Alzaidi, Suhair Alshehri, and Seyed M. Buhari. DroidRista: a highly precise static data flow analysis framework for Android applications. International Journal of Information Security, 19(5): 523–536, October 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00471-w.

Al-ani:2023:PSI

[AaBZS23] Ruqayah Al-ani, Thar Baker, Bo Zhou, and Qi Shi. Privacy and safety improvement of VANET data via a safety-related privacy scheme. International Journal of Information Security, 22(4): 763–783, August 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00662-6.

Asad:2024:PRA

[AAG24] H. Asad, S. Adhikari, and Ilir Gashi. A perspectiveretrospective analysis of diversity in signature-based open-source network intrusion detection systems. International Journal of Information Security, 23 (2):1331–1346, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00794-9.

Ahmad:2023:TPB

[AAM23] Hira Ahmad, Muhammad Ahtazaz Ahsan, and Adnan Noor Mian. Trends in publishing blockchain surveys: a bibliometric perspective. International Journal of Information Security, 22 (2):511–523, April 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00653-z.

Al-Aqrabi:2024:DAI

[AAMH^{+}24] Hussain Al-Aqrabi, Ahmed M.

Manasrah, Richard Hill, Mohammed Q. Shatnawi, Mohammad Sh Daoud, and Hoda Alkhzaimi. Dynamic authentication for intelligent sensor clouds in the Internet of Things. International Journal of Information Security, 23 (3):2003–2021, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00829-9.

Abdussami:2022:LLA

[AAV22] Mohammad Abdussami, Ruhul Amin, and Satyanarayana Vollala. LASSI: a lightweight authenticated key agreement protocol for fog-enabled IoT deployment. International Journal of Information Security, 21(6):1373–1387, December 2022. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00619-1.

Alawneh:2023:UIU

[AAZAA23] Luay Alawneh, Mohammad Al-Zinati, and Mahmoud Al-Ayyoub. User identification using deep learning and human activity mobile sensor data. International Journal of Information Security, 22(1): 289–301, February 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00640-4.

Agostini:2022:BBM

[AB22] E. Agostini and M. Bernaschi. BitCracker: BitLocker meets GPUs. International Journal of Information Security, 21(5): 1005–1018, October 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00589-4.

Ayodele:2024:SDM

[AB24] Believe Ayodele and Victor Buttigieg. SDN as a defence mechanism: a comprehensive survey. International Journal of Information Security, 23(1): 141–185, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00764-1.

Abbadi:2013:FET

[Abb13]  Imad M. Abbadi.  A

framework for establishing trust in cloud provenance. International Journal of Information Security, 12(2):111–128, April 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0179-0.

Aliasgari:2017:SCH

[ABB17] Mehrdad Aliasgari, Marina Blanton, and Fattaneh Bayatbabolghani. Secure computation of hidden Markov models and secure floating-point arithmetic in the malicious model. International Journal of Information Security, 16(6): 577–601, November 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0350-0.

Alfaro:2008:CAC

[ABCC08] J. G. Alfaro, N. Boulahia-Cuppens, and F. Cuppens. Complete analysis of configuration rules to guarantee reliable network security policies. International Journal of Information Security, 7(2):103–122, April 2008. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        007-0045-7.

Ates:2012:WHI

[ABFL12] Mikaël Ates, Francesco Buccafurri, Jacques Fayolle, and Gianluca Lax. A warning on how to implement anonymous credential protocols into the information card framework. International Journal of Information Security, 11(1):33–40, February 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        011-0150-5.

Aimeur:2008:LPP

[ABFO08] Esma Aïmeur, Gilles Brassard, José M. Fernandez, and Flavien Serge Mani Onana. Alambic: a privacy-preserving recommender system for electronic commerce. International Journal of Information Security, 7(5): 307–334, October 2008. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0049-3.

Agrawal:2022:SAE

[ABK22] Ankit Agrawal, Ashutosh Bhatia, and Rekha Kaushik. A survey on analyzing encrypted network traffic of mobile devices. International Journal of Information Security, 21(4): 873–915, August 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00581-y.

Andreeva:2012:SAS

[ABM^{+}12] Elena Andreeva, Andrey Bogdanov, Bart Mennink, Bart Preneel, and Christian Rechberger. On security arguments of the second round SHA-3 candidates. International Journal of Information Security, 11(2):103–120, April 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0156-7.

Albeshri:2014:EGI

[ABN14] Aiiad Albeshri, Colin Boyd, and Juan González Nieto. Enhanced Geo-Proof: improved geographic assurance for data in the cloud. International Journal of Information Security, 13 (2):191–198, April 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0217-6.

Amine:2024:SEV

[ABPR24] Oussama Amine, Karim Baghery, Zaira Pindado, and Carla Ràfols. Simulation extractable versions of Groth’s zk-SNARK revisited. International Journal of Information Security, 23(1):431–445, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00750-7.

Abbes:2016:DFC

[ABR16] Tarek Abbes, Adel Bouhoula, and Michaël Rusinowitch. Detection of firewall configuration errors with updatable tree. International Journal of Information Security, 15(3):301–317, June 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

com/article/10.1007/s10207-
015-0290-0.

Armando:2008:SBM

[AC08] Alessandro Armando and Luca Compagna. SATbased model-checking for security protocols analysis. International Journal of Information Security, 7(1):3–32, January 2008. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0041-y.

Alabrah:2014:ESC

[ACB14] Amerah Alabrah, Jeffrey Cashion, and Mostafa Bassiouni. Enhancing security of cookie-based sessions in mobile networks using sparse caching. International Journal of Information Security, 13(4): 355–366, August 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0223-8.

Azkia:2015:DPA

[ACBC^{+}15] Hanieh Azkia, Nora Cuppens-Boulahia, Frédéric Cuppens, Gouenou Coatrieux, and Said Oulmakhzoune. Deployment of a posteriori access control using IHE ATNA. International Journal of Information Security, 14(5): 471–483, October 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        014-0265-6.

Ahmad:2024:IFC

[ACCS24] Zubair Ahmad, Stefano Calzavara, Samuele Casarin, and Ben Stock. Information flow control for comparative privacy analyses. International Journal of Information Security, 23(5):3199–3216, October 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        024-00886-0.

Albertini:2017:EAC

[ACF17] Davide Alberto Albertini, Barbara Carminati, and Elena Ferrari. An extended access control mechanism exploiting data dependencies. International Journal of Information Security, 16(1):75–89, February 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        016-0322-4.

Abe:2013:DTA

[ACHO13] Masayuki Abe, Sherman S. M. Chow, Kristiyan Haralambiev, and Miyako Ohkubo. Double-trapdoor anonymous tags for traceable signatures. International Journal of Information Security, 12 (1):19–31, February 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0184-3.

Autexier:2005:PSI

[ACM05] Serge Autexier, Iliano Cervesato, and Heiko Mantel. Preface to the special issue of selected papers from FCS/VERIFY 2002. International Journal of Information Security, 4(1–2):1, February 2005. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
004-0053-9;   http://
link.springer.com/content/
pdf/10.1007/s10207-004-
0053-9.pdf.

Armando:2015:FMA

[ACMV15] Alessandro Armando, Gabriele Costa, Alessio Merlo, and Luca Verderame. Formal modeling and automatic enforcement of Bring Your Own Device policies. International Journal of Information Security, 14(2):123–140, April 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        014-0252-y.

Alecakir:2021:ATI

[ACS21] Huseyin Alecakir, Burcu Can, and Sevil Sen. Attention: there is an inconsistency between Android permissions and application metadata! International Journal of Information Security, 20(6): 797–815, December 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        020-00536-1.

Alcaraz:2023:OST

[ACT23] Cristina Alcaraz, Jesus Cumplido, and Alicia Triviño. OCPP in the spotlight: threats and countermeasures for electric vehicle charging infrastructures 4.0. International Journal of Information Security, 22(5):1395–1421, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00698-8.

Aldini:2008:EMI

[AD08] Alessandro Aldini and Alessandra Di Pierro. Estimating the maximum information leakage. International Journal of Information Security, 7(3): 219–242, June 2008. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0050-x.

Ali:2024:FIS

[ADHN24] Rao Faizan Ali, P. D. D.

Dominic, Sadaf Hina, and Sheraz Naseer. Fostering information security policies compliance with ISA-95-based framework: an empirical study of oil and gas employees. International Journal of Information Security, 23(2): 1197–1213, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00786-9.

Akbarzadeh:2024:TSA

[AEHS24] Aida Akbarzadeh, Laszlo Erdodi, Siv Hilde Houmb, and Tore Geir Soltvedt. Two-stage advanced persistent threat (APT) attack on an IEC 61850 power grid substation. International Journal of Information Security, 23(4): 2739–2758, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        024-00856-6.

Alzahrani:2023:DAC

[AFA^{+}23] Bander Alzahrani, Nikos Fotiou, Aiiad Albeshri, Abdullah Almuhaimeed, and Khalid Alsubhi. Distributed access control for information-centric networking architectures using verifiable credentials. International Journal of Information Security, 22 (2):467–478, April 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00649-9.

Ariyadasa:2024:SRL

[AFF24] Subhash Ariyadasa, Shantha Fernando, and Subha Fernando. SmartiPhish: a reinforcement learningbased intelligent antiphishing solution to detect spoofed website attacks. International Journal of Information Security, 23 (2):1055–1076, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00778-9.

Amro:2023:CRM

[AG23] Ahmed Amro and Vasileios Gkioulos. Cyber risk management for autonomous passenger ships using threatinformed defense-in-depth. International Journal of Information Security, 22 (1):249–288, February 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00638-y.

Anagnostakis:2007:CDR

[AGIK07] Kostas G. Anagnostakis, Michael B. Greenwald, Sotiris Ioannidis, and Angelos D. Keromytis. COV-ERAGE: detecting and reacting to worm epidemics using cooperation and validation. International Journal of Information Security, 6(6):361–378, October 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0032-z.

Alkhateeb:2024:SRT

[AGL24] Ehab Alkhateeb, Ali Ghorbani, and Arash Habibi Lashkari. A survey on run-time packers and mitigation techniques. International Journal of Information Security, 23 (2):887–913, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.

                                        com/article/10.1007/s10207-
                                        023-00759-y.

Ahmim:2022:SAT

[AGZA22] Ilyes Ahmim, Nacira Ghoualmi-Zine, and Marwa Ahmim. Security analysis on “Three-factor authentication protocol using physical unclonable function for IoV”. International Journal of Information Security, 21(5): 1019–1026, October 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00595-6.

Aniello:2021:ABT

[AHC^{+}21] Leonardo Aniello, Basel Halak, Peter Chai, Riddhi Dhall, Mircea Mihalea, and Adrian Wilczynski. Anti-BlUFf: towards counterfeit mitigation in IC supply chains using blockchain and PUF. International Journal of Information Security, 20(3):445–460, June 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        020-00513-8;   http://
                                        link.springer.com/content/
                                        pdf/10.1007/s10207-020-
                                        00513-8.pdf.

Abdelkreem:2024:FEI

[AHT24] Eslam Abdelkreem, Sherif Hussein, and Ashraf Tammam. Feature engineering impact on position falsification attacks detection in vehicular ad-hoc network. International Journal of Information Security, 23(3):1939–1961, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00830-2.

An:2023:HEL

[AHX^{+}23] Qin An, Wilson Cheong Hin Hong, XiaoShu Xu, Yunfeng Zhang, and Kimberly Kolletar-Zhu. How education level influences Internet security knowledge, behaviour, and attitude: a comparison among undergraduates, postgraduates and working graduates. International Journal of Information Security, 22(2):305–317, April 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00637-z.

Ayed:2018:ADS

[AICC18] Samiha Ayed, Muhammad Sabir Idrees, Nora Cuppens, and Frederic Cuppens. Achieving dynamicity in security policies enforcement using aspects. International Journal of Information Security, 17(1): 83–103, February 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        016-0357-6.

Adnan:2024:ISE

[AIJM24] Muhammad Adnan, Muhammad Osama Imam, Muhammad Furqan Javed, and Iqbal Murtza. Improving spam email classification accuracy using ensemble techniques: a stacking approach. International Journal of Information Security, 23(1): 505–517, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00756-1.

An:2009:UIC

[AJC^{+}09] Xiangdong An, Dawn Jutla, Nick Cercone, Charnyote Pluempitiwiriyawej, and Hai Wang. Uncertain inference control in privacy protection. International Journal of Information Security, 8(6): 423–431, December 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.

com/article/10.1007/s10207-
009-0088-z.

Akbarzadeh:2023:DBS

[AK23] Aida Akbarzadeh and Sokratis K. Katsikas. Dependencybased security risk assessment for cyber-physical systems. International Journal of Information Security, 22(3):563–578, June 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00608-4.

Amutha:2024:ESQ

[AK24a] M. Amutha and K. R.

Kavitha. Enhancing security in QCA-based circuits using optimal key gate placement. International Journal of Information Security, 23(3): 2395–2405, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00842-y.

Anjum:2024:HST

[AK24b] Anjum and Rahul Katarya.

Hate speech, toxicity detection in online social media: a recent survey of state of the art and opportunities. International Journal of Information Security, 23(1):577–608, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00755-2.

Arikan:2024:ASC

[AKA24] Süleyman Muhammed Arikan, Aynur Koçak, and Mustafa Alkan. Automating shareable cyber threat intelligence production for closed source software vulnerabilities: a deep learning based detection system. International Journal of Information Security, 23(5): 3135–3151, October 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        024-00882-4.

Anagnostopoulos:2016:NFM

[AKG16] Marios Anagnostopoulos, Georgios Kambourakis, and Stefanos Gritzalis. New facets of mobile botnet: architecture and evaluation. International Journal of Information Security, 15(5): 455–473, October 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        015-0310-0.

Anada:2020:KUP

[AKMW20] Hiroaki Anada, Akira Kanaoka, Natsume Matsuzaki, and Yohei Watanabe. Key-updatable publickey encryption with keyword search (or: How to realize PEKS with efficient key updates for IoT environments). International Journal of Information Security, 19 (1):15–38, February 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00441-2.

Abbasi:2024:TMI

[AKR24] Shirin Abbasi, Navid Khaledian, and Amir Masoud Rahmani. Trust management in the internet of vehicles: a systematic literature review of blockchain integration. International Journal of Information Security, 23(4): 3065–3088, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00878-0.

Andriotis:2023:BDU

[AKT23] Panagiotis Andriotis, Myles Kirby, and Atsuhiro Takasu. Bu-Dash: a universal and dynamic graphical password scheme (extended version). International Journal of Information Security, 22 (2):381–401, April 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00642-2.

Adepu:2020:ASG

[AKZM20] Sridhar Adepu, Nandha Kumar Kandasamy, Jianying Zhou, and Aditya Mathur. Attacks on smart grid: power supply interruption and malicious power generation. International Journal of Information Security, 19(2):189–211, April 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        019-00452-z.

Atallah:2005:SOS

[AL05] Mikhail J. Atallah and Jiangtao Li. Secure outsourcing of sequence comparisons. International Journal of Information Security, 4(4): 277–287, October 2005. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        005-0070-3.

Alawatugoda:2017:GCM

[Ala17] Janaka Alawatugoda. Generic construction of an eCKsecure key exchange protocol in the standard model. International Journal of Information Security, 16 (5):541–557, October 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0346-9.

Alghamdi:2022:NTA

[Alg22] Saleh A. Alghamdi. Novel trust-aware intrusion detection and prevention system for 5G MANET-Cloud. International Journal of Information Security, 21(3):469–488, June 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
020-00531-6.

Alharbi:2024:FTL

[Alh24] Afnan A. Alharbi. Federated transfer learning for attack detection for Internet of Medical Things. International Journal of Information Security, 23 (1):81–100, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00805-9.

Almoqbil:2024:ADE

[Alm24] Abdullah Hamad N. Almoqbil. Anomaly detection for early ransomware and spyware warning in nuclear power plant systems based on FusionGuard. International Journal of Information Security, 23 (3):2377–2394, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00841-z.

Arbit:2015:IPK

[ALOW15] Alex Arbit, Yoel Livne, Yossef Oren, and Avishai Wool. Implementing public-key cryptography on passive RFID tags is practical. International Journal of Information Security, 14(1):85–99, February 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0236-y.

Asghar:2013:CCH

[ALPW13] Hassan Jameel Asghar, Shujun Li, Josef Pieprzyk, and Huaxiong Wang. Cryptanalysis of the convex hull click human identification protocol. International Journal of Information Security, 12(2):83–96, April 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

com/article/10.1007/s10207-
012-0161-x.

Alqahtani:2024:SBR

[Alq24] Sultan S. Alqahtani. Security bug reports classification using fasttext. International Journal of Information Security, 23 (2):1347–1358, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00793-w.

Ahmad:2023:PCS

[AM23] Suhail Ahmad and Ajaz Hussain Mir. Protection of centralized SDN control plane from high-rate Packet–In messages. International Journal of Information Security, 22(5): 1197–1206, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00685-z.

Ali:2018:ZMN

[AMLH18] Syed Taha Ali, Patrick McCorry, Peter Hyun-Jeen Lee, and Feng Hao. ZombieCoin 2.0: managing next-generation botnets using Bitcoin. International Journal of Information Security, 17(4): 411–422, August 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        017-0379-8.

Andreeva:2012:PFG

[AMP12] Elena Andreeva, Bart Mennink, and Bart Preneel. The parazoa family: generalizing the sponge hash functions. International Journal of Information Security, 11(3):149–165, June 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        012-0157-6.

Arapinis:2017:APM

[AMRR17] Myrto Arapinis, Loretta Ilaria Mancini, Eike Ritter, and Mark Dermot Ryan. Analysis of privacy in mobile telephony systems. International Journal of Information Security, 16(5): 491–523, October 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        016-0338-9;   http://
                                        link.springer.com/content/
                                        pdf/10.1007/s10207-016-
                                        0338-9.pdf.

Ahmad:2022:EAC

[AMZ22] Tahir Ahmad, Umberto Morelli, and Nicola Zannone. Extending access control in AWS IoT through event-driven functions: an experimental evaluation using a smart lock system. International Journal of Information Security, 21(2):379– 408, April 2022. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00558-3.

Alidoosti:2023:BLC

[ANN23] Mitra Alidoosti, Alireza Nowroozi, and Ahmad Nickabadi. Business-layer client-side racer: dynamic security testing of the web application against clientside race condition in the business layer. International Journal of Information Security, 22(4): 1029–1054, August 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00671-5.

Anonymous:2011:P

[Ano11a] Anonymous. Preface. International Journal of Information Security, 10(2): 61, June 2011. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-

                                        011-0134-5;   http://
                                        link.springer.com/content/
                                        pdf/10.1007/s10207-011-
                                        0134-5.pdf.

Anonymous:2011:SIS

[Ano11b] Anonymous. Special issue on “SCADA and control system security”. International Journal of Information Security, 10 (2):135–136, June 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        011-0128-3.

Anonymous:2014:SCC

[Ano14] Anonymous. Security in cloud computing. International Journal of Information Security, 13(2): 95–96, April 2014. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        014-0232-2.

Anonymous:2022:TS

[Ano22] Anonymous. Track for surveys. International Journal of Information Security, 21(1):159, February 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        021-00546-7.

Ahmed:2012:MGE

[ANS^{+}12] Irfan Ahmed, Martin Naedele, Bradley Schatz, Ryoichi Sasaki, and Andrew West. Message from the guest editors. International Journal of Information Security, 11(4): 213, August 2012. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0168-3;   http://
link.springer.com/content/
pdf/10.1007/s10207-012-
0168-3.pdf.

Armour:2022:ASA

[AP22] Marcel Armour and Bertram Poettering. Algorithm substitution attacks against receivers. International Journal of Information Security, 21(5):1027–1050, October 2022. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00596-5.

Al-Riyami:2006:EFE

[ARMLS06] S. S. Al-Riyami, J. Malone-Lee, and N. P. Smart. Escrow-free encryption supporting cryptographic workflow. International Journal of Information Security, 5(4):217–229, October 2006. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        006-0002-x.

Al-Shehari:2023:RRA

[ASA23] Taher Al-Shehari and Rakan A. Alsowail. Random resampling algorithms for addressing the imbalanced dataset classes in insider threat detection. International Journal of Information Security, 22 (3):611–629, June 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00651-1.

Al-Saleh:2015:IDC

[ASAAS15] Mohammed I. Al-Saleh, Fatima M. AbuHjeela, and Ziad A. Al-Sharif. Investigating the detection capabilities of antiviruses under concurrent attacks. International Journal of Information Security, 14 (4):387–396, August 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        014-0261-x.

Andelic:2024:RPS

[AŠC24] Nikola Andelić, Sandi Baressi Šegota, and Zlatan Car. Robust password security: a genetic programming approach with imbalanced dataset handling. International Journal of Information Security, 23(3):1761–1786, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00814-2.

Abie:2004:DDR

[ASF04] Habtamu Abie, Pål Spilling, and Bent Foyn. A distributed digital rights management model for secure information-distribution systems. International Journal of Information Security, 3(2):113–128, November 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL http://

link.springer.com/article/
10.1007/s10207-004-0058-
4.

Asfahani:2024:POR

[Asf24] Ahmed M. Asfahani. Perceptions of organizational responsibility for cybersecurity in Saudi Arabia: a moderated mediation analysis. International Journal of Information Security, 23(4):2515–2530, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.

                                        com/article/10.1007/s10207-
                                        024-00859-3.

Ayub:2024:RSI

[ASFG24] Md. Ahsan Ayub, Ambareen Siraj, Bobby Filar, and Maanak Gupta. RWArmor: a staticinformed dynamic analysis approach for early detection of cryptographic windows ransomware. International Journal of Information Security, 23(1): 533–556, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00758-z.

Adamos:2024:EAR

[ASKG24] Konstantinos Adamos, George Stergiopoulos, Michalis Karamousadakis, and Dimitris Gritzalis. Enhancing attack resilience of cyber-physical systems through state dependency graph models. International Journal of Information Security, 23(1): 187–198, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00731-w.

Almaatouq:2016:IIL

[ASN^{+}16] Abdullah Almaatouq, Erez Shmueli, Mariam Nouh, Ahmad Alabdulkareem, Vivek K. Singh, Mansour Alsaleh, Abdulrahman Alarifi, Anas Alfaris, and Alex ‘Sandy’ Pentland. If it looks like a spammer and behaves like a spammer, it must be a spammer: analysis and detection of microblogging spam accounts. International Journal of Information Security, 15(5): 475–491, October 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0321-5.

Alhaj:2023:EAS

[ASZ^{+}23] Taqwa Ahmed Alhaj, Maheyzah Md Siraj, Anazida Zainal, Inshirah Idris, Anjum Nazir, Fatin Elhaj, and Tasneem Darwish. An effective attack scenario construction model based on identification of attack steps and stages. International Journal of Information Security, 22(5): 1481–1496, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00701-2.

Adewole:2022:DDL

[AT22] Kayode S. Adewole and Vicenç Torra. DFTMicroagg: a dual-level anonymization algorithm for smart grid data. International Journal of Information Security, 21 (6):1299–1321, December 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00612-8.

Abbas:2024:RFF

[ATN^{+}24] Sohail Abbas, Manar Abu Talib, Qassim Nasir, Sally Idhis, Mariam Alaboudi, and Ali Mohamed. Radio frequency fingerprinting techniques for device identification: a survey. International Journal of Information Security, 23 (2):1389–1427, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00801-z.

Abdulganiyu:2023:SLR

[ATS23] Oluwadamilare Harazeem Abdulganiyu, Taha Ait Tchakoucht, and Yakub Kayode Saheed. A systematic literature review for network intrusion detection system (IDS). International Journal of Information Security, 22(5): 1125–1162, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00682-2.

Aufner:2020:ISG

[Auf20] Peter Aufner. The IoT security gap: a look down into the valley between threat models and their implementation. International Journal of Information Security, 19 (1):3–14, February 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00445-y;   http://
link.springer.com/content/
pdf/10.1007/s10207-019-
00445-y.pdf.

Ali:2023:ENI

[AuHD^{+}23] Muhammad Ali, Mansoor ul Haque, Muhammad Hanif Durad, Anila Usman, Syed Muhammad Mohsin, Hana Mujlid, and Carsten Maple. Effective network intrusion detection using stacking-based ensemble approach. International Journal of Information Security, 22 (6):1781–1798, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00718-7.

Antunes:2017:DVT

[AV17] Nuno Antunes and Marco Vieira. Designing vulnerability testing tools for web services: approach, components, and tools. International Journal of Information Security, 16 (4):435–457, August 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        016-0334-0.

Asha:2023:DFD

[AVM23] S. Asha, P. Vinod, and Varun G. Menon. A defensive framework for deepfake detection under adversarial settings using temporal and spatial features. International Journal of Information Security, 22(5): 1371–1382, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00695-x.

Alsaleh:2013:EAA

[AvO13] Mansour Alsaleh and P. C. van Oorschot. Evaluation in the absence of absolute ground truth: toward reliable evaluation methodology for scan detectors. International Journal of Information Security, 12 (2):97–110, April 2013.

CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0178-1.

Akiyama:2018:HDC

[AYHK18] Mitsuaki Akiyama, Takeshi Yagi, Takeo Hariu, and Youki Kadobayashi. HoneyCirculator: distributing credential honeytoken for introspection of web-based attack cycle. International Journal of Information Security, 17 (2):135–151, April 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0361-5;   http://
link.springer.com/content/
pdf/10.1007/s10207-017-
0361-5.pdf.

Angafor:2024:SRO

[AYM24] Giddeon Njamngang Angafor, Iryna Yevseyeva, and Leandros Maglaras. Securing the remote office: reducing cyber risks to remote working through regular security awareness education campaigns. International Journal of Information Security, 23(3): 1679–1693, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.

                                        com/article/10.1007/s10207-
                                        023-00809-5.

Aamir:2019:DAD

[AZ19] Muhammad Aamir and Syed Mustafa Ali Zaidi. DDoS attack detection with feature engineering and machine learning: the framework and performance evaluation. International Journal of Information Security, 18(6): 761–785, December 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        019-00434-1.

Amraoui:2022:ABD

[AZ22] Noureddine Amraoui and Belhassen Zouari. Anomalous behavior detectionbased approach for authenticating smart home system users. International Journal of Information Security, 21(3):611–636, June 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        021-00571-6.

Aliabadian:2024:SCG

[AZM24] Amir Aliabadian, Mohammadreza Zahabi, and Majid Mobini. Spatial de-correlation of generated keys from wireless channels using adversarial deep learning. International Journal of Information Security, 23(3): 2063–2073, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00831-1.

Ali:2024:RSP

[AZS24] Auwal Shehu Ali, Zarul Fitri Zaaba, and Manmeet Mahinderjit Singh. The rise of “security and privacy”: bibliometric analysis of computer privacy research. International Journal of Information Security, 23 (2):863–885, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00761-4.

Blanton:2016:POS

[BA16] Marina Blanton and Everaldo Aguiar. Private and oblivious set and multiset operations. International Journal of Information Security, 15(5): 493–518, October 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
015-0301-1.

Bilgen:2023:AEA

[BAB23] Melike Burakgazi Bilgen, Osman Abul, and Kemal Bicakci. Authenticationenabled attribute-based access control for smart homes. International Journal of Information Security, 22(2):479–495, April 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00639-x.

Chikha:2016:BBA

[BABB16] Randa Jabeur Ben Chikha, Tarek Abbes, Wassim Ben Chikha, and Adel Bouhoula. Behavior-based approach to detect spam over IP telephony attacks. International Journal of Information Security, 15 (2):131–143, April 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        015-0281-1.

Browne:2024:SRR

[BAC24] Thomas Oakley Browne, Mohammad Abedin, and Mohammad Jabed Morshed Chowdhury. A systematic review on research utilising artificial intelligence for open source intelligence (OS-INT) applications. International Journal of Information Security, 23(4): 2911–2938, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00868-2.

Baek:2010:RAR

[Bae10] Yoo-Jin Baek. Regular 2^{w}-ary right-to-left exponentiation algorithm with very efficient DPA and FA countermeasures. International Journal of Information Security, 9(5): 363–370, October 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
010-0118-x.

Bhandari:2024:IAE

[BAG^{+}24] Guru Prasad Bhandari, Gebremariam Assres, Nikola Gavric, Andrii Shalaginov, and Tor-Morten Grønli. IoTvulCode: AI-enabled vulnerability detection in software products designed for IoT applications. International Journal of Information Security, 23(4): 2677–2690, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00848-6.

Biskup:2004:CQE

[BB04a] Joachim Biskup and Piero Bonatti. Controlled query evaluation for enforcing confidentiality in complete information systems. International Journal of Information Security, 3 (1):14–27, October 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        004-0032-1.

Blakley:2004:ASN

[BB04b] B. Blakley and G. R. Blakley. All sail, no anchor II: Acceptable high-end PKI. International Journal of Information Security, 2(2):66–77, January 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        003-0029-1.

Bajic:2022:ABN

[BB22] Alexander Bajic and Georg T. Becker. Automated benchmark network diversification for realistic attack simulation with application to moving target defense. International Journal of Information Security, 21(2):253–278, April 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00552-9.

BenAttia:2020:UHT

[BBB20] Hasiba Ben Attia, Laid Kahloul Saber Benharzallah, and Samir Bourekkache. Using hierarchical timed coloured Petri nets in the formal study of TRBAC security policies. International Journal of Information Security, 19 (2):163–187, April 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00448-9. See author

name correction [BKBB20].

Bansal:2024:MPP

[BBG24] Vidhi Bansal, Niyati Baliyan, and Mohona Ghosh. MLChain: a privacy-preserving model learning framework using blockchain. International Journal of Information Security, 23(1): 649–677, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00754-3.

Bhale:2024:HID

[BBN24] Pradeepkumar Bhale, Santosh Biswas, and Sukumar Nandi. A hybrid IDS for detection and mitigation of sinkhole attack in 6LoWPAN networks. International Journal of Information Security, 23 (2):915–934, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00763-2.

Bindel:2018:CAA

[BBR18] Nina Bindel, Johannes Buchmann, and Susanne Rieß. Comparing apples with apples: performance analysis of latticebased authenticated key exchange protocols. International Journal of Information Security, 17(6): 701–718, November 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0397-6.

Bayly:2010:FBS

[BCA^{+}10] Duncan Bayly, Maurice Castro, Arathi Arakala, Jason Jeffers, and Kathy Horadam. Fractional biometrics: safeguarding privacy in biometric applications. International Journal of Information Security, 9(1):69–82, February 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
009-0096-z.

Bresson:2013:LLS

[BCD^{+}13] Emmanuel Bresson, Dario Catalano, Mario Di Raimondo, Dario Fiore, and Rosario Gennaro. Offline/on-line signatures revisited: a general unifying paradigm, efficient threshold variants and experimental results. International Journal of Information Security, 12(6): 439–465, November 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0200-2.

Burgess:2004:GTM

[BCEM04] Mark Burgess, Geoffrey Canright, and Kenth Engø-Monsen. A graph-theoretical model of computer security. International Journal of Information Security, 3(2):70–85, November 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
004-0044-x.

Boyd:2017:AAK

[BCF^{+}17] Colin Boyd, Cas Cremers, Michèle Feltz, Kenneth G. Paterson, Bertram Poettering, and Douglas Stebila. ASICS: authenticated key exchange security incorporating certification systems. International Journal of Information Security, 16 (2):151–171, April 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
015-0312-y.

Bowen:2023:BHD

[BCGL23] Brandon Bowen, Anitha Chennamaneni, Ana Goulart, and Daisy Lin. BLoC-Net: a hybrid, datasetindependent intrusion detection system using deep learning. International Journal of Information Security, 22(4):893–917, August 2023. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00663-5.

Backes:2011:CSS

[BCJ^{+}11] Michael Backes, Iliano Cervesato, Aaron D. Jaggard, Andre Scedrov, and Joe-Kai Tsay. Cryptographically sound security proofs for basic and public-key Kerberos. International Journal of Information Security, 10 (2):107–134, June 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
011-0125-6.

Brickell:2009:SSN

[BCL09] Ernie Brickell, Liqun Chen, and Jiangtao Li. Simplified security notions of direct anonymous attestation and a concrete scheme from pairings. International Journal of Information Security, 8(5): 315–330, October 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
009-0076-3.

Beauquier:2013:SPE

[BCL13] Danièle Beauquier, Joëlle Cohen, and Ruggero Lanotte. Security policies enforcement using finite and pushdown edit automata. International Journal of Information Security, 12 (4):319–336, August 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0195-8.

Blundo:2001:PDI

[BDD01] Carlo Blundo, Paolo D’Arco, and Alfredo De Santis. A t-private k-database information retrieval scheme.

International Journal of Information Security, 1 (1):64–68, August 2001. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s102070100005.

Bartoletti:2004:SIS

[BDF04] Massimo Bartoletti, Pierpaolo Degano, and Gian Luigi Ferrari. Stack inspection and secure program transformations. International Journal of Information Security, 2(3– 4):187–217, August 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        004-0038-8.

Berger:2021:WTC

[BDG21] Harel Berger, Amit Z. Dvir, and Moti Geva. A wrinkle in time: a case study in DNS poisoning. International Journal of Information Security, 20 (3):313–329, June 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        020-00502-x.

Brama:2023:ENN

[BDG23] Haya Brama, Lihi Dery, and Tal Grinshpoun. Evaluation of neural networks defenses and attacks using NDCG and reciprocal rank metrics. International Journal of Information Security, 22 (2):525–540, April 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00652-0.

Buhan:2010:ERC

[BDH^{+}10] Ileana Buhan, Jeroen Doumen, Pieter Hartel, Qian Tang, and Raymond Veldhuis. Embedding renewable cryptographic keys into noisy data. International Journal of Information Security, 9(3):193–208, June 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
010-0103-4.

Backes:2008:CRS

[BDHK08] Michael Backes, Markus Dürmuth, Dennis Hofheinz, and Ralf Küsters. Conditional reactive simulatability. International Journal of Information Security, 7(2):155–169, April 2008. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0046-6.

Bertino:2015:SPE

[BDHZ15] Elisa Bertino, Robert H. Deng, Xinyi Huang, and Jianying Zhou. Security and privacy of electronic health information systems. International Journal of Information Security, 14(6): 485–486, November 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        015-0303-z;   http://
                                        link.springer.com/content/
                                        pdf/10.1007/s10207-015-
                                        0303-z.pdf.

Bernardi:2019:DMD

[BDMM19] Mario Luca Bernardi, Marta Cimitile Damiano Distante, Fabio Martinelli, and Francesco Mercaldo. Dynamic malware detection and phylogeny analysis using process mining. International Journal of Information Security, 18(3):257–284, June 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        018-0415-3.

Bertoni:2014:SCS

[BDPV14] Guido Bertoni, Joan Daemen, Michaël Peeters, and Gilles Van Assche. Sufficient conditions for sound tree and sequential hashing modes. International Journal of Information Security, 13(4):335–353, August 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0220-y.

Barhoun:2023:TMD

[BEd23] Rabie Barhoun and Maryam Ed-daibouni. Trust modeling in a distributed collaborative environment: application to a collaborative healthcare system. International Journal of Information Security, 22 (6):1585–1604, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00707-w.

Bella:2010:PGA

[Bel10] Giampaolo Bella. The principle of guarantee availability for security protocol analysis. International Journal of Information Security, 9(2):83–97, April 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
009-0097-y.

Bat-Erdene:2017:EAC

[BEPL^{+}17] Munkhbayar Bat-Erdene, Hyundo Park, Hongzhe Li, Heejo Lee, and Mahn-Soo Choi. Entropy analysis to classify unknown packing algorithms for malware detection. International Journal of Information Security, 16(3):227–248, June 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0330-4.

Baig:2023:PPC

[BEY23] Ahmed Fraz Baig, Sigurd Eskeland, and Bian Yang. Privacy-preserving continuous authentication using behavioral biometrics. International Journal of Information Security, 22 (6):1833–1847, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00721-y.

Boteanu:2013:CSQ

[BF13] Daniel Boteanu and José M. Fernandez. A comprehensive study of queue management as a DoS counter-measure. International Journal of Information Security, 12(5): 347–382, October 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

com/article/10.1007/s10207-
013-0197-6.

Bernhard:2013:AAU

[BFG^{+}13] D. Bernhard, G. Fuchsbauer, E. Ghadafi, N. P. Smart, and B. Warinschi. Anonymous attestation with user-controlled linkability. International Journal of Information Security, 12(3):219–249, June 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0191-z.

Bertino:2003:CBF

[BFP03] Elisa Bertino, Elena Ferrari, and Andrea Perego. Content-based filtering of Web documents: the MaX system and the EUFOR-BIA project. International Journal of Information Security, 2(1): 45–58, November 2003. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
003-0024-6.

Bertino:2007:SSP

[BFPP07] Elisa Bertino, Elena Ferrari, Federica Paci, and Loredana Parasiliti Provenza. A system for securing pushbased distribution of XML documents. International Journal of Information Security, 6(4):255–284, July 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        007-0020-3.

Brzuska:2013:LMR

[BFS^{+}13] C. Brzuska, M. Fischlin, N. P. Smart, B. Warinschi, and S. C. Williams. Less is more: relaxed yet composable security notions for key exchange. International Journal of Information Security, 12 (4):267–297, August 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        013-0192-y.

Bracciali:2008:SFM

[BFT08] Andrea Bracciali, Gianluigi Ferrari, and Emilio Tuosto. A symbolic framework for multi-faceted security protocol analysis. International Journal of Information Security, 7 (1):55–84, January 2008. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        007-0043-9.

Balopoulos:2008:SIP

[BGK08]  Theodoros  Balopoulos,

Stefanos Gritzalis, and Sokratis K. Katsikas. Specifying and implementing privacy-preserving cryptographic protocols. International Journal of Information Security, 7(6): 395–420, November 2008. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
008-0057-y.

Bagheri:2012:SFP

[BGKZ12] Nasour Bagheri, Praveen Gauravaram, Lars R. Knudsen, and Erik Zenner. The suffix-free-prefix-free hash function construction and its indifferentiability security analysis. International Journal of Information Security, 11(6):419–434, November 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0175-4.

Backes:2007:P

[BGP07a] Michael Backes, Stefanos Gritzalis, and Bart Preneel. Preface. International Journal of Information Security, 6(6): 359–360, October 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

                                        com/article/10.1007/s10207-
                                        007-0034-x.

Blundo:2007:LRC

[BGP07b] Carlo Blundo, Clemente Galdi, and Giuseppe Persiano. Low-randomness constant-round private XOR computations. International Journal of Information Security, 6(1):15– 26, January 2007. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        006-0007-5.

Ben-Ghorbel-Talbi:2010:DME

[BGTCCBB10] Meriam Ben-Ghorbel-Talbi, Frédéric Cuppens, Nora Cuppens-Boulahia, and Adel Bouhoula. A delegation model for extended RBAC. International Journal of Information Security, 9(3):209–236, June 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        010-0104-3.

Baradaran:2023:UBS

[BHKM23] Sara Baradaran, Mahdi Heidari, Ali Kamali, and Maryam Mouzarani. A unit-based symbolic execution method for detecting memory corruption vulnerabilities in executable codes. International Journal of Information Security, 22(5): 1277–1290, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00691-1.

Benarfa:2021:CED

[BHL^{+}21] Abdelmadjid Benarfa, Muhammad Hassan, Eleonora Losiouk, Alberto Compagno, Mohamed Bachir Yagoubi, and Mauro Conti. ChoK-IFA+: an early detection and mitigation approach against interest flooding attacks in NDN. International Journal of Information Security, 20 (3):269–285, June 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00500-z.

Bi:2021:MPA

[BHZ^{+}21] Kun Bi, Dezhi Han, Guichen Zhang, Kuan-Ching Li, and Aniello Castiglione. K maximum probability attack paths generation algorithm for target nodes in networked systems. International Journal of Information Security, 20(4):535–551, August 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        020-00517-4.

Bella:2015:SIS

[BJ15] Giampaolo Bella and Helge Janicke. Special issue on the Security Track at the ACM Symposium on Applied Computing 2013. International Journal of Information Security, 14 (2):101–102, April 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        015-0280-2;   http://
                                        link.springer.com/content/
                                        pdf/10.1007/s10207-015-
                                        0280-2.pdf.

Brown:2016:API

[BJ16] Christopher W. Brown and Michael Jenkins. Analyzing proposals for improving authentication on the TLS- /SSL-protected Web. International Journal of Information Security, 15(6): 621–635, November 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        016-0316-2.

BenAttia:2020:CUH

[BKBB20] Hasiba Ben Attia, Laid Kahloul, Saber Benharzallah, and Samir Bourekkache.

Correction to: Using Hierarchical Timed Coloured Petri Nets in the formal study of TRBAC security policies. International Journal of Information Security, 19(2):241, April 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00454-x;   http://
link.springer.com/content/
pdf/10.1007/s10207-019-

00454-x.pdf. See [BBB20].

Bryans:2008:OGT

[BKMR08] Jeremy W. Bryans, Maciej Koutny, Laurent Mazaré, and Peter Y. A. Ryan. Opacity generalised to transition systems. International Journal of Information Security, 7(6): 421–435, November 2008. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
008-0058-x.

Bouzida:2011:CAB

[BLM11] Yacine Bouzida, Luigi Logrippo, and Serge Mankovski. Concrete- and abstractbased access control. International Journal of Information Security, 10(4): 223–238, August 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        011-0138-1.

Bandara:2023:OPP

[BLS^{+}23] Eranga Bandara, Xueping Liang, Sachin Shetty, Ravi Mukkamala, Peter Foytik, Nalin Ranasinghe, and Kasun De Zoysa. Octopus: privacy preserving peer-to-peer transactions system with InterPlanetary file system (IPFS). International Journal of Information Security, 22 (3):591–609, June 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00650-2.

Boyd:2005:PSI

[BM05] Colin Boyd and Wenbo Mao. Preface to the special issue on ISC 2003. International Journal of Information Security, 4 (4):227, October 2005. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        004-0066-4.

Bielova:2011:DYR

[BM11] Nataliia Bielova and Fabio Massacci. Do you really mean what you actually enforced? International Journal of Information Security, 10(4): 239–254, August 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
011-0137-2.

Barik:2024:AAD

[BMFS24] Kousik Barik, Sanjay Misra, and Luis Fernandez-Sanz. Adversarial attack detection framework based on optimized weighted conditional stepwise adversarial network. International Journal of Information Security, 23(3): 2353–2376, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00844-w.

Bella:2005:OVS

[BMP05] Giampaolo Bella, Fabio Massacci, and Lawrence C. Paulson. An overview of the verification of SET. International Journal of Information Security, 4(1– 2):17–28, February 2005. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
004-0047-7.

Bates:2014:DCR

[BMP^{+}14] Adam Bates, Benjamin Mood, Joe Pletcher, Hannah Pruse, Masoud Valafar, and Kevin Butler. On detecting co-resident cloud instances using network flow watermarking techniques. International Journal of Information Security, 13(2):171–189, April 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0210-0.

Basin:2005:OSM

[BMV05] David Basin, Sebastian Mödersheim, and Luca Viganò. OFMC: a symbolic model checker for security protocols. International Journal of Information Security, 4(3):181–208, June 2005. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
004-0055-7.

Buchholtz:2004:CCF

[BNN04] Mikael Buchholtz, Hanne Riis Nielson, and Flemming Nielson. A calculus for control flow analysis of security protocols. International Journal of Information Security, 2(3– 4):145–167, August 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

com/article/10.1007/s10207-
004-0036-x.

Bogdanov:2012:HPS

[BNTW12] Dan Bogdanov, Margus Niitsoo, Tomas Toft, and Jan Willemson. Highperformance secure multiparty computation for data mining applications. International Journal of Information Security, 11(6): 403–418, November 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0177-2.

Backes:2004:CPN

[BP04] Michael Backes and Birgit Pfitzmann. Computational probabilistic noninterference. International Journal of Information Security, 3(1):42–60, October 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
004-0039-7.

Backes:2008:LBU

[BP08] Michael Backes and Birgit Pfitzmann. Limits of the BRSIM/UC soundness of Dolev–Yao-style XOR. International Journal of Information Security, 7(1):33–54, January 2008. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        007-0040-z.

Backes:2005:RSS

[BPW05a] Michael Backes, Birgit Pfitzmann, and Michael Waidner. Reactively secure signature schemes. International Journal of Information Security, 4(4): 242–252, October 2005. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        004-0062-8.

Backes:2005:SAS

[BPW05b] Michael Backes, Birgit Pfitzmann, and Michael Waidner. Symmetric authentication in a simulatable Dolev–Yao-style cryptographic library. International Journal of Information Security, 4(3): 135–154, June 2005. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        004-0056-6.

Bernardini:2017:MRP

[BR17] Riccardo Bernardini and Roberto Rinaldo. Making random permutations from physically unclonable constants. International Journal of Information Security, 16(3):249–261, June 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0324-2.

Bernardini:2018:GES

[BR18] Riccardo Bernardini and Roberto Rinaldo. Generalized Elias schemes for efficient harvesting of truly random bits. International Journal of Information Security, 17(1): 67–81, February 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0358-5.

Bernardini:2020:ASS

[BR20] Riccardo Bernardini and Roberto Rinaldo. Analysis of some simple stabilizers for physically obfuscated keys. International Journal of Information Security, 19 (5):547–565, October 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00473-8.

Bhaskara:2023:CEA

[BR23] Srividya Bhaskara and Santosh Singh Rathore. Causal effect analysisbased intrusion detection system for IoT applications. International Journal of Information Security, 22(4):931–946, August 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00674-2.

Brandt:2006:HOF

[Bra06] Felix Brandt. How to obtain full privacy in auctions. International Journal of Information Security, 5(4):201–216, October 2006. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        006-0001-y.

Braeken:2022:PKV

[Bra22]  An Braeken.  Pub-

lic key versus symmetric key cryptography in client–server authentication protocols. International Journal of Information Security, 21(1): 103–114, February 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        021-00543-w.

Bohli:2006:KSA

[BRS06] Jens-Matthias Bohli, Stefan Röhrich, and Rainer Steinwandt. Key substitution attacks revisited: Taking into account malicious signers. International Journal of Information Security, 5(1):30–36, January 2006. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
005-0071-2.

Baldwin:2005:ESA

[BS05] Adrian Baldwin and Simon Shiu. Enabling shared audit data. International Journal of Information Security, 4(4): 263–276, October 2005. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
004-0061-9.

Borges:2021:EPP

[BS21] Ricard Borges and Francesc Sebé. An efficient privacypreserving pay-by-phone system for regulated parking areas. International Journal of Information Security, 20(5): 715–727, October 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00527-2.

Bhattacharjya:2022:EVS

[BS22] Sairath Bhattacharjya and Hossein Saiedian. Establishing and validating secured keys for IoT devices: using P3 connection model on a cloud-based architecture. International Journal of Information Security, 21(3):427–436, June 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        021-00562-7.

Borah:2024:UCT

[BSBK24] Parthajit Borah, Upasana Sarmah, D. K. Bhattacharyya, and J. K. Kalita. Unmasking the common traits: an ensemble approach for effective malware detection. International Journal of Information Security, 23(4): 2547–2557, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        024-00854-8.

Baumgarten:2011:CSH

[BSCZ11] Alex Baumgarten, Michael Steffen, Matthew Clausman, and Joseph Zambreno. A case study in hardware Trojan design and implementation. International Journal of Information Security, 10 (1):1–14, February 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
010-0115-0.

Bao:2020:LSP

[BSK^{+}20] Zijian Bao, Wenbo Shi, Saru Kumari, Zhi yin Kong, and Chien-Ming Chen. Lockmix: a secure and privacy-preserving mix service for Bitcoin anonymity. International Journal of Information Security, 19 (3):311–321, June 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00459-6.

Balikcioglu:2023:MCD

[BSK^{+}23] Pinar G. Balikcioglu, Melih Sirlanci, Ozge A. Kucuk, Bulut Ulukapi, Ramazan K. Turkmen, and Cengiz Acarturk. Malicious code detection in Android: the role of sequence characteristics and disassembling methods. International Journal of Information Security, 22(1): 107–118, February 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.

                                        com/article/10.1007/s10207-
                                        022-00626-2.

Bonte:2021:THM

[BST21] Charlotte Bonte, Nigel P. Smart, and Titouan Tanguy. Thresholdizing HashEd-DSA: MPC to the rescue. International Journal of Information Security, 20(6):879–894, December 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        021-00539-6.

Borges:2022:AUE

[BSV22] Ricard Borges, Francesc Sebé, and Magda Valls. An anonymous and unlinkable electronic toll collection system. International Journal of Information Security, 21(5): 1151–1162, October 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00604-8.

Bononi:2007:IDS

[BT07] Luciano Bononi and Carlo Tacconi. Intrusion detection for secure clustering and routing in Mobile Multi-hop Wireless Networks. International Journal of Information Security, 6(6):379–392, October 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0035-9.

Bajaj:2024:NAN

[BV24] Ashish Bajaj and Dinesh Kumar Vishwakarma. Non-alpha-num: a novel architecture for generating adversarial examples for bypassing NLPbased clickbait detection mechanisms. International Journal of Information Security, 23(4): 2711–2737, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00861-9.

Bohli:2007:SGK

[BVS07] Jens-Matthias Bohli, María Isabel González Vasco, and Rainer Steinwandt. Secure group key establishment revisited. International Journal of Information Security, 6(4):243–254, July 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0018-x.

Biskup:2008:KSI

[BW08] Joachim Biskup and Torben Weibert. Keeping secrets in incomplete databases. International Journal of Information Security, 7(3):199–217, June 2008. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        007-0037-7.

Baek:2003:ZSP

[BZ03] Joonsang Baek and Yuliang Zheng. Zheng and Seberry’s public key encryption scheme revisited. International Journal of Information Security, 2(1): 37–44, November 2003. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        003-0023-7.

Yehuda:2020:PAR

[BZ20] Raz Ben Yehuda and Nezer Jacob Zaidenberg. Protection against reverse engineering in ARM. International Journal of Information Security, 19 (1):39–51, February 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        019-00450-1.

Bai:2005:SAP

[BZV05] Yun Bai, Yan Zhang, and Vijay Varadharajan. On the sequence of authorization policy transformations. International Journal of Information Security, 4(1–2): 120–131, February 2005. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
004-0069-1.

Cevik:2022:LAG

[CAS22] Beyza Cevik, Nur Altiparmak, and Sevil Sen. Lib2Desc: automatic generation of security-centric Android app descriptions using third-party libraries. International Journal of Information Security, 21 (5):1107–1125, October 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00601-x.

Corbett:2008:PCW

[CBC08] Cherita L. Corbett, Raheem A. Beyah, and John A. Copeland. Passive classification of wireless NICs during active scanning. International Journal of Information Security, 7(5):335–348, October 2008. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0053-7.

Cheng:2020:SAA

[CBRY20] Peng Cheng, Ibrahim Ethem Bagci, Utz Roedig, and Jeff Yan. SonarSnoop: active acoustic side-channel attacks. International Journal of Information Security, 19(2):213–228, April 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        019-00449-8;   http://
                                        link.springer.com/content/
                                        pdf/10.1007/s10207-019-
                                        00449-8.pdf.

Chen:2010:HSS

[CC10] Haiyong Chen and Hailiang Chen. A hybrid scheme for securing fingerprint templates. International Journal of Information Security, 9(5):353–361, October 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        010-0114-1.

Chen:2012:DHP

[CC12] Liqun Chen and Yu Chen. The n-Diffie–Hellman problem and multiple-key encryption. International Journal of Information Security, 11(5):305–320, October 2012. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0171-8.

Cuppens:2008:MCS

[CCB08] Frédéric Cuppens and Nora Cuppens-Boulahia. Modeling contextual security policies. International Journal of Information Security, 7(4): 285–305, August 2008. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0051-9.

Cederquist:2007:ABC

[CCD^{+}07] J. G. Cederquist, R. Corin, M. A. C. Dekker, S. Etalle, J. I. den Hartog, and G. Lenzini. Audit-based compliance control. International Journal of Information Security, 6(2– 3):133–151, March 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0017-y.

Chen:2007:IBK

[CCS07] L. Chen, Z. Cheng, and N. P. Smart. Identitybased key agreement protocols from pairings. International Journal of Information Security, 6(4): 213–241, July 2007. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        006-0011-9.

Catalano:2013:FNI

[CDF^{+}13] Dario Catalano, Mario Di Raimondo, Dario Fiore, Rosario Gennaro, and Orazio Puglisi. Fully non-interactive onion routing with forward secrecy. International Journal of Information Security, 12 (1):33–47, February 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        012-0185-2.

Camtepe:2022:ABC

[CDP22] Seyit Camtepe, Jarek Duda, and Josef Pieprzyk. ANS-based compression and encryption with 128bit security. International Journal of Information Security, 21(5): 1051–1067, October 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00597-4.

Carminati:2003:MAC

[CF03] Barbara Carminati and Elena Ferrari. Management of access control policies for XML document sources. International Journal of Information Security, 1(4):236–260, July 2003. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
003-0020-x.

Clarke:2007:AMP

[CF07] N. L. Clarke and S. M.

Furnell. Authenticating mobile phone users using keystroke analysis. International Journal of Information Security, 6 (1):1–14, January 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
006-0006-6.

Chiasson:2009:UID

[CFBvO09] Sonia Chiasson, Alain Forget, Robert Biddle, and P. C. van Oorschot. User interface design affects security: patterns in clickbased graphical passwords. International Journal of Information Security, 8(6): 387–398, December 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
009-0080-7.

Catalano:2017:CAO

[CFG17] Dario Catalano, Dario Fiore, and Rosario Gennaro. A certificateless approach to onion routing. International Journal of Information Security, 16(3):327–343, June 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0337-x.

Catuogno:2014:ATF

[CG14] Luigi Catuogno and Clemente Galdi. Analysis of a two-factor graphical password scheme. International Journal of Information Security, 13(5): 421–437, October 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0228-y.

Cianciullo:2021:OCD

[CG21] Louis Cianciullo and Hossein Ghodosi. Outsourced cheating detection for secret sharing. International Journal of Information Security, 20(6): 871–878, December 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
021-00538-7.

Chowdhury:2023:PLT

[CG23] Nabin Chowdhury and Vasileios Gkioulos. A personalized learning theorybased cyber-security training exercise. International Journal of Information Security, 22(6):1531– 1546, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00704-z.

Coker:2011:PRA

[CGL^{+}11] George Coker, Joshua Guttman, Peter Loscocco, Amy Herzog, Jonathan Millen, Brian O’Hanlon, John Ramsdell, Ariel Segall, Justin Sheehy, and Brian Sniffen. Principles of remote attestation. International Journal of Information Security, 10(2):63–81, June 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
011-0124-7.

Cremers:2016:III

[CH16] Cas Cremers and Marko Horvat. Improving the ISO/IEC 11770 standard for key management techniques. International Journal of Information Security, 15(6):659–673, November 2016. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
015-0306-9;   http://
link.springer.com/content/
pdf/10.1007/s10207-015-
0306-9.pdf.

Cui:2024:DCV

[CHC24] Beibei Cui, Wei He, and Yan Cui. A dynamic C-V2X anonymous authentication and group key agreement protocol. International Journal of Information Security, 23(4): 2977–2989, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00876-2.

Camacho:2012:SAC

[CHKO12] Philippe Camacho, Alejandro Hevia, Marcos Kiwi, and Roberto Opazo. Strong accumulators from collision-resistant hashing. International Journal of Information Security, 11 (5):349–363, October 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0169-2.

Chen:2018:MGE

[CHM18] Liqun Chen, Jinguang Han, and Chris Mitchell. Message from the Guest Editors. International Journal of Information Security, 17(5):491–492, October 2018. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
018-0416-2.

Chen:2021:PAS

[CHMS21] Liqun Chen, Kaibin Huang, Mark Manulis, and Venkkatesh Sekar. Password-authenticated searchable encryption. International Journal of Information Security, 20(5): 675–693, October 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00524-5;   http://
link.springer.com/content/
pdf/10.1007/s10207-020-
00524-5.pdf.

Chen:2016:SOK

[CHZ16] Yu Chen, Qiong Huang, and Zongyang Zhang. Sakai–Ohgishi–Kasahara identity-based non-interactive key exchange revisited and more. International Journal of Information Security, 15(1):15–33, February 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        015-0274-0.

Chang:2019:CTM

[CJMS19] Donghoon Chang, Arpan Jati, Sweta Mishra, and Somitra Kumar Sanadhya. Cryptanalytic timememory trade-off for password hashing schemes. International Journal of Information Security, 18 (2):163–180, April 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        018-0405-5.

Crampton:2008:DRB

[CK08] Jason Crampton and Hemanth Khambhammettu. Delegation in role-based access control. International Journal of Information Security, 7(2):123– 136, April 2008. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        007-0044-8.

Cai:2023:CRT

[CK23a] Feiyang Cai and Xenofon Koutsoukos. Correction: Real-time detection of deception attacks in cyber-physical systems.

International Journal of Information Security, 22 (5):1383, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-

023-00696-w. See [CK23b].

Cai:2023:RTD

[CK23b] Feiyang Cai and Xenofon Koutsoukos. Real-time detection of deception attacks in cyber-physical systems. International Journal of Information Security, 22 (5):1099–1114, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00677-z. See correc-

tion [CK23a].

Chanal:2024:BBD

[CK24] Poornima M. Chanal and Mahabaleshwar S. Kakkasageri. Blockchain-based data integrity framework for Internet of Things. International Journal of Information Security, 23(1): 519–532, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00719-6.

Chatzoglou:2023:RQA

[CKKK23] Efstratios Chatzoglou, Vasileios Kouliaridis, Georgios Karopoulos, and Georgios Kambourakis. Revisiting QUIC attacks: a comprehensive review on QUIC security and a hands-on study. International Journal of Information Security, 22 (2):347–365, April 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00630-6.

Chew:2024:RTS

[CKPM24] Christopher Jun Wen Chew, Vimal Kumar, Panos Patros, and Robi Malik. Real-time system call-based ransomware detection. International Journal of Information Security, 23(3):1839–1858, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL https://

link.springer.com/article/
10.1007/s10207-024-00819-
x.

Chatterjee:2021:SEW

[CKS21] Sanjit Chatterjee, Manish Kesarwani, and Akash Shah. Secure and efficient wildcard search over encrypted data. International Journal of Information Security, 20 (2):199–244, April 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
020-00492-w.

Cai:2019:STP

[CKW19] Yixian Cai, George Karakostas, and Alan Wassyng. Secure and trusted partial grey-box verification. International Journal of Information Security, 18(6): 677–700, December 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00431-4.

Chen:2008:IID

[CL08] Pei-Te Chen and Chi-Sung Laih. IDSIC: an intrusion detection system with identification capability. International Journal of Information Security, 7(3):185–197, June 2008. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0024-z.

Cheng:2009:CKA

[CL09] Jiin-Chiou Cheng and Chi-Sung Laih. Conference key agreement protocol with non-interactive fault-tolerance over broadcast network. International Journal of Information Security, 8(1):37– 48, February 2009. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        008-0062-1.

Chen:2013:AMA

[CL13] Yu-Shian Chen and Chin-Laung Lei. Aggregate message authentication codes (AMACs) with on-the-fly verification. International Journal of Information Security, 12(6): 495–504, November 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        013-0202-0.

Chen:2023:EIS

[CLG23] Yingchun Chen, Jinguo Li, and Naiwang Guo. Efficient and interpretable SRU combined with Tab-Net for network intrusion detection in the big data environment. International Journal of Information Security, 22(3):679–689, June 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00656-w.

Chen:2024:IEC

[CLO^{+}24] Tzer-Long Chen, Chia-Hui Liu, Ya-Hui Ou, Yao-Min Huang, and Zhen-Yu Wu. An improved and efficient coercion-resistant measure for electronic voting system. International Journal of Information Security, 23(4):2637–2654, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00852-w.

Crampton:2011:UFC

[CLPP11] Jason Crampton, Hoon Wei Lim, Kenneth G. Paterson, and Geraint Price. Userfriendly and certificate-free grid security infrastructure. International Journal of Information Security, 10(3):137–153, June 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
011-0123-8.

Chang:2011:EHS

[CLW^{+}11] Ee-Chien Chang, Liming Lu, Yongzheng Wu, Roland H. C. Yap, and Jie Yu. Enhancing host security using external environment sensors. International Journal of Information Security, 10 (5):285–299, October 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

                                        com/article/10.1007/s10207-
                                        011-0130-9.

Chen:2016:MGE

[CM16] Liqun Chen and Chris Mitchell. Message from the guest editors. International Journal of Information Security, 15(6): 573–574, November 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        016-0353-x.

Chen:2024:MAU

[CMH^{+}24] Chun-Fu Chen, Bill Moriarty, Shaohan Hu, Sean Moran, Marco Pistoia, Vincenzo Piuri, and Pierangela Samarati. Model-agnostic utility-preserving biometric information anonymization. International Journal of Information Security, 23(4):2809–2826, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        024-00862-8.

Casassa-Mont:2015:TSI

[CMMPS15] Marco Casassa-Mont, Ilaria Matteucci, Marinella Petrocchi, and Marco Luca Sbodio. Towards safer information sharing in the cloud. International Journal of Information Security, 14(4):319–334, August 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0258-5.

Cimitile:2018:TNM

[CMN^{+}18] Aniello Cimitile, Francesco Mercaldo, Vittoria Nardone, Antonella Santone, and Corrado Aaron Visaggio. Talos: no more ransomware victims with formal methods. International Journal of Information Security, 17(6): 719–738, November 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0398-5.

Chakrabarti:2006:KPD

[CMR06]  Dibyendu  Chakrabarti,

Subhamoy Maitra, and Bimal Roy. A key predistribution scheme for wireless sensor networks: merging blocks in combinatorial design. International Journal of Information Security, 5(2):105–114, April 2006. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
006-0085-4.

Chakraborty:2010:CDB

[CMS10] Anindya Chakraborty, Arun K.

Majumdar, and Shamik Sural. A column dependencybased approach for static and dynamic recovery of databases from malicious transactions. International Journal of Information Security, 9(1):51– 67, February 2010. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
009-0095-0.

Chadwick:2009:ASX

[CON09] David W. Chadwick, Sassa Otenko, and Tuan Anh Nguyen. Adding support to XACML for multidomain user to user dynamic delegation of authority. International Journal of Information Security, 8(2):137–152, April 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
008-0073-y.

Chakravarty:2015:DAE

[CPPK15] Sambuddho Chakravarty, Georgios Portokalidis, Michalis Polychronakis, and Angelos D. Keromytis. Detection and analysis of eavesdropping in anonymous communication networks. International Journal of Information Security, 14(3):205–220, June 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0256-7.

Casas-Roma:2020:DGD

[CR20] Jordi Casas-Roma. DUEF-GA: data utility and privacy evaluation framework for graph anonymization. International Journal of Information Security, 19 (4):465–478, August 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00469-4.

Cheimonidis:2024:DVS

[CR24] Pavlos Cheimonidis and Konstantinos Rantos. Dynamic vulnerability severity calculator for industrial control systems. International Journal of Information Security, 23(4): 2655–2676, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00858-4.

Celdran:2023:IBB

[CSC^{+}23] Alberto Huertas Celdrán, Pedro Miguel Sánchez Sánchez, Miguel Azorín Castillo, Gérôme Bovet, Gregorio Martínez Pérez, and Burkhard Stiller. Intelligent and behavioralbased detection of malware in IoT spectrum sensors. International Journal of Information Security, 22(3):541–561, June 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00602-w.

Chimuco:2023:SCB

[CSL^{+}23] Francisco T. Chimuco, João B. F. Sequeiros, Carolina Galvão Lopes, Tiago M. C. Simões, Mário M. Freire, and Pedro R. M. Inácio. Secure cloud-based mobile apps: attack taxonomy, requirements, mechanisms, tests and automation. International Journal of Information Security, 22(4):833–867, August 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00669-z.

Chillara:2024:DSM

[CSM^{+}24] Anil Kumar Chillara, Paresh Saxena, Rajib Ranjan Maiti, Manik Gupta, Raghu Kondapalli, Zhichao Zhang, and Krishnakumar Kesavan. Deceiving supervised machine learning models via adversarial data poisoning attacks: a case study with USB keyboards. International Journal of Information Security, 23(3):2043–2061, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00834-y.

Chimuco:2024:EDD

[CSS^{+}24] Francisco T. Chimuco, Joāo B. F. Sequeiros, Tiago M. C. Simões, Mário M. Freire, and Pedro R. M. Inácio. Expediting the design and development of secure cloudbased mobile apps. International Journal of Information Security, 23(4): 3043–3064, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00880-6.

Celdran:2016:RPP

[CTM^{+}16] Alberto Huertas Celdrán, Ginés Dólera Tormo, Félix Gómez Mármol, Manuel Gil Pérez, and Gregorio Martínez Pérez. Resolving privacypreserving relationships over outsourced encrypted data storages. International Journal of Information Security, 15 (2):195–209, April 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        015-0283-z.

Chakraborty:2024:BBA

[CTN24] Partha Sarathi Chakraborty, Somanath Tripathy, and Sanjeet Kumar Nayak. BASPED: Blockchain assisted searchable public key encryption over outsourced data. International Journal of Information Security, 23(1):487–503, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00753-4.

Chiba:2018:DTA

[CYA^{+}18] Daiki Chiba, Takeshi Yagi, Mitsuaki Akiyama, Toshiki Shibahara, Tatsuya Mori, and Shigeki Goto. Domain-Profiler: toward accurate and early discovery of domain names abused in future. International Journal of Information Security, 17(6):661–680, November 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        017-0396-7;   http://
                                        link.springer.com/content/
                                        pdf/10.1007/s10207-017-
                                        0396-7.pdf.

Cook:2009:EBC

[CYK09] Debra L. Cook, Moti Yung, and Angelos D. Keromytis. Elastic block ciphers: method, security and instantiations. International Journal of Information Security, 8 (3):211–231, June 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
008-0075-9.

Crescini:2006:PSD

[CZ06] Vino Fernando Crescini and Yan Zhang. PolicyUpdater: a system for dynamic access control. International Journal of Information Security, 5 (3):145–165, July 2006. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
005-0078-8.

Deldar:2021:EST

[DA21] Fatemeh Deldar and Mahdi Abadi. Enhancing spatial and temporal utilities in differentially private moving objects database release. International Journal of Information Security, 20(4):511–533, August 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        020-00516-5.

deAlbuquerque:2010:FVA

[dAKdG10] João Porto de Albuquerque, Heiko Krumm, and Paulo Lício de Geus. Formal validation of automated policy refinement in the management of network security systems. International Journal of Information Security, 9 (2):99–125, April 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        010-0101-6.

Danezis:2007:BFM

[Dan07] George Danezis. Breaking four mix-related schemes based on Universal Reencryption. International Journal of Information Security, 6(6): 393–402, October 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        007-0033-y.

Das:2012:RKE

[Das12] Ashok Kumar Das. A random key establishment scheme for multi-phase deployment in large-scale distributed sensor networks. International Journal of Information Security, 11 (3):189–211, June 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0162-9.

Dawson:2004:PSI

[Daw04]  E. Dawson.  Preface

to the special issue on PKI. International Journal of Information Security, 2(2):65, January 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
003-0031-7;   http://
link.springer.com/content/
pdf/10.1007/s10207-003-
0031-7.pdf.

Dupasquier:2010:AIL

[DBMS10] Benoît Dupasquier, Stefan Burschka, Kieran McLaughlin, and Sakir Sezer. Analysis of information leakage from encrypted Skype conversations. International Journal of Information Security, 9(5):313–325, October 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
010-0111-4.

DArco:2013:TTR

[DdP13] Paolo D’Arco and Angel Perez del Pozo. Toward tracing and revoking schemes secure against collusion and any form of secret information leakage. International Journal of Information Security, 12(1):1–17, February 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0186-1.

Damiani:2002:SSS

[DDPS02] E. Damiani, S. De Capitani di Vimercati, S. Paraboschi, and P. Samarati. Securing SOAP e-services. International Journal of Information Security, 1(2): 100–115, February 2002. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s102070100009.

Dyer:2019:PHE

[DDX19] James Dyer, Martin Dyer, and Jie Xu. Practical homomorphic encryption over the integers for secure computation in the cloud. International Journal of Information Security, 18(5): 549–579, October 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00427-0.

DArco:2022:GWP

[DDZ22] P. D’Arco, R. De Prisco, and R. Zaccagnino. Gossamer: weaknesses and performance. International Journal of Information Security, 21(3):669–687, June 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00575-2.

Dent:2008:SCE

[Den08] Alexander W. Dent. A survey of certificateless encryption schemes and security models. International Journal of Information Security, 7(5): 349–377, October 2008. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
008-0055-0.

Desoky:2009:LLB

[Des09] Abdelrahman Desoky. Listega: list-based steganography methodology. International Journal of Information Security, 8(4): 247–261, August 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
009-0079-0.

Domingo-Ferrer:2018:DGS

[DFBJR18] Josep Domingo-Ferrer, Alberto Blanco-Justicia, and Carla Ràfols. Dynamic group size accreditation and group discounts preserving anonymity. International Journal of Information Security, 17 (3):243–260, June 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0368-y.

Degabriele:2016:UPC

[DFF^{+}16] Jean Paul Degabriele, Victoria Fehr, Marc Fischlin, Tommaso Gagliardoni, Felix Günther, Giorgia Azzurra Marson, Arno Mittelbach, and Kenneth G. Paterson. Unpicking PLAID: a cryptographic analysis of an ISO-standards-track authentication protocol. International Journal of Information Security, 15(6): 637–657, November 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
015-0309-6.

Duessel:2017:DZD

[DGF^{+}17] Patrick Duessel, Christian Gehl, Ulrich Flegel, Sven Dietrich, and Michael Meier. Detecting zeroday attacks using contextaware anomaly detection at the application-layer. International Journal of Information Security, 16(5): 475–490, October 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0344-y.

Draper-Gil:2013:OFE

[DGZFGH13] G. Draper-Gil, J. Zhou, J. L. Ferrer-Gomila, and M. F. Hinarejos. An optimistic fair exchange protocol with active intermediaries. International Journal of Information Security, 12(4):299–318, August 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0194-9.

Daza:2004:PUI

[DHS04] Vanesa Daza, Javier Herranz, and Germán Sáez. Protocols useful on the Internet from distributed signature schemes. International Journal of Information Security, 3(2): 61–69, November 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
004-0043-y.

DiPierro:2011:PTC

[DHW11] Alessandra Di Pierro, Chris Hankin, and Herbert Wiklicky. Probabilistic timing covert channels: to close or not to close? International Journal of Information Security, 10(2):83–106, June 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        010-0107-0.

Damgaard:2010:GPP

[DJN10] Ivan Damgård, Mads Jurik, and Jesper Buus Nielsen. A generalization of Paillier’s public-key system with applications to electronic voting. International Journal of Information Security, 9(6): 371–385, December 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        010-0119-9.

Dhingra:2024:RLM

[DJS^{+}24] Bhavya Dhingra, Vidhi Jain, Deepak Kumar Sharma, Koyel Datta Gupta, and Deepika Kukreja. RLET: a lightweight model for ubiquitous multi-class intrusion detection in sustainable and secured smart environment. International Journal of Information Security, 23(1): 315–330, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00739-2.

Dinesh:2024:EET

[DK24] K. Dinesh and S. V.

N. Santhosh Kumar. Energyefficient trust-aware secured neuro-fuzzy clustering with sparrow search optimization in wireless sensor network. International Journal of Information Security, 23(1): 199–223, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00737-4.

Dolzhenko:2015:MRE

[DLR15] Egor Dolzhenko, Jay Ligatti, and Srikar Reddy. Modeling runtime enforcement with mandatory results automata. International Journal of Information Security, 14 (1):47–60, February 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0239-8.

DeFrancesco:2007:ILS

[DM07] Nicoletta De Francesco and Luca Martini. Instructionlevel security typing by abstract interpretation. International Journal of Information Security, 6(2– 3):85–106, March 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        007-0015-0.

DeKeulenaer:2016:LTS

[DMDD16] Ronald De Keulenaer, Jonas Maebe, Koen De Bosschere, and Bjorn De Sutter. Link-time smart card code hardening. International Journal of Information Security, 15 (2):111–130, April 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        015-0282-0.

DeCristofaro:2013:PDC

[DMP13] Emiliano De Cristofaro, Mark Manulis, and Bertram Poettering. Private discovery of common social contacts. International Journal of Information Security, 12(1):49–65, February 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.

com/article/10.1007/s10207-
012-0183-4.

Dimitrakos:2007:GEP

[DMRS07] Theo Dimitrakos, Fabio Martinelli, Peter Y. A. Ryan, and Steve Schneider. Guest Editors’ preface. International Journal of Information Security, 6(2–3):65–66, March 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0012-3.

Dasgupta:2019:DIN

[DNF^{+}19] Dipankar Dasgupta, Abhijit Kumar Nag, Denise Ferebee, Sanjib Kumar Saha, Kul Prasad Subedi, Arunava Roy, Alvaro Madero, Abel Sanchez, and John R. Williams. Design and implementation of Negative Authentication System. International Journal of Information Security, 18 (1):23–48, February 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0395-8.

Dasgupta:2024:IDP

[DPMC24] Rupshali Dasgupta, Meghabriti Pramanik, Pabitra Mitra, and Dipanwita Roy Chowdhury. Intrusion detection for power grid: a review. International Journal of Information Security, 23(2):1317–1329, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL https://

link.springer.com/article/
10.1007/s10207-023-00789-
6.

Dewri:2012:OSH

[DRPW12] Rinku Dewri, Indrajit Ray, Nayot Poolsappasit, and Darrell Whitley. Optimal security hardening on attack tree models of networks: a cost-benefit analysis. International Journal of Information Security, 11(3):167–188, June 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0160-y.

Delicata:2007:AAV

[DS07] Rob Delicata and Steve Schneider. An algebraic approach to the verification of a class of Diffie– Hellman protocols. International Journal of Information Security, 6(2– 3):183–196, March 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0013-2.

Das:2019:DSI

[DSB19] Debasish Das, Utpal Sharma, and D. K. Bhattacharyya. Defeating SQL injection attack in authentication security: an experimental study. International Journal of Information Security, 18(1):1–22, February 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0393-x.

Faria:2019:DAA

[dSFK19] Gerson de Souza Faria and Hae Yong Kim. Differential audio analysis: a new side-channel attack on PIN pads. International Journal of Information Security, 18(1):73–84, February 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
018-0403-7.

Diaz-Santiago:2016:CST

[DSRHC16] Sandra Díaz-Santiago, Lil María Rodríguez-Henríquez, and Debrup Chakraborty. A cryptographic study of tokenization systems. International Journal of Information Security, 15(4): 413–432, August 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        015-0313-x.

Ding:2006:CNT

[DSY06] Jintai Ding, Dieter Schmidt, and Zhijun Yin. Cryptanalysis of the new TTS scheme in CHES 2004. International Journal of Information Security, 5(4): 231–240, October 2006. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        006-0003-9.

Deepa:2018:BBD

[DTK^{+}18] G. Deepa, P. Santhi Thilagam, Furqan Ahmed Khan, Amit Praseed, Alwyn R. Pais, and Nushafreen Palsetia. Blackbox detection of XQuery injection and parameter tampering vulnerabilities in web applications. International Journal of Information Security, 17(1): 105–120, February 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        016-0359-4.

Degano:2008:P

[DV08] Pierpaolo Degano and Luca Viganò. Preface. International Journal of Information Security, 7(1):

1, January 2008. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0038-6.

Dawson:2002:CCP

[DVB02] Ed Dawson, Kapali Viswanathan, and Colin Boyd. Compliant cryptologic protocols. International Journal of Information Security, 1(3): 189–202, November 2002. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
002-0016-y.

Dwivedi:2024:IJI

[Dwi24] Rahul Dwivedi. International journal of information security: a bibliometric study, 2007–2023. International Journal of Information Security, 23 (3):2159–2187, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00840-0.

Din:2023:TSM

[DWU^{+}23] Nizamud Din, Abdul Waheed, Shamsher Ullah, Noor Ul Amin, Gautam Srivastava, Farhan Ullah, and Jerry Chun-Wei Lin.

A typology of secure multicast communication over 5 G/6 G networks. International Journal of Information Security, 22(4): 1055–1073, August 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00678-y.

Ding:2010:NHA

[DYDW10] Xuhua Ding, Yanjiang Yang, Robert H. Deng, and Shuhong Wang. A new hardware-assisted PIR with O(n) shuffle cost. International Journal of Information Security, 9 (4):237–252, August 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
010-0105-2.

Deng:2018:SPT

[DZW^{+}18] Hua Deng, Yunya Zhou, Qianhong Wu, Bo Qin, and Jianwei Liu. Secure pay-TV for chained hotels. International Journal of Information Security, 17(1):33–42, February 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0354-9.

Esponda:2007:PDP

[EAH^{+}07] Fernando Esponda, Elena S.

Ackley, Paul Helman, Haixia Jia, and Stephanie Forrest. Protecting data privacy through hard-to-reverse negative databases. International Journal of Information Security, 6(6): 403–415, October 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0030-1.

Ekstedt:2023:YAC

[EAM^{+}23] Mathias Ekstedt, Zeeshan Afzal, Preetam Mukherjee, Simon Hacks, and Robert Lagerström. Yet another cybersecurity risk assessment framework. International Journal of Information Security, 22 (6):1713–1729, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00713-y.

Ebrahimpour:2024:AIT

[EB24] Elham Ebrahimpour and Shahram Babaie. Authentication in Internet of Things, protocols, attacks, and open issues: a systematic literature review. International Journal of Information Security, 23 (3):1583–1602, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00806-8.

ElHassani:2015:ION

[EEB^{+}15] Abdeljebar Ameziane El Hassani, Anas Abou El Kalam, Adel Bouhoula, Ryma Abassi, and Abdellah Ait Ouahman. Integrity-OrBAC: a new model to preserve Critical Infrastructures integrity. International Journal of Information Security, 14(4): 367–385, August 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0254-9.

ElMendili:2024:EDM

[EFB^{+}24] Fatna El Mendili, Mohammed Fattah, Nisrine Berros, Youness Filaly, and Younès El Bouzekri El Idrissi. Enhancing detection of malicious profiles and spam tweets with an automated honeypot framework powered by deep learning. International Journal of Information Security, 23(2): 1359–1388, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00796-7.

Esponda:2009:NRI

[EFH09] Fernando Esponda, Stephanie Forrest, and Paul Helman. Negative representations of information. International Journal of Information Security, 8(5): 331–345, October 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
009-0078-1.

ElSalamouny:2018:ONF

[EG18] Ehab ElSalamouny and Sébastien Gambs. Optimal noise functions for location privacy on continuous regions. International Journal of Information Security, 17(6): 613–630, November 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0384-y.

Egners:2015:MOW

[EHM15] André Egners, Patrick Herrmann, and Ulrike Meyer. Multi-operator wireless mesh networks secured by an all-encompassing security architecture. International Journal of Information Security, 14 (2):169–186, April 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0244-y.

Emura:2014:GSI

[EHSS14] Keita Emura, Goichiro Hanaoka, Yusuke Sakai, and Jacob C. N. Schuldt. Group signature implies public-key encryption with non-interactive opening. International Journal of Information Security, 13 (1):51–62, February 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0204-y.

Enomoto:2024:EMC

[EKY^{+}24] Shuhei Enomoto, Hiroki Kuzuno, Hiroshi Yamada, Yoshiaki Shiraishi, and Masakatu Morii. Early mitigation of CPU-optimized ransomware using monitoring encryption instructions. International Journal of Information Security, 23(5):3393–3413, October 2024. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00892-2.

Elmiger:2024:STG

[ELPB24] Marius Elmiger, Mouad Lemoudden, Nikolaos Pitropakis, and William J. Buchanan. Start thinking in graphs: using graphs to address critical attack paths in a Microsoft cloud tenant. International Journal of Information Security, 23(1): 467–485, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00751-6.

Esfahani:2017:EHM

[EMRN17] Alireza Esfahani, Georgios Mantas, Jonathan Rodriguez, and José Carlos Neves. An efficient homomorphic MAC-based scheme against data and tag pollution attacks in network coding-enabled wireless networks. International Journal of Information Security, 16(6): 627–639, November 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0351-z.

Amin:2024:BBM

[EOC^{+}24] Habib El Amin, Lina Oueidat, Maroun Chamoun, Abed Ellatif Samhat, and Antoine Feghali. Blockchainbased multi-organizational cyber risk management framework for collaborative environments. International Journal of Information Security, 23(2): 1231–1249, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00788-7.

Empl:2024:GIV

[ESSP24] Philip Empl, Daniel Schlette, Lukas Stöger, and Günther Pernul. Generating ICS vulnerability playbooks with open standards. International Journal of Information Security, 23(2): 1215–1230, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00760-5.

Estevez-Tapiador:2008:BRE

[ETAHCR08] Juan M. Estevez-Tapiador, Almudena Alcaide, Julio C. Hernandez-Castro, and Arturo Ribagorda. Bayesian rational exchange. International Journal of Information Security, 7(1): 85–100, January 2008. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

com/article/10.1007/s10207-
007-0039-5.

Elberri:2024:CDS

[ETRLG24] Mustafa Ahmed Elberri, Ümit Tokeşer, Javad Rahebi, and Jose Manuel Lopez-Guede. A cyber defense system against phishing attacks with deep learning game theory and LSTM–CNN with African vulture optimization algorithm (AVOA). International Journal of Information Security, 23(4): 2583–2606, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00851-x.

Elmufti:2009:MWS

[EWR^{+}09] Kalid Elmufti, Dasun Weerasinghe, M. Rajarajan, Veselin Rakocevic, Sanowar Khan, and John A. MacDonald. Mobile Web services authentication using SAML and 3GPP generic bootstrapping architecture. International Journal of Information Security, 8(2): 77–87, April 2009. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
008-0065-y.

Erdodi:2022:AWM

[EZ22] László Erdődi and Fabio Massimo Zennaro. The Agent Web Model: modeling web hacking for reinforcement learning. International Journal of Information Security, 21 (2):293–309, April 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        021-00554-7.

El-Zawawy:2021:DLN

[EZLC21] Mohamed A. El-Zawawy, Eleonora Losiouk, and Mauro Conti. Do not let Next-Intent Vulnerability be your next nightmare: type system-based approach to detect it in Android apps. International Journal of Information Security, 20 (1):39–58, February 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        020-00491-x.

Far:2024:GFB

[FAH24] Saeed Banaeian Far, Maryam Rajabzadeh Asaar, and Afrooz Haghbin. A generic framework for blockchainassisted on-chain auditing for off-chain storage. International Journal of Information Security, 23 (3):2407–2435, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00846-8.

Farhat:2023:CMD

[FAMMZ23] Saida Farhat, Manel Abdelkader, Amel Meddeb-Makhlouf, and Faouzi Zarai. CADS-ML/DL: efficient cloud-based multiattack detection system. International Journal of Information Security, 22 (6):1989–2013, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00729-4.

Ferraris:2021:TMP

[FBFGEM21] Davide Ferraris, Daniel Bastos, Carmen Fernandez-Gago, and Fadi El-Moussa. A trust model for popular smart home devices. International Journal of Information Security, 20 (4):571–587, August 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00519-2.

Frati:2024:CTH

[FDS^{+}24] Fulvio Frati, Georgiana Darau, Nikos Salamanos, Pantelitsa Leonidou, Costas Iordanou, Dimitris Plachouris, Efstratios Syrmas, Evangelos Floros, George Nikitakis, George Spanoudakis, Konstantinos Kalais, Stella Tsichlaki, Ernesto Damiani, George C. Kagadis, Jihane Najar, and Michael Sirivianos. Cybersecurity training and healthcare: the AERAS approach. International Journal of Information Security, 23(2): 1527–1539, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00802-y.

Ferraris:2020:TTR

[FFG20] Davide Ferraris and Carmen Fernandez-Gago. TrUStAPIS: a trust requirements elicitation method for IoT. International Journal of Information Security, 19(1): 111–127, February 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00438-x.

Ferrer-Gomila:2024:AFC

[FGH24] Josep-Lluís Ferrer-Gomila and M. Francisca Hinarejos. Abuse-freeness in contract signing: a blockchainbased proposal. International Journal of Information Security, 23(3): 1963–1974, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00818-y.

Fiore:2012:RBS

[FGS12] D. Fiore, R. Gennaro, and N. P. Smart. Relations between the security models for certificateless encryption and ID-based key agreement. International Journal of Information Security, 11 (1):1–22, February 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
011-0149-y.

Faust:2018:OPM

[FHV18] Sebastian Faust, Carmit Hazay, and Daniele Venturi. Outsourced pattern matching. International Journal of Information Security, 17(3):327–346, June 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0374-0.

Feng:2024:SAI

[FJZZ24] Tao Feng, Wentao Jin, Minghu Zhang, and Lu Zheng. Security assessment and improvement of smart grid NIKE protocol. International Journal of Information Security, 23(1): 411–429, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00746-3.

Faghani:2019:MBM

[FN19] Mohammad R. Faghani and Uyen T. Nguyen. Mobile botnets meet social networks: design and analysis of a new type of botnet. International Journal of Information Security, 18(4):423–449, August 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
018-0412-6.

Fong:2008:DCC

[Fon08] Philip W. L. Fong. Discretionary capability confinement. International Journal of Information Security, 7(2):137–154, April 2008. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0047-5.

Farao:2024:ICI

[FPP^{+}24] Aristeidis Farao, Georgios Paparis, Sakshyam Panda, Emmanouil Panaousis, Apostolis Zarras, and Christos Xenakis. INCHAIN: a cyber insurance architecture with smart contracts and self-sovereign identity on top of blockchain. International Journal of Information Security, 23(1): 347–371, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00741-8.

Frattolillo:2018:WPE

[Fra18] Franco Frattolillo. Watermarking protocols: an excursus to motivate a new approach. International Journal of Information Security, 17(5): 587–601, October 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0386-9.

Farras:2019:PSP

[FRG19] Oriol Farràs and Jordi Ribes-González. Provably secure public-key encryption with conjunctive and subset keyword search. International Journal of Information Security, 18(5):

533–548, October 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        018-00426-7.

Fernandes:2014:SIC

[FSG^{+}14] Diogo A. B. Fernandes, Liliana F. B. Soares, João V. Gomes, Mário M. Freire, and Pedro R. M. Inácio. Security issues in cloud environments: a survey. International Journal of Information Security, 13 (2):113–170, April 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        013-0208-7.

Fan:2020:SHB

[FTS^{+}20] Chun-I Fan, Yi-Fan Tseng, Hui-Po Su, Ruei-Hau Hsu, and Hiroaki Kikuchi. Secure hierarchical Bitcoin wallet scheme against privilege escalation attacks. International Journal of Information Security, 19 (3):245–255, June 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        019-00476-5.

Fu:2022:CGP

[FYF22] Bingxue Fu, Xing Yu, and Tao Feng. CT-GCN: a phishing identification model for blockchain cryptocurrency transactions. International Journal of Information Security, 21 (6):1223–1232, December 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00606-6.

Faisal:2021:SAT

[FZ21] Abu Faisal and Mohammad Zulkernine. A secure architecture for TCP/ UDP-based cloud communications. International Journal of Information Security, 20(2):161– 179, April 2021. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
020-00511-w.

Gandhi:2023:TUS

[GA23] Kapilan Kulayan Arumugam Gandhi and Chamundeswari Arumugam. Toward a unified and secure approach for extraction of forensic digital evidence from an IoT device. International Journal of Information Security, 22 (2):417–431, April 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00645-z.

Gimenez-Aguilar:2023:MUB

[GAdFGM23] Mar Gimenez-Aguilar, Jose Maria de Fuentes, and Lorena Gonzalez-Manzano. Malicious uses of blockchains by malware: from the analysis to Smart–Zephyrus. International Journal of Information Security, 22(5): 1445–1480, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00700-3.

Guo:2024:CAS

[GBC24] Yuejun Guo, Seifeddine Bettaieb, and Fran Casino. A comprehensive analysis on software vulnerability detection datasets: trends, challenges, and road ahead. International Journal of Information Security, 23 (5):3311–3327, October 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        024-00888-y.

Granadillo:2014:RBC

[GBDJ14] Gustavo Gonzalez Granadillo, Malek Belhaouane, Hervé Debar, and Grégoire Jacob. RORI-based countermeasure selection using the OrBAC formalism. International Journal of Information Security, 13(1):63–79, February 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0207-8.

Gruber:2018:UTB

[GBG18] A. Gruber and I. Ben-Gal.

Using targeted Bayesian network learning for suspect identification in communication networks. International Journal of Information Security, 17 (2):169–181, April 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0362-4.

Gao:2019:LBD

[GCH^{+}19] Wen Gao, Liqun Chen, Yupu Hu, Christopher J. P. Newton, Baocang Wang, and Jiangshan Chen. Lattice-based deniable ring signatures. International Journal of Information Security, 18 (3):355–370, June 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
018-0417-1;   http://
link.springer.com/content/

                                        pdf/10.1007/s10207-018-
                                        0417-1.pdf.

Guerra-Casanova:2012:AMD

[GCSÁBdSS12] J. Guerra-Casanova, C. Sánchez-Ávila, G. Bailador, and A. de Santos Sierra. Authentication in mobile devices through hand gesture recognition. International Journal of Information Security, 11(2):65–83, April 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        012-0154-9.

Goenka:2024:CSP

[GCT24] Richa Goenka, Meenu Chawla, and Namita Tiwari. A comprehensive survey of phishing: mediums, intended targets, attack and defence techniques and a novel taxonomy. International Journal of Information Security, 23(2):819–848, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00768-x.

Gong:2022:TDR

[GDA22] Qian Gong, Phil DeMar, and Mine Altunay. ThunderSecure: deploying realtime intrusion detection for 100G research networks by leveraging stream-based features and one-class classification network. International Journal of Information Security, 21(4): 799–812, August 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00584-9.

Garcia:2014:WLS

[GdKGV14] Flavio D. Garcia, Gerhard de Koning Gans, and Roel Verdult. Wirelessly lockpicking a smart card reader. International Journal of Information Security, 13(5): 403–420, October 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0234-0.

Goncalves:2024:OGC

[GdSdC24] Ricardo Martins Gonçalves, Miguel Mira da Silva, and Paulo Rupino da Cunha. Olympus: a GDPR compliant blockchain system. International Journal of Information Security, 23 (2):1021–1036, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00782-z.

Ghorbel:2022:APP

[GGJ22] Amal Ghorbel, Mahmoud Ghorbel, and Mohamed Jmaiel. Accountable privacy preserving attributebased access control for cloud services enforced using blockchain. International Journal of Information Security, 21 (3):489–508, June 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        021-00565-4.

Granese:2021:EMP

[GGP21] Federica Granese, Daniele Gorla, and Catuscia Palamidessi. Enhanced models for privacy and utility in continuoustime diffusion networks. International Journal of Information Security, 20 (5):763–782, October 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        020-00530-7.

Guttman:2005:RAN

[GH05] Joshua D. Guttman and Amy L. Herzog. Rigorous automated network security management. International Journal of Information Security, 4(1– 2):29–48, February 2005. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
004-0052-x.

Gulyas:2019:HIA

[GI19] Gábor György Gulyás and Sándor Imre. Hiding information against structural re-identification. International Journal of Information Security, 18 (2):125–139, April 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
018-0400-x.

Gritzalis:2012:FAR

[GKBS12] Dimitris Gritzalis, Panagiotis Katsaros, Stylianos Basagiannis, and Yannis Soupionis. Formal analysis for robust anti-SPIT protection using model checking. International Journal of Information Security, 11(2):121–135, April 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0159-4.

Gauravaram:2010:HFU

[GKKT10] Praveen Gauravaram, John Kelsey, Lars R. Knudsen, and Søren S. Thomsen. On hash functions using checksums. International Journal of Information Security, 9(2):137–151, April 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        009-0100-7.

Grining:2019:PPP

[GKS19] Krzysztof Grining, Marek Klonowski, and Piotr Syga. On practical privacypreserving fault-tolerant data aggregation. International Journal of Information Security, 18 (3):285–304, June 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        018-0413-5.

Gerasimou:2024:SPS

[GL24] Stylianos Gerasimou and Konstantinos Limniotis. A study on privacy and security aspects of personalised apps. International Journal of Information Security, 23(5): 3217–3239, October 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        024-00887-z.

Gorrieri:2004:AAT

[GLMS^{+}04] Roberto Gorrieri, Ruggero Lanotte, Andrea Maggiolo-Schettini, Fabio Martinelli, Simone Tini, and Enrico Tronci. Automated analysis of timed security: a case study on web privacy. International Journal of Information Security, 2(3– 4):168–186, August 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
004-0037-9.

Garra:2019:RAB

[GLMS19] Ricard Garra, Dominik Leibenger, Josep M. Miret, and Francesc Sebé. Repairing an aggregation-based smart metering system. International Journal of Information Security, 18(5): 637–646, October 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00435-0.

Gurgens:2003:ACP

[GLP03] Sigrid Gürgens, Javier Lopez, and René Peralta. Analysis of e-commerce protocols: Adapting a traditional technique. International Journal of Information Security, 2(1): 21–36, November 2003. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

                                        com/article/10.1007/s10207-
                                        003-0021-9.

Gonzalez-Manzano:2023:TCA

[GMdFLR23] Lorena González-Manzano, José M. de Fuentes, Flavio Lombardi, and Cristina Ramos. A technical characterization of APTs by leveraging public resources. International Journal of Information Security, 22 (6):1567–1584, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00706-x.

Gomez:2021:CCM

[GMGM21] Juan Manuel Castelo Gómez, Javier Carrillo Mondéjar, José Roldán Gómez, and José Luis Martínez Martínez. A context-centered methodology for IoT forensic investigations. International Journal of Information Security, 20(5): 647–673, October 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        020-00523-6.

Gouglidis:2014:SPV

[GMH14] Antonios Gouglidis, Ioannis Mavridis, and Vincent C. Hu. Security policy verification for multidomains in cloud systems.

International Journal of Information Security, 13 (2):97–111, April 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0205-x.

Gruner:2023:ACS

[GMLM23] Andreas Grüner, Alexander Mühle, Niko Lockenvitz, and Christoph Meinel. Analyzing and comparing the security of selfsovereign identity management systems through threat modeling. International Journal of Information Security, 22(5): 1231–1248, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00688-w.

Galindo:2005:FOH

[GMMV05] David Galindo, Sebastià Martín, Paz Morillo, and Jorge L. Villar. Fujisaki– Okamoto hybrid encryption revisited. International Journal of Information Security, 4(4): 228–241, October 2005. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
004-0042-z.

Giorgini:2006:RET

[GMMZ06] Paolo Giorgini, Fabio Massacci, John Mylopoulos, and Nicola Zannone. Requirements engineering for trust management: model, methodology, and reasoning. International Journal of Information Security, 5(4):257–274, October 2006. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        006-0005-7.

Gollman:2001:E

[GMO01] Dieter Gollman, Catherine A. Meadows, and Eiji Okamoto. Editorial. International Journal of Information Security, 1(1):1–2, August 2001. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s102070100004.

Geiselmann:2003:AIP

[GMS03] Willi Geiselmann, Willi Meier, and Rainer Steinwandt. An attack on the isomorphisms of polynomials problem with one secret. International Journal of Information Security, 2(1):59–64, November 2003. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.

com/article/10.1007/s10207-
003-0025-5.

Guardiola-Muzquiz:2023:SCS

[GMSS23] Gorka Guardiola-Múzquiz and Enrique Soriano-Salvador. SealFSv2: combining storage-based and ratcheting for tamperevident logging. International Journal of Information Security, 22 (2):447–466, April 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00643-1.

Gao:2014:TTR

[GNS14] Weizheng Gao, Kashi Neupane, and Rainer Steinwandt. Tuning a two-round group key agreement. International Journal of Information Security, 13(5): 467–476, October 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0225-6.

Garcia:2011:SID

[GOBdlC11] Sergio Sánchez García, Ana Gómez Oliva, Emilia Pérez Belleboni, and Iván Pau de la Cruz. Solving identity delegation problem in the e-government environment. International Journal of Information Security, 10(6): 351–372, November 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        011-0140-7.

Gollmann:2008:E

[Gol08] Dieter Gollmann. Editorial. International Journal of Information Security, 7(2):101, April 2008. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        007-0052-8.

Golic:2012:NAM

[Gol12] Jovan Dj. Golić. A new authentication model for ad hoc networks. International Journal of Information Security, 11(5): 333–347, October 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        012-0167-4.

Gunther:2017:LMT

[GP17] Felix Günther and Bertram Poettering. Linkable message tagging: solving the key distribution problem of signature schemes. International Journal of Information Security, 16 (3):281–297, June 2017.

CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0327-z.

Grontas:2023:AEP

[GP23] Panagiotis Grontas and Aris Pagourtzis. Anonymity and everlasting privacy in electronic voting. International Journal of Information Security, 22(4): 819–832, August 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00666-2.

Garg:2017:NBD

[GPS17] Shree Garg, Sateesh K.

Peddoju, and Anil K. Sarje. Network-based detection of Android malicious apps. International Journal of Information Security, 16(4): 385–400, August 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0343-z.

Gritzalis:2006:PKI

[Gri06] Stefanos Gritzalis. Public Key Infrastructure: Research and applications. International Journal of Information Security, 5 (1):1–2, January 2006. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        005-0075-y.

Gurgens:2005:SFN

[GRV05] Sigrid Gürgens, Carsten Rudolph, and Holger Vogt. On the security of fair nonrepudiation protocols. International Journal of Information Security, 4(4): 253–262, October 2005. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        004-0063-7.

Gomez:2024:IEC

[GRV24] Juan Manuel Castelo Gómez and Sergio Ruiz-Villafranca. Integrating the edge computing paradigm into the development of IoT forensic methodologies. International Journal of Information Security, 23(2): 1093–1116, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00776-x.

Giffhorn:2015:NAL

[GS15] Dennis Giffhorn and Gregor Snelting. A new algorithm for low-deterministic security. International Journal of Information Security, 14(3):263–287, June 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0257-6.

Gonzalez-Serrano:2018:SML

[GSAMCA18] Francisco-Javier González-Serrano, Adrián Amor-Martín, and Jorge Casamayón-Antón. Supervised machine learning using encrypted training data. International Journal of Information Security, 17(4): 365–377, August 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0381-1.

Glisson:2011:ERW

[GSM^{+}11] William Bradley Glisson, Tim Storer, Gavin Mayall, Iain Moug, and George Grispos. Electronic retention: what does your mobile phone reveal about you? International Journal of Information Security, 10(6):337–349, November 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
011-0144-3.

Gritti:2016:BED

[GSP^{+}16] Clémentine Gritti, Willy Susilo, Thomas Plantard, Kaitai Liang, and Duncan S. Wong. Broadcast encryption with dealership. International Journal of Information Security, 15(3):271–283, June 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        015-0285-x.

Goldschlag:2010:THB

[GSS10] David M. Goldschlag, Stuart G. Stubblebine, and Paul F. Syverson. Temporarily hidden bit commitment and lottery applications. International Journal of Information Security, 9(1):33– 50, February 2010. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        009-0094-1.

Gyorffy:2011:TBG

[GTM11] John Charles Gyorffy, Andrew F. Tappenden, and James Miller. Tokenbased graphical password authentication. International Journal of Information Security, 10(6): 321–336, November 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
011-0147-0.

Geron:2009:CCR

[GW09] Erel Geron and Avishai Wool. CRUST: cryptographic remote untrusted storage without public keys. International Journal of Information Security, 8(5):357–377, October 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
009-0081-6.

Gritzalis:2007:PMO

[GYL^{+}07] S. Gritzalis, A. N. Yannacopoulos, C. Lambrinoudakis, P. Hatzopoulos, and S. K. Katsikas. A probabilistic model for optimal insurance contracts against security risks and privacy violation in IT outsourcing environments. International Journal of Information Security, 6 (4):197–211, July 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
006-0010-x.

Guan:2023:AHO

[GZH^{+}23] Zhenyu Guan, Lixin Zhang, Bohan Huang, Bihe Zhao, and Song Bian. Adaptive hyperparameter optimization for black-box adversarial attack. International Journal of Information Security, 22 (6):1765–1779, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00716-9.

Haider:2024:CEF

[HAAI24] Raja Zeeshan Haider, Baber Aslam, Haider Abbas, and Zafar Iqbal. C2-Eye: framework for detecting command and control (C2) connection of supply chain attacks. International Journal of Information Security, 23(4): 2531–2545, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00850-y.

Hale:2020:UMA

[Hal20] Britta Hale. User-mediated authentication protocols and unforgeability in key collision. International Journal of Information Security, 19(6):609–621, December 2020. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

com/article/10.1007/s10207-
019-00479-2.

Hamdi:2023:FLB

[Ham23]  Najet Hamdi.  Feder-

ated learning-based intrusion detection system for Internet of Things. International Journal of Information Security, 22 (6):1937–1948, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00727-6.

Hilario:2024:GAP

[HAS^{+}24] Eric Hilario, Sami Azam, Jawahar Sundaram, Khwaja Imran Mohammed, and Bharanidharan Shanmugam. Generative AI for pentesting: the good, the bad, the ugly. International Journal of Information Security, 23(3):2075–2097, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00835-x.

Hoglund:2024:APK

[HBF^{+}24] Joel Höglund, Simon Bouget, Martin Furuhed, John Preuß Mattsson, Göran Selander, and Shahid Raza. AutoPKI: public key infrastructure for IoT with automated trust transfer. International Journal of Information Security, 23 (3):1859–1875, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00825-z.

Hadziosmanovic:2012:LMA

[HBH12] Dina Hadziosmanović, Damiano Bolzoni, and Pieter H. Hartel. A log mining approach for process monitoring in SCADA. International Journal of Information Security, 11(4): 231–251, August 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0163-8;   http://
link.springer.com/content/
pdf/10.1007/s10207-012-
0163-8.pdf.

Hadjadj:2022:OPF

[HBK22] Taha Elamine Hadjadj, Adel Bouhoula, and Riadh Ksantini. Optimization of parallel firewalls filtering rules. International Journal of Information Security, 21(2):323–340, April 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00557-4.

Han:2010:CBI

[HC10] Shui-Hua Han and Chao-Hsien Chu. Contentbased image authentication: current status, issues, and challenges. International Journal of Information Security, 9(1): 19–32, February 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
009-0093-2.

Hoang:2015:GAM

[HCN15] Thang Hoang, Deokjai Choi, and Thuc Nguyen. Gait authentication on mobile phone using biometric cryptosystem and fuzzy commitment scheme. International Journal of Information Security, 14(6): 549–560, November 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
015-0273-1.

Hamdi:2024:VAA

[HFA24] Ahlem Hamdi, Lamia Fourati, and Samiha Ayed. Vulnerabilities and attacks assessments in blockchain 1.0, 2.0 and 3.0: tools, analysis and countermeasures. International Journal of Information Security, 23(2):713–757, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00765-0.

Heluany:2024:RDT

[HG24] Jessica B. Heluany and Vasileios Gkioulos. A review on digital twins for power generation and distribution. International Journal of Information Security, 23(2): 1171–1195, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00784-x.

Hamzenejadi:2023:MBD

[HGH23] Sajad Hamzenejadi, Mahdieh Ghazvini, and Seyedamiryousef Hosseini. Mobile botnet detection: a comprehensive survey. International Journal of Information Security, 22(1): 137–175, February 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00624-4.

Hu:2016:EWS

[HH16] Changhui Hu and Lidong Han. Efficient wildcard search over encrypted data. International Journal of Information Security, 15 (5):539–547, October 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
015-0302-0.

Hamila:2024:ESF

[HHSS24] Firas Hamila, Mohammad Hamad, Daniel Costa Salgado, and Sebastian Steinhorst. Enhancing security in Fiat–Shamir transformation-based noninteractive zero-knowledge protocols for IoT authentication. International Journal of Information Security, 23(2): 1131–1148, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00779-8.

Hinarejos:2019:DPE

[HIDFGHR19] M. Francisca Hinarejos, Andreu-Pere Isern-Deyà, Josep-Lluís Ferrer-Gomila, and Llorenç Huguet-Rotger. Deployment and performance evaluation of mobile multicoupon solutions. International Journal of Information Security, 18(1): 101–124, February 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

                                        com/article/10.1007/s10207-
                                        018-0404-6;   http://
                                        link.springer.com/content/
                                        pdf/10.1007/s10207-018-
                                        0404-6.pdf.

Hasegawa:2009:PFC

[HIST09] Shingo Hasegawa, Shuji Isobe, Hiroki Shizuya, and Katsuhiro Tashiro. On the pseudo-freeness and the CDH assumption. International Journal of Information Security, 8(5): 347–355, October 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        009-0087-0.

Hadavi:2015:SSS

[HJDC15] Mohammad Ali Hadavi, Rasool Jalili, Ernesto Damiani, and Stelvio Cimato. Security and searchability in secret sharing-based data outsourcing. International Journal of Information Security, 14(6):513–529, November 2015. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        015-0277-x.

Hiemenz:2019:DSS

[HK19] Benedikt Hiemenz and Michel Krämer. Dynamic searchable symmetric encryption for storing geospatial data in the cloud. International Journal of Information Security, 18 (3):333–354, June 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
018-0414-4.

Huszti:2022:SPB

[HKO22] Andrea Huszti, Szabolcs Kovács, and Norbert Oláh. Scalable, password-based and threshold authentication for smart homes. International Journal of Information Security, 21 (4):707–723, August 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00578-7.

Hopcroft:2004:ASA

[HL04] Philippa Hopcroft and Gavin Lowe. Analysing a stream authentication protocol using model checking. International Journal of Information Security, 3(1):2–13, October 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
004-0040-1.

Hatzivasilis:2024:SIM

[HLA^{+}24] George Hatzivasilis, Eftychia Lakka, Manos Athanatos, Sotiris Ioannidis, Grigoris Kalogiannis, Manolis Chatzimpyrros, George Spanoudakis, Spyros Papastergiou, Stylianos Karagiannis, Andreas Alexopoulos, Dimitry Amelin, and Stephan Kiefer. Swarmintelligence for the modern ICT ecosystems. International Journal of Information Security, 23(4): 2951–2975, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        024-00869-1.

Han:2015:MAU

[HLKI15] Kyoung Soo Han, Jae Hyun Lim, Boojoong Kang, and Eul Gyu Im. Malware analysis using visualized images and entropy graphs. International Journal of Information Security, 14 (1):1–14, February 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        014-0242-0.

Ham:2018:IYP

[HLS18] HyoungMin Ham, JongHyup Lee, and JooSeok Song. Improved yoking proof protocols for preserving anonymity. International Journal of Information Security, 17(4): 379–393, August 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0383-z.

Hougaard:2022:ALO

[HM22] Hector B. Hougaard and Atsuko Miyaji. Authenticated logarithmic-order supersingular isogeny group key exchange. International Journal of Information Security, 21 (2):207–221, April 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00549-4.

Hitchcock:2004:ESM

[HMCD04] Yvonne Hitchcock, Paul Montague, Gary Carter, and Ed Dawson. The efficiency of solving multiple discrete logarithm problems and the implications for the security of fixed elliptic curves. International Journal of Information Security, 3(2): 86–98, November 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

                                        com/article/10.1007/s10207-
                                        004-0045-9.

Herranz:2014:SEA

[HN14] Javier Herranz and Jordi Nin. Secure and efficient anonymization of distributed confidential databases. International Journal of Information Security, 13(6):497–512, November 2014. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        014-0237-x.

Hannousse:2023:DLM

[HNHY23] Abdelhakim Hannousse, Mohamed Cherif Nait-Hamoud, and Salima Yahiouche. A deep learner model for multi-language webshell detection. International Journal of Information Security, 22 (1):47–61, February 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00615-5.

Harn:2009:EIB

[HRL09] Lein Harn, Jian Ren, and Changlu Lin. Efficient identity-based GQ multisignatures. International Journal of Information Security, 8(3):205–210, June 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
008-0072-z.

Holzl:2020:DDA

[HRMM20] Michael Hölzl, Michael Roland, Omid Mir, and René Mayrhofer. Disposable dynamic accumulators: toward practical privacy-preserving mobile eIDs with scalable revocation. International Journal of Information Security, 19(4):401–417, August 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00458-7;   http://
link.springer.com/content/
pdf/10.1007/s10207-019-
00458-7.pdf.

Hammer:2009:FSC

[HS09] Christian Hammer and Gregor Snelting. Flowsensitive, context-sensitive, and object-sensitive information flow control based on program dependence graphs. International Journal of Information Security, 8(6): 399–422, December 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
009-0086-1.

Halevi:2015:KAS

[HS15] Tzipora Halevi and Nitesh Saxena. Keyboard acoustic side channel attacks: exploring realistic and security-sensitive scenarios. International Journal of Information Security, 14 (5):443–456, October 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        014-0264-7.

Haverkamp:2024:EMC

[HS24] Indy Haverkamp and Dipti K. Sarmah. Evaluating the merits and constraints of cryptography-steganography fusion: a systematic analysis. International Journal of Information Security, 23(4):2607–2635, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        024-00853-9.

Hayata:2024:PIR

[HSHM24] Junichiro Hayata, Jacob C. N. Schuldt, Goichiro Hanaoka, and Kanta Matsuura. On private information retrieval supporting range queries. International Journal of Information Security, 23(1): 629–647, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00743-6.

Haugli-Sandvik:2024:MDM

[HSLB24] Marie Haugli-Sandvik, Mass Soldal Lund, and Frøy Birte Bjørneseth. Maritime decision-makers and cyber security: deck officers’ perception of cyber risks towards IT and OT systems. International Journal of Information Security, 23(3):1721–1739, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00810-y.

Huang:2008:SUD

[HSMW08] Xinyi Huang, Willy Susilo, Yi Mu, and Wei Wu. Secure universal designated verifier signature without random oracles. International Journal of Information Security, 7 (3):171–183, June 2008. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0021-2.

Han:2012:NCO

[HSMY12] Jinguang Han, Willy Susilo, Yi Mu, and Jun Yan. New constructions of OSBE schemes and their applications in oblivious access control. International Journal of Information Security, 11(6): 389–401, November 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0176-3.

Hanley:2011:UTD

[HTM11] Neil Hanley, Michael Tunstall, and William P. Marnane. Using templates to distinguish multiplications from squaring operations. International Journal of Information Security, 10(4):255–266, August 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
011-0135-4.

Huber:2012:PSS

[Hub12] Michael Huber. Perfect secrecy systems immune to spoofing attacks. International Journal of Information Security, 11(4): 281–289, August 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0166-5.

Higgins:2023:CPR

[HXTP23] Martin Higgins, Wangkun Xu, Fei Teng, and Thomas Parisini. Cyber-physical risk assessment for false data injection attacks considering moving target defences. International Journal of Information Security, 22(3):579–589, June 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00621-7.

Huang:2011:ESD

[HYWS11] Qiong Huang, Guomin Yang, Duncan S. Wong, and Willy Susilo. Efficient strong designated verifier signature schemes without random oracle or with non-delegatability. International Journal of Information Security, 10(6): 373–385, November 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
011-0146-1.

Huang:2012:NEO

[HYWS12] Qiong Huang, Guomin Yang, Duncan S. Wong, and Willy Susilo. A new efficient optimistic fair exchange protocol without random oracles. International Journal of Information Security, 11 (1):53–63, February 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
011-0152-3.

Huang:2017:MAP

[HZL^{+}17] Cheng-Ta Huang, Yu-Hong Zhang, Li-Chiun Lin, Wei-Jen Wang, and Shiuh-Jeng Wang. Mutual authentications to parties with QR-code applications in mobile systems. International Journal of Information Security, 16(5): 525–540, October 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0349-6.

Isern-Deya:2015:PUG

[IDHRPCMP15] Andreu Pere Isern-Deyà, Llorenç Huguet-Rotger, M. Magdalena Payeras-Capellà, and Macià Mut-Puigserver. On the practicability of using group signatures on mobile devices: implementation and performance analysis on the Android platform. International Journal of Information Security, 14(4): 335–345, August 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0259-4.

Itakura:2002:PPI

[IHNT02] Yukio Itakura, Masaki Hashiyada, Toshio Nagashima, and Shigeo Tsujii. Proposal on personal identifiers generated from the STR information of DNA. International Journal of Information Security, 1(3): 149–160, November 2002. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
002-0013-1.

Inayoshi:2022:VUT

[IKS22] Hiroki Inayoshi, Shohei Kakei, and Shoichi Saito. Value-utilized taint propagation: toward precise detection of apps’ information flows across Android API calls. International Journal of Information Security, 21(5): 1127–1149, October 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00603-9.

Imamura:2018:IAA

[IMI18] Kazuya Imamura, Kazuhiko Minematsu, and Tetsu Iwata. Integrity analysis of authenticated encryption based on stream ciphers. International Journal of Information Security, 17 (5):493–511, October 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0378-9.

Ince:2024:EPD

[Inc24] Kenan Ince. Exploring the potential of deep learning and machine learning techniques for randomness analysis to enhance security on IoT. International Journal of Information Security, 23(2): 1117–1130, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00783-y.

Ibnugraha:2021:RMD

[INS21] Prajna Deshanta Ibnugraha, Lukito Edi Nugroho, and Paulus Insap Santosa. Risk model development for information security in organization environment based on business perspectives. International Journal of Information Security, 20(1): 113–126, February 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

com/article/10.1007/s10207-
020-00495-7.

Imamura:2021:WAM

[IOU^{+}21] Yuta Imamura, Rintaro Orito, Hiroyuki Uekawa, Kritsana Chaikaew, Pattara Leelaprute, Masaya Sato, and Toshihiro Yamauchi. Web access monitoring mechanism via Android WebView for threat analysis. International Journal of Information Security, 20(6): 833–847, December 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00534-3;   http://
link.springer.com/content/
pdf/10.1007/s10207-020-
00534-3.pdf.

Ibrahim:2023:CSS

[IS23] Nasir Ibrahim and Harin Sellahewa. A cross-setting study of user unlocking behaviour in a graphical authentication scheme: a case study on Android Pattern Unlock. International Journal of Information Security, 22 (6):1849–1863, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00722-x.

Itakura:2005:PMB

[IT05] Yukio Itakura and Shigeo Tsujii. Proposal on a multifactor biometric authentication method based on cryptosystem keys containing biometric signatures. International Journal of Information Security, 4(4):288–296, October 2005. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        004-0065-5.

Imamoto:2008:AEC

[IZS08] Kenji Imamoto, Jianying Zhou, and Kouichi Sakurai. Achieving evenhandedness in certified email system for contract signing. International Journal of Information Security, 7(6): 383–394, November 2008. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        008-0056-z.

Josephlal:2021:EIR

[JAYZ21] Edwin Franco Myloth Josephlal, Sridhar Adepu, Zheng Yang, and Jianying Zhou. Enabling isolation and recovery in PLC redundancy framework of metro train systems. International Journal of Information Security, 20(6):

783–795, December 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00529-0.

Jouini:2021:QAS

[JBK21] Mouna Jouini, Latifa Ben Arfa Rabai, and Ridha Khedri. A quantitative assessment of security risks based on a multifaceted classification approach. International Journal of Information Security, 20(4): 493–510, August 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00515-6.

Jia:2018:ERH

[JCL^{+}18] Hongyong Jia, Yue Chen, Julong Lan, Kaixiang Huang, and Jun Wang. Efficient revocable hierarchical identity-based encryption using cryptographic accumulators. International Journal of Information Security, 17(4): 477–490, August 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0387-8.

James:2015:AIP

[JG15] Joshua I. James and Pavel Gladyshev. Automated inference of past action instances in digital investigations. International Journal of Information Security, 14(3):249–261, June 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        014-0249-6.

Jovanovikj:2014:CMS

[JGK14] Vladimir Jovanovikj, Dusan Gabrijelcic, and Tomaz Klobucar. A conceptual model of security context. International Journal of Information Security, 13(6):571–581, November 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        014-0229-x.

Ji:2021:EVA

[JJJ21] Soo-Yeon Ji, Bong-Keun Jeong, and Dong Hyun Jeong. Evaluating visualization approaches to detect abnormal activities in network traffic data. International Journal of Information Security, 20 (3):331–345, June 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00504-9.

Junaid:2024:CBR

[JJQ24] Kashif Junaid, Muhammad Umar Janjua, and Junaid Qadir. A compliancebased ranking of certificate authorities using probabilistic approaches. International Journal of Information Security, 23(4): 2881–2910, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00867-3.

Jeet:2022:SIP

[JK22] Rabari Jeet and P. Arun Raj Kumar. A survey on interest packet flooding attacks and its countermeasures in named data networking. International Journal of Information Security, 21(5):1163–1187, October 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00591-w.

Jin:2024:FID

[JKLJ24] Philgeun Jin, Namjun Kim, Sangjin Lee, and Doowon Jeong. Forensic investigation of the dark web on the Tor network: pathway toward the surface web. International Journal of Information Security, 23(1):331–346, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00745-4.

Joh:2017:PSV

[JM17] HyunChul Joh and Yashwant K. Malaiya. Periodicity in software vulnerability discovery, patching and exploitation. International Journal of Information Security, 16(6): 673–690, November 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        016-0345-x.

Johnson:2001:ECD

[JMV01] Don Johnson, Alfred Menezes, and Scott Vanstone. The Elliptic Curve Digital Signature Algorithm (ECDSA). International Journal of Information Security, 1 (1):36–63, August 2001. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s102070100002.

Jiang:2018:CPA

[JSMG18a] Yinhao Jiang, Willy Susilo, Yi Mu, and Fuchun Guo. Ciphertext-policy attribute-based encryption supporting access policy update and its extension with preserved attributes. International Journal of Information Security, 17 (5):533–548, October 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0388-7.

Jiang:2018:FCP

[JSMG18b] Yinhao Jiang, Willy Susilo, Yi Mu, and Fuchun Guo. Flexible ciphertext-policy attribute-based encryption supporting ANDgate and threshold with short ciphertexts. International Journal of Information Security, 17(4): 463–475, August 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0376-y.

Jia:2021:CNI

[JT21] Huiwen Jia and Chunming Tang. Cryptanalysis of a non-interactive deniable ring signature scheme. International Journal of Information Security, 20(1): 103–112, February 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

                                        com/article/10.1007/s10207-
                                        020-00497-5.

Jain:2024:IDL

[JT24] Vikas Kumar Jain and Meenakshi Tripathi. An integrated deep learning model for Ethereum smart contract vulnerability detection. International Journal of Information Security, 23(1): 557–575, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00752-5.

Joudaki:2019:ETS

[JTV19] Zeinab Joudaki, Julie Thorpe, and Miguel Vargas Martin. Enhanced Tacit Secrets: System-assigned passwords you can’t write down, but don’t need to. International Journal of Information Security, 18 (2):239–255, April 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        018-0408-2.

Jajodia:2011:MGE

[JZ11] Sushil Jajodia and Jianying Zhou. Message from the Guest Editors. International Journal of Information Security, 10(5): 267–268, October 2011.

CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
011-0131-8;   http://
link.springer.com/content/
pdf/10.1007/s10207-011-
0131-8.pdf.

Kirubavathi:2018:SAD

[KA18] G. Kirubavathi and R. Anitha.

Structural analysis and detection of Android botnets using machine learning techniques. International Journal of Information Security, 17 (2):153–167, April 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0363-3.

Karimi:2016:UAA

[KAC16] Vahid R. Karimi, Paulo S. C. Alencar, and Donald D. Cowan. A uniform approach for access control and business models with explicit rule realization. International Journal of Information Security, 15(2):145–171, April 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
015-0275-z.

Karimi:2017:FMA

[KAC17] Vahid R. Karimi, Paulo S. C. Alencar, and Donald D. Cowan. A formal modeling and analysis approach for access control rules, policies, and their combinations. International Journal of Information Security, 16 (1):43–74, February 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        016-0314-4.

Kifouche:2022:DIN

[KAK22] Abdenour Kifouche, Mohamed Salah Azzaz, and Remy Kocik. Design and implementation of a new lightweight chaosbased cryptosystem to secure IoT communications. International Journal of Information Security, 21 (6):1247–1262, December 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00609-3.

Kaur:2023:PPP

[KAK23] Jasleen Kaur, Alka Agrawal, and Raees Ahmad Khan. P2ADF: a privacy-preserving attack detection framework in fog-IoT environment. International Journal of Information Security, 22(4):749–762, August 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00661-7.

Kakisim:2024:DLA

[Kak24] Arzu Gorgulu Kakisim.

A deep learning approach based on multi-view consensus for SQL injection detection. International Journal of Information Security, 23(2): 1541–1556, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00791-y.

Kundu:2013:PPA

[KB13] Ashish Kundu and Elisa Bertino. Privacy-preserving authentication of trees and graphs. International Journal of Information Security, 12(6): 467–494, November 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0198-5.

Kumar:2022:SML

[KB22] Raju Kumar and Aruna Bhat. A study of machine learning-based models for detection, control, and mitigation of cyberbullying in online social media. International Journal of Information Security, 21 (6):1409–1431, December 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00600-y.

Kuppusamy:2023:APS

[KB23] K. S. Kuppusamy and G. Balayogi. Accessible password strength assessment method for visually challenged users. International Journal of Information Security, 22 (6):1731–1741, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00714-x.

Knight:2007:IJI

[KBH07] Scott Knight, Scott Buffett, and Patrick C. K. Hung. The International Journal of Information Security special issue on privacy, security and trust technologies and e-business services. International Journal of Information Security, 6(5): 285–286, September 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0036-8.

Karl:2022:DNI

[KBJ22] Ryan Karl, Hannah Burchfield, and Taeho Jung. Developing non-interactive MPC with trusted hardware for enhanced security. International Journal of Information Security, 21(4):777–797, August 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00583-w.

Kim:2022:MPD

[KBY22] Jinsung Kim, Younghoon Ban, and Jeong Hyun Yi. MAPAS: a practical deep learning-based Android malware detection system. International Journal of Information Security, 21 (4):725–738, August 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00579-6.

Kwarteng:2024:MTM

[KC24] Emmanuel Kwarteng and Mumin Cebe. MEDICAL-HARM: A threat modeling designed for modern medical devices and a comprehensive study on effectiveness, user satisfaction, and security perspectives. International Journal of Information Security, 23(3):2225–2268, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00826-y.

Kotzanikolaou:2017:BAR

[KCB17] Panayiotis Kotzanikolaou, George Chatzisofroniou, and Mike Burmester. Broadcast anonymous routing (BAR): scalable real-time anonymous communication. International Journal of Information Security, 16(3):313–326, June 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0318-0.

Ko:2020:FDD

[KCB20] Ili Ko, Desmond Chambers, and Enda Barrett. Feature dynamic deep learning approach for DDoS mitigation within the ISP domain. International Journal of Information Security, 19 (1):53–70, February 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00453-y.

Kuzlu:2023:ASM

[KCC^{+}23] Murat Kuzlu, Ferhat Ozgur Catak, Umit Cali, Evren Catak, and Ozgur Guler. Adversarial security mitigations of mmWave beamforming prediction models using defensive distillation and adversarial retraining. International Journal of Information Security, 22(2):319–332, April 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00644-0.

Kozakevicius:2015:UQS

[KCM^{+}15] Alice Kozakevicius, Cristian Cappo, Bruno A. Mozzaquatro, Raul Ceretta Nunes, and Christian E. Schaerer. URL query string anomaly sensor designed with the bidimensional Haar wavelet transform. International Journal of Information Security, 14 (6):561–581, November 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
015-0276-y.

Khan:2022:PAK

[KDM22] Haibat Khan, Benjamin Dowling, and Keith M. Martin. Pragmatic authenticated key agreement for IEEE Std 802.15.6. International Journal of Information Security, 21 (3):577–595, June 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00570-7.

Kulah:2019:SAD

[KDYS19] Yusuf Kulah, Berkay Dincer, Cemal Yilmaz, and Erkay Savas. SpyDetector: an approach for detecting side-channel attacks at runtime. International Journal of Information Security, 18(4): 393–422, August 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
018-0411-7.

Kekul:2024:EVM

[KEA24] Hakan Kekül, Burhan Ergen, and Halil Arslan. Estimating vulnerability metrics with word embedding and multiclass classification methods. International Journal of Information Security, 23(1): 247–270, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00734-7.

Kate:2011:GCB

[KG11] Aniket Kate and Ian Goldberg. Generalizing cryptosystems based on the subset sum problem. International Journal of Information Security, 10 (3):189–199, June 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
011-0129-2.

Kozina:2009:MIW

[KGG09] Mario Kozina, Marin Golub, and Stjepan Gros. A method for identifying Web applications. International Journal of Information Security, 8(6): 455–467, December 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
009-0092-3.

Kishore:2022:SPD

[KIP22] R. Kishore, I. Ioannou, and A. Pitsillides. A security protocol for D2D communications in 5G networks using elliptic curve cryptography. International Journal of Information Security, 21(6):1389– 1408, December 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00620-8.

Kim:2014:EVE

[KJ14] Kee Sung Kim and Ik Rae Jeong. Efficient verifiably encrypted signatures from lattices. International Journal of Information Security, 13(4): 305–314, August 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        014-0226-0.

Kong:2011:SSA

[KJG^{+}11] Deguang Kong, Yoon-Chan Jhi, Tao Gong, Sencun Zhu, Peng Liu, and Hongsheng Xi. SAS: semantics aware signature generation for polymorphic worm detection. International Journal of Information Security, 10(5): 269–283, October 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        011-0132-7.

Kim:2017:MES

[KJS17] Minchul Kim, Younghoon Jung, and Junghwan Song. A modified exhaustive search on a password system using SHA-1. International Journal of Information Security, 16 (3):263–269, June 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0332-2.

Kananizadeh:2017:DDP

[KK17] Shahrzad Kananizadeh and Kirill Kononenko. Development of dynamic protection against timing channels. International Journal of Information Security, 16(6):641–651, November 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0356-7.

Kazoleas:2022:NMR

[KK22] Ioannis Kazoleas and Panagiotis Karampelas. A novel malicious remote administration tool using stealth and self-defense techniques. International Journal of Information Security, 21 (2):357–378, April 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00559-2.

Kolias:2017:TDS

[KKK17] Constantinos Kolias, Vasilis Kolias, and Georgios Kambourakis. TermID: a distributed swarm intelligencebased approach for wireless intrusion detection. International Journal of Information Security, 16 (4):401–416, August 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0335-z.

Kim:2022:DAD

[KKK22a] Hee Yeon Kim, Ji Hoon Kim, and Kyounggon Kim. DAPP: automatic detection and analysis of prototype pollution vulnerability in Node.js modules. International Journal of Information Security, 21(1):1–23, February 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
020-00537-0.

Koike:2022:IOH

[KKK22b] Kazuki Koike, Ryotaro Kobayashi, and Masahiko Katoh. IoT-oriented high-efficient anti-malware hardware focusing on time series metadata extractable from inside a processor core. International Journal of Information Security, 21(4):1–19, August 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00577-0.

Klinkoff:2007:ENS

[KKKV07] Patrick Klinkoff, Engin Kirda, Christopher Kruegel, and Giovanni Vigna. Extending .NET security to unmanaged code. International Journal of Information Security, 6(6): 417–428, October 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0031-0.

Kalaria:2024:ACA

[KKR^{+}24] Rudri Kalaria, A. S. M.

Kayes, Wenny Rahayu, Eric Pardede, and Ahmad Salehi Shahraki. Adaptive context-aware access control for IoT environments leveraging fog computing. International Journal of Information Security, 23(4): 3089–3107, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00866-4.

Kim:2023:ERS

[KKY^{+}23] Moon-Seok Kim, Sungho Kim, Sang-Kyung Yoo, Bong-Soo Lee, Ji-Man Yu, Il-Woong Tcho, and Yang-Kyu Choi. Error reduction of SRAM-based physically unclonable function for chip authentication. International Journal of Information Security, 22(5): 1087–1098, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00668-0.

Koshutanski:2009:EGS

[KLMM09] Hristo Koshutanski, Aliaksandr Lazouski, Fabio Martinelli, and Paolo Mori. Enhancing grid security by fine-grained behavioral control and negotiationbased authorization. International Journal of Information Security, 8(4): 291–314, August 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
009-0083-4.

Kim:2021:ISD

[KLPL21] Hyoseung Kim, Kwangsu Lee, Jong Hwan Park, and Dong Hoon Lee. Improving the security of direct anonymous attestation under host corruptions. International Journal of Information Security, 20(4):475–492, August 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00507-6.

Khan:2021:EAB

[KLZ^{+}21] Fawad Khan, Hui Li, Yinghui Zhang, Haider Abbas, and Tahreem Yaqoob. Efficient attribute-based encryption with repeated attributes optimization. International Journal of Information Security, 20 (3):431–444, June 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00505-8.

Kremer:2003:FMP

[KM03] Steve Kremer and Olivier Markowitch. Fair multiparty non-repudiation protocols. International Journal of Information Security, 1(4):223–235, July 2003. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
003-0019-3.

Kopf:2007:TTU

[KM07] Boris Köpf and Heiko Mantel. Transformational typing and unification for automatically correcting insecure programs. International Journal of Information Security, 6(2– 3):107–131, March 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        007-0016-z.

Knudsen:2010:CEA

[KM10] Lars R. Knudsen and Charlotte V. Miolane. Counting equations in algebraic attacks on block ciphers. International Journal of Information Security, 9 (2):127–135, April 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        009-0099-9.

Komano:2022:CBS

[KM22] Yuichi Komano and Takaaki Mizuki. Coin-based secure computations. International Journal of Information Security, 21(4): 833–846, August 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00585-8.

Kasamatsu:2016:TSE

[KME^{+}16] Kohei Kasamatsu, Takahiro Matsuda, Keita Emura, Nuttapong Attrapadung, Goichiro Hanaoka, and Hideki Imai. Time-specific encryption from forwardsecure encryption: generic and direct constructions. International Journal of Information Security, 15 (5):549–571, October 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
015-0304-y.

Korzhik:2003:HAB

[KML03] Valery Korzhik and Guillermo Morales-Luna. Hybrid authentication based on noisy channels. International Journal of Information Security, 1(4):203–210, July 2003. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
002-0017-x.

Katsantonis:2023:CRD

[KMMG23] M. N. Katsantonis, A. Manikas, I. Mavridis, and D. Gritzalis. Cyber range design framework for cyber security education and training. International Journal of Information Security, 22(4):1005–1027, August 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00680-4.

Kuper:2009:GXS

[KMR09] Gabriel Kuper, Fabio Massacci, and Nataliya Rassadko. Generalized XML security views. International Journal of Information Security, 8(3):173– 203, June 2009. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        008-0074-x.

Khorasgani:2024:AIP

[KMW24] Hossein Abedi Khorasgani, Noman Mohammed, and Yang Wang. Attribute inference privacy protection for pre-trained models. International Journal of Information Security, 23(3):2269–2285, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        024-00839-7.

Krukow:2007:TS

[KN07] Karl Krukow and Mogens Nielsen. Trust structures. International Journal of Information Security, 6(2– 3):153–181, March 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        007-0014-1.

Kurek:2016:TBC

[KNL16] Tytus Kurek, Marcin Niemiec, and Artur Lason. Taking back control of privacy: a novel framework for preserving cloudbased firewall policy confidentiality. International Journal of Information Security, 15(3):235–250, June 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
015-0292-y;   http://
link.springer.com/content/
pdf/10.1007/s10207-015-
0292-y.pdf.

Karagiannis:2024:MDP

[KNM^{+}24] Stylianos Karagiannis, Christoforos Ntantogian, Emmanouil Magkos, Aggeliki Tsohou, and Luís Landeiro Ribeiro. Mastering data privacy: leveraging K-anonymity for robust health data sharing. International Journal of Information Security, 23 (3):2189–2201, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00838-8.

Kunang:2024:EEI

[KNSS24] Yesi Novaria Kunang, Siti Nurmaini, Deris Stiawan, and Bhakti Yudho Suprapto. An end-to-end intrusion detection system with IoT dataset using deep learning with unsupervised feature extraction. International Journal of Information Security, 23(3):1619–1648, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00807-7.

Kim:2002:NEC

[KO02] S. Kim and H. Oh. A new electronic check system with reusable refunds. International Journal of Information Security, 1(3): 175–188, November 2002. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        002-0015-z.

Khayati:2016:PPP

[KOSU16] Leyli Javid Khayati, Cengiz Orencik, Erkay Savas, and Berkant Ustaoğlu. A practical privacy-preserving targeted advertising scheme for IPTV users. International Journal of Information Security, 15(4): 335–360, August 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        015-0296-7.

Kontaxis:2012:MID

[KPM12] Georgios Kontaxis, Michalis Polychronakis, and Evangelos P. Markatos. Minimizing information disclosure to third parties in social login platforms. International Journal of Information Security, 11(5): 321–332, October 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0173-6.

Kaur:2023:IAL

[KSB23] Sarabjeet Kaur, Amanpreet Kaur Sandhu, and Abhinav Bhandari. Investigation of application layer DDoS attacks in legacy and software-defined networks: a comprehensive review. International Journal of Information Security, 22 (6):1949–1988, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00728-5.

Kundu:2010:DID

[KSM10] Amlan Kundu, Shamik Sural, and A. K. Majumdar. Database intrusion detection using sequence alignment. International Journal of Information Security, 9(3):179–191, June 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        010-0102-5.

Konstantinou:2007:EGS

[KSZ07] Elisavet Konstantinou, Yannis C. Stamatiou, and Christos Zaroliagis. Efficient generation of secure elliptic curves. International Journal of Information Security, 6(1):47– 63, January 2007. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        006-0009-3.

Kiraz:2016:EVA

[KU16] Mehmet Sabir Kiraz and Osmanbey Uzunkol. Efficient and verifiable algorithms for secure outsourcing of cryptographic computations. International Journal of Information Security, 15(5): 519–537, October 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        015-0308-7.

Kudo:2002:PPB

[Kud02] Michiharu Kudo. PBAC: Provision-based access control model. International Journal of Information Security, 1(2): 116–130, February 2002. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s102070100010.

Kusters:2005:DCP

[Küs05] Ralf Küsters. On the decidability of cryptographic protocols with open-ended data structures. International Journal of Information Security, 4(1– 2):49–70, February 2005. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
004-0050-z.

Kamm:2015:SFP

[KW15] Liina Kamm and Jan Willemson. Secure floating point arithmetic and private satellite collision analysis. International Journal of Information Security, 14(6):531–548, November 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0271-8;   http://
link.springer.com/content/
pdf/10.1007/s10207-014-
0271-8.pdf.

Kluczniak:2019:MDA

[KWCK19] Kamil Kluczniak, Jianfeng Wang, Xiaofeng Chen, and Miroslaw Kutylowski. Multi-device anonymous authentication. International Journal of Information Security, 18 (2):181–197, April 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
018-0406-4;   http://
link.springer.com/content/
pdf/10.1007/s10207-018-
0406-4.pdf.

Kuzuno:2024:MPE

[KY24] Hiroki Kuzuno and Toshihiro Yamauchi. Mitigation of privilege escalation attack using kernel data relocation mechanism. International Journal of Information Security, 23(5): 3351–3367, October 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00890-4.

Kuo:2018:DRA

[KYH18] Tsung-Min Kuo, Sung-Ming Yen, and Meng-Che Han. Dynamic reversed accumulator. International Journal of Information Security, 17 (2):183–191, April 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0360-6.

Landwehr:2001:CS

[Lan01] Carl E. Landwehr. Computer security. International Journal of Information Security, 1(1):3– 13, August 2001. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s102070100003.

Ligatti:2005:EAE

[LBW05] Jay Ligatti, Lujo Bauer, and David Walker. Edit automata: enforcement mechanisms for run-time security policies. International Journal of Information Security, 4(1– 2):2–16, February 2005. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
004-0046-8.

Liu:2010:EOO

[LBZ^{+}10] Joseph K. Liu, Joonsang Baek, Jianying Zhou, Yanjiang Yang, and Jun Wen Wong. Efficient online/ offline identity-based signature for wireless sensor network. International Journal of Information Security, 9(4):287–296, August 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
010-0109-y.

Laih:2004:GVR

[LC04] C. S. Laih and K. Y. Chen. Generating visible RSA public keys for PKI. International Journal of Information Security, 2(2):103–109, January 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
003-0030-8.

Lian:2014:PSC

[LCL14] Bin Lian, Gongliang Chen, and Jianhua Li. Provably secure e-cash system with practical and efficient complete tracing. International Journal of Information Security, 13(3):271–289, June 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0240-2.

Liu:2016:STP

[LCL16] Liang Liu, Xiaofeng Chen, and Wenjing Lou. Secure three-party computational protocols for triangle area. International Journal of Information Security, 15(1):1–13, February 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
015-0284-y.

Li:2014:AAM

[LCPD14] Fudong Li, Nathan Clarke, Maria Papadaki, and Paul Dowland. Active authentication for mobile devices utilising behaviour profiling. International Journal of Information Security, 13(3):229–244, June 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0209-6.

Luo:2018:ASI

[LcSCL^{+}18] Ying Luo, Sen ching S. Cheung, Riccardo Lazzeretti, Tommaso Pignata, and Mauro Barni. Anonymous subject identification and privacy information management in video surveillance. International Journal of Information Security, 17(3):261–278, June 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0380-2.

Lauer:2007:BOT

[LD07] Thomas W. Lauer and Xiaodong Deng. Building online trust through privacy practices. International Journal of Information Security, 6(5): 323–331, September 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0028-8.

Li:2017:WOO

[LD17] Lichun Li and Anwitaman Datta. Write-only oblivious RAM-based privacypreserved access of outsourced data. International Journal of Information Security, 16(1): 23–42, February 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0329-x;   http://
link.springer.com/content/
pdf/10.1007/s10207-016-
0329-x.pdf.

Leszczyna:2024:ABP

[Les24] Rafał Leszczyna. Activitybased payments: alternative (anonymous) online payment model. International Journal of Information Security, 23(3): 1741–1759, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00816-0.

Levin:2007:WSL

[Lev07] Avner Levin. Is workplace surveillance legal in Canada? International Journal of Information Security, 6(5): 313–321, September 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0026-x.

Li:2015:TMM

[LH15] Tao Li and Aiqun Hu.

Trusted mobile model based on DTE technology. International Journal of Information Security, 14 (5):457–469, October 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0266-5.

Luong:2023:BML

[LH23] Ngoc T. Luong and Doan Hoang. BAPRP: a machine learning approach to blackhole attacks prevention routing protocol in vehicular ad hoc networks. International Journal of Information Security, 22 (6):1547–1566, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.

com/article/10.1007/s10207-
023-00705-y.

Lagraa:2024:RGB

[LHS^{+}24] Sofiane Lagraa, Martin Husák, Hamida Seba, Satyanarayana Vuppala, Radu State, and Moussa Ouedraogo. A review on graph-based approaches for network security monitoring and botnet detection. International Journal of Information Security, 23(1):119–140, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00742-7.

Liu:2007:ICR

[LI07] Jinshan Liu and Valérie Issarny. An incentive compatible reputation mechanism for ubiquitous computing environments. International Journal of Information Security, 6(5): 297–311, September 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0029-7.

Lin:2015:RNR

[Lin15] Han-Yu Lin. RPCAE: a novel revocable proxy convertible authenticated encryption scheme. International Journal of Information Security, 14(5): 431–441, October 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0269-2.

Lee:2009:RTB

[LKH09] Sangho Lee, Jong Kim, and Sung Je Hong. Redistributing time-based rights between consumer devices for content sharing in DRM system. International Journal of Information Security, 8(4): 263–273, August 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
009-0082-5.

Liao:2014:POC

[LL14] Qi Liao and Zhen Li. Portfolio optimization of computer and mobile botnets. International Journal of Information Security, 13 (1):1–14, February 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0206-9.

Leontiadis:2021:SCR

[LL21] Iraklis Leontiadis and Ming Li. Secure and collusionresistant data aggregation from convertible tags. International Journal of Information Security, 20 (1):1–20, February 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        019-00485-4.

Lee:2022:SPM

[LL22] Junwon Lee and Heejo Lee. An SSH predictive model using machine learning with web proxy session logs. International Journal of Information Security, 21(2):311–322, April 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        021-00555-6.

Lanet:2018:WTM

[LLBL18] Jean-Louis Lanet, Hélène Le Bouder, Mohammed Benattou, and Axel Legay. When time meets test. International Journal of Information Security, 17 (4):395–409, August 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        017-0371-3.

Lyvas:2022:EIS

[LLG22] Christos Lyvas, Costas Lambrinoudakis, and Dimitris Geneiatakis. IntentAuth: Securing Android’s intentbased inter-process communication. International Journal of Information Security, 21(5): 973–982, October 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00592-9.

Lee:2021:BBM

[LLH21] Tian-Fu Lee, Hong-Ze Li, and Yi-Pei Hsieh. A blockchain-based medical data preservation scheme for telecare medical information systems. International Journal of Information Security, 20(4): 589–601, August 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00521-8.

Liu:2016:PCC

[LLW^{+}16] Weiran Liu, Jianwei Liu, Qianhong Wu, Bo Qin, and Yan Li. Practical chosen-ciphertext secure Hierarchical Identity-Based Broadcast Encryption. International Journal of Information Security, 15(1):35–50, February 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        015-0287-8.

Li:2009:DSA

[LLWY09] Jiangtao Li, Ninghui Li, XiaoFeng Wang, and Ting Yu. Denial of service attacks and defenses in decentralized trust management. International Journal of Information Security, 8(2):89–101, April 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        008-0068-8.

Li:2006:USS

[LM06] Ninghui Li and John C. Mitchell. Understanding SPKI/SDSI using firstorder logic. International Journal of Information Security, 5(1):48–64, January 2006. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        005-0073-0.

Li:2017:RRS

[LMD17] Lichun Li, Michael Militzer, and Anwitaman Datta. rPIR: ramp secret sharing-based communicationefficient private information retrieval. International Journal of Information Security, 16(6): 603–625, November 2017.

CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0347-8.

Lai:2017:EIB

[LMG17] Jianchang Lai, Yi Mu, and Fuchun Guo. Efficient identity-based online/ offline encryption and signcryption with short ciphertext. International Journal of Information Security, 16(3):299–311, June 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0320-6.

Lopez:2004:PDB

[LMMO04] Javier Lopez, Antonio Maña, Jose A. Montenegro, and Juan J. Ortega. PKI design based on the use of on-line certification authorities. International Journal of Information Security, 2(2):91–102, January 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
003-0027-3.

Lioy:2006:PPP

[LMMP06] Antonio Lioy, Marius Marian, Natalia Moltchanova, and Massimiliano Pala. PKI past, present and future. International Journal of Information Security, 5(1):18–29, January 2006. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
005-0077-9.

Lazouski:2017:SDU

[LMMS17] Aliaksandr Lazouski, Fabio Martinelli, Paolo Mori, and Andrea Saracino. Stateful data usage control for Android mobile devices. International Journal of Information Security, 16 (4):345–369, August 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0336-y.

Lyvas:2022:MME

[LNX22] Christos Lyvas, Christoforos Ntantogian, and Christos Xenakis. [m]allotROPism: a metamorphic engine for malicious software variation development. International Journal of Information Security, 21 (1):61–78, February 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00541-y.

Lopriore:2018:ARM

[Lop18] Lanfranco Lopriore. Access right management by extended password capabilities. International Journal of Information Security, 17 (5):603–612, October 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0390-0.

Lim:2011:IBC

[LP11] Hoon Wei Lim and Kenneth G. Paterson. Identitybased cryptography for grid security. International Journal of Information Security, 10(1): 15–32, February 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
010-0116-z.

Lin:2010:ECE

[LRB^{+}10] Dan Lin, Prathima Rao, Elisa Bertino, Ninghui Li, and Jorge Lobo. EXAM: a comprehensive environment for the analysis of access control policies. International Journal of Information Security, 9(4): 253–273, August 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

                                        com/article/10.1007/s10207-
                                        010-0106-1.

Lovrencic:2023:MCA

[LS23] Rudolf Lovrencić and Dejan Skvorc. Multi-cloud applications: data and code fragmentation for improved security. International Journal of Information Security, 22(3):713–721, June 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00658-8.

Lygerou:2022:DHI

[LSG22] Irini Lygerou, Shreyas Srinivasa, and Dimitris Gritzalis. A decentralized honeypot for IoT protocols based on Android devices. International Journal of Information Security, 21 (6):1211–1222, December 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00605-7. See correc-

tion [LSV^{+}23].

Lopez:2023:AAA

[LSR^{+}23] Christian López, Jesús Solano, Esteban Rivera, Lizzy Tengana, Johana Florez-Lozano, Alejandra Castelblanco, and Martín Ochoa. Adversarial attacks against mouse- and keyboard-based biometric authentication: blackbox versus domain-specific techniques. International Journal of Information Security, 22(6):1665–1685, December 2023. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00711-0.

Lygerou:2023:CDH

[LSV^{+}23] Irini Lygerou, Shreyas Srinivasa, Emmanouil Vasilomanolakis, George Stergiopoulos, and Dimitris Gritzalis. Correction to: A decentralized honeypot for IoT Protocols based on Android devices. International Journal of Information Security, 22 (1):303, February 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-

022-00628-0. See [LSG22].

Lee:2014:AGU

[LSWW14] Ming-Feng Lee, Nigel P.

Smart, Bogdan Warinschi, and Gaven J. Watson. Anonymity guarantees of the UMTS/LTE authentication and connection protocol. International Journal of Information Security, 13(6): 513–527, November 2014.

CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        014-0231-3.

Levshun:2023:TMA

[LTC23] Dmitry Levshun, Olga Tushkanova, and Andrey Chechulin. Two-model active learning approach for inappropriate information classification in social networks. International Journal of Information Security, 22(6):1921– 1936, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00726-7.

Lu:2009:RKR

[Lu09] Jiqiang Lu. Related-key rectangle attack on 36 rounds of the XTEA block cipher. International Journal of Information Security, 8(1):1–11, February 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        008-0059-9.

Lin:2010:MTE

[LV10] Ching Lin and Vijay Varadharajan. MobileTrust: a trust enhanced security architecture for mobile agent systems. International Journal of Information Security, 9(3): 153–178, June 2010. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
009-0098-x.

Laszka:2018:GTA

[LVK18] Aron Laszka, Yevgeniy Vorobeychik, and Xenofon Koutsoukos. A game-theoretic approach for integrity assurance in resource-bounded systems. International Journal of Information Security, 17 (2):221–242, April 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0364-2.

Li:2021:TBB

[LWA21] Wenjuan Li, Yu Wang, and Man Ho Au. Toward a blockchain-based framework for challengebased collaborative intrusion detection. International Journal of Information Security, 20 (2):127–139, April 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
020-00488-6.

Liu:2021:NEM

[LWL^{+}21] Pengfei Liu, Weiping Wang, Xi Luo, Haodong Wang, and Chushu Liu. NSDroid: efficient multi-classification of Android malware using neighborhood signature in local function call graphs. International Journal of Information Security, 20(1):59–71, February 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        020-00489-5.

Li:2023:BEC

[LWL23] Wenjuan Li, Yu Wang, and Jin Li. A blockchainenabled collaborative intrusion detection framework for SDN-assisted cyber-physical systems. International Journal of Information Security, 22(5): 1219–1230, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00687-x.

Li:2018:IVC

[LZQ^{+}18] Yuxi Li, Fucai Zhou, Yuhai Qin, Muqing Lin, and Zifeng Xu. Integrityverifiable conjunctive keyword searchable encryption in cloud storage. International Journal of Information Security, 17(5): 549–568, October 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0394-9.

Mandal:2021:ATC

[Man21] Mriganka Mandal. Anonymity in traceable cloud data broadcast system with simultaneous individual messaging. International Journal of Information Security, 20(3):405–430, June 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00512-9.

Manasrah:2020:KPD

[MARK20] Ahmed M. Manasrah, Areej R. AL-Rabadi, and Najib A. Kofahi. Key pre-distribution approach using block LU decomposition in wireless sensor network. International Journal of Information Security, 19(5): 579–596, October 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00477-4.

Masmoudi:2024:TAP

[MAZS24] Mariam Masmoudi, Ikram Amous, Corinne Amel Zayani, and Florence Sèdes. Trust attack prevention based on spark-blockchain in social IoT: a survey. International Journal of Information Security, 23(5): 3179–3198, October 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00885-1.

Malhi:2016:PPA

[MB16] Avleen Malhi and Shalini Batra. Privacy-preserving authentication framework using Bloom filter for secure vehicular communications. International Journal of Information Security, 15(4):433–453, August 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
015-0299-4.

Mohammadi:2023:NHH

[MB23] Shahriar Mohammadi and Mehdi Babagoli. A novel hybrid hunger games algorithm for intrusion detection systems based on nonlinear regression modeling. International Journal of Information Security, 22(5):1177–1195, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00684-0.

Manulis:2021:CSN

[MBH^{+}21] M. Manulis, C. P. Bridges, R. Harrison, V. Sekar, and A. Davis. Cyber security in new space. International Journal of Information Security, 20(3):287–311, June 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00503-w;   http://
link.springer.com/content/
pdf/10.1007/s10207-020-
00503-w.pdf.

Mehrnezhad:2017:PCI

[MBHT17] Maryam Mehrnezhad, Abbas Ghaemi Bafghi, Ahad Harati, and Ehsan Toreini. PiSHi: click the images and I tell if you are a human. International Journal of Information Security, 16(2):133–149, April 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
015-0311-z;   http://
link.springer.com/content/
pdf/10.1007/s10207-015-
0311-z.pdf.

Martinez-Balleste:2018:DIS

[MBRPS18] Antoni Martínez-Ballesté, Hatem Rashwan, Domenec Puig, and Agusti Solanas. Design and implementation of a secure and trustworthy platform for privacyaware video surveillance. International Journal of Information Security, 17 (3):279–290, June 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0370-4.

Marconi:2011:CPE

[MCD11] Luciana Marconi, Mauro Conti, and Roberto Di Pietro. CASSANDRA: a probabilistic, efficient, and privacy-preserving solution to compute set intersection. International Journal of Information Security, 10 (5):301–319, October 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
011-0133-6.

McHugh:2001:IID

[McH01]  John McHugh.  Intru-

sion and intrusion detection. International Journal of Information Security, 1(1):14–35, August 2001. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s102070100001.

Madabhushi:2023:SAD

[MD23a] Srinidhi Madabhushi and Rinku Dewri. A survey of anomaly detection methods for power grids. International Journal of Information Security, 22 (6):1799–1832, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00720-z.

Mishra:2023:SSU

[MD23b] Kailash Chandra Mishra and Subrata Dutta. A simple and secure user authentication scheme using Map Street View with usability analysis based on ISO/IEC 25022. International Journal of Information Security, 22(2):403–415, April 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00636-0.

Menezes:2023:BSC

[MdAN23] Leonardo Dias Menezes, Luciano Vieira de Araújo, and Marislei Nishijima. Blockchain and smart contract architecture for notaries services under civil law: a Brazilian experience. International Journal of Information Security, 22(4):869–880, August 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00673-3.

Melo:2022:IAI

[MdMF22] Roberto Vasconcelos Melo, Douglas D. J. de Macedo, and Mauricio Martinuzzi Fiorenza. ISM-AC: an immune security model based on alert correlation and software-defined networking. International Journal of Information Security, 21(2):191–205, April 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00550-x.

Martina:2015:ATM

[MdSC^{+}15] Jean Everson Martina, Eduardo dos Santos, Marcelo Carlomagno Carlos, Geraint Price, and Ricardo Felipe Custódio. An adaptive threat model for security ceremonies. International Journal of Information Security, 14 (2):103–121, April 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0253-x.

Mohamed:2023:DSB

[ME23] Safa Mohamed and Ridha Ejbali. Deep SARSAbased reinforcement learning approach for anomaly network intrusion detection system. International Journal of Information Security, 22(1): 235–247, February 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00634-2.

Munoz:2004:CRS

[MFES04] Jose L. Muñoz, Jordi Forne, Oscar Esparza, and Miguel Soriano. Certificate revocation system implementation based on the Merkle hash tree. International Journal of Information Security, 2(2): 110–124, January 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
003-0026-4.

Makhlouf:2019:SAS

[MG19] Amel Meddeb Makhlouf and Mohsen Guizani. SE-AOMDV: secure and efficient AOMDV routing protocol for vehicular communications. International Journal of Information Security, 18(5):

665–676, October 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        019-00436-z.

Mittal:2023:NTL

[MG23] Shweta Mittal and Mohona Ghosh. A novel two-level secure access control approach for blockchain platform in healthcare. International Journal of Information Security, 22(4): 799–817, August 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00664-4.

Mawla:2024:MAM

[MGAS24] Tanjila Mawla, Maanak Gupta, Safwa Ameer, and Ravi Sandhu. The ACAC_{D} model for mutable activity control and chain of dependencies in smart and connected systems. International Journal of Information Security, 23(5): 3283–3310, October 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        024-00881-5.

Marco-Gisbert:2019:SES

[MGRR19] Héctor Marco-Gisbert and Ismael Ripoll-Ripoll. SSPFA: effective stack smashing protection for Android OS. International Journal of Information Security, 18 (4):519–532, August 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
018-00425-8;   http://
link.springer.com/content/
pdf/10.1007/s10207-018-
00425-8.pdf.

Malatras:2017:EUI

[MGV17] Apostolos Malatras, Dimitris Geneiatakis, and Ioannis Vakalis. On the efficiency of user identification: a system-based approach. International Journal of Information Security, 16(6):653–671, November 2017. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0340-2;   http://
link.springer.com/content/
pdf/10.1007/s10207-016-
0340-2.pdf.

Mimura:2022:ANT

[MI22] Mamoru Mimura and Ryo Ito. Applying NLP techniques to malware detection in a practical environment. International Journal of Information Security, 21(2):279–291, April 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        021-00553-8.

Mighan:2021:NSI

[MK21] Soosan Naderi Mighan and Mohsen Kahani. A novel scalable intrusion detection system based on deep learning. International Journal of Information Security, 20(3):387–403, June 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        020-00508-5.

Mazumder:2023:CCB

[MKS^{+}23] Md. Mumtahin Habib Ullah Mazumder, Md. Eusha Kadir, Sadia Sharmin, Md. Shariful Islam, and Muhammad Mahbub Alam. cFEM: a cluster based feature extraction method for network intrusion detection. International Journal of Information Security, 22(5):1355–1369, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00694-y.

Mousazadeh:2014:RGA

[ML14] Mousa Mousazadeh and Behrouz Tork Ladani. Randomized gossip algorithms under attack. International Journal of Information Security, 13(4): 391–402, August 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0221-x.

Mann:2017:TFA

[ML17] Christopher Mann and Daniel Loebenberger. Twofactor authentication for the Bitcoin protocol. International Journal of Information Security, 16 (2):213–226, April 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0325-1.

Mohammadi:2023:HML

[MLC23] Reza Mohammadi, Chhagan Lal, and Mauro Conti. HTTPScout: a machine learning based countermeasure for HTTP flood attacks in SDN. International Journal of Information Security, 22 (2):367–379, April 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00641-3.

Medina-Lopez:2021:SAD

[MLCQ21] Cristóbal Medina-López, L. G. Casado, and Yuansong Qiao. An SDN approach to detect targeted attacks in P2P fully connected overlays. International Journal of Information Security, 20 (2):245–255, April 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        020-00499-3.

Miao:2016:MDU

[MLCS16] Qiguang Miao, Jiachen Liu, Ying Cao, and Jianfeng Song. Malware detection using bilayer behavior abstraction and improved one-class support vector machines. International Journal of Information Security, 15(4): 361–379, August 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        015-0297-6.

Mesbah:2019:REJ

[MLM19] Abdelhak Mesbah, Jean-Louis Lanet, and Mohamed Mezghiche. Reverse engineering Java Card and vulnerability exploitation: a shortcut to ROM. International Journal of Information Security, 18(1): 85–100, February 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
018-0401-9.

Mana:2004:FSE

[MLO^{+}04] Antonio Maña, Javier Lopez, Juan J. Ortega, Ernesto Pimentel, and Jose M. Troya. A framework for secure execution of software. International Journal of Information Security, 3(2): 99–112, November 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
004-0048-6.

Meng:2020:ECB

[MLYL20] Weizhi Meng, Wenjuan Li, Laurence T. Yang, and Peng Li. Enhancing challenge-based collaborative intrusion detection networks against insider attacks using blockchain. International Journal of Information Security, 19 (3):279–290, June 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00462-x.

Micale:2024:CAB

[MMF^{+}24] Davide Micale, Ilaria Matteucci, Florian Fenzl, Roland Rieke, and Giuseppe Patanè. A context-aware on-board intrusion detection system for smart vehicles. International Journal of Information Security, 23(3):2203–2223, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        024-00821-3.

Mateu:2016:HAV

[MMS16] Víctor Mateu, Josep M. Miret, and Francesc Sebé. A hybrid approach to vector-based homomorphic tallying remote voting. International Journal of Information Security, 15 (2):211–221, April 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        015-0279-8.

Melki:2020:LMF

[MNC20] Reem Melki, Hassan N. Noura, and Ali Chehab. Lightweight multi-factor mutual authentication protocol for IoT devices. International Journal of Information Security, 19(6): 679–694, December 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00484-5.

Moen:2024:SBA

[MOA^{+}24a] Ivar Moen, Aybars Oruc, Ahmed Amro, Vasileios Gkioulos, and Georgios Kavallieratos. Surveybased analysis of cybersecurity awareness of Turkish seafarers. International Journal of Information Security, 23(5): 3153–3178, October 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00884-2.

Mughaid:2024:ICA

[MOA^{+}24b] Ala Mughaid, Ibrahim Obeidat, Laith Abualigah, Shadi Alzubi, Mohammad Sh. Daoud, and Hazem Migdady. Intelligent cybersecurity approach for data protection in cloud computing based Internet of Things. International Journal of Information Security, 23 (3):2123–2137, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00832-0.

Mohammadi:2023:CBO

[Moh23]  Reza Mohammadi.  A

comprehensive blockchainoriented secure framework for SDN/Fog-based IoUT. International Journal of Information Security, 22 (5):1163–1175, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00683-1.

Martina:2015:VMB

[MP15] Jean Everson Martina and Lawrence Charles Paulson. Verifying multicastbased security protocols using the inductive method. International Journal of Information Security, 14 (2):187–204, April 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        014-0251-z.

Mennink:2016:EPH

[MP16] Bart Mennink and Bart Preneel. Efficient parallelizable hashing using small non-compressing primitives. International Journal of Information Security, 15(3):285–300, June 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.

com/article/10.1007/s10207-
015-0288-7.

Munonye:2022:MLA

[MP22] Kindson Munonye and Martinek Péter. Machine learning approach to vulnerability detection in OAuth 2.0 authentication and authorization flow. International Journal of Information Security, 21(2):223–237, April 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00551-w.

Murvay:2021:SCA

[MPG21] Pal-Stefan Murvay, Lucian Popa, and Bogdan Groza. Securing the controller area network with covert voltage channels. International Journal of Information Security, 20(6): 817–831, December 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00532-5.

Mavrogiannopoulos:2014:TSK

[MPP14] Nikos Mavrogiannopoulos, Andreas Pashalidis, and Bart Preneel. Toward a secure Kerberos key exchange with smart cards. International Journal of Information Security, 13 (3):217–228, June 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0213-x.

Menges:2021:DDI

[MPP21] Florian Menges, Benedikt Putz, and Günther Pernul. DEALER: decentralized incentives for threat intelligence reporting and exchange. International Journal of Information Security, 20(5): 741–761, October 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00528-1;   http://
link.springer.com/content/
pdf/10.1007/s10207-020-
00528-1.pdf.

MacKenzie:2010:PAK

[MPS10] Philip MacKenzie, Sarvar Patel, and Ram Swaminathan. Passwordauthenticated key exchange based on RSA. International Journal of Information Security, 9(6): 387–410, December 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
010-0120-3.

Manulis:2014:PAI

[MPS14] Mark Manulis, Bertram Poettering, and Douglas Stebila. Plaintext awareness in identity-based key encapsulation. International Journal of Information Security, 13(1): 25–49, February 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0218-5.

MacKenzie:2003:NCD

[MR03] Philip MacKenzie and Michael K. Reiter. Networked cryptographic devices resilient to capture. International Journal of Information Security, 2 (1):1–20, November 2003. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
003-0022-8.

MacKenzie:2004:TPG

[MR04] Philip MacKenzie and Michael K. Reiter. Twoparty generation of DSA signatures. International Journal of Information Security, 2(3–4): 218–239, August 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

                                        com/article/10.1007/s10207-
                                        004-0041-0.

Maddu:2024:NID

[MR24] Mamatha Maddu and Yamarthi Narasimha Rao. Network intrusion detection and mitigation in SDN using deep learning models. International Journal of Information Security, 23(2):849–862, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00771-2.

Mangipudi:2024:PPM

[MRCK24] Easwar Vivek Mangipudi, Krutarth Rao, Jeremy Clark, and Aniket Kate. Pepal: Penalizing multimedia breaches and partial leakages. International Journal of Information Security, 23(1): 447–465, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00744-5.

Molla:2023:ESS

[MRG23] Eirini Molla, Panagiotis Rizomiliotis, and Stefanos Gritzalis. Efficient searchable symmetric encryption supporting range queries. International Journal of Information Security, 22 (4):785–798, August 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00667-1.

Monrose:2002:PHB

[MRW02] Fabian Monrose, Michael K.

Reiter, and Susanne Wetzel. Password hardening based on keystroke dynamics. International Journal of Information Security, 1(2):69–83, February 2002. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s102070100006.

Mashatan:2009:ITC

[MS09] Atefeh Mashatan and Douglas R. Stinson. Interactive two-channel message authentication based on Interactive–Collision Resistant hash functions. International Journal of Information Security, 8(1): 49–60, February 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
008-0063-0.

Miret:2011:CAH

[MS11] Josep M. Miret and Francesc Sebé. Cryptanalysis of an ad-hoc cryptosystem for mix-based e-voting robust against relation attacks. International Journal of Information Security, 10 (6):387–389, November 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
011-0145-2.

Mizuki:2014:FCB

[MS14] Takaaki Mizuki and Hiroki Shizuya. A formalization of card-based cryptographic protocols via abstract machine. International Journal of Information Security, 13(1):15–23, February 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0219-4.

Mustafa:2015:UIA

[MS15] Tanveer Mustafa and Karsten Sohr. Understanding the implemented access control policy of Android system services with slicing and extended static checking. International Journal of Information Security, 14(4):347–366, August 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0260-y.

Morales-Sandoval:2018:PBC

[MSGCDPSS18] Miguel Morales-Sandoval, Jose Luis Gonzalez-Compean, Arturo Diaz-Perez, and Victor J. Sosa-Sosa. A pairing-based cryptographic approach for data security in the cloud. International Journal of Information Security, 17(4): 441–461, August 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0375-z.

Manulis:2016:SMP

[MSKD16] Mark Manulis, Douglas Stebila, Franziskus Kiefer, and Nick Denham. Secure modular password authentication for the web using channel bindings. International Journal of Information Security, 15(6): 597–620, November 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0348-7;   http://
link.springer.com/content/
pdf/10.1007/s10207-016-
0348-7.pdf.

Mitropoulos:2020:PAC

[MSKS20] Dimitris Mitropoulos, Thodoris Sotiropoulos, Nikos Koutsovasilis, and Diomidis Spinellis. PDGuard: an architecture for the control and secure processing of personal data. International Journal of Information Security, 19(4):479–498, August 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00468-5.

Mizuki:2002:CCF

[MSN02] Takaaki Mizuki, Hiroki Shizuya, and Takao Nishizeki. A complete characterization of a family of key exchange protocols. International Journal of Information Security, 1(2): 131–142, February 2002. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s102070100011.

Marmol:2013:PEA

[MSP^{+}13] Félix Gómez Mármol, Christoph Sorge, Ronald Petrlic, Osman Ugus, Dirk Westhoff, and Gregorio Martínez Pérez. Privacy-enhanced architecture for smart metering. International Journal of Information Security, 12(2):67–82, April 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0181-6.

Martinez:2024:MCP

[MSSO^{+}24] Ferney Martínez, Luis Enrique Sànchez, Antonio Santos-Olmo, David G. Rosado, and Eduardo Fernàndez-Medina. Maritime cybersecurity: protecting digital seas. International Journal of Information Security, 23(2): 1429–1457, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00800-0.

Mvah:2024:GGT

[MTD^{+}24] Fabrice Mvah, Vianney Kengne Tchendji, Clémentin Tayou Djamegni, Ahmed H. Anwar, Deepak K. Tosh, and Charles Kamhoua. GaTeBaSep: game theorybased security protocol against ARP spoofing attacks in software-defined networks. International Journal of Information Security, 23(1):373–387, February 2024. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00749-0.

Mehrnezhad:2018:SPM

[MTSH18] Maryam Mehrnezhad, Ehsan Toreini, Siamak F. Shahandashti, and Feng Hao.

Stealing PINs via mobile sensors: actual risk versus user perception. International Journal of Information Security, 17 (3):291–313, June 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0369-x;   http://
link.springer.com/content/
pdf/10.1007/s10207-017-
0369-x.pdf.

McEvoy:2014:ANT

[MTW^{+}14] Robert P. McEvoy, Michael Tunstall, Claire Whelan, Colin C. Murphy, and William P. Marnane. Allor-Nothing Transforms as a countermeasure to differential side-channel analysis. International Journal of Information Security, 13(3):291–304, June 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0212-y.

Murakami:2018:OAA

[MU18] Keisuke Murakami and Takeaki Uno. Optimization algorithm for kanonymization of datasets with low information loss. International Journal of Information Security, 17 (6):631–644, November 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0392-y.

Miyahara:2021:ECB

[MUiH^{+}21] Daiki Miyahara, Itaru Ueda, Yu ichi Hayashi, Takaaki Mizuki, and Hideaki Sone. Evaluating cardbased protocols in terms of execution time. International Journal of Information Security, 20(5): 729–740, October 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00525-4;   http://
link.springer.com/content/
pdf/10.1007/s10207-020-
00525-4.pdf.

Madhushanie:2024:BFS

[MVA24a] Nadisha Madhushanie, Sugandima Vidanagamachchi, and Nalin Arachchilage. BAflag: a self-prevention mechanism of selfish mining attacks in blockchain technology. International Journal of Information Security, 23(4): 2783–2792, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.

                                        com/article/10.1007/s10207-
                                        024-00857-5.

Madhushanie:2024:SMA

[MVA24b] Nadisha Madhushanie, Sugandima Vidanagamachchi, and Nalin Arachchilage. Selfish mining attack in blockchain: a systematic literature review. International Journal of Information Security, 23(3): 2333–2351, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        024-00849-5.

Mayer:2006:OFA

[MWZ06] Alain Mayer, Avishai Wool, and Elisha Ziskind. Offline firewall analysis. International Journal of Information Security, 5 (3):125–144, July 2006. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        005-0074-z.

Mu:2014:RBD

[MYLZ14] Chengpo Mu, Meng Yu, Yingjiu Li, and Wanyu Zang. Risk balance defense approach against intrusions for network server. International Journal of Information Security, 13 (3):255–269, June 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0214-9.

Nunez:2014:BPP

[NA14] David Nuñez and Isaac Agudo. BlindIdM: a privacy-preserving approach for identity management as a service. International Journal of Information Security, 13 (2):199–215, April 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0230-4.

Navajas-Adan:2024:PDA

[NABGJS^{+}24] Joaquín Navajas-Adán, Eulàlia Badia-Gelabert, Laura Jiménez-Saurina, Ma. Jesús Marijuán-Martín, and Rafael Mayo-García. Perceptions and dilemmas around cyber-security in a Spanish research center after a cyber-attack. International Journal of Information Security, 23 (3):2315–2331, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00847-7.

Nali:2006:HTB

[NAM06] Deholo Nali, Carlisle Adams, and Ali Miri. Hierarchical time-based information release. International Journal of Information Security, 5(2):92–104, April 2006. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
006-0084-5.

Nasereddin:2024:NAD

[NAQ24] Mohammed Nasereddin and Raad Al-Qassas. A new approach for detecting process injection attacks using memory analysis. International Journal of Information Security, 23(3):2099–2121, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00836-w.

Ngamboe:2021:RAC

[NBA^{+}21] Mikaëla Ngamboé, Paul Berthier, Nader Ammari, Katia Dyrda, and José M. Fernandez. Risk assessment of cyber-attacks on telemetry-enabled cardiac implantable electronic devices (CIED). International Journal of Information Security, 20(4): 621–645, August 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

com/article/10.1007/s10207-
020-00522-7;   http://
link.springer.com/content/
pdf/10.1007/s10207-020-
00522-7.pdf.

Naim:2023:MWI

[NCBG23] Or Naim, Doron Cohen, and Irad Ben-Gal. Malicious website identification using design attribute learning. International Journal of Information Security, 22(5): 1207–1217, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00686-y.

Nandanwar:2024:TBI

[NK24] Himanshu Nandanwar and Rahul Katarya. TL-BILSTM IoT: transfer learning model for prediction of intrusion detection system in IoT environment. International Journal of Information Security, 23 (2):1251–1277, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00787-8.

Niktabe:2024:DCP

[NLS24] Sepideh Niktabe, Arash Habibi Lashkari, and Dilli Prasad Sharma. Detection, characterization, and profiling DoH malicious traffic using statistical pattern recognition. International Journal of Information Security, 23 (2):1293–1316, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00790-z.

Niebuhr:2012:SPS

[NMBB12] Robert Niebuhr, Mohammed Meziani, Stanislav Bulygin, and Johannes Buchmann. Selecting parameters for secure McEliece-based cryptosystems. International Journal of Information Security, 11(3):137–147, June 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
011-0153-2.

Narita:2024:SDP

[NMH^{+}24] Jun Narita, Takao Murakami, Hideitsu Hino, Masakatsu Nishigaki, and Tetsushi Ohki. Synthesizing differentially private location traces including co-locations. International Journal of Information Security, 23(1): 389–410, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.

com/article/10.1007/s10207-
023-00740-9.

Nguyen:2020:NGB

[NNL20] Huy-Trung Nguyen, Quoc-Dung Ngo, and Van-Hoang Le. A novel graph-based approach for IoT botnet detection. International Journal of Information Security, 19(5): 567–577, October 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00475-6.

Naor:2010:ETR

[NP10] Moni Naor and Benny Pinkas. Efficient trace and revoke schemes. International Journal of Information Security, 9(6): 411–424, December 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
010-0121-2.

Nappa:2015:MDI

[NRC15] Antonio Nappa, M. Zubair Rafique, and Juan Caballero. The MALICIA dataset: identification and analysis of drive-by download operations. International Journal of Information Security, 14 (1):15–33, February 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        014-0248-7.

Nguyen:2006:VSF

[NSNK06] Lan Nguyen, Rei Safavi-Naini, and Kaoru Kurosawa. Verifiable shuffles: a formal model and a paillierbased three-round construction with provable security. International Journal of Information Security, 5(4):241–255, October 2006. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        006-0004-8.

Narasimha:2009:PPR

[NST09] M. Narasimha, J. Solis, and G. Tsudik. Privacypreserving revocation checking. International Journal of Information Security, 8(1):61–75, February 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        008-0064-z.

Nayak:2020:SSE

[NT20] Sanjeet Kumar Nayak and Somanath Tripathy. SEDS: secure and efficient serveraided data deduplication scheme for cloud storage. International Journal of Information Security, 19 (2):229–240, April 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00455-w.

Nuida:2012:SCS

[Nui12] Koji Nuida. Short collusionsecure fingerprint codes against three pirates. International Journal of Information Security, 11 (2):85–102, April 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0155-8.

Nieto:2002:KRC

[NVB^{+}02] J. M. Gonzáles Nieto, K. Viswanathan, C. Boyd, A. Clark, and E. Dawson. Key recovery for the commercial environment. International Journal of Information Security, 1(3): 161–174, November 2002. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
002-0014-0.

Ogunseyi:2023:SRP

[OAJ23] Taiwo Blessing Ogunseyi, Cossi Blaise Avoussoukpo, and Yiqiang Jiang. A systematic review of privacy techniques in recommendation systems. International Journal of Information Security, 22 (6):1651–1664, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00710-1.

Olivero:2024:SMS

[OBDM^{+}24] Miguel Angel Olivero, Antonia Bertolino, Francisco José Dominguez-Mayo, María José Escalona, and Ilaria Matteucci. A systematic mapping study on security for systems of systems. International Journal of Information Security, 23(2):787– 817, April 2024. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00757-0.

Ozawa:2020:SIM

[OBH^{+}20] Seiichi Ozawa, Tao Ban, Naoki Hashimoto, Junji Nakazato, and Jumpei Shimamura. A study of IoT malware activities using association rule learning for darknet sensor data. International Journal of Information Security, 19 (1):83–92, February 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00439-w.

Oruc:2024:MCS

[OCG24] Aybars Oruc, Nabin Chowdhury, and Vasileios Gkioulos. A modular cyber security training programme for the maritime domain. International Journal of Information Security, 23 (2):1477–1512, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00799-4.

Ofte:2024:AOG

[Oft24] Håvard Jakobsen Ofte. The awareness of operators: a goal-directed task analysis in SOCs for critical infrastructure. International Journal of Information Security, 23(5): 3253–3282, October 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00872-6.

Ogata:2020:SAS

[OO20] Wakaha Ogata and Takaaki Otemori. Security analysis of secure kNN and ranked keyword search over encrypted data. International Journal of Information Security, 19(4): 419–425, August 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00461-y.

Otokwala:2024:OCF

[OPK24] Uneneibotejit Otokwala, Andrei Petrovski, and Harsha Kalutarage. Optimized common features selection and deep-autoencoder (OCFSDA) for lightweight intrusion detection in Internet of Things. International Journal of Information Security, 23(4): 2559–2581, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00855-7.

Orooji:2023:FAD

[ORK23] Marmar Orooji, Seyedeh Shaghayegh Rabbanian, and Gerald M. Knapp. Flexible adversary disclosure risk measure for identity and attribute disclosure attacks. International Journal of Information Security, 22 (3):631–645, June 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00654-y.

Orencik:2016:MKS

[OSSK16] Cengiz Orencik, Ayse Selcuk, Erkay Savas, and Murat Kantarcioğlu. Multikeyword search over encrypted data with scoring and search pattern obfuscation. International Journal of Information Security, 15(3):251–269, June 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
015-0294-9.

Okeya:2006:SAC

[OT06] Katsuyuki Okeya and Tsuyoshi Takagi. Security analysis of CRT-based cryptosystems. International Journal of Information Security, 5(3):177– 185, July 2006. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
005-0080-1.

Papanikolaou:2023:ANT

[PAI^{+}23] Alexandros Papanikolaou, Aggelos Alevizopoulos, Christos Ilioudis, Konstantinos Demertzis, and Konstantinos Rantos. An autoML network traffic analyzer for cyber threat detection. International Journal of Information Security, 22(5):

1511–1530, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00703-0.

Prantl:2023:RSG

[PBI^{+}23] Thomas Prantl, André Bauer, Lukas Iffländer, Christian Krupitzer, and Samuel Kounev. Recommendation of secure group communication schemes using multi-objective optimization. International Journal of Information Security, 22(5): 1291–1332, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00692-0.

Patsakis:2019:HID

[PC19] Constantinos Patsakis and Fran Casino. Hydras and IPFS: a decentralised playground for malware. International Journal of Information Security, 18(6): 787–799, December 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        019-00443-0.

Park:2022:EVC

[PCK22a] Jai Hyun Park, Jung Hee Cheon, and Dongwoo Kim. Efficient verifiable computation over quotient polynomial rings. International Journal of Information Security, 21(5): 953–971, October 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00590-x.

Potteiger:2022:DSR

[PCK22b] Bradley Potteiger, Feiyang Cai, and Xenofon Koutsoukos. Data space randomization for securing cyber-physical systems. International Journal of Information Security, 21 (3):597–610, June 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00568-1.

Pirbhulal:2024:ICB

[PCSA24] Sandeep Pirbhulal, Sabarathinam Chockalingam, Ankur Shukla, and Habtamu Abie. IoT cybersecurity in 5G and beyond: a systematic literature review. International Journal of Information Security, 23(4): 2827–2879, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.

                                        com/article/10.1007/s10207-
                                        024-00865-5.

Parish:2022:PGU

[PCT22] Zach Parish, Connor Cushing, and Julie Thorpe. Password guessers under a microscope: an in-depth analysis to inform deployments. International Journal of Information Security, 21(2):409–425, April 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        021-00560-9.

Parida:2021:PMC

[PD21] Trushna Parida and Suvrojit Das. PageDumper: a mechanism to collect page table manipulation information at run-time. International Journal of Information Security, 20 (4):603–619, August 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        020-00520-9.

Peng:2011:MOS

[PDB11] Kun Peng, Ed Dawson, and Feng Bao. Modification and optimisation of a shuffling scheme: stronger security, formal analysis and higher efficiency. International Journal of Information Security, 10 (1):33–47, February 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
010-0117-y.

Patil:2020:DVA

[PDM20] Rajendra Patil, Harsha Dudeja, and Chirag Modi. Designing in-VM-assisted lightweight agent-based malware detection framework for securing virtual machines in cloud computing. International Journal of Information Security, 19(2):147–162, April 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00447-w.

Peng:2011:GEC

[Pen11] Kun Peng. A general and efficient countermeasure to relation attacks in mix-based e-voting. International Journal of Information Security, 10 (1):49–60, February 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
010-0122-1.

Peng:2012:TDA

[Pen12]  Kun Peng.  Threshold

distributed access control with public verification: a practical application of PVSS. International Journal of Information Security, 11(1):23–31, February 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
011-0151-4.

Peng:2013:SAH

[Pen13]  Kun Peng.  A shuffle

to achieve high efficiency through pre-computation and batch verification. International Journal of Information Security, 12(4): 337–345, August 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0193-x.

Pereniguez-Garcia:2013:KPA

[PGMLK^{+}13] F. Pereñíguez-García, R. Marín-López, G. Kambourakis, A. Ruiz-Martínez, S. Gritzalis, and A. F. Skarmeta-Gómez. KAMU: providing advanced user privacy in Kerberos multidomain scenarios. International Journal of Information Security, 12(6): 505–525, November 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

com/article/10.1007/s10207-
013-0201-1.

Pericas-Gornals:2022:HPB

[PGMPPC22] Rosa Pericàs-Gornals, Macià Mut-Puigserver, and M. Magdalena Payeras-Capellà. Highly private blockchainbased management system for digital COVID-19 certificates. International Journal of Information Security, 21(5): 1069–1090, October 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00598-3.

Patruni:2024:PMP

[PH24a] Muralidhara Rao Patruni and Abdul Gaffar Humayun. PPAM-mIoMT: a privacy-preserving authentication with device verification for securing healthcare systems in 5G networks. International Journal of Information Security, 23(1):679–698, February 2024. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00762-3.

Patruni:2024:PCP

[PH24b] Muralidhara Rao Patruni and Abdul Gaffar Humayun. Publisher correction: PPAM-mIoMT: a privacy-preserving authentication with device verification for securing healthcare systems in 5G networks. International Journal of Information Security, 23(2):1019, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00775-y.

Prantl:2024:BHI

[PHE^{+}24] Thomas Prantl, Lukas Horn, Simon Engel, Lukas Iffländer, Lukas Beierlieb, Christian Krupitzer, André Bauer, Mansi Sakarvadia, Ian Foster, and Samuel Kounev. De Bello Homomorphico: Investigation of the extensibility of the OpenFHE library with basic mathematical functions by means of common approaches using the example of the CKKS cryptosystem. International Journal of Information Security, 23 (2):1149–1169, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00781-0.

Park:2022:EDP

[PHS22] Cheolhee Park, Dowon Hong, and Changho Seo. Evaluating differentially private decision tree model over model inversion attack. International Journal of Information Security, 21(3):1–14, June 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00564-5.

Pais:2010:NPR

[PJ10] Alwyn R. Pais and Shankar Joshi. A new probabilistic rekeying method for secure multicast groups. International Journal of Information Security, 9 (4):275–286, August 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
010-0108-z.

Pandey:2023:EMC

[PKHS23] Saurabh Pandey, Nitesh Kumar, Anand Handa, and Sandeep Kumar Shukla. Evading malware classifiers using RL agent with action-mask. International Journal of Information Security, 22(6):1743– 1763, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00715-w.

Plaga:2009:BKS

[Pla09] Rainer Plaga. Biometric keys: suitable use cases and achievable information content. International Journal of Information Security, 8(6): 447–454, December 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        009-0090-5.

Pitolli:2021:MAF

[PLA^{+}21] Gregorio Pitolli, Giuseppe Laurenza, Leonardo Aniello, Leonardo Querzoni, and Roberto Baldoni. MalFamAware: automatic family identification and malware classification through online clustering. International Journal of Information Security, 20(3):371–386, June 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        020-00509-4.

Pandey:2023:PSN

[PM23] Pankaj Pandey and Nishchol Mishra. Phish-Sight: a new approach for phishing detection using dominant colors on web pages and machine learning. International Journal of Information Security, 22(4): 881–891, August 2023.

CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00672-4.

Panda:2023:SAP

[PMDS23] Suryakanta Panda, Samrat Mondal, Ashok Kumar Das, and Willy Susilo. Secure access privilege delegation using attributebased encryption. International Journal of Information Security, 22(5): 1261–1276, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00690-2.

Perez-Mendez:2012:CLS

[PMPGMLLM12] Alejandro Pérez-Méndez, Fernando Pereñíguez-García, Rafael Marín-López, and Gabriel López-Millán. A cross-layer SSO solution for federating access to kerberized services in the eduroam/DAMe network. International Journal of Information Security, 11 (6):365–388, November 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0174-5.

Patel:2024:CED

[PMW24] N. D. Patel, B. M. Mehtre, and Rajeev Wankar. A computationally efficient dimensionality reduction and attack classification approach for network intrusion detection. International Journal of Information Security, 23(3): 2457–2487, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00792-x.

Pilz:2020:SAS

[PNG^{+}20] M. Pilz, F. Baghaei Naeini, K. Grammont, C. Smagghe, M. Davis, J.-C. Nebel, L. Al-Fagih, and E. Pfluegel. Security attacks on smart grid scheduling and their defences: a game-theoretic approach. International Journal of Information Security, 19(4):427–443, August 2020. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        019-00460-z;   http://
                                        link.springer.com/content/
                                        pdf/10.1007/s10207-019-
                                        00460-z.pdf.

Petrean:2024:RFE

[PP24a] Diana-Elena Petrean and Rodica Potolea. Random forest evaluation using multi-key homomorphic encryption and lookup tables. International Journal of Information Security, 23(3):2023–2041, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00823-1.

Pise:2024:PAV

[PP24b] Rohini G. Pise and Sonali Patil. Pioneering automated vulnerability detection for smart contracts in blockchain using KEVM: Guardian ADR-GAN. International Journal of Information Security, 23(3):1805–1819, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00817-z.

Patsakis:2024:CCS

[PPAHC24] Constantinos Patsakis, Eugenia Politou, Efthimios Alepis, and Julio Hernandez-Castro. Cashing out crypto: state of practice in ransom payments. International Journal of Information Security, 23 (2):699–712, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00766-z.

Pitropakis:2015:BRP

[PPL15] Nikolaos Pitropakis, Aggelos Pikrakis, and Costas Lambrinoudakis. Behaviour reflects personality: detecting co-residence attacks on Xen-based cloud environments. International Journal of Information Security, 14(4): 299–305, August 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        014-0255-8.

Phan:2013:ACB

[PPSS13] Duong-Hieu Phan, David Pointcheval, Siamak F. Shahandashti, and Mario Strefler. Adaptive CCA broadcast encryption with constant-size secret keys and ciphertexts. International Journal of Information Security, 12(4): 251–265, August 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        013-0190-0.

Priami:2004:PSI

[Pri04] Corrado Priami. Preface to the special issue on Security in Global Computing. International Journal of Information Security, 2(3–4):125, August 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
004-0034-z;   http://
link.springer.com/content/
pdf/10.1007/s10207-004-
0034-z.pdf.

Poettering:2017:DAP

[PS17] Bertram Poettering and Douglas Stebila. Doubleauthentication-preventing signatures. International Journal of Information Security, 16(1):1– 22, February 2017. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
015-0307-8.

Picazo-Sanchez:2023:GIP

[PSA23] Pablo Picazo-Sanchez and Magnus Almgren. Gridchain: an investigation of privacy for the future local distribution grid. International Journal of Information Security, 22 (1):29–46, February 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00622-6.

Perez-Sola:2019:DSP

[PSDSNAHJ19] Cristina Pérez-Solà, Sergi Delgado-Segura, Guillermo Navarro-Arribas, and Jordi Herrera-Joancomartí. Doublespending prevention for Bitcoin zero-confirmation transactions. International Journal of Information Security, 18(4): 451–463, August 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        018-0422-4.

Picazo-Sanchez:2022:CEC

[PSOMS22] Pablo Picazo-Sanchez, Lara Ortiz-Martin, and Andrei Sabelfeld. Are chrome extensions compliant with the spirit of least privilege? International Journal of Information Security, 21 (6):1283–1297, December 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00610-w.

Picazo-Sanchez:2020:AYP

[PSTS20] Pablo Picazo-Sanchez, Juan Tapiador, and Gerardo Schneider. After you, please: browser extensions order attacks and countermeasures. International Journal of Information Security, 19(6):

623–638, December 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00481-8;   http://
link.springer.com/content/
pdf/10.1007/s10207-019-
00481-8.pdf.

Prajisha:2022:EID

[PV22] C. Prajisha and A. R. Vasudevan. An efficient intrusion detection system for MQTT-IoT using enhanced chaotic salp swarm algorithm and LightGBM. International Journal of Information Security, 21 (6):1263–1282, December 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00611-9.

Paya:2024:SNS

[PVGG24] Antonio Paya, Vicente-García, and Alberto Gómez. Securesdp: a novel softwaredefined perimeter implementation for enhanced network security and scalability. International Journal of Information Security, 23(4):2793–2808, August 2024. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.

                                        com/article/10.1007/s10207-
                                        024-00863-7.

Qin:2009:SSS

[QDW09] Huawang Qin, Yuewei Dai, and Zhiquan Wang. A secret sharing scheme based on (t, n) threshold and adversary structure. International Journal of Information Security, 8(5): 379–385, October 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        009-0085-2.

Qin:2015:FAB

[QDW^{+}15] Bo Qin, Hua Deng, Qianhong Wu, Josep Domingo-Ferrer, David Naccache, and Yunya Zhou. Flexible attribute-based encryption applicable to secure e-healthcare records. International Journal of Information Security, 14(6): 499–511, November 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        014-0272-7.

Qian:2024:OID

[QLH^{+}24] Guangyu Qian, Jinyuan Li, Wei He, Wei Zhang, and You Cao. An online intrusion detection method for industrial control systems based on extended belief rule base. International Journal of Information Security, 23(4): 2491–2514, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00845-9.

Quinn:2009:AAE

[QLOW09] Karl Quinn, David Lewis, Declan O’Sullivan, and Vincent P. Wade. An analysis of accuracy experiments carried out over of a multi-faceted model of trust. International Journal of Information Security, 8(2):103–119, April 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
008-0069-7.

Qian:2015:PPP

[QLZH15] Huiling Qian, Jiguo Li, Yichen Zhang, and Jinguang Han. Privacypreserving personal health record using multi-authority attribute-based encryption with revocation. International Journal of Information Security, 14(6): 487–497, November 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

                                        com/article/10.1007/s10207-
                                        014-0270-9.

Quincozes:2024:SIA

[QQK^{+}24] Vagner E. Quincozes, Silvio E. Quincozes, Juliano F. Kazienko, Simone Gama, Omar Cheikhrouhou, and Anis Koubaa. A survey on IoT application layer protocols, security challenges, and the role of explainable AI in IoT (XAIoT). International Journal of Information Security, 23(3): 1975–2002, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        024-00828-w.

Rajasekaran:2024:FBA

[RA24] Arun Sekar Rajasekaran and M. Azees. A fogbased anonymous authentication scheme with location privacy for wireless body area network with FPGA implementation. International Journal of Information Security, 23(1):1–13, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00717-8.

Raad:2016:PSR

[RAC16] Elie Raad, Bechara Al Bouna, and Richard Chbeir. Preventing sensitive relationships disclosure for better social media preservation. International Journal of Information Security, 15(2):173–194, April 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
015-0278-9.

Rastogi:2023:RFN

[RB23] Shubhangi Rastogi and Divya Bansal. A review on fake news detection 3T’s: typology, time of detection, taxonomies. International Journal of Information Security, 22(1): 177–212, February 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00625-3.

Rodeh:2002:UAT

[RBD02] Ohad Rodeh, Kenneth P.

Birman, and Danny Dolev. Using AVL trees for faulttolerant group key management. International Journal of Information Security, 1(2):84–99, February 2002. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s102070100008.

Riecker:2015:LEC

[RBEH15] Michael Riecker, Sebastian Biedermann, Rachid El Bansarkhani, and Matthias Hollick. Lightweight energy consumption-based intrusion detection system for wireless sensor networks. International Journal of Information Security, 14 (2):155–167, April 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        014-0241-1.

Ryu:2024:DAA

[RC24] Gwonsang Ryu and Daeseon Choi. Detection of adversarial attacks based on differences in image entropy. International Journal of Information Security, 23(1): 299–314, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00735-6.

Rekeraho:2024:CCI

[RCC^{+}24] Alexandre Rekeraho, Daniel Tudor Cotfas, Petru Adrian Cotfas, Titus Constantin Bălan, Emmanuel Tuyishime, and Rebecca Acheampong. Cybersecurity challenges in IoT-based smart renewable energy. International Journal of Information Security, 23(1): 101–117, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00732-9.

Rao:2016:EAB

[RD16] Y. Sreenivasa Rao and Ratna Dutta. Efficient attribute-based signature and signcryption realizing expressive access structures. International Journal of Information Security, 15(1): 81–109, February 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
015-0289-6.

Rial:2018:PPS

[RDK18] Alfredo Rial, George Danezis, and Markulf Kohlweiss. Privacy-preserving smart metering revisited. International Journal of Information Security, 17 (1):1–31, February 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0355-8.

Rao:2024:CDM

[RE24] Dimmiti Srinivasa Rao and Ajith Jubilson Emerson. Cyberattack defense mechanism using deep learning techniques in softwaredefined networks. International Journal of Information Security, 23(2): 1279–1291, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00785-w.

Rizomiliotis:2013:SAP

[RG13] Panagiotis Rizomiliotis and Stefanos Gritzalis. On the security of AUTH, a provably secure authentication protocol based on the subspace LPN problem. International Journal of Information Security, 12(2):151–154, April 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0188-z.

Rowe:2016:MPS

[RGL16] Paul D. Rowe, Joshua D. Guttman, and Moses D. Liskov. Measuring protocol strength with security goals. International Journal of Information Security, 15(6): 575–596, November 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0319-z.

Ruiz-Heras:2017:AAB

[RHGTSC17] A. Ruiz-Heras, P. García-Teodoro, and L. Sánchez-Casado. ADroid: anomalybased detection of malicious events in Android platforms. International Journal of Information Security, 16(4): 371–384, August 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0333-1.

Ragab-Hassen:2017:KMS

[RHL17] Hani Ragab-Hassen and Esma Lounes. A key management scheme evaluation using Markov processes. International Journal of Information Security, 16(3):271–280, June 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0323-3;   http://
link.springer.com/content/
pdf/10.1007/s10207-016-
0323-3.pdf.

Ribeiro:2018:SRH

[RLEM18] Carlos Ribeiro, Herbert Leitold, Simon Esposito, and David Mitzam.

STORK: a real, heterogeneous, large-scale eID management system. International Journal of Information Security, 17(5): 569–585, October 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0385-x.

Reaves:2012:OVT

[RM12] Bradley Reaves and Thomas Morris. An open virtual testbed for industrial control system security research. International Journal of Information Security, 11(4):215–229, August 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0164-7.

Rebollo-Monedero:2013:MPA

[RMPADF13] David Rebollo-Monedero, Javier Parra-Arnau, Claudia Diaz, and Jordi Forné. On the measurement of privacy as an attacker’s estimation error. International Journal of Information Security, 12 (2):129–149, April 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0182-5.

Ros-Martin:2019:SND

[RMSCR19] Miguel Ros-Martín, Julián Salas, and Jordi Casas-Roma. Scalable nondeterministic clusteringbased k-anonymization for rich networks. International Journal of Information Security, 18 (2):219–238, April 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
018-0409-1.

Roelse:2011:DST

[Roe11a] Peter Roelse. Dynamic subtree tracing and its application in pay-TV systems. International Journal of Information Security, 10(3):173–187, June 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-

011-0126-5. See erratum [Roe11b].

Roelse:2011:EDS

[Roe11b] Peter Roelse. Erratum to: Dynamic subtree tracing and its application in pay-TV systems. International Journal of Information Security, 10 (6):391, November 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        011-0141-6;   http://
                                        link.springer.com/content/
                                        pdf/10.1007/s10207-011-

0141-6.pdf. See [Roe11a].

Rovida:2023:FAH

[Rov23]  Lorenzo Rovida.  Fast

but approximate homomorphic k-means based on masking technique. International Journal of Information Security, 22 (6):1605–1619, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00708-9.

Rezvani:2019:AXP

[RRI^{+}19] Mohsen Rezvani, David Rajaratnam, Aleksandar Ignjatovic, Maurice Pagnucco, and Sanjay Jha. Analyzing XACML policies using answer set programming. International Journal of Information Security, 18(4):465–479, August 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        018-0421-5.

Roy:2018:DSC

[RS18] Sangita Roy and Ashok Singh Sairam. Distributed star coloring of network for IP traceback. International Journal of Information Security, 17(3):315–326, June 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0366-0.

Rastegari:2019:CDV

[RSD19] Parvin Rastegari, Willy Susilo, and Mohammad Dakhilalian. Certificateless designated verifier signature revisited: achieving a concrete scheme in the standard model. International Journal of Information Security, 18(5): 619–635, October 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00430-5.

Rabie:2024:CFP

[RSH^{+}24a] Osama Bassam J. Rabie, Shitharth Selvarajan, Tawfiq Hasanin, Gouse Baig Mohammed, Abddulrhman M. Alshareef, and Mueen Uddin. Correction: A full privacy-preserving distributed batch-based certificateless aggregate signature authentication scheme for healthcare wearable wireless medical sensor networks (HWMSNs). International Journal of Information Security, 23(3): 2489, June 2024. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-

023-00798-5. See [RSH^{+}24b].

Rabie:2024:FPP

[RSH^{+}24b] Osama Bassam J. Rabie, Shitharth Selvarajan, Tawfiq Hasanin, Gouse Baig Mohammed, Abddulrhman M. Alshareef, and Mueen Uddin. A full privacypreserving distributed batchbased certificate-less aggregate signature authentication scheme for healthcare wearable wireless medical sensor networks (HWM-SNs). International Journal of Information Security, 23(1):51–80, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00748-1. See correc-

tion [RSH^{+}24a].

Rahal:2024:BBM

[RSK^{+}24] Hakima Rym Rahal, Sihem Slatnia, Okba Kazar, Ezedin Barka, and Saad Harous. Blockchain-based multi-diagnosis deep learning application for various diseases classification. International Journal of Information Security, 23 (1):15–30, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00733-8.

Resende:2019:BMI

[RSMA19] João S. Resende, Patrícia R.

Sousa, Rolando Martins, and Luís Antunes. Breaking MPC implementations through compression. International Journal of Information Security, 18(4): 505–518, August 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
018-0424-2.

Rasori:2024:UAF

[RSMT24] Marco Rasori, Andrea Saracino, Paolo Mori, and Marco Tiloca. Using the ACE framework to enforce access and usage control with notifications of revoked access rights. International Journal of Information Security, 23(5): 3109–3133, October 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00877-1.

Rashwan:2016:UTP

[RSPMB16] Hatem A. Rashwan, Agusti Solanas, Domènec Puig, and Antoni Martínez-Ballesté. Understanding trust in privacy-aware video surveillance systems. International Journal of Information Security, 15 (3):225–234, June 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
015-0286-9.

Rajic:2023:EWA

[RSV23] Branislav Rajić, Zarko Stanisavljević, and Pavle Vuletić. Early web application attack detection using network traffic analysis. International Journal of Information Security, 22(1):77–91, February 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00627-1.

Rathore:2023:RDL

[RT23] Nemi Chandra Rathore and Somanath Tripathy. Restricting data-leakage using fine-grained access control on OSN objects. International Journal of Information Security, 22 (1):93–106, February 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00629-z.

Russell:2004:TBR

[Rus04]  Selwyn Russell.  The-

ory and benefits of recursive certificate structures. International Journal of Information Security, 2(2):78–90, January 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
003-0028-2.

Ruan:2003:FGB

[RV03] Chun Ruan and Vijay Varadharajan. A formal graph based framework for supporting authorization delegations and conflict resolutions. International Journal of Information Security, 1(4):211–222, July 2003. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
003-0018-4.

Riesco:2019:LCT

[RV19] R. Riesco and V. A. Villagrá. Leveraging cyber threat intelligence for a dynamic risk framework. International Journal of Information Security, 18(6): 715–739, December 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

com/article/10.1007/s10207-
019-00433-2.

Sharma:2024:CRP

[SA24] Yash Sharma and Anshul Arora. A comprehensive review on permissions-based Android malware detection. International Journal of Information Security, 23(3):1877–1912, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00822-2.

Saifuzzaman:2022:SLR

[SAC22] Munshi Saifuzzaman, Tajkia Nuri Ananna, and Farida Chowdhury. A systematic literature review on wearable health data publishing under differential privacy. International Journal of Information Security, 21(4):847–872, August 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00576-1.

Syrmakesis:2022:CRA

[SAH22] Andrew D. Syrmakesis, Cristina Alcaraz, and Nikos D. Hatziargyriou. Classifying resilience approaches for protecting smart grids against cyber threats. International Journal of Information Security, 21(5): 1189–1210, October 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00594-7.

Shaikh:2017:DCM

[SAL17] Riaz Ahmed Shaikh, Kamel Adi, and Luigi Logrippo. A data classification method for inconsistency and incompleteness detection in access control policy sets. International Journal of Information Security, 16 (1):91–113, February 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0317-1.

Saad:2024:NDP

[SAM^{+}24] Redhwan M. A. Saad, Mohammed Anbar, Selvakumar Manickam, Samir I. Shaheen, and Iznan H. Hasbullah. Neighbor discovery protocol anomalybased detection system using neural network algorithm. International Journal of Information Security, 23(3):1787–1803, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00815-1.

Sourour:2009:ESD

[SAT09] Meharouech Sourour, Bouhoula Adel, and Abbes Tarek. Ensuring security in depth based on heterogeneous network security technologies. International Journal of Information Security, 8(4):233–246, August 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        009-0077-2.

Sattath:2020:IQB

[Sat20] Or Sattath. On the insecurity of quantum Bitcoin mining. International Journal of Information Security, 19(3):291–302, June 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        020-00493-9.

Shay:2009:CST

[SB09] Richard Shay and Elisa Bertino. A comprehensive simulation tool for the analysis of password policies. International Journal of Information Security, 8(4):275–289, August 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        009-0084-3.

Saikia:2014:PHF

[SB14] Navajit Saikia and Prabin K. Bora. Perceptual hash function for scalable video. International Journal of Information Security, 13(1):81–93, February 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0211-z.

Spence:2022:SBC

[SB22] Aaron Spence and Shaun Bangay. Security beyond cybersecurity: side-channel attacks against non-cyber systems and their countermeasures. International Journal of Information Security, 21(3):437–453, June 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00563-6.

Samadani:2019:SPM

[SBB19] Mohammad Hasan Samadani, Mehdi Berenjkoob, and Marina Blanton. Secure pattern matching based on bit parallelism. International Journal of Information Security, 18 (3):371–391, June 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

                                        com/article/10.1007/s10207-
                                        018-0410-8.

Schlette:2021:MVC

[SBCP21] Daniel Schlette, Fabian Böhm, Marco Caselli, and Günther Pernul. Measuring and visualizing cyber threat intelligence quality. International Journal of Information Security, 20(1):21–38, February 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        020-00490-y;   http://
                                        link.springer.com/content/
                                        pdf/10.1007/s10207-020-
                                        00490-y.pdf.

Seth:2023:LDD

[SBD23] Abhay Deep Seth, Santosh Biswas, and Amit Kumar Dhar. LDES: detector design for version number attack detection using linear temporal logic based on discrete event system. International Journal of Information Security, 22(4):961–985, August 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00665-3.

Skopik:2022:SDA

[SBG22] Florian Skopik, Arndt Bonitz, and Günter Göhler.

From scattered data to actionable knowledge: flexible cyber security reporting in the military domain. International Journal of Information Security, 21 (6):1323–1347, December 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00613-7.

Sah:2023:IDS

[SBS23] Gulab Sah, Subhasish Banerjee, and Sweety Singh. Intrusion detection system over real-time data traffic using machine learning methods with feature selection approaches. International Journal of Information Security, 22 (1):1–27, February 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00616-4.

Saikam:2024:EAB

[SC24] Jalaiah Saikam and Koteswararao Ch. An ensemble approachbased intrusion detection system utilizing ISHO– HBA and SE–ResNet152. International Journal of Information Security, 23 (2):1037–1054, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00777-w.

Scaife:2018:OSR

[SCL^{+}18] Nolen Scaife, Henry Carter, Lyrissa Lidsky, Rachael L. Jones, and Patrick Traynor. OnionDNS: a seizureresistant top-level domain. International Journal of Information Security, 17 (6):645–660, November 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        017-0391-z.

Solano:2021:CBB

[SCO21] Jesus Solano, Luis Camacho, and Martín Ochoa. Combining behavioral biometrics and session context analytics to enhance risk-based static authentication in web applications. International Journal of Information Security, 20(2):181–197, April 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        020-00510-x.

Shah:2024:CHR

[SCY^{+}24] Rajiv Shah, Deniz Cemiloglu, Cagatay Yucel, Raian Ali, and Vasilis Katos. Is cyber hygiene a remedy to IPTV infringement? A study of online streaming behaviours and cyber security practices. International Journal of Information Security, 23(3): 1913–1926, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00824-0.

Stergiopoulos:2022:AAA

[SDG22] George Stergiopoulos, Panagiotis Dedousis, and Dimitris Gritzalis. Automatic analysis of attack graphs for risk mitigation and prioritization on large-scale and complex networks in Industry 4.0. International Journal of Information Security, 21 (1):37–59, February 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
020-00533-4.

Skoric:2016:FBR

[SdHZ16] Boris Skorić, Sebastiaan J. A. de Hoogh, and Nicola Zannone. Flow-based reputation with uncertainty: evidence-based subjective logic. International Journal of Information Security, 15(4):381–402, August 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        015-0298-5;   http://
                                        link.springer.com/content/
                                        pdf/10.1007/s10207-015-
                                        0298-5.pdf.

Sowjanya:2020:ECC

[SDR20] K. Sowjanya, Mou Dasgupta, and Sangram Ray. An elliptic curve cryptography based enhanced anonymous authentication protocol for wearable health monitoring systems. International Journal of Information Security, 19(1): 129–146, February 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        019-00464-9.

Sangari:2023:MRD

[SDW23] Seema Sangari, Eric Dallal, and Michael Whitman. Modeling reporting delays in cyber incidents: an industry-level comparison. International Journal of Information Security, 22(1):63–76, February 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00623-5.

Sen:2014:UIW

[Sen14] Sevil Sen. Using instanceweighted naive Bayes for adapting concept drift in masquerade detection. International Journal of Information Security, 13(6): 583–590, November 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0238-9.

Sepehrzadeh:2023:MIT

[Sep23] Hamed Sepehrzadeh. A method for insider threat assessment by modeling the internal employee interactions. International Journal of Information Security, 22(5): 1385–1393, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00697-9.

Seo:2018:AOF

[SEXY18] Jae Hong Seo, Keita Emura, Keita Xagawa, and Kazuki Yoneyama. Accumulable optimistic fair exchange from verifiably encrypted homomorphic signatures. International Journal of Information Security, 17 (2):193–220, April 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

                                        com/article/10.1007/s10207-
                                        017-0367-z.

Sommervoll:2024:SAA

[SEZ24] Åvald Åslaugson Sommervoll, László Erdődi, and Fabio Massimo Zennaro. Simulating all archetypes of SQL injection vulnerability exploitation using reinforcement learning agents. International Journal of Information Security, 23(1): 225–246, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00738-3.

Singh:2017:RUA

[SF17] Ankit Singh and Hervais C. Simo Fhom. Restricted usage of anonymous credentials in vehicular ad hoc networks for misbehavior detection. International Journal of Information Security, 16 (2):195–211, April 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        016-0328-y.

Smyrlis:2024:RRA

[SFB^{+}24] Michail Smyrlis, Evangelos Floros, Ioannis Basdekis, Dumitru-Bogdan Prelipcean, Aristeidis Sotiropoulos, Herve Debar, Apostolis Zarras, and George Spanoudakis. RAMA: a risk assessment solution for healthcare organizations. International Journal of Information Security, 23(3):1821–1838, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00820-4.

She:2024:MPC

[SFT^{+}24] Wei She, Haige Feng, Zhao Tian, Xinpeng Rong, Tianxiang Ma, and Wei Liu. Multi-party codebook distribution strategy based on secret sharing. International Journal of Information Security, 23(4): 3029–3042, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00873-5.

Shah:2022:MLO

[SGC22] Ankit Shah, Rajesh Ganesan, and Hasan Cam. Maintaining the level of operational effectiveness of a CSOC under adverse conditions. International Journal of Information Security, 21(3):637–651, June 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        021-00573-4.

Steinwandt:2002:APB

[SGE02] Rainer Steinwandt, Willi Geiselmann, and Regine Endsuleit. Attacking a polynomial-based cryptosystem: Polly cracker. International Journal of Information Security, 1(3): 143–148, November 2002. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        002-0012-2.

Shah:2019:MEF

[SGJ19] Ankit Shah, Rajesh Ganesan, and Sushil Jajodia. A methodology for ensuring fair allocation of CSOC effort for alert investigation. International Journal of Information Security, 18(2):199–218, April 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        018-0407-3.

Shah:2018:MMM

[SGJC18] Ankit Shah, Rajesh Ganesan, Sushil Jajodia, and Hasan Cam. A methodology to measure and monitor level of operational effectiveness of a CSOC.

International Journal of Information Security, 17 (2):121–134, April 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0365-1.

Saini:2019:YCS

[SGLC19] Anil Saini, Manoj Singh Gaur, Vijay Laxmi, and Mauro Conti. You click, I steal: analyzing and detecting click hijacking attacks in web pages. International Journal of Information Security, 18(4): 481–504, August 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
018-0423-3.

Sharma:2023:NAD

[SGSS23] Amit Sharma, Brij B.

Gupta, Awadhesh Kumar Singh, and V. K. Saraswat. A novel approach for detection of APT malware using multi-dimensional hybrid Bayesian belief network. International Journal of Information Security, 22(1):119–135, February 2023. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.

                                        com/article/10.1007/s10207-
                                        022-00631-5.

Shayesteh:2020:TMS

[SHA20] Behshid Shayesteh, Vesal Hakami, and Ahmad Akbari. A trust management scheme for IoT-enabled environmental health/accessibility monitoring services. International Journal of Information Security, 19(1): 93–110, February 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        019-00446-x.

Shuhan:2024:DIF

[SHD^{+}24] Mirza Kamrul Bashar Shuhan, Syed Md. Hasnayeen, Tanmoy Krishna Das, Md. Nazmus Sakib, and Md Sadek Ferdous. Decentralised identity federations using blockchain. International Journal of Information Security, 23(4): 2759–2782, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        024-00864-6.

Suren:2023:PPA

[SHOL23] Emre Süren, Fredrik Heiding, Johannes Olegård, and Robert Lagerström. PatrIoT: practical and agile threat research for IoT. International Journal of Information Security, 22(1): 213–233, February 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00633-3.

Song:2023:NFS

[SHW23] Guangjia Song, Jianhua Hu, and Hui Wang. A novel frame switching model based on virtual MAC in SDN. International Journal of Information Security, 22(3):723–736, June 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00659-7.

Silvestri:2024:CTA

[SIA^{+}24] Stefano Silvestri, Shareful Islam, Dmitry Amelin, Gabriele Weiler, Spyridon Papastergiou, and Mario Ciampi. Cyber threat assessment and management for securing healthcare ecosystems using natural language processing. International Journal of Information Security, 23 (1):31–50, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00769-w.

Sajedi:2009:SSB

[SJ09] Hedieh Sajedi and Mansour Jamzad. Secure steganography based on embedding capacity. International Journal of Information Security, 8(6): 433–445, December 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        009-0089-y.

Sajedi:2010:UCT

[SJ10] Hedieh Sajedi and Mansour Jamzad. Using contourlet transform and cover selection for secure steganography. International Journal of Information Security, 9(5): 337–352, October 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        010-0112-3.

Sidiroglou:2006:ETD

[SK06] Stelios Sidiroglou and Angelos D. Keromytis. Execution transactions for defending against software failures: use and evaluation. International Journal of Information Security, 5(2):77–91, April 2006. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
006-0083-6.

Shin:2014:AAA

[SK14] Sooyeon Shin and Taekyoung Kwon. AAnA: Anonymous authentication and authorization based on short traceable signatures. International Journal of Information Security, 13 (5):477–495, October 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0227-z.

Salini:2016:EPA

[SK16] P. Salini and S. Kanmani.

Effectiveness and performance analysis of modeloriented security requirements engineering to elicit security requirements: a systematic solution for developing secure software systems. International Journal of Information Security, 15(3):319–334, June 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
015-0305-x.

Smiliotopoulos:2023:DLM

[SKB23] Christos Smiliotopoulos, Georgios Kambourakis, and Konstantia Barbatsalou. On the detection of lateral movement through supervised machine learning and an open-source tool to create turnkey datasets from Sysmon logs. International Journal of Information Security, 22 (6):1893–1919, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00725-8.

Sneha:2023:RRT

[SKH^{+}23] M. Sneha, A. Keerthan Kumar, Nikhil V. Hegde, A. S. Anish, and G. Shobha. RADS: a real-time anomaly detection model for softwaredefined networks using machine learning. International Journal of Information Security, 22 (6):1881–1891, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00724-9.

Sisaat:2017:STM

[SKK^{+}17] Khamphao Sisaat, Surin Kittitornkun, Hiroaki Kikuchi, Chaxiong Yukonhiatou, Masato Terada, and Hiroshi Ishii. A spatiotemporal malware and country clustering algorithm: 2012 IIJ MITF case study. International Journal of Information Security, 16(5): 459–473, October 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0342-0.

Seo:2020:ERL

[SKLP20] Minhye Seo, Suhri Kim, Dong Hoon Lee, and Jong Hwan Park. EM-BLEM: (R)LWE-based key encapsulation with a new multi-bit encoding method. International Journal of Information Security, 19 (4):383–399, August 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00456-9.

Shukla:2023:QFS

[SKY23] Ankur Shukla, Basel Katt, and Muhammad Mudassar Yamin. A quantitative framework for security assurance evaluation and selection of cloud services: a case study. International Journal of Information Security, 22 (6):1621–1650, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00709-8.

Lhee:2005:DFB

[sLC05] Kyung suk Lhee and Steve J. Chapin. Detection of file-based race conditions. International Journal of Information Security, 4(1–2): 105–119, February 2005. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        004-0068-2.

Sarhan:2023:ZSM

[SLGP23] Mohanad Sarhan, Siamak Layeghy, Marcus Gallagher, and Marius Portmann. From zero-shot machine learning to zeroday attack detection. International Journal of Information Security, 22(4): 947–959, August 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00676-0.

Skoric:2010:FDS

[SM10] Boris Skorić and Marc X. Makkes. Flowchart description of security primitives for controlled physical unclonable functions. International Journal of Information Security, 9(5): 327–335, October 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
010-0113-2.

Saheed:2024:VGW

[SM24] Yakub Kayode Saheed and Sanjay Misra. A voting gray wolf optimizer-based ensemble learning models for intrusion detection in the Internet of Things. International Journal of Information Security, 23 (3):1557–1581, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00803-x.

Smith:2004:OAP

[Smi04] Sean W. Smith. Outbound authentication for programmable secure coprocessors. International Journal of Information Security, 3(1):28–41, October 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
004-0033-0.

Sole:2012:EMT

[SMMN12] Marc Solé, Victor Muntés-Mulero, and Jordi Nin. Efficient microaggregation techniques for large numerical data volumes. International Journal of Information Security, 11(4):

253–267, August 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        012-0158-5.

Shukla:2024:EBA

[SMV24] Sanjeev Shukla, Manoj Misra, and Gaurav Varshney. Email bombing attack detection and mitigation using machine learning. International Journal of Information Security, 23(4):2939–2949, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        024-00871-7.

Snekkenes:2005:PSI

[Sne05] Einar Snekkenes. Preface to the special issue on ES-ORICS 2003. International Journal of Information Security, 4(3):133–134, June 2005. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        004-0054-8.

Stasinopoulos:2019:CAE

[SNX19] Anastasios Stasinopoulos, Christoforos Ntantogian, and Christos Xenakis. Commix: automating evaluation and exploitation of command injection vulnerabilities in Web applications. International Journal of Information Security, 18(1):49–72, February 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
018-0399-z.

Stevanovic:2017:MIC

[SPDR17] Matija Stevanovic, Jens Myrup Pedersen, Alessandro D’Alconzo, and Stefan Ruehrup. A method for identifying compromised clients based on DNS traffic analysis. International Journal of Information Security, 16 (2):115–132, April 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0331-3.

Schreuders:2013:FBA

[SPM13] Z. Cliffe Schreuders, Christian Payne, and Tanya McGill. The functionalitybased application confinement model. International Journal of Information Security, 12(5): 393–422, October 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0199-4.

Sicari:2021:ABE

[SRD^{+}21] Sabrina Sicari, Alessandra Rizzardi, Gianluca Dini, Pericle Perazzo, Michele La Manna, and Alberto Coen-Porisini. Attribute-based encryption and sticky policies for data access control in a smart home scenario: a comparison on networked smart object middleware. International Journal of Information Security, 20 (5):695–713, October 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        020-00526-3;   http://
                                        link.springer.com/content/
                                        pdf/10.1007/s10207-020-
                                        00526-3.pdf.

Sethi:2020:CAR

[SRK^{+}20] Kamalakanta Sethi, E. Sai Rupesh, Rahul Kumar, Padmalochan Bera, and Y. Venu Madhav. A context-aware robust intrusion detection system: a reinforcement learningbased approach. International Journal of Information Security, 19(6): 657–678, December 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        019-00482-7.

Swati:2023:DAD

[SRSM23] Swati, Sangita Roy, Jawar Singh, and Jimson Mathew. Design and analysis of DDoS mitigating network architecture. International Journal of Information Security, 22(2):333– 345, April 2023. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00635-1.

Serjantov:2005:PAA

[SS05a] Andrei Serjantov and Peter Sewell. Passive-attack analysis for connectionbased anonymity systems. International Journal of Information Security, 4 (3):172–180, June 2005. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
004-0059-3.

Skalka:2005:SUB

[SS05b] Christian Skalka and Scott Smith. Static use-based object confinement. International Journal of Information Security, 4(1– 2):87–104, February 2005. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

com/article/10.1007/s10207-
004-0049-5.

Sampangi:2017:HSR

[SS17] Raghav V. Sampangi and Srinivas Sampalli. HiveSec: security in resource-constrained wireless networks inspired by beehives and bee swarms. International Journal of Information Security, 16(4):417–433, August 2017. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0341-1.

Shameli-Sendi:2014:ACA

[SSD14] Alireza Shameli-Sendi and Michel Dagenais. ARITO: Cyber-attack response system using accurate risk impact tolerance. International Journal of Information Security, 13(4): 367–390, August 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0222-9.

Spreitzenbarth:2015:MSC

[SSE^{+}15] Michael Spreitzenbarth, Thomas Schreck, Florian Echtler, Daniel Arp, and Johannes Hoffmann. Mobile-Sandbox: combining static and dynamic analysis with machinelearning techniques. International Journal of Information Security, 14 (2):141–153, April 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0250-0.

Seifi:2015:ATA

[SSFB15] Younes Seifi, Suriadi Suriadi, Ernest Foo, and Colin Boyd. Analysis of two authorization protocols using Colored Petri Nets. International Journal of Information Security, 14 (3):221–247, June 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0243-z.

Soltani:2022:CBD

[SSJ22] Mahdi Soltani, Mahdi Jafari Siavoshani, and Amir Hossein Jahangir. A contentbased deep intrusion detection system. International Journal of Information Security, 21(3):547–562, June 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00567-2.

Sheikhalishahi:2022:PPD

[SSL22a] Mina Sheikhalishahi, Andrea Saracino, and Antonio La Marra. Privacy preserving data sharing and analysis for edgebased architectures. International Journal of Information Security, 21(1): 79–101, February 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00542-x.

Shin:2022:NSS

[SSL22b] Hansub Shin, Sungyong Sim, and Younho Lee. A new smart smudge attack using CNN. International Journal of Information Security, 21 (1):25–36, February 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00540-z.

Sheikhalishahi:2020:DWD

[SSM^{+}20] Mina Sheikhalishahi, Andrea Saracino, Fabio Martinelli, Antonio La Marra, Mohammed Mejri, and Nadia Tawbi. Digital waste disposal: an automated framework for analysis of spam emails. International Journal of Information Security, 19(5): 499–522, October 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00470-x.

Silva:2015:RMP

[SSN15] Helber Silva, Aldri Santos, and Michele Nogueira. Routing management for performance and security tradeoff in wireless mesh networks. International Journal of Information Security, 14 (1):35–46, February 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0246-9.

Sepahi:2014:LBC

[SSP14] Reza Sepahi, Ron Steinfeld, and Josef Pieprzyk. Lattice-based certificateless public-key encryption in the standard model. International Journal of Information Security, 13 (4):315–333, August 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0215-8.

Singh:2022:DAI

[SSR22] Ram Govind Singh, Ananya Shrivastava, and Sushmita Ruj. A digital asset inheritance model to convey online persona posthumously. International Journal of Information Security, 21(5): 983–1003, October 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00593-8.

Singh:2024:BBB

[SSSJ24] Upendra Singh, Sumit Kumar Sharma, Mukul Shukla, and Preeti Jha. Blockchainbased BATMAN protocol using mobile ad hoc network (MANET) with an ensemble algorithm. International Journal of Information Security, 23 (3):1667–1677, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00804-w.

Singh:2022:IAM

[SSV22] Jaya Singh, Ayush Sinha, and Om Prakash Vyas. Insider attack mitigation in a smart metering infrastructure using reputation score and blockchain technology. International Journal of Information Security, 21(3):527–546, June 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.

com/article/10.1007/s10207-
021-00561-8.

Susil:2016:SSA

[SSVC16] Petr Susil, Pouyan Sepehrdad, Serge Vaudenay, and Nicolas Courtois. On selection of samples in algebraic attacks and a new technique to find hidden low degree equations. International Journal of Information Security, 15 (1):51–65, February 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
015-0295-8.

Soumelidou:2024:VET

[ST24] Aikaterini Soumelidou and Aggeliki Tsohou. Validation and extension of two domain-specific information privacy competency models. International Journal of Information Security, 23(3): 2437–2455, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00843-x.

Spathoulas:2021:UHE

[STD21] Georgios Spathoulas, Georgios Theodoridis, and Georgios-Paraskevas Damiris. Using homomorphic encryption for privacy-preserving clustering of intrusion detection alerts. International Journal of Information Security, 20 (3):347–370, June 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00506-7.

Shi:2023:SAM

[STW^{+}23] Sibo Shi, Shengwei Tian, Bo Wang, Tiejun Zhou, and Guanxin Chen. SFCG-Droid: Android malware detection based on sensitive function call graph. International Journal of Information Security, 22(5): 1115–1124, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00679-x.

Saxena:2011:DRE

[SV11] Nitesh Saxena and Jonathan Voris. Data remanence effects on memory-based entropy collection for RFID systems. International Journal of Information Security, 10(4):213–222, August 2011. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
011-0139-0.

Sermpinis:2021:DDT

[SVKV21] Thomas Sermpinis, George Vlahavas, Konstantinos Karasavvas, and Athena Vakali. DeTRACT: a decentralized, transparent, immutable and open PKI certificate framework. International Journal of Information Security, 20(4): 553–570, August 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00518-3.

Sun:2020:PDI

[SXZC20] Lixue Sun, Chunxiang Xu, Yuan Zhang, and Kefei Chen. Public data integrity auditing without homomorphic authenticators from indistinguishability obfuscation. International Journal of Information Security, 19(6): 711–720, December 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00486-8.

Sheidani:2024:PPT

[SZ24] Sorour Sheidani and Alireza Zarei. Privacy-preserving two-party computation of line segment intersection. International Journal of Information Security, 23 (5):3415–3432, October 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        024-00895-z.

Taheri:2020:CTU

[TBGB20] Sona Taheri, Adil M. Bagirov, Iqbal Gondal, and Simon Brown. Cyberattack triage using incremental clustering for intrusion detection systems. International Journal of Information Security, 19 (5):597–607, October 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        019-00478-3.

Tuna:2023:TNH

[TCE23] Omer Faruk Tuna, Ferhat Ozgur Catak, and M. Taner Eskil. TENET: a new hybrid network architecture for adversarial defense. International Journal of Information Security, 22(4): 987–1004, August 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00675-1.

Takahashi:2020:MGE

[TCS^{+}20] Takeshi Takahashi, Rodrigo Roman Castro, Bilhanan Silverajan, Ryan K. L. Ko, and Said Tabet. Message from the guest editors. International Journal of Information Security, 19(1):1–2, February 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00472-9;   http://
link.springer.com/content/
pdf/10.1007/s10207-019-
00472-9.pdf.

Tommasi:2022:BMB

[TCT22] Franco Tommasi, Christian Catalano, and Ivan Taurino. Browser-in-the-Middle (BitM) attack. International Journal of Information Security, 21 (2):179–189, April 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00548-5.

Tsohou:2023:CIS

[TDGL23] Aggeliki Tsohou, Vasiliki Diamantopoulou, Stefanos Gritzalis, and Costas Lambrinoudakis. Cyber insurance: state of the art, trends and future directions. International Journal of Information Security, 22(3):737–748, June 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00660-8.

Tiloca:2020:GRB

[TDRR20] Marco Tiloca, Gianluca Dini, Kiki Rizki, and Shahid Raza. Group rekeying based on member join history. International Journal of Information Security, 19(4): 343–381, August 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        019-00451-0.

Trostle:2005:TIS

[TG05] Jonathan Trostle and Bill Gossman. Techniques for improving the security and manageability of IPsec policy. International Journal of Information Security, 4(3):209–226, June 2005. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        004-0064-6.

Torra:2022:PFP

[TGNA22] Vicenç Torra, Edgar Galván, and Guillermo Navarro-Arribas. PSO + FL = PAASO: particle swarm optimization + federated learning = privacy-aware agent swarm optimization.

International Journal of Information Security, 21 (6):1349–1359, December 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00614-6.

Tiloca:2017:IRD

[TGS17] Marco Tiloca, Christian Gehrmann, and Ludwig Seitz. On improving resistance to Denial of Service and key provisioning scalability of the DTLS handshake. International Journal of Information Security, 16(2):173–193, April 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0326-0.

Tiplea:2022:NPD

[Tip22] Ferucio Laurentiu Tiplea.

Narrow privacy and desynchronization in Vaudenay’s RFID model. International Journal of Information Security, 21(3):563–575, June 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00569-0.

Tyagi:2024:EUA

[TK24] Gaurav Tyagi and Rahul Kumar. An efficient user authentication and key agreement scheme for wireless sensor networks using physically unclonable function. International Journal of Information Security, 23(2):935–962, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00770-3.

Takase:2020:PIE

[TKKO20] Hayate Takase, Ryotaro Kobayashi, Masahiko Kato, and Ren Ohmura. A prototype implementation and evaluation of the malware detection mechanism for IoT devices using the processor information. International Journal of Information Security, 19 (1):71–81, February 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00437-y.

Tian:2009:LSN

[TLX09] Daxin Tian, Yanheng Liu, and Yang Xiang. Largescale network intrusion detection based on distributed learning algorithm. International Journal of Information Security, 8(1):25–35, February 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
008-0061-2.

Takahashi:2019:SSF

[TMM^{+}19] Kenta Takahashi, Takahiro Matsuda, Takao Murakami, Goichiro Hanaoka, and Masakatsu Nishigaki. Signature schemes with a fuzzy private key. International Journal of Information Security, 18(5): 581–617, October 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00428-z;   http://
link.springer.com/content/
pdf/10.1007/s10207-019-
00428-z.pdf.

Tormo:2013:DAI

[TMP13] Ginés Dólera Tormo, Gabriel López Millán, and Gregorio Martínez Pérez. Definition of an advanced identity management infrastructure. International Journal of Information Security, 12 (3):173–200, June 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0189-y.

Toreini:2024:FSF

[TMvM24] Ehsan Toreini, Maryam Mehrnezhad, and Aad van Moorsel. Fairness as a Service (FaaS): verifiable and privacy-preserving fairness auditing of machine learning systems. International Journal of Information Security, 23(2):981– 997, April 2024. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00774-z.

Torra:2023:ADR

[TNA23] Vicenç Torra and Guillermo Navarro-Arribas. Attribute disclosure risk for k-anonymity: the case of numerical data. International Journal of Information Security, 22 (6):2015–2024, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00730-x.

Tonicelli:2015:ITS

[TND^{+}15] Rafael Tonicelli, Anderson C. A. Nascimento, Rafael Dowsley, Jörn Müller-Quade, Hideki Imai, Goichiro Hanaoka, and Akira Otsuka. Information-theoretically secure oblivious polynomial evaluation in the commodity-based model. International Journal of Information Security, 14 (1):73–84, February 2015.

CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0247-8.

Torra:2020:RDP

[Tor20] Vicenç Torra. Random dictatorship for privacypreserving social choice. International Journal of Information Security, 19 (5):537–545, October 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00474-7;   http://
link.springer.com/content/
pdf/10.1007/s10207-019-
00474-7.pdf.

Triakosia:2024:SOM

[TRT^{+}24] Aikaterini Triakosia, Panagiotis Rizomiliotis, Cecilia Tonelli, Fabio Federici, and Valerio Senni. Secure outsourcing of manufacturing compliance checks. International Journal of Information Security, 23(1): 609–627, February 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00747-2.

Tavizi:2020:LPS

[TS20] Tina Tavizi and Mehdi Shajari. A language and a pattern system for temporal property specification: advanced metering infrastructure case study. International Journal of Information Security, 19(6): 695–710, December 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00483-6.

Templ:2022:SOM

[TS22] Matthias Templ and Murat Sariyar. A systematic overview on methods to protect sensitive data provided for various analyses. International Journal of Information Security, 21 (6):1233–1246, December 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00607-5.

Toreini:2019:DEW

[TSMH19] Ehsan Toreini, Siamak F. Shahandashti, Maryam Mehrnezhad, and Feng Hao. DOMtegrity: ensuring web page integrity against malicious browser extensions. International Journal of Information Security, 18(6):801–814, December 2019. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00442-1;   http://
link.springer.com/content/
pdf/10.1007/s10207-019-
00442-1.pdf.

Talukder:2024:MWM

[TSU^{+}24] Md. Alamin Talukder, Selina Sharmin, Md Ashraf Uddin, Md Manowarul Islam, and Sunil Aryal. MLSTL-WSN: machine learning-based intrusion detection using SMOTE-Tomek in WSNs. International Journal of Information Security, 23(3): 2139–2158, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00833-z.

Tan:2022:CPQ

[TSZ22] Teik Guan Tan, Pawel Szalachowski, and Jianying Zhou. Challenges of postquantum digital signing in real-world applications: a survey. International Journal of Information Security, 21(4):937–952, August 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00587-6.

Tsunoo:2006:ICA

[TTS^{+}06] Yukiyasu Tsunoo, Etsuko Tsujihara, Maki Shigeri, Hiroyasu Kubo, and Kazuhiko Minematsu. Improving cache attacks by considering cipher structure. International Journal of Information Security, 5(3):166–176, July 2006. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
005-0079-7.

Tzavara:2024:TEC

[TV24] Vasiliki Tzavara and Savvas Vassiliadis. Tracing the evolution of cyber resilience: a historical and conceptual review. International Journal of Information Security, 23(3):1695–1719, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00811-x.

Tartary:2008:CAM

[TWP08] Christophe Tartary, Huaxiong Wang, and Josef Pieprzyk. A coding approach to the multicast stream authentication problem. International Journal of Information Security, 7(4): 265–283, August 2008. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0048-4.

Teller:2004:UAC

[TZH04] David Teller, Pascal Zimmer, and Daniel Hirschkoff. Using ambients to control resources. International Journal of Information Security, 2(3– 4):126–144, August 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
004-0035-y.

Ullah:2023:PTA

[UBK23] Imdad Ullah, Roksana Boreli, and Salil S. Kanhere. Privacy in targeted advertising on mobile devices: a survey. International Journal of Information Security, 22 (3):647–678, June 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00655-x.

Ukeje:2024:ISP

[UGP24] Ndukwe Ukeje, Jairo Gutierrez, and Krassie Petrova. Information security and privacy challenges of cloud computing for government adoption: a systematic review. International Journal of Information Security, 23(2): 1459–1475, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00797-6.

Upreti:2024:TML

[ULEY24] Ramesh Upreti, Pedro G. Lind, Ahmed Elmokashfi, and Anis Yazidi. Trustworthy machine learning in the context of security and privacy. International Journal of Information Security, 23(3):2287–2314, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        024-00813-3.

Ueda:2020:SIR

[UMN^{+}20] Itaru Ueda, Daiki Miyahara, Akihiro Nishimura, Yu ichi Hayashi, Takaaki Mizuki, and Hideaki Sone. Secure implementations of a random bisection cut. International Journal of Information Security, 19 (4):445–452, August 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        019-00463-w.

Ustaoglu:2011:IIB

[Ust11] Berkant Ustaoğlu. Integrating identity-based and certificate-based authenticated key exchange protocols. International Journal of Information Security, 10(4):201–212, August 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
011-0136-3.

Vajda:2016:ATA

[Vaj16] István Vajda. On the analysis of time-aware protocols in universal composability framework. International Journal of Information Security, 15(4): 403–412, August 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
015-0300-2.

Vasilellis:2024:GST

[VBAG24] Efstratios Vasilellis, Vasileios Botsos, Argiro Anagnostopoulou, and Dimitris Gritzalis. Gaming the system: tetromino-based covert channel and its impact on mobile security. International Journal of Information Security, 23(4): 3007–3027, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.

                                        com/article/10.1007/s10207-
                                        024-00875-3.

VandenBroeck:2021:OIS

[VCD21] Jens Van den Broeck, Bart Coppens, and Bjorn De Sutter. Obfuscated integration of software protections. International Journal of Information Security, 20(1): 73–101, February 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        020-00494-8.

Vishwakarma:2024:BBB

[VD24] Lokendra Vishwakarma and Debasis Das. BLISS: blockchain-based integrated security system for Internet of Things (IoT) applications. International Journal of Information Security, 23(3):1649–1665, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00808-6.

Veeningen:2014:DMC

[VdWZ14] Meilof Veeningen, Benne de Weger, and Nicola Zannone. Data minimisation in communication protocols: a formal analysis framework and application to identity management. International Journal of Information Security, 13(6): 529–569, November 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0235-z.

Vasilellis:2024:PPI

[VGG24] Efstratios Vasilellis, Grigoris Gkionis, and Dimitris Gritzalis. Press play, install malware: a study of rhythm game-based malware dropping. International Journal of Information Security, 23(5): 3369–3391, October 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00893-1.

Vora:2019:KBP

[VH19] Aishwarya Vipul Vora and Saumya Hegde. Keywordbased private searching on cloud data along with keyword association and dissociation using cuckoo filter. International Journal of Information Security, 18(3):305–319, June 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
018-0418-0.

Villalon-Huerta:2024:SCC

[VHRRMG24] Antonio Villalón-Huerta, Ismael Ripoll-Ripoll, and Hector Marco-Gisbert. A survey and characterization of Close Access Cyberspace Operations. International Journal of Information Security, 23 (2):963–980, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00772-1.

Vasserman:2009:IKN

[VHT09] Eugene Y. Vasserman, Nicholas Hopper, and James Tyra. Silent-Knock: practical, provably undetectable authentication. International Journal of Information Security, 8(2):121–135, April 2009. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        008-0070-1.

Vrakas:2013:IDP

[VL13] Nikos Vrakas and Costas Lambrinoudakis. An intrusion detection and prevention system for IMS and VoIP services. International Journal of Information Security, 12 (3):201–217, June 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0187-0.

Vitel:2024:SVL

[VLGL24] Silviu Viţel, Marilena Lupaşcu, Dragoş Teodor Gavriluţ, and Henri Luchian. Short- versus long-term performance of detection models for obfuscated MSOffice-embedded malware. International Journal of Information Security, 23(1):271–297, February 2024. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00736-5.

Velasquez:2023:SRS

[VMCR23] Juan Miguel López Velásquez, Sergio Mauricio Martínez Monterrubio, Luis Enrique Sánchez Crespo, and David Garcia Rosado. Systematic review of SIEM technology: SIEM–SC birth. International Journal of Information Security, 22(3):691–711, June 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00657-9.

vonOheimb:2005:ASM

[vOLW05] David von Oheimb, Volkmar Lotz, and Georg Walter. Analyzing SLE 88 memory management security using Interacting State Machines. International Journal of Information Security, 4(3):155–171, June 2005. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        004-0057-5.

vanOorschot:2006:MSD

[vORM06] Paul C. van Oorschot, Jean-Marc Robert, and Miguel Vargas Martin. A monitoring system for detecting repeated packets with applications to computer worms. International Journal of Information Security, 5(3):186–199, July 2006. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        006-0081-8.

Vasiliadis:2015:GAM

[VPI15] Giorgos Vasiliadis, Michalis Polychronakis, and Sotiris Ioannidis. GPU-assisted malware. International Journal of Information Security, 14(3):289–297, June 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0262-9.

Vanrenen:2006:DSM

[VSM06] Gabriel Vanrenen, Sean Smith, and John Marchesini. Distributing securitymediated PKI. International Journal of Information Security, 5(1):3– 17, January 2006. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
005-0076-x.

Vlachos:2022:SOS

[VSN22] Vasileios Vlachos, Yannis C. Stamatiou, and Sotiris Nikoletseas. The SAINT observatory subsystem: an open-source intelligence tool for uncovering cybersecurity threats. International Journal of Information Security, 21 (5):1091–1106, October 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00599-2.

Valenzuela:2015:MAO

[VSR15] Michael Valenzuela, Ferenc Szidarovszky, and Jerzy Rozenblit. A multiresolution approach for optimal defense against random attacks. International Journal of Information Security, 14(1):61–72, February 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        014-0245-x.

Wu:2009:IDV

[WAB^{+}09] Yu-Sung Wu, Vinita Apte, Saurabh Bagchi, Sachin Garg, and Navjot Singh. Intrusion detection in voice over IP environments. International Journal of Information Security, 8(3): 153–172, June 2009. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        008-0071-0.

Waidner:2004:P

[Wai04] Michael Waidner. Preface. International Journal of Information Security, 3(1):1, October 2004. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        004-0051-y;   http://
                                        link.springer.com/content/
                                        pdf/10.1007/s10207-004-
                                        0051-y.pdf.

Wu:2020:CIB

[WCS20] Chunhui Wu, Xiaofeng Chen, and Willy Susilo.

Concise ID-based mercurial functional commitments and applications to zero-knowledge sets. International Journal of Information Security, 19(4): 453–464, August 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00466-7.

Wang:2023:ANM

[WGH23] Runzheng Wang, Jian Gao, and Shuhua Huang. AI-HGAT: a novel method of malware detection and homology analysis using assembly instruction heterogeneous graph. International Journal of Information Security, 22(5): 1423–1443, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00699-7.

Wang:2013:USM

[WGMB13] Pu Wang, Marta C.

González, Ronaldo Menezes, and Albert-László Barabási. Understanding the spread of malicious mobile-phone programs and their damage potential. International Journal of Information Security, 12(5): 383–392, October 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        013-0203-z.

Wangen:2018:FEI

[WHS18] Gaute Wangen, Christoffer Hallstensen, and Einar Snekkenes. A framework for estimating information security risk assessment method completeness. International Journal of Information Security, 17(6): 681–699, November 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        017-0382-0;   http://
                                        link.springer.com/content/
                                        pdf/10.1007/s10207-017-
                                        0382-0.pdf.

Wardana:2024:CID

[WKWS24] Aulia Arif Wardana, Grzegorz Kołaczek, Arkadiusz Warzyński, and Parman Sukarno. Collaborative intrusion detection using weighted ensemble averaging deep neural network for coordinated attack detection in heterogeneous network. International Journal of Information Security, 23(5): 3329–3349, October 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.

com/article/10.1007/s10207-
024-00891-3.

Wu:2014:SSP

[WLLW14] Tzong-Sun Wu, Ming-Lun Lee, Han-Yu Lin, and Chao-Yuan Wang. Shoulder-surfing-proof graphical password authentication scheme. International Journal of Information Security, 13(3):245–254, June 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0216-7.

Wijesundara:2024:SEF

[WLT^{+}24] W. M. A. B. Wijesundara, Joong-Sun Lee, Dara Tith, Eleni Aloupogianni, Hiroyuki Suzuki, and Takashi Obi. Securityenhanced firmware management scheme for smart home IoT devices using distributed ledger technologies. International Journal of Information Security, 23(3):1927–1937, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00827-x.

Lye:2005:GSN

[wLW05] Kong wei Lye and Jeannette M. Wing. Game strategies in network security. International Journal of Information Security, 4(1–2):71–86, February 2005. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        004-0060-x.

Wu:2019:TPP

[WMS^{+}19] Ge Wu, Yi Mu, Willy Susilo, Fuchun Guo, and Futai Zhang. Threshold privacy-preserving cloud auditing with multiple uploaders. International Journal of Information Security, 18(3):321–331, June 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        018-0420-6.

Wang:2018:VEC

[WPD18] Yujue Wang, HweeHwa Pang, and Robert H. Deng. Verifiably encrypted cascade-instantiable blank signatures to secure progressive decision management. International Journal of Information Security, 17(3):347–363, June 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        017-0372-2.

Wang:2008:MLF

[WR08] XiaoFeng Wang and Michael K.

Reiter. A multi-layer framework for puzzle-based denial-of-service defense. International Journal of Information Security, 7 (4):243–263, August 2008. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0042-x.

Wei:2015:TPE

[WR15] Lei Wei and Michael K.

Reiter. Toward practical encrypted email that supports private, regularexpression searches. International Journal of Information Security, 14(5): 397–416, October 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0268-3;   http://
link.springer.com/content/
pdf/10.1007/s10207-014-
0268-3.pdf.

Wang:2016:SSE

[WT16] Yuyu Wang and Keisuke Tanaka. Strongly simulationextractable leakage-resilient NIZK. International Journal of Information Security, 15(1):67–79, February 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

                                        com/article/10.1007/s10207-
                                        015-0291-z.

Wang:2021:SLR

[WT21] Yuntao Wang and Tsuyoshi Takagi. Studying lattice reduction algorithms improved by quick reordering technique. International Journal of Information Security, 20 (2):257–268, April 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        020-00501-y.

Wu:2007:RFT

[WW07] Zhengping Wu and Alfred C. Weaver. Requirements of federated trust management for serviceoriented architectures. International Journal of Information Security, 6(5): 287–296, September 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        007-0027-9.

Wang:2023:DAM

[WWZ^{+}23] Kai Wang, Jinxia Wu, Tianqing Zhu, Wei Ren, and Ying Hong. Defense against membership inference attack in graph neural networks through graph perturbation. International Journal of Information Security, 22 (2):497–509, April 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
022-00646-y.

Wang:2012:SGN

[WYL^{+}12] Yuanzhuo Wang, Min Yu, Jingyuan Li, Kun Meng, Chuang Lin, and Xueqi Cheng. Stochastic game net and applications in security analysis for enterprise network. International Journal of Information Security, 11 (1):41–52, February 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
011-0148-z.

Wang:2007:HDA

[WZ07] Shujing Wang and Yan Zhang. Handling distributed authorization with delegation through answer set programming. International Journal of Information Security, 6 (1):27–46, January 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
006-0008-4.

Wang:2024:FSL

[WZZ^{+}24] Yahui Wang, Zhiyong Zhang, Kejing Zhao, Peng Wang, and Ruirui Wu. A few-shot learning based method for industrial internet intrusion detection. International Journal of Information Security, 23(5): 3241–3252, October 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        024-00889-x.

Xin:2012:RAA

[XCW^{+}12] Zhi Xin, Huiyu Chen, Xinche Wang, Peng Liu, Sencun Zhu, Bing Mao, and Li Xie. Replacement attacks: automatically evading behaviorbased software birthmark. International Journal of Information Security, 11 (5):293–304, October 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        012-0170-9.

Xu:2013:VBP

[XSA13] Wenjuan Xu, Mohamed Shehab, and Gail-Joon Ahn. Visualization-based policy analysis for SELinux: framework and user study. International Journal of Information Security, 12 (3):155–171, June 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0180-7.

Xiao:2024:AAT

[XT24] Chong Xiao and Ming Tang. Acnn: arbitrary trace attacks based on leakage area detection. International Journal of Information Security, 23(4): 2991–3006, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00874-4.

Xun:2024:LCA

[XYZT24] Peng Xun, Ziyu Yang, Haoyang Zhu, and Zhu Tang. Locating collaborative attack targets based on physical invariants toward cyber-physical systems. International Journal of Information Security, 23(2):999–1017, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00773-0.

Xu:2021:SLR

[XZ21] Jia Xu and Jianying Zhou.

Strong leakage-resilient encryption: enhancing data confidentiality by hiding partial ciphertext. International Journal of Information Security, 20 (2):141–159, April 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
020-00487-7.

Xu:2024:GAE

[XZ24] Lei Xu and Junhai Zhai. Generating adversarial examples with collaborative generative models. International Journal of Information Security, 23(2): 1077–1091, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00780-1.

Yildirim:2022:MIT

[YA22] Metehan Yildirim and Emin Anarim. Mitigating insider threat by profiling users based on mouse usage pattern: ensemble learning and frequency domain analysis. International Journal of Information Security, 21 (2):239–251, April 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00544-9.

Yu:2015:EPR

[YAM^{+}15] Yong Yu, Man Ho Au, Yi Mu, Shaohua Tang, Jian Ren, Willy Susilo, and Liju Dong. Enhanced privacy of a remote data integrity-checking protocol for secure cloud storage. International Journal of Information Security, 14(4):307–318, August 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
014-0263-8.

Yankson:2021:CIP

[Yan21] Benjamin Yankson. Continuous improvement process (CIP)-based privacypreserving framework for smart connected toys. International Journal of Information Security, 20(6): 849–869, December 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00535-2.

Yamauchi:2021:AKO

[YAY^{+}21] Toshihiro Yamauchi, Yohei Akao, Ryota Yoshitani, Yuichi Nakamura, and Masaki Hashimoto. Additional kernel observer: privilege escalation attack prevention mechanism focusing on system call privilege changes. International Journal of Information Security, 20(4): 461–473, August 2021. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00514-7;   http://
link.springer.com/content/
pdf/10.1007/s10207-020-
00514-7.pdf.

Yichiet:2022:SAL

[YKP22] Aun Yichiet, Yen-Min Jasmina Khaw, and Vasaki Ponnusamy. A semanticaware log generation method for network activities. International Journal of Information Security, 21 (2):161–177, April 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00547-6.

Yu:2019:UUP

[YL19] Xiaoying Yu and Qi Liao. Understanding user passwords through password prefix and postfix (P3) graph analysis and visualization. International Journal of Information Security, 18(5): 647–663, October 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

com/article/10.1007/s10207-
019-00432-3.

Yohan:2020:FSB

[YL20] Alexander Yohan and Nai-Wei Lo. FOTB: a secure blockchain-based firmware update framework for IoT environment. International Journal of Information Security, 19(3):257–278, June 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00467-6.

Yang:2018:NSS

[YLL^{+}18] Zheng Yang, Chao Liu, Wanping Liu, Daigu Zhang, and Song Luo. A new strong security model for stateful authenticated group key exchange. International Journal of Information Security, 17(4): 423–440, August 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0373-1.

Yildirim:2019:EUI

[YM19] M. Yildirim and I. Mackie.

Encouraging users to improve password security and memorability. International Journal of Information Security, 18(6): 741–759, December 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        019-00429-y;   http://
                                        link.springer.com/content/
                                        pdf/10.1007/s10207-019-
                                        00429-y.pdf.

Yaacoub:2022:RCS

[YNC22] Jean-Paul A. Yaacoub, Hassan N. Noura, and Ali Chehab. Robotics cyber security: vulnerabilities, attacks, countermeasures, and recommendations. International Journal of Information Security, 21(1): 115–158, February 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        021-00545-8.

Yoneyama:2018:FMR

[Yon18] Kazuki Yoneyama. Formal modeling of random oracle programmability and verification of signature unforgeability using task-PIOAs. International Journal of Information Security, 17(1):43–66, February 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        016-0352-y.

Young:2006:CEU

[You06] Adam L. Young. Cryptoviral extortion using Microsoft’s crypto API. International Journal of Information Security, 5(2): 67–76, April 2006. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
006-0082-7.

Yoshino:2009:BMM

[YOV09] Masayuki Yoshino, Katsuyuki Okeya, and Camille Vuillaume. Bipartite modular multiplication with twice the bit-length of multipliers. International Journal of Information Security, 8(1):13– 23, February 2009. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
008-0060-3.

Yeo:2006:SWE

[YP06] Gary S.-W. Yeo and Raphael C.-W. Phan. On the security of the Win-RAR encryption feature. International Journal of Information Security, 5 (2):115–123, April 2006. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
006-0086-3.

Yaseen:2012:ITM

[YP12] Qussai Yaseen and Brajendra Panda. Insider threat mitigation: preventing unauthorized knowledge acquisition. International Journal of Information Security, 11(4): 269–280, August 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        012-0165-6.

Yasmin:2014:PSP

[YRW14] Rehana Yasmin, Eike Ritter, and Guilin Wang. Provable security of a pairing-free one-pass authenticated key establishment protocol for wireless sensor networks. International Journal of Information Security, 13(5): 453–465, October 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        013-0224-7.

Yeh:2020:SIS

[YSD^{+}20] Kuo-Hui Yeh, Chunhua Su, Robert H. Deng, Moti Yung, and Miroslaw Kutylowski. Special issue on security and privacy of blockchain technologies. International Journal of Information Security, 19 (3):243–244, June 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
020-00496-6;   http://
link.springer.com/content/
pdf/10.1007/s10207-020-
00496-6.pdf.

Yuen:2010:HCI

[YSM10] Tsz Hon Yuen, Willy Susilo, and Yi Mu. How to construct identity-based signatures without the key escrow problem. International Journal of Information Security, 9(4): 297–311, August 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
010-0110-5.

Yang:2016:MGE

[YSM16] Guomin Yang, Willy Susilo, and Yi Mu. Message from the Guest Editors. International Journal of Information Security, 15 (2):223–224, April 2016. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
016-0315-3;   http://
link.springer.com/content/
pdf/10.1007/s10207-016-
0315-3.pdf.

Yang:2022:SAI

[YWW22] Xiaodong Yang, Wenjia Wang, and Caifen Wang. Security analysis and improvement of a privacypreserving authentication scheme in VANET. International Journal of Information Security, 21 (6):1361–1371, December 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00617-3.

Yang:2023:AEA

[YYH^{+}23] Zhen Yang, Shisong Yang, Yunbo Huang, José-Fernán Martínez, Lourdes López, and Yuwen Chen. AAIA: an efficient aggregation scheme against inverting attack for federated learning. International Journal of Information Security, 22(4):919–930, August 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00670-6.

Yoneyama:2018:MCK

[YYK^{+}18] Kazuki Yoneyama, Reo Yoshida, Yuto Kawahara, Tetsutaro Kobayashi, Hitoshi Fuji, and Tomohide Yamamoto. Multi-cast key distribution: scalable, dynamic and provably secure construction. International Journal of Information Security, 17(5): 513–532, October 2018. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
017-0389-6.

Yousaf:2024:SSM

[YZ24] Awais Yousaf and Jianying Zhou. From sinking to saving: MITRE ATT&CK and D3FEND frameworks for maritime cybersecurity. International Journal of Information Security, 23(3):1603–1618, June 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
024-00812-4.

Zeghida:2023:SMP

[ZBC23] Hayette Zeghida, Mehdi Boulaiche, and Ramdane Chikh. Securing MQTT protocol for IoT environment using IDS based on ensemble learning. International Journal of Information Security, 22(4): 1075–1086, August 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00681-3.

Zhou:2006:MTI

[ZBD06] Jianying Zhou, Feng Bao, and Robert Deng. Minimizing TTP’s involvement in signature validation. International Journal of Information Security, 5(1):37–47, January 2006. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        005-0072-1.

Zabihimayvan:2022:FLR

[ZD22] Mahdieh Zabihimayvan and Derek Doran. A first look at references from the dark to the surface web world: a case study in Tor. International Journal of Information Security, 21(4):739–755, August 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00580-z.

Zenitani:2022:MOC

[Zen22] Kengo Zenitani. A multiobjective cost-benefit optimization algorithm for network hardening. International Journal of Information Security, 21(4): 813–832, August 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.

com/article/10.1007/s10207-
022-00586-7.

Zhang:2007:FBT

[ZGC07] Jie Zhang, Ali A. Ghorbani, and Robin Cohen. A familiarity-based trust model for effective selection of sellers in multiagent e-commerce systems. International Journal of Information Security, 6(5): 333–344, September 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0025-y.

Zulkernine:2007:ISS

[ZGK07] Mohammad Zulkernine, Mathews Graves, and Muhammad Umair Ahmed Khan. Integrating software specifications into intrusion detection. International Journal of Information Security, 6(5): 345–357, September 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0023-0.

Zhang:2023:ISF

[ZGRS23] Denghui Zhang, Zhaoquan Gu, Lijing Ren, and Muhammad Shafiq. An interpretability security framework for intelligent decision support systems based on saliency map. International Journal of Information Security, 22(5): 1249–1260, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00689-9.

Zhang:2023:AOS

[Zha23] Yin Sheng Zhang. Analysis of OSPU security effect and data assembly verification under semi-network OS architecture. International Journal of Information Security, 22(5): 1497–1509, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00702-1.

Zhang:2022:DTA

[ZHZ22] Xiaokuan Zhang, Jihun Hamm, and Yinqian Zhang. Defeating traffic analysis via differential privacy: a case study on streaming traffic. International Journal of Information Security, 21(3):689–706, June 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00574-3.

Zouhri:2024:EIF

[ZIR24] Houssam Zouhri, Ali Idri, and Ahmed Ratnani. Evaluating the impact of filterbased feature selection in intrusion detection systems. International Journal of Information Security, 23(2):759–785, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00767-y.

Zulfiqar:2022:TAR

[ZJS22] Maryam Zulfiqar, Muhammad Umar Janjua, and Jack W. Stokes. Tracking adoption of revocation and cryptographic features in X.509 certificates. International Journal of Information Security, 21 (3):653–668, June 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00572-5.

Zhou:2023:IDA

[ZKP^{+}23] Hanxun Zhou, Longyu Kang, Hong Pan, Guo Wei, and Yong Feng. An intrusion detection approach based on incremental long short-term memory. International Journal of Information Security, 22(2):433–446, April 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        022-00632-4.

Zacharis:2024:IAD

[ZKP24] Alexandros Zacharis, Vasilios Katos, and Constantinos Patsakis. Integrating AI-driven threat intelligence and forecasting in the cyber security exercise content generation lifecycle. International Journal of Information Security, 23(4):2691–2710, August 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        024-00860-w.

Zhou:2006:P

[ZL06] Jianying Zhou and Javier Lopez. Preface. International Journal of Information Security, 5(2): 65–66, April 2006. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        006-0087-2.

Zhang:2019:FGA

[ZLGZ19] Tao Zhang, Wang Hao Lee, Mingyuan Gao, and Jianying Zhou. File Guard: automatic format-based media file sanitization. International Journal of Information Security, 18(6): 701–713, December 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00440-3.

Zhou:2020:IFV

[ZLJW20] Yuanjian Zhou, Yining Liu, Chengshun Jiang, and Shulan Wang. An improved FOO voting scheme using blockchain. International Journal of Information Security, 19(3):303–310, June 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00457-8.

Zhou:2012:MGE

[ZLL12] Jianying Zhou, Xuejia Lai, and Hui Li. Message from the Guest Editors. International Journal of Information Security, 11(5): 291–292, October 2012. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
012-0172-7;   http://
link.springer.com/content/
pdf/10.1007/s10207-012-
0172-7.pdf.

Zhang:2020:FAB

[ZLZL20] Yanting Zhang, Jianwei Liu, Zongyang Zhang, and Weiran Liu. FDCO: attribute-based fast data cloud-outsourcing scheme for mobile devices. International Journal of Information Security, 19(6): 639–656, December 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
019-00480-9.

Zheng:2007:DSL

[ZM07] Lantian Zheng and Andrew C. Myers. Dynamic security labels and static information flow control. International Journal of Information Security, 6(2–3):67–84, March 2007. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
007-0019-9.

Zia:2022:SIE

[ZMS22] Unsub Zia, Mark McCartney, and Ali Sajjad. Survey on image encryption techniques using chaotic maps in spatial, transform and spatiotemporal domains. International Journal of Information Security, 21(4):917–935, August 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.

com/article/10.1007/s10207-
022-00588-5.

Zakerzadeh:2013:DSA

[ZO13] Hessam Zakerzadeh and Sylvia L. Osborn. Delaysensitive approaches for anonymizing numerical streaming data. International Journal of Information Security, 12(5): 423–437, October 2013. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
013-0196-7.

Zacharis:2023:AAA

[ZP23] Alexandros Zacharis and Constantinos Patsakis. AiCEF: an AI-assisted cyber exercise content generation framework using named entity recognition. International Journal of Information Security, 22(5): 1333–1354, October 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00693-z.

Zuo:2024:VDB

[ZR24] Fei Zuo and Junghwan Rhee. Vulnerability discovery based on source code patch commit mining: a systematic literature review. International Journal of Information Security, 23(2): 1513–1526, April 2024. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00795-8.

Zhang:2014:DFA

[ZRJ14] Meng Zhang, Anand Raghunathan, and Niraj K. Jha. A defense framework against malware and vulnerability exploits. International Journal of Information Security, 13(5): 439–452, October 2014. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        014-0233-1.

Zyout:2023:MCA

[ZSN23] Mo’ath Zyout, Raed Shatnawi, and Hassan Najadat. Malware classification approaches utilizing binary and text encoding of permissions. International Journal of Information Security, 22 (6):1687–1712, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10207-
                                        023-00712-z.

Zhang:2022:RLA

[ZTG22] Lu Zhang, Arie Taal, and Paola Grosso. A risk-level assessment system based on the STRIDE/DREAD model for digital data marketplaces. International Journal of Information Security, 21(3):509–525, June 2022. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00566-3.

Ziza:2023:DED

[ZTV23] Kristijan Ziza, Predrag Tadić, and Pavle Vuletić. DNS exfiltration detection in the presence of adversarial attacks and modified exfiltrator behaviour. International Journal of Information Security, 22 (6):1865–1880, December 2023. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
023-00723-w.

Zhou:2015:GCR

[ZVH15] Lan Zhou, Vijay Varadharajan, and Michael Hitchens. Generic constructions for role-based encryption. International Journal of Information Security, 14(5): 417–430, October 2015. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.

                                        com/article/10.1007/s10207-
                                        014-0267-4.

Zhang:2017:CIB

[ZWQ^{+}17] Lei Zhang, Qianhong Wu, Bo Qin, Hua Deng, Jiangtao Li, Jianwei Liu, and Wenchang Shi. Certificateless and identity-based authenticated asymmetric group key agreement. International Journal of Information Security, 16(5): 559–576, October 2017. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        016-0339-8.

Zhang:2020:CBE

[ZWX20] Shufan Zhang, Lili Wang, and Hu Xiong. Chaintegrity: blockchain-enabled large-scale e-voting system with robustness and universal verifiability. International Journal of Information Security, 19 (3):323–341, June 2020. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        019-00465-8.

Zhu:2011:SLA

[ZXZ^{+}11] Wen Tao Zhu, Yang Xiang, Jianying Zhou, Robert H. Deng, and Feng Bao. Secure localization with attack detection in wireless sensor networks. International Journal of Information Security, 10 (3):155–171, June 2011. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
011-0127-4.

Zhang:2022:THG

[ZXZ22] Zheng Zhang, Shaohao Xie, and Fangguo Zhang. Topology-hiding garbled circuits without universal circuits. International Journal of Information Security, 21(2):341– 356, April 2022. CO-DEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

https://link.springer.
com/article/10.1007/s10207-
021-00556-5.

Zheng:2019:IDR

[ZZG19] Jian-Wu Zheng, Jing Zhao, and Xin-Ping Guan. Identifier discrimination: realizing selective-ID HIBE with authorized delegation and dedicated encryption privacy. International Journal of Information Security, 18 (2):141–162, April 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

http://link.springer.
com/article/10.1007/s10207-
018-0402-8.

Zhao:2008:NAP

[ZZH08] Chang-An Zhao, Fangguo Zhang, and Jiwu Huang. A note on the Ate pairing. International Journal of Information Security, 7(6): 379–382, November 2008. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        008-0054-1.

Zheng:2010:AAB

[ZZW^{+}10] Ruijuan Zheng, Mingchuan Zhang, Qingtao Wu, Shibao Sun, and Jiexin Pu. Analysis and application of Bio-Inspired Multi-Net Security Model. International Journal of Information Security, 9(1):1–17, February 2010. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/s10207-
                                        009-0091-4.